A Complete Bibliography of IEEE Transactions on Computers (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.18

Title word cross-reference

(k, K) [NC03]. (m, k) [JS09]. (S_{t}b/EC − S_{b}ED) [UF03]. (t, k) [AS03c, CCC07]. (t; k) [CCC06]. 1 [EVDN05, Tze04]. 1/2 [HT00]. 2 [AVZ08, Biq05, Cha06, LS09b, MYL^{+}01, TP02b, Wu03]. 2^{k} [FFLTM09]. 2^{m} [AQ08]. 2^{n} + 1 [EVN04a, VEN02, EVDN05]. 2^{n} − (2^{n−2} + 1) [PBB07]. 2^{n} − 1 [EVN04b, KNE^{+}00, PB07b]. 2^{n} ± 1 [EVN03]. 3 [AVZ08, CMD05, GT06, KPDS01, LA05, LLP09, NKY08, PLK^{+}03, PL09b, TK00]. 4 [KK09, LA03a]. >= 4 [FDBS05b]. ^{2} [Gho01]. 4 [OMM02]. _{p} [LJ01]. C + AB^{2} [WG00]. ηT [BBD^{+}08b]. F (2^{m}) [NWA07]. F (2^{n}) [Gol02]. GF (2^{m}) [CCL^{+}09, CHC05, PDS07, RMH04]. IF_{2} [CO09]. k [BB03a, CDLS06, IO03, KYY^{+}03, KNS01, SAOKM01]. log_{3}P [CGS07]. log_{n} P [CGS07]. M [KF04, MT06, MBCP07]. m(≥ 2) [EB09]. GF(2^{k}) [Sun04]. GF(2^{m}) [AP00, BBGM08, Cil09, DH05, EL02, FH06, GMQS02, HRM09, HW00, KWP06, LLL01, LHJL05, RMPJ08, RMH02, RHMLL08, TYT01, WG00, WWSH04, YS03]. GF(2^{n}) [ADMRK02, FD05, KOL02, SBPV07]. GF(q^{n}) [GS03]. n [BB03a, IO03, KYY^{+}03, KNS01, PR04b, SAOKM01, Tze04]. N × N [Sue09]. O(log n) [LS04c, SK04]. t [AB00, BET07, CLTH04]. t/k [FL05]. Z_{m} [BET07].

-approximating [CDLS06]. -ary [BB03a, EB09, IO03, KYY^{+}03, KNS01, MT06, SAOKM01, KF04]. -color [TP02b]. -connected [CLTH04]. -cube [IO03, KYY^{+}03]. -cubes [BB03a, KNS01, SAOKM01]. -detection [PR04b]. -diagnosability [FL05]. -diagnosable [AS03c]. -Diagnosis [CCC06, CCC07]. -EC [AB00]. -EC/AUED [AB00]. -Level [Cha06]. -out-of- [Tze04]. -track [HT00]. -Unidirectional [BET07].

1 [OBB^{+}02].

5 [BP07].

60-year [Ano06f]. 61499 [YRVS09].

754R [CHA^{+}09b].

802.11 [LLP09, LLC03]. 802.16 [NH06].

A-combined [MM04]. Abort [KLS01]. Abort-oriented [KLS01]. aborting [LKS03]. abstraction [KR01, PEP06]. Accelerating [BMR04, KDM^{+}09, LM00, ZBI^{+}07]. Acceleration [JMH02, CZM05]. Accelerator [SC06]. Access [CX07, FAL06, GMSC09, JDZ07, LCL07, MK07, SZZ09, BW03, CMD05, ICM03, MWA^{+}00, ZBC02]. Accesses [BP09, FPL^{+}08, LKS03, SE05]. Accrual [BWR^{+}07, WRJ07, CRJ06, LWRJ06, WRJL05]. Accumulate [DT05]. Accumulation [PKCD09]. accumulations [LM00]. Accumulator [NKSG09, Voy05]. Accumulator-Based [NKSG09]. accumulator-generated [Voy05]. Accuracy [MR06a, SM09, YSLH07]. Accurate [CGS07, EP09, Gra09, HSS^{+}08, Gho01, LG01, XV04]. achieve [RPH01]. achieved [PR02b]. Achieving [OS03, GNF^{+}06]. Achitectures [Ano07i]. Acknowledgment [Bal04]. acknowledgments [ABG04]. Acoustic [lFB07, HW07, SMN07]. Across [ZDS^{+}07, CK05, RSM^{+}05]. Action [MNR08]. actions [XRR^{+}02]. Active [CLE^{+}07, SMS07, KCHS04, ZC05b]. activities [LWRJ06]. Actor [LSS09]. Ad [Ano04h, WCYR07, WDY08, BB03b, CCS^{+}04, DGZA03, LWF03, TNS03, WD04]. adaptation [HSU03, IM06, Thi00]. Adaptations [BHS09]. Adapting [CU02]. Adaptive [ACPP08, BDK^{+}02, FML03, HUN07, JVG07, KSL08, LL08, LLLD06, MY07, MWK^{+}09, SKK^{+}09, THC^{+}08, TNS03, WGZ^{+}08, ZJ08, ASB03, CNM^{+}02, Hie04, KP00, KL03a, KL04, KNS01, LLS02a, RBC^{+}03, RKM05, Uht05, XLN05, GCNS08]. adaptivity [DNVG05]. Add [BDL09, LB04]. Addendum [KL04]. Adder [CS09a, Dad07, PBB07, EVN03, Hia02, NMLE00, VEN02]. Adders [CUT00, HAH08, KNE^{+}00, KS05, COP^{+}06, DN05, EVN04a, UK01, VNBE03]. adding [YLH05]. Addition [CCY00, CLV05, PBB07, PB07b, WSTJ09, ADMRK02, CV00, LM00, PGK01, SE04, YJ03]. Addition/Subtraction [CCY00]. additional [Tsu00]. Address [HAAvdG06, MKP06, OVB^{+}06, SHR08, SZ05]. Address-Value [MKP06]. Addressability [CL09]. Addressable [KPP09, MDJM05, ZIPL00]. addresses [MH01]. ADir [LJ01]. adjacent [PR00a]. Adjustment [ZV06, KH00]. Admission [WFP08, CP02, Fan03b, HH02, LMM03, Mey01]. advance [BMR04]. Advanced [AFM^{+}06, HL09, SM09, BBK^{+}03, EGP03, YW06]. Adversarial [GSS08]. AES [HV06, MAD03, SK02]. AES-Proposal [SK02]. affecting [Fav06, MFR00]. Affinity [BFP^{+}06]. Affinity-driven [BFP^{+}06]. after [MS02]. against [GSH^{+}08, ML08, MRL06, MG08, SXZ^{+}06, WHZ09, YJ00b, YKLM03]. AGAMOS [ACS^{+}09]. agent [EGP03, PS03]. agent-based [EGP03]. aggregate [AVVY05]. Aggregation [LKTT08]. aggressive [LLVA01]. Agreement [FMRR07, AK00, KPT04, Tze02a, Tze02b]. aided [KKP00, SC05]. Algebra [BS08, ZP08a, ZP08b]. Algebraic [Kap09, WT00, YF09]. Algorithm [Bol09, BDK^{+}05, DZ06, LN07, LHC^{+}08, MY07, MBCP07, PEB04, SL06, SC06, Sum08, WCKD04, WH06, Yan08, YJ06, BDBR05, CCMS02, DM00a, Gho01, HSW03, JZ05, KT05, LWRJ06, dALB03, LWG01, Low00, Mey01, MB03, MO06, OPZ00, PY05a, PC02a, RB05, Saa04, SV06, SP02, SZ05, Sun05, TK00, TYT01, TK03, Thi00, YJ03, YOC^{+}01, ZWST03, ZV02]. algorithm-based [MB03, YOC^{+}01]. algorithm-specific [BDBR05]. Algorithmic [KDM^{+}09, Has01]. Algorithmic/Architectural [KDM^{+}09]. Algorithms [Ano06e, Ano06c, Ano06d, BBD^{+}08b, BM08b, BGH07, CD09, ES00, FH06, FFLTM09, GDM07, KS08, KDM^{+}09, KB08b, LOP07, MAA^{+}08, NZ07, RS08b, UBWF08, XHLC08, XS07, Ali02, BFG01, BGJ^{+}05, CBP00, CLF05, DV04, DJM00, FML03, FFS02, Fro00, GCI06, HW00, JD06, JS06, JMH02, KI04, KYL02, LR04, Ngu05, RL04, RM06, SS06, SHK06, WT00, WWSH04, WRJL05, WK06, XTX06]. Aliasing [Had05, BDG03, LRJ00, MGZ06]. all-one [CHC05, LLL01]. all-pairs [MO06]. all-port [WT00, YW01]. all-to-all [YW01]. allocated [Tsu00]. Allocation [HUN07, SZL08, XQ08, ABP^{+}05, CCK00, CC02, DSV05, ET03, HSH01, HR02, HLLR00, KYY^{+}03, KL03b, LLC02a, LP00, MVG^{+}04, RL04, ST04, UK01, WZX05, YJMS05, YNOJ02, YD02, Zha03]. almost [CM01]. ALR [GCNS08]. Alternating [CXP06]. Alternative [VAZSR07, WH06]. alternatives [FFS02]. ALU [Voy08]. ALU-Based [Voy08]. Analysis [Ano07f, ACPP08, Ayd07, AFM^{+}06, CPRS07, CMS08, DPZ07, DTHS09, Gir06, Hey03, HW07, KL09, KSL08, Lev07, LLW07, Lo02, ML08, MKS03, MG08, PB07a, PRB09, RS08a, SKJ07, SSST06, TB06, XPGP06, XYR^{+}09, XHLC08, ZB09, Ali02, BBK^{+}03, BBB03, BB04, Car04, CWC02, CMCJ04, DNVG05, Fan02, Fan03a, Fav06, Has01, KDB^{+}05, KM03, LRJ00, LVLL06, LMS04, LP00, LLC00, LG01, LGK01, MDS02, MSM02, MKAP05, PCR01, SMBY05, SMMMK03, TB03, WPP05, XP04, XSMH04, YFKB04, YWV05, YECV02, YBB00, ZWST03, ZBS^{+}04]. Analytical [CGS07, SAOKM01, XV04]. Analytically [Gro04]. Analyzing [BF08, HKM09, DH06]. and/or [CL06]. anniversaries [Ano06f]. Annual [Ano04b, Ano05a, Ano06b, Ano08a]. Anomaly [MT02]. Anonymous [Wan04]. Antennas [ICRSR^{+}09]. antilogarithmic [AS03b]. Antisequential [BB05]. Aperiodic [SL06, ASL04, LMM00]. aperture [VPR04]. Application [HL05, Imr07, RK06, XLSF07, AAS00, BO03, BB03b, GL02, HR02, KKP00, KS00a, NAH02, SSST06, TN00]. Applications [ACV05, Ano07d, BM08a, BG08, CLE^{+}07, CCdS04, GSS08, JPSR07, KGB07, LL08, LZ07, MK07, NTA08, RC06, WHT09, XQ08, YRF08, ZSS08, AS00, BGL^{+}03, CD04, CWM01, DK04, FFS02, GL05, KGM^{+}05, LMM03, LJ05, ML00, NS03, Raj00, SLL^{+}00, Sma03, SMBY05, TSC^{+}00, XQ06, XSMH04]. Applied [KGG08]. Applying [FLM^{+}03, JLJ07]. Approach [ACS^{+}09, BF08, BKM07, CNG^{+}09, CKC^{+}08, Dad07, FH07b, HAH08, ORM07, OMG07, SKG09, WAU^{+}08, XM07, YRVS09, ZVT09, ABD^{+}04, CC02, FS00, GHP03, HAK05, HS04, HML00, KLY^{+}05, LML01, LS04a, MM04, MG02, MKBG00, PEP06, RK03, RG05]. approaches [Has01, TNS03]. approximants [CL01]. Approximate [Fio08, AK00, CSS02]. Approximating [SC07b, ZR07, CDLS06]. Approximation [POMB05, SS06, XS07, CLW^{+}03, XTX06]. Approximations [LCLV08, KDB^{+}05]. AQuA [RBC^{+}03]. AQUILA [HCC^{+}00]. Arbitrary [BNRB09, BM08c, PGVB08, WRJ07, FWCL06, Had05]. arbitrary-shaped [FWCL06]. Arbitration [ASMD07]. Arc [XWC^{+}08]. Architectural [KCHS04, KDM^{+}09, PPB^{+}07, MTB^{+}01, RVJ^{+}01, TJB03]. Architecture [ANPS07, Ano09c, AH08, Cha06, CS08a, CBC07, CRS09, CWM01, DT05, GBD07, GT06, HSLN08, KGG08, KC07a, LN07, Lee09, Moo07, MWK^{+}09, OM07, PDS07, PL09a, PEB04, PBLM08, RG09, SZGS09, SFRV09, SKK^{+}09, SH09, VPG^{+}08, WWH^{+}07, WYMG09, AP00, BDBR05, Con00, EFX^{+}04, EGP03, FZM00, GL05, IB00, KTK06, KGA01, KOL02, LPS00, LCKR03, MAD03, MG02, PLK^{+}03, PP06, RMB05, RCA01, RBC^{+}03, SSC03, Sav05, SLT01, TM04, TK03, VC02, XS02, YLH05]. Architectures [Ano06e, Ano06c, Ano06d, Ano07g, Ano07d, BSH09, DH05, FN09, GDM07, LG06, LG09, NWA07, PB07b, RLJ^{+}09, RMH06, SSA09, SC07a, ST08, SK02, TPB^{+}08, Car03, GYA^{+}03, HW00, KPDS01, LKF03, LLVA01, MRS00, MKBG00, PGJ^{+}05, PEP06, RMH04, RM06, SKS04, San06, SD06, VW05, YS03, YL06b, ZK01]. Area [BG05, CH07c, HV06, Kim09, NKY08, PKM00, SLS04, WGZ^{+}08, Hia02, PDS04, ZS00]. Area-Efficient [CH07c, NKY08, PKM00]. Area-throughput [HV06]. Area-time [SLS04]. Argument [BDL09]. Arithmetic [Ano07c, Ano07e, Ano09d, BIN06, BSH09, BBD^{+}08b, CCSK00a, CCSK00b, CHA^{+}09b, CLV05, DLM09, FFLTM09, HAH08, Kah04, KK00b, KMMS09, LM08, RC06, SBPV07, SS00, SB05, VVSA07, GMQS02, GS03, GPS05, HL05, NYC05, PS04, RZ04, Red03, SAJ02, UK01, YS01]. ARQ [ABT07, EB09, SSCG06]. arrangement [HLTH04]. Array [DK04, FS07, HE05, KOA07, AAvdG03, Alw06, FDBS05a, FDBS05b, KCKC05, LP01b, LLC02b, LCL03, SXWL04, YY01]. Arrays [BP09, EP09, FS07, JSW07, TFH07, AKS^{+}03, HPS02, HT00, Low00, PGPZ00, San03, SCZ01, Tsu00, WG00]. Arrival [WRJ07]. Articles [Ano07a]. ary [BB03a, EB09, IO03, KF04, KYY^{+}03, KNS01, MT06, MYL^{+}01, SAOKM01]. ASIC [CH07c]. ASICs [CBP00]. Assessing [EP09, WTL04]. assignment [AS00, CBP00, LSV00, dALB03, OR00, RTD00]. Assisted [NASK^{+}08]. Associative [CNG^{+}09, Sez04, Zha05]. Assurance [PH08]. assured [SHK06]. asymmetric [KF04]. Asynchronous [AM07, FMRR07, MRL06, San03, SM09, YGL07, CBPC01, Fet03, Gho01, HR02, HL02, JR02, PHA06, RL04, Sto00, TB03, TM04, WAB^{+}02]. Atlas [CWM01, SW05]. ATM [SMSM01]. atomic [XRR^{+}02]. atomicity [CMCJ04]. attached [Par04]. Attack [Lev07, PV06, GCD00]. Attacks [Gir06, GSH^{+}08, MG08, OVB^{+}06, SMS07, Has01, MDS02, SXZ^{+}06, XL03]. Attribute [FAL06]. Attribute-Based [FAL06]. Auction [PKCD07]. audit [YECV02]. AUED [AB00]. augmenting [LY02b]. authentication [Wan04]. Author [Ano00b, Ano01b, Ano02a]. Authors [Ano04i, Ano04j]. autocorrelation [KSA03]. Automata [CS09a, HB01a, VOL08, Bao04, DCCS01, OR00, TSP00]. Automata-based [HB01a, DCCS01]. Automated [CZM05, KDM^{+}09, AAS00]. Automatic [ACCL06, BM08a, CH07c, KHB02, LG01, MJ02, NR04, SLT01, VVSA07, LCCA02]. automaton [AB05]. Autonomic [APV09]. Availability [BGB08, QX08, SP07, Kum00, XL03]. Availability-Aware [QX08]. AVD [MKP06]. Average [BSM05, Ker08, KGB05]. avoid [PR06]. Aware [AZ09, BP09, GCN^{+}09, HSS^{+}08, KLCV08, LSLsK07, LJS^{+}07, MWK^{+}09, NBAR08, QX08, SL06, Xie08, ZW08, ZX07, AMMMA04, BDBR05, Cao02, KM06a, LP06a, Pal05a, PSZS04, ZZ05].

B [BBH06, LS05]. B-Cubing [BBH06]. B-tree [LS05]. Backup [ZVT09]. Bag [LZ07]. Bag-of-Tasks [LZ07]. Balanced [LHC^{+}08, NHSC07, MT06, Res01, YB03]. Balancing [HL03, ZV06, HSY00, Kum00]. Band [ZSXZ07, LLC00]. Bandwidth [OMG07, WWH^{+}07, CHH^{+}03, CYL01, HSH01, LY01, LY02b, RG05, ST04, WO01]. Bandwidth-Guaranteed [OMG07]. banked [SE05, TA05]. barrier [CHG00, MYL^{+}01]. base [VC02]. Based [ANPS07, ACS^{+}09, Ano04d, Ano05f, Ano06e, Ano06c, Ano06d, AGPP09, BP07, CS08a, CBM07, CS09b, CKS^{+}08, CMAB09, FN09, FAL06, GCN^{+}09, GDM07, Har06, Has00, HAH08, HW07, JP07, JSW^{+}06, KPJ^{+}09, KS08, KB08b, KWP06, LA05, LG06, LV07, LSBV08, LBP08, LP09, LHC^{+}08, LS09a, MAA^{+}08, Moo07, NS09, NKSG09, NH06, NTA08, PV06, PDS07, RS08a, SBPV07, SZGS09, SCK06, SC06, Voy08, VPG^{+}08, WYYZ06, WZL08, WBW08, WWC06, WH06, XCF07, XZP09, XW08, Xie08, YRF08, ZMM07, AS00, ACCL06, AFR02, AB05, AMMMA01, BK00, BB05, BO05, BD05, BCGG00, BGL^{+}03, CBT05, COP^{+}06, CK05, Che04, CH06, CP02, CN05, CD03, CNM^{+}02, CPN^{+}06, CHG00, CS00, CN03, DSV05, DCCS01, DM00a, EGP03, ET01, ET03, EKK04, Fan03a, GKM03, GHP03, GBHL06, GL02, GPS05, HKA01, HB01a]. based [HT00, HMR02, KTK06, KSL05, KMPE02, KBK03, KPGX05, KNS01, KPDS02, KM00, KH00, KYL02, KKK03, LRJ00, LLS02a, LLVC04, LMS02, Lev07, LJ01, LCC02, LCL03, LLA04, LHLL06, LJ04, MRY06, MEB01, MB03, MYL^{+}01, MG02, NR04, NYC05, PKCD07, PY05a, PC02a, PC07b, PGK01, Pie02a, PR02a, PR02c, PR04d, RMB05, RHK03, Ros04, RY05, RB05, SCZ01, SAKR03, SM00, SLZ05a, Ste02, SV06, TK03, THWH01, VD05, WB03, WFMSW00, Wu03, XXF03, XP04, XTX06, XRR^{+}02, XN06, YHIO02, YWV05, YECV02, YJ00b, YOC^{+}01, ZWST03, ZIPL00, ZCW^{+}06, PLP08]. Bases [FH07c, RMH06, DHH^{+}06, Gol02, RMH03a, RM06]. basic [HL00]. Basis [CCL^{+}09, FH06, HL09, PCH06, Wu08, GPS05, OKLC00, RMH03b, RMH04, RMH05, SK01, Sun05, TYT01, Wu02a]. Batch [CL06]. batching [AWY01]. battery [BBM^{+}03, MS03]. Baugh [VT09]. Bayesian [CLF05, KI04]. BC [FL05]. BCD [Dad07]. BCH [CS09b]. BDD [iM02, ZWST03]. BDD-based [ZWST03]. be [TMD05, YJ00b, AK09]. Beating [BSN^{+}06]. before [YJ00b]. behavior [AAvdG03, XV04]. Behavioral [LRJ00]. Benchmark [MK07, JPEJ06, KS00b]. benchmarks [YL06b]. bends [BDD00]. benefit [HR02]. Benes [Kan05]. bent [BCV01]. Berger [Pie02a]. best [LR04, SB01]. best-effort [LR04]. Better [KP00, ZR07]. between [CKS^{+}06, FH06]. beyond [RMH03a, TTA^{+}02]. BGP [DTHS09]. BICS [Lo02]. BICS-only [Lo02]. bidirectional [LPAM04]. Binary [CHA^{+}09b, Dad07, FH07b, FH07c, HKM09, JPSR07, JY00, Ker08, Kor09, MP09, MDJM05, SKG09, BSM05, DV04, EAGS01, GD03, PKM00, PGK01, RK05, SLS04]. Binary-to-Multidigit [MDJM05]. Biometric [lFB07]. Biometrics [HAD06]. BIP [BS08]. Bipartite [KT08]. bird [DSK00]. bisecting [Saa04]. Bisection [TK07, Jha03]. BIST [BO05, CP03, CH06, LLW07, LC04, NTA08, VNBE03, Voy08, VPG^{+}08, XCF07]. BIST-Based [NTA08]. Bit [Cil09, DH05, HRM09, HL09, IST06, LLL01, LV07, PCH06, RMPJ08, RHMLL08, Wu02a, Wu08, Yan08, CHC05, FD05, LHJL05, RMH04, vdGT03]. bit-oriented [vdGT03]. Bit-Parallel [Cil09, HRM09, PCH06, RMPJ08, RHMLL08, Wu08, IST06, LLL01, Wu02a, CHC05, FD05, LHJL05]. Bit-Serial [HRM09]. Bit-Width [LV07]. bits [UF03].

Block [PS06b, WBW08, XYR^{+}09, YRVS09, CNM^{+}02, CPN^{+}06, Hey03, Wan03, WBW08]. Block-Level [XYR^{+}09]. blocks [EVN03, FWC02, HML00, HL00, SAKR03, SWCC00]. Blue [MRP^{+}08]. BlueStars [PBC03]. Bluetooth [LCCA02, PBC03]. Boolean [DN06, DM00a, ASM06, ARSM07, BCK09, NASR04, SM03, WFMSW00]. Boost [MRM07, FLM^{+}03]. Boosting [ML01]. Booth [EVN04b, YJ00a]. Bottlenecks [TJB03]. Bound [BNRB09, CX07, MR01, TK07, TMD05, ASL04, BBB03, CMK03, LG01, MEB01, PR04d, SW00]. bound-based [MEB01]. Bound-set [TMD05]. Boundary [LL09, SCZ01]. Bounded [HA05, KB08a, ABG04, PCR01]. Bounded-latency [HA05]. bounded-length [ABG04]. Bounding [GL02]. Bounds [CMS08, Tal05, WRJ07, AE00, Bar04, Car02, DV04, HLM00, LSP04, Par03]. box [FLW03, LVLL06]. boxes [FWCL06]. BR [HXW09]. BR-Tree [HXW09]. Branch [KJM^{+}09, LJS^{+}07, MGZ06, MEB01, PSZS04, SL04]. breaking [ASM06]. bridge [SAKR03]. bridging [Thi00]. Broadband [NH06]. broadcast [ABP^{+}05, DPIK05, LLC02a, LLS02a, LLSC02, PC02b, RDH^{+}01, TNS03, WD04, YW01, YNOJ02]. Broadcasting [Bal04, ICRSR^{+}09, KB08a, ABG04, PS05, WT00]. broadside [XN06]. Buffer [BGH07, CBP00, JDZ07, KPJ^{+}09, LSLsK07, QD08, TIVYL09, ZJ08, CC02, JBV^{+}05, JZ05, KPEG04, SXZ^{+}06, WH03]. Buffered [Mha09, PY09, ZY06]. Buffering [SGB08, LPAM04]. Buffers [KSL08, FLM^{+}03, GVMC^{+}06, JJ06]. Build [Ano08c, Ano08b, Ano08d, Ano09b]. Building [HSU03, FN04]. Built [PR02a, SMS07, PR02c]. Built-In [SMS07, PR02a, PR02c]. Bullet [Ano08f, Ano08g, Ano09h]. Burrows [BK00, YHIO02]. Burst [AM07, UF05]. Burst-Mode [AM07]. Bus [KLCV08, SAYN09, CHG00, DV04, HHTH00, HPS02, IC02, LKF03, LP01b, MFR00]. bus-based [CHG00]. bush [DSK00]. BWT [BB05]. BWT-based [BB05]. bypass [Tsu00]. Bypassing [KS08]. byte [UF03]. Byzantine [BCdSFL09]. Byzantine-Resilient [BCdSFL09].

Cache

[BP07, BFR01, BGH07, CX07, CS08b, JD06, JDZ07, KOHC03, KS08, KGM^{+}05, LSLsK07, MKP06, MWK^{+}09, PA08, RG09, SZZ09, WHZ09, WYMG09, ZXL02, ZDS^{+}07, ZJ08, ZL07, Cao02, CPN^{+}06, DNVG05, GS00b, GSG05, HK00b, IB00, JZ05, KSL05, LY02b, LwJKW03, LJ01, MH01, ML00, NC01, SKS04, SE05, WO01, XSMH04, XV04, Zha05, ZZ05, RLPV05]. Cache-memory [BFR01]. Caches [Kim09, QD08, TIVYL09, BWTE04, KHRR02, MH01, ZZZ04]. Caching [ACMM07, SX09, XW08, ZJ08, BB03b, KV02, KRP05, KHM01, TC02]. Caching-Based [XW08]. Calculating [AQ08]. Calculation [BRC08, KLS09, ESE05, JSD01]. Calculations [DLM09, Had05]. calculi [LVS01]. calculus [LML01]. Calibration [MRM07]. Call [Ano04e, Ano04c, Ano04d, Ano05b, Ano05c, Ano05d, Ano05e, Ano06e, Ano06c, Ano06d, Ano07c, Ano07e, Ano07f, Ano07g, Ano07i, Ano07j, Ano07d, Ano07h, Ano09d, Ano09c, WFP08, Fan03b, HH02, TL02]. Call-for-Papers [Ano09d]. calls [DLBS03]. CAM [LW00, Ray06]. Can [LLJA07, PV03]. Can-Follow [LLJA07]. Canonical [CKA06]. Cantor [WPP05]. capability [HSY00, ZY06]. capacity [OR00, Ray06, Tal05]. card [BGL^{+}03, MDS02]. Cards [MG08]. Career [Ano08c, Ano08b, Ano08d, Ano09b]. Carry [CUT00, TPT06, Sav05, UK01]. carry-free [Sav05]. Carry-Lookahead [CUT00]. Carry-save [TPT06]. carry-save-adders [UK01]. Cascades [SNB07]. Cascading [ING09]. Case [BBGM08, CC07, PCG07, YD09, ESE05]. Cases [SLZ05b]. Causal [Bal04, ABG04]. Causality [CK05]. Causality-based [CK05]. CC [BW03, IB00]. CC-NUMA [BW03, IB00]. CDMA [CMD05, KL03b]. Celebrates [Ano06f]. cell [AAvdG03, AK09]. Cellular [CS09a, DCCS01, VOL08, Bao04, Lin01, TSP00, Wan04, YJMS05]. Centered [BKM07]. Centric [CNG^{+}09, ZC05b]. CeRA [EKK04]. Certification [Ano09e]. CFD [CKS^{+}08]. CFD-Based [CKS^{+}08]. chains [Li05, NAH02]. challenges [FN04, RKFTF03]. Chandra [HMR02]. changes [LO04]. Channel [EB09, KSL08, ABP^{+}05, CRJ06, CMCJ04, SM00, SLZ05a, YJMS05]. channel-based [SM00]. Channels [HY08, ABP^{+}05, LLC02a, SW00, Tal05, YNOJ02]. Chaotic [WLS09]. Characteristic [BIN06, BBD^{+}08b, MMPT08, GPS05]. characteristics [JPEJ06]. Characterization [WHZ09, BCV01, PSZS04, RVJ^{+}01]. characterizations [Biq05]. Charge [CLV05, KZP05, VAZSR07]. Charge-recovery [KZP05]. Charge-Voltage [VAZSR07]. Chebyshev [Li04]. check [YOC^{+}01]. Checker [MTHA08, ORM05, Pie02a]. checkers [Pie02b]. Checking [HU09, HCL07, OZ06, WH06, YJ00b, HU02, HU06, HCC^{+}00, MFR00, Pie02a, SXZ^{+}06]. checkpoint [LML01]. Checkpointed [EBPG06]. Checkpointing [KC07a]. Checkpoints [ABLP07]. Chief [Ano07n, Lom08a, Pra06b]. Chinese [BG08, CO09, YKLM03]. Chip [Ano04c, ACPP08, CKC^{+}08, FP09, FPL^{+}08, GE09, KSL08, LMV^{+}08, LHC^{+}08, LML06, Mar08, MRM07, PPB^{+}07, SBC08, VLP^{+}08, WHZ09, ZW08, BBR^{+}06, CC03, CWM01, GRV05, HBH05, IC02, ICM03, LP06a, MPP^{+}05, PGJ^{+}05, SW05, SZ05, TP02a, WO01, ZZZ04]. chip-level [HBH05]. chip-multiprocessor [CWM01, SW05]. Chips [Ano07f, Ano07d, HMM06]. Cholla [BHS09]. chronological [PR02b]. chunking [LLS02a]. Cipher [PS06b, Hey03]. Ciphers [BKM07, Hey03]. Circuit [GSS08, MG08, CBPC01, HSY00, PR04c, XP04]. Circuits [KLS09, Lev07, MRL06, NPB07, SM09, SB07, VVSA07, Ano00c, CDLS06, DBB00, DM00b, Fuj00b, GMQS02, GS03, KHP00, LPAM04, MSMS04, MSM04, NC03, NAH02, PR00c, PR00b, PR02a, PR02c, PR04d, PR06, SMBY05, TM00, UK01, XXF03]. Circular [ALB00]. claim [Jha03]. Class [CXP06, Cil09, DH05, PBB07, UBWF08, WB03, ZMM07, BMO04, CJ01, Fuj00b, KF04, KYL02, Pie02b, UF03, Wu02b]. Class-based [WB03]. Classes [Wu08, LHJL05]. Classification [Cha09a, FN09, LS08, LS09b, DN06, ZL04, ZCW^{+}06]. Client [WX06, KD02, KKK03]. Client-Perceived [WX06]. client-server [KD02]. client/server [KKK03]. Clock [MRM07, BG05, KL03a, KL04, LR06, MDM04, Uht05]. Clocks [Ksh07]. Closed [CMS08, LY02a, CSS02, CL01]. cluster [BBP^{+}01, DZZ00, IO03, KLY^{+}05]. Clustered [ACS^{+}09, ESE05, JBV^{+}05, GSG05]. Clustering [HWW07, Kar06]. Clusters [ZMY08, AN05, XQ06]. CMOS [AS03a, HAAvdG06, PGPZ00]. CMP [BG05]. Co. [GCN^{+}09]. coarse [KBK03, RTT05]. coarse-grain [RTT05]. Coconut [AK09]. Code [BD05, HPH^{+}04, MTHA08, BMO04, ORM05, Pie02a, RSM^{+}05, SLT01]. Codecs [CPRS07]. Codes [ABT07, ABA07, BET07, CS09b, FMRR07, FS07, Jha13, KPP09, SS09b, SZZ09, Yan08, AB00, CC03, CHL01, FDBS05a, FDBS05b, KF04, MT06, Ngu05, Pie02b, Red03, UF03, UF05, YL06a, YB03].

Coding [BCL07, HX08, KKN07, RK06, WHT09, HL05, NYC05]. Coefficients [JP07]. coemulation [CK06]. cognizance [DSK00]. coherence [KOHC03, LJ01]. Cohorts [KC07b, Jia04]. collaboration [MPAS03]. Collaborative [CRS09, MRP^{+}08, XQ08, CSR04]. Collision [ZW08]. Collision-Free [ZW08]. Collusive [LH09]. color [GL03, TP02b]. Coloring [XWC^{+}08, Tsu00]. Column [JSW07, LLW07]. Comb [NWA07]. Combination [LA03a]. combinational [Fuj00b, MSM04, Pie02b, PR04d]. combinations [Ste02]. Combinatorial [PKCD07]. Combined [NASK^{+}08, MM04]. Combining [EB09, HAD06, Car03]. command [MG02]. Comment [Bal04, CJ01, FWC02, Zac06, DN06, DM00a, Fal03]. Comments [CS08a, CLF05, FH07a, Jha13, JLZ^{+}09, KL08, KC07b, Lee12, Pie02a, WZ07, XWC^{+}08, Zha03]. Commercial [JLJ07, CP02]. commitment [KPDS02]. commodity [MBF^{+}04]. Common [GMSC09, RY05, WK03]. Communication [ABLP07, CGS07, CWC07, CS04, FSL07, GCN^{+}09, HY08, LG09, NASK^{+}08, RC06, SMN07, UBWF08, VNM07, VLP^{+}08, AF05, AE00, CK06, KPT04, MFC02, NS03, PV03, RTD00, TB03, XTX06, YFKB04, Yan02]. Communication-Aware [GCN^{+}09]. Communications [NHSC07, HSH01]. Compact [LC04]. Compacting [YJ06]. compaction [BO03, BDG03, Had05, PR00c, PR02b, PR04d]. Compactors [WBW08]. comparative [NASR04]. comparators [PKEG04b]. Comparison [ACK^{+}03, BG08, CT09, ES00, HC08a, HC08b, YT07, CLTH04, LTTH04, PKR04]. compatibility [CS00]. competitive [BC01b]. compilation [Car03, SL04]. compiled [KHB02]. Compiler [BLAA01, HL00, LP01a, LCKR03, ZL04]. compiling [DK04]. Complete [KLT07, LKTT08, CM01, PDS04].

Completeness [CC07]. completion [LZ06]. Complex [CL06, KN08]. Complexities [NS09]. Complexity [ABLP07, DKT07, DJM00, FH07b, FH07c, KPEG04, MBR^{+}09, MMPT08, RHMLL08, SFRV09, VLP^{+}08, YD09, AKS^{+}03, CHC05, EL02, Fuj00b, GMQS02, JWF01, KB03, LHJL05, RMH04, RMH05, San06, Sun04, WWSH04]. Complexity-Effective [MBR^{+}09, SFRV09, KPEG04]. compliant [EP00, NMLE00]. component [KCKC05]. component-oriented [KCKC05]. Components [PL09b, PKG06, ZWST03]. Composable [MPAS03]. Composing [BHS09]. composite [OKLC00, SSK03, Sun05]. Composition [CCC07, CCC06, JWF01, LTTH04]. compressed [ASF^{+}01, BFR01, TM05, WB03]. compressing [YHIO02]. Compression [KKN07, KGB07, RK07, TSS08, AT05, BK00, BB05, BO03, BMO04, BGJ^{+}05, CC03, HL05, JHZ01, LJ05, LW00, LCL03, LJ04, MK06, NYC05, PKR04]. Compressor [BR09]. Computation [Bol09, CY09, CCY00, HCK09, Kah04, ML08, NKY08, PB02, PEB04, RHMLL08, SG07, VNM07, Ali02, BC01a, BC01b, MR01, Mis06, RY05, SSS05, SLL^{+}00]. computation-dags [RY05]. computation-intensive [SLL^{+}00]. Computational [BC01a, DHS08, FH07c, ZVT09]. Computationally [Car02, PLP08]. computations [AHS06, Li04, MWA^{+}00, Ros04]. Computer [Ano04f, Ano04g, Ano06f, Ano06h, Ano06g, Ano07a, Ano07c, Ano07e, Ano07k, Ano07l, Ano07m, Ano09d, Ano09c, Ano09g, Ano09f, KKP00, KK00b, KMMS09, LA05, LH08, SB05, AN05, KS00b, Par04, Sma03, WS01, YLH05, YL06b]. Computers [FPFP06, HS04, KKKB05, MS03, Sma03, Pra06b]. Computing [Ano05c, Ano05d, Ano05e, Ano07a, Ano08c, Ano08d, Ano09g, APV09, BDD00, BBD^{+}08b, CLE^{+}07, CAK04, CD09, CLW^{+}03, DLL07, JP07, LL08, MBG08, Mud05, PPB^{+}07, SKG09, XT02, YF09, Zac06, ZP08b, AP00, Cao02, CS04, DZZ00, GSW02, KZP05, MRY06, OBB^{+}02, Pal05a, Ros04, RY05, SKS04, SLG02, TK00, VC02]. Concepts [LMV^{+}08]. concerning [BFG03]. concerns [RK03]. Concurrency [LwLH02, LLJA07, KM00, KLS01, KKK03]. Concurrent [AM07, BP09, BSH09, BO03, Che05, CCL^{+}09, JLZ^{+}09, JSW^{+}06, LMW07, RN04, Sez04, VPG^{+}08, CS04, LO04, Lo02]. Condition [IM06]. Conditional [KJM^{+}09, LTCH05]. conditioned [BFG01]. conference [Tze02a, Tze02b, Yan02]. conference-key [Tze02a, Tze02b]. Confidence [gLMK07]. Configurable [CH07c, HTKL08, LM00, NYC05]. Configuration [TEG09, WGZ^{+}08, CRL00, HJ01]. Configurations [ZDS^{+}07]. Configuring [PBC03]. Conflict [SE05, DSK00, KSL05, KGM^{+}05, MGZ06, XXF03]. Conflict-free [SE05]. conformance [DU04, RTD00]. conforms [NMLE00]. congestion [GHP03]. Connected [WDY08, CKBF03, CLTH04, CWC02, HS04, WLD06, YD02, Zha03]. connectedness [CU02]. connection [Mey01, XLN05]. Connections [XWC^{+}08]. Connectivity [AD08, BD05, HLTH04, YW04b, ZC05b]. connectivity-centric [ZC05b]. Connectors [BS08]. Connectors-Structuring [BS08]. Conscious [ABF^{+}07, KRCB01, KGM^{+}05, LLVA01]. Consensus [WCYR07, GR04, HL02, HMR02, IM06]. Conserve [WYZ08]. Conserving [WZL08]. considerations [CDV^{+}05, DZZ00, TP04]. Considering [DLT07]. consistency [GS00b, KLA^{+}03, LP01a]. Consistent [JDZ07, WCKD04]. Consolidated [GCN^{+}09]. Constant [PGK01, BT05]. Constant-time [PGK01]. Constants [BM08c]. Constrained [Has00, MLB^{+}09, WLS09, WZ07, BF06, LP00, LS02, PXK^{+}02, WZSP04]. constraint [LHLL06, RPH01]. Constraints [ARSM07, GCS08, gLMK07, YRF08, HL01, LWRJ06, MKS03, NASR04]. Constructing [KLT07, LCD02, LKTT08, SSK03, AB00, Sun04]. Construction [KS07, BO05, RMH02]. Consumption [ACMM07, BSPG08, GCNS08, LSLsK07, BCF^{+}03, BMO04, HAK05]. Container [GBD07]. Containment [SS09a, TTA^{+}02]. Content [KPP09, LH09, ET03, GCI06, HA05, ZIPL00]. Context [LOP07, MP03, Rho03]. Contextual [HW07]. Contiguity [BP07]. contiguous [LP00]. continuity [RPH01]. Continuous [KF01, OR00, PXK^{+}02, SAJ02]. contraction [SXWL04]. Control [AGG06, BVM07, CKC^{+}08, FAL06, GG01, HUN07, HASL07, KM07, LJS^{+}07, LLJA07, MLB^{+}09, WFP08, WCLK09, AAS00, CBS02, CMD05, Che05, CCS^{+}04, CP02, CBPC01, Fan03b, GHP03, HH02, Kan05, KM00, KLS01, KKK03, LMM03, LwLH02, Mey01, OM04, PNRP04, TA05, Uht05, ZBC02]. Control-flow [GG01]. control-optimal [Kan05]. Controllable [WCLK09]. Controlled [CLV05]. Controller [RLJ^{+}09, LRJ00, ZFP^{+}01, ZS00]. controller/datapaths [LRJ00]. Controllers [OZ06, FS00]. controlling [DKV^{+}01]. Conventional [WZL08]. Convergence [ZSS08]. Convergence-Guaranteed [ZSS08]. Conversion [HWW07, MDJM05, PGK01, SSK03, Sun05, YW04b, ZY06]. conversions [PB04]. Converter [RDH^{+}01, AS03a, AS03b, CN03]. Converting [RW08]. convolution [Ata01]. convolutional [Red03]. Cooperative [GS09, SZL08, NS03]. Coordinated [JDZ07, SX09, TC02, XRR^{+}02]. Coordinating [BHS09]. COPACOBANA [GKN^{+}08]. Coprocessors [CWZL08]. CoPTUA [WCKD04]. copy [KMPE02]. CORBA [FN04]. CORDIC [ALB00, AVZ08, LA05, Sum08]. CORDIC-Based [LA05]. Core [GBD07, LG09, KTK06, MP03, NYC05, SP02, XN06]. core-based [KTK06, XN06]. Cores [CH07c, GMSC09, RK07, CH06, HMM06]. Correct [Bol09, CM01]. Correcting [FMRR07, HX08, Kor09, KPP09, SZZ09, KF04, Red03, UF03, UF05]. Correction [Ano00c, Ano01c, Ano05f, BCL07, CCL^{+}09, HU09, JLZ^{+}09, LMW07, Sum08, COP^{+}06]. Corrections [CCSK00b, Tze02a]. Correctly [BM08b, BM08c, BMR04]. Correlation [DLT07, ML00]. Coset [XWC^{+}08]. cosimulation [CK06]. cosine [Biq05, LPS00]. Cost [BWR^{+}07, CCY00, Kah04, KLCV08, KJM^{+}09, LKTT08, LLVA01, XLSF07, XS02, BO05, CMCJ04, DZZ00, JPJ^{+}04, LJ01, LG01, ORM05, Ray06, TTA^{+}02, TP04]. cost-bound [LG01]. Cost-conscious [LLVA01]. Cost-Effective [KLCV08, XS02, BO05, DZZ00, JPJ^{+}04, LJ01]. Costs [ZVT09, CRL00, JD06, LA03b]. Coteries [KC07b, Jia04]. COTS [AFR02]. Counter [HAH08, KS08, KJM^{+}09, SSA09, GBHL06]. Counter-Based [KS08, GBHL06]. counterexample [Jha03]. counterflow [CD04]. Countermeasure [ML08]. Countermeasures [GSH^{+}08, Has01]. counters [ABG04]. counting [ACCL06, Hie04]. coupled [THWH01]. Course [Ano06h]. Cover [KLT07]. Coverage [AD08, AFM^{+}06, BF08, CKC^{+}08, KJD02, KLS09, LWF03, CIQC02, CP03, PR04c, ZC05b]. coverage- [ZC05b]. Covers [PKL09]. CPRS [LLW07]. CPU [BGB08, GCN^{+}09, RTT05]. CRC [CPR03, KB08b, SSS05]. CRCs [Ngu09].

Credentials [FAL06]. Crisis [GR07].

Critical

[HCL07, KLT07, SC06, FFS02, XQ06].

CRM [SM00]. Cross

[AFM^{+}06, WWH^{+}07, CS00]. Cross-Domain [WWH^{+}07]. cross-generation [CS00]. Cross-Product [AFM^{+}06]. Crossbar [Mha09, PY09, HKA01]. crossbar-based [HKA01]. Crosstalk [HY08, HR06, MFR00]. CRT [CN03]. CRT-based [CN03]. Cryptanalysis [Ano07j, Ano07h, Bao04, GKN^{+}08, PS06b, SGK08, RSQL03, YJ00b, YKLM03]. Crypto [HAD06]. Cryptographic [GSS08, BGL^{+}03, KP03, PS04, ST03]. cryptographically [SM03]. Cryptography [Ano04e, Ano07j, Ano07h, BBGM08, BK06, BKM07, KWP06, PV06, SGK08, GPS05]. Cryptoprocessor [SBPV07]. cryptosystem [Bao04]. Cryptosystems [Has00, SKG09, Has01, Mis06, SSST06, WPP05]. CSDA [Ano09e]. CTAM [Ray06]. Cube [PBLM08, HC06, IO03, KYY^{+}03]. cubes [BB03a, KNS01, SAOKM01]. Cubing [BBH06]. culling [CPN^{+}06]. Current [MTHA08, BBM^{+}03]. Curve [AH08, BBGM08, KWP06, LSBV08, LM08, SBPV07, Has01, Mis06, ST03, SSST06, WPP05]. Curve-Based [KWP06, SBPV07]. Curves [DJJ^{+}08, HKM09, Lee12, VOT08, ADMRK02]. Custom [CD04, LV09, CZM05]. Customized [WMZH02]. cut [HSY00]. cut-through [HSY00]. Cutting [VPR04]. Cycle [ZBC02]. Cycle-time [ZBC02]. Cycles [ABLP07, BB03a]. Cyclic [Yan08, Sch01, UF05]. Cyclo [SGB08]. Cyclo-Static [SGB08].

D [AVZ08, Biq05, GT06, KPDS01, KK09, LA05, LLP09, LS09b, MYL^{+}01, NKY08, PLK^{+}03, PL09b, TK00, Wu03]. D-Integrated [PL09b]. dags [MRY06, RY05]. Data [BR09, CX07, CKA06, CKC^{+}08, DLT07, DMS^{+}09, GE08, Gol06, HXW09, KR01, KGG08, KKN07, LLC02a, LX09, LP07, LKTT08, LLLD06, ML08, MP09, PH08, PC07a, RK06, RK07, TIVYL09, TP02b, WYMG09, XLSF07, XYR^{+}09, Xie08, XHLC08, ZBI^{+}07, ZL07, AT05, AVVY05, AFPS01, ABP^{+}05, BK00, BB05, BO03, BDK^{+}02, CC03, CBP00, DNVG05, DK04, GCI06, GSG05, HL05, HJS04, ICM03, KF04, KRP05, KGM^{+}05, KM00, KL03b, LLS02a, LKS03, LW00, LCC02, LCL03, ML00, OMM02, PCL06, PC02b, SC05, SLL^{+}00, SXWL04, SLG02, TP02a, WK03, XV04, YHIO02, YNOJ02, YOC^{+}01, Zha05, ZXL02]. data-check [YOC^{+}01]. Data-Distribution [KGG08]. data-flow [WK03]. Data-Independent [RK07]. data-intensive [DK04]. data-parallel [SLL^{+}00]. database [DSK00, JLN04, KLA^{+}03, LCCA02]. Databases [Imr07, AHS06, KD02, KLS01, KKK03, LwLH02, Lin01]. Dataflow [SGB08, KGA01, TM04]. Datapath [OZ06, PKG06]. datapaths [LRJ00, PKEG04b]. DCT [YS01]. DDoS [MRP^{+}08]. Deactivation [KG07]. Deactivation/Reactivation [KG07]. Deadline [BRC08, MAA^{+}08, BF06]. Deadline-Based [MAA^{+}08]. deadline-constrained [BF06]. Deadlines [BNRB09, XR04]. Deadlock [LCC06, LMF^{+}08, Sch01, XZP09, SW00, Wu03]. Deadlock-Free [LMF^{+}08, XZP09, Sch01, SW00, Wu03]. Deallocation [EBPG06]. Debug [HMM06, HBH05]. Debugging [YJ06, ACCL06]. decentralized [CKBF03]. Decimal [CHA^{+}09b, Dad07, EHS09, JK09, KS05, WSTJ09]. Decision [JSD01, Ker08, BSM05, GD03, KSA03]. Decoders [CL09, HAAvdG06, RS08a]. Decoding [CS09b, SFRV09, UF05, YB03].

Decomposition

[FWCL06, MR06a, LY02a, NR04, WRJL05]. decompressor [WB03]. decoupled [RCA01, SSC03]. Decoupling [VAZSR07, MH01, PG01]. defect [MP03]. defect-tolerant [MP03]. Defective [PB07a]. Defects [CMAB09, AAvdG03]. Defense [MRP^{+}08]. Deferrable [XHLC08]. Defined [MBG08, CHC05, LLL01]. definition [PRO03]. degradable [Low00]. degradation [COP^{+}06]. degree [Zha02]. Delay [HAAvdG06, HASL07, HDQK09, KLS09, SE04, SC07b, WLS09, XQ08, CT03, Fav06, FZM00, HR06, LY01, LPAM04, LS02, MFR00, TM00]. Delay-Constrained [WLS09]. delay-insensitive [LPAM04]. Delay-optimized [SE04]. delayed [LM00]. delays [AF05]. Delivery [KB08a, LP07, LH09, ET01, ET03]. Delta [MKP06]. Demand [CS08b, LMW07, AWY01, GL02]. demanding [WKS^{+}05]. denial [XL03]. Deno [CKBF03]. denormalized [SST05]. Dense [RM09, BBP^{+}01, MMRT06]. Density [FS07, EKK04]. Dependability [AFR02, ING09, MR01, MB01, YBB00]. Dependable [Ano09c, PGVB08, PBWB00, RBC^{+}03]. Dependence [DLT07, MS02]. dependencies [Sch01]. Dependent [LLWS08, MKP06, ZXL02]. Deploying [KLT07]. Deployment [MY07, XS07]. Depth [CV08, ALMN05]. Derivation [RMPJ08]. derived [DM00b]. Deriving [XR04]. Describing [PS06b]. description [PR00a]. descriptions [Gho01]. Design [Ano04c, Ano04d, AB05, CDV^{+}05, CS09a, CNM^{+}02, CH07c, GSS08, GE08, GE09, GBD07, Har06, HE05, HSS^{+}08, IB00, JSW^{+}06, KG06, KF01, LH08, LMS04, LML06, OZ06, PCG07, Pie02b, RZ04, SKJ07, SD06, SHYV06, SLZ05a, SMSM01, SMBY05, TIVYL09, WCYR07, XYR^{+}09, YFKB04, YJ00a, YJ06, ZV06, ZZZ04, ZSXZ07, ZW08, ZVT09, ZS00, ABD^{+}04, BFP^{+}06, CRL00, CP03, EVN03, EGP03, EP00, FLW03, FWCL06, FCB04, KGS00, KKP00, LP06a, LW00, LS05, MS03, MPP^{+}05, MSM02, MSM04, NC01, PGJ^{+}05, PSZS04, RB05, SK04, VEN02, WB03, WO01, XXF03, XP04, YOC^{+}01, ZP01]. designers [LKF03]. Designing [HAH08, LRB01, MRL06, YW04b, GS03, KLY^{+}05]. Designs [CKA06, WSTJ09, ZP08a, ZP08b, BO03, CT05, FLW03, GSG05, HCC^{+}00, KPEG04, PDS04, Pie02a, WWSH04, XS02]. detailed [NASR04]. detect [Fal03]. Detecting [BET07, HML00, KR04, Red03, UF03]. Detection [AM07, BSH09, BKM07, CRS09, CCL^{+}09, CMAB09, DTHS09, JSW^{+}06, Ksh07, LCC06, RMH06, SBAB00, SM09, AFPS01, BBR^{+}06, BBK^{+}03, Car04, CK05, CLF05, Che05, CNM^{+}02, CSR04, Fet03, GSA06, GRV05, KI04, Lo02, LDH06, MT02, MFR00, MB03, Ngu05, OMM02, PR02a, PR04b, PR06, RN04, SLS04, Ste02, TGKL03, YECV02, YW06]. detection-based [CNM^{+}02]. detectors [CTA02a, CTA02b, HMR02, LFA04]. Determine [CT09, WLD06]. determinism [PBWB00]. Deterministic [VNBE03, BO05, HPS02, HBH05, PY05b, ZV02]. Development [WAU^{+}08, XRR^{+}02]. device [HPH^{+}04, MG02, OPZ00]. Devices [ABF^{+}07, KPJ^{+}09, KC07a, LBP08, PLP08, KF04]. DFGs [HB01a]. DFT [XN06]. DIA [SFRV09]. Diagnosability [CLTH04, CT09, FL05, LTTH04, LTCH05]. Diagnosable [HC08a, HC08b, YT07, AS03c]. Diagnosis [Ano04e, BK06, CC07, CCC06, CCC07, CT09, HC08a, HC08b, Li05, LLW07, BO05, CCMS02, CLTH04, CM01, FML03, KC01, KP00, LTTH04, LC04, MFR00, SC04, Thi00]. diagnostic [NR04]. Diagonal [TK07, Jha03]. diagonals [CHL01]. diagram [CHH^{+}00, JSD01]. Diagrams [CKA06, HAH08, Ker08, WAU^{+}08, BSM05, GD03, KSA03]. Diameter [Mel07, XWC^{+}08]. Diarization [PAW07]. dictionaries [LC04]. Differential [MG08, PRB09, SSCG06]. Differentiating [MT09, CM02]. differentiation [WZX05]. Digit [ALMN05, JY00, Kor05, KWP06, LN07, MBCP07, PBLM08, COP^{+}06, CV00, PB04, SLS04]. digit-based [COP^{+}06]. Digit-by-Digit [MBCP07, PBLM08]. Digit-Recurrence [LN07, ALMN05]. Digital [Ano07a, Ano07k, Ano07l, Gol06]. digits [KM06b, SAJ02]. Dilated [RW08]. Dimensional [AD08, FP09, Che04, CHL01, LPS00, TSP00]. dimensioning [CT03]. Diminished [VEN02, EVDN05]. diminished- [EVDN05]. Diminished-one [VEN02]. Direct [PH08, DNVG05, GNF^{+}06]. Direct-Voting [PH08]. Directed [Par00, dOBNL09, CC03, SSC03]. Directional [CLLL09, ICRSR^{+}09]. directory [LJ01]. directory-based [LJ01]. Discharge [BBM^{+}03]. discipline [LMS02, LMS04]. Discovery [LSS09]. Discrete [FFLTM09, TN00, Biq05, KLA^{+}03, LPS00, ST04]. Discriminant [Bol09]. discriminate [BCGG00]. Disjoint [PKL09, BB03a, LCD02]. Disk [BP09, FS07, HH04, SZGS09, TB06, TFH07, THC^{+}08, WZL08, WYZ08, AKS^{+}03, DRC05, FDBS05a, FDBS05b, San03, YY01]. Disk-Based [WZL08]. Disk-Scheduling [THC^{+}08]. Disks [EP09, MKA03]. Disproving [WK03]. disseminating [BDK^{+}02]. Dissemination [PC07a, Fuj00a, SC05]. Distance [Ano06h, CXP06, GCS08, LSS09, CHH^{+}00]. Distance-Sensitive [LSS09]. Distant [PAW07]. Distributed [AM08, CLE^{+}07, CGS07, CLF05, DHS08, GSG05, GVB^{+}09, HH02, KI04, RLJ^{+}09, SKK^{+}09, VNM07, WCLK09, YF09, ZV06, ZCWZ08, ABD^{+}04, CIQC02, CHG00, GL05, Gho01, HSY00, HR02, HL02, HL01, JR02, KHM03, KGS00, LC02, LCC02, LDH06, MG02, DS00, RG05, RL04, RBC^{+}03, RPH01, Sto00, UB03, WD04, WRJL05, XL03, YJMS05, YS01, ZCW^{+}06, ZC05b]. Distributing [KHRR02]. Distribution [DHS08, JVG07, KGG08, HA05, MDM04, PvST02]. distributions [HT00]. diverse [OMM02]. Diversity [EB09, Fav06, MSM02, MSM04]. divider [JPJ^{+}04]. dividers [ALMN05, GS00a]. Divisible [JVG07]. Division [CBC07, DH05, EIM^{+}00, Kor05, LA03a, LN07, LMV^{+}08, PB02, BMR04, KT05, ML01, Par03, WG00, WWSH04, YS03]. divisor [BMR04, Par03]. DMP [YY01]. documents [PvST02]. Domain [WWH^{+}07, CZM05, SA02]. domain-specific [CZM05]. Dominating [WDY08, WLD06]. Dot [CS09a, VOL08]. Double [BR09, FS07, KGB07, LL09, ML08, PB02, VLG06, CHH^{+}00, LWG01]. Double-Data-Rate [ML08]. double-loop [CHH^{+}00, LWG01]. Double-Precision [BR09]. Double-residue [VLG06]. Downlink [ZYZ09]. DPPC [ZCW^{+}06]. DPPC-RE [ZCW^{+}06]. DPR [RL04]. Draco [MLB^{+}09]. DRAM [DKV^{+}01, GVMC^{+}06, VAZSR07]. DRAM/SRAM [GVMC^{+}06]. DRAMs [AAvdG03, CJDM01, HWI^{+}02]. drawings [BDD00]. DRES [CWZL08]. Driven [WAU^{+}08, ASL04, BFP^{+}06, CBP00, Gho01, KDB^{+}05, KHB02]. DSM [CHH^{+}03, MKBG00]. DSPs [BD05]. Dual [JSM09, SC07a, WYYZ06, ZCWZ08, Gol02, HvdG02, HHTH00, LLC00, ST03, SMBY05]. dual-bus [HHTH00]. dual-field [ST03]. Dual-Homing [WYYZ06]. dual-port [HvdG02]. dual-rail [SMBY05]. Dual-Speed [SC07a]. DualFS [PCG07]. due [AF05]. duplex [Fav06, YFKB04]. duplicated [OMM02]. duplication [PC02a]. Durations [MNR08]. during [MDM04, NAH02, PKR04]. Dwell [LKSS06, KCKC05]. Dynamic [AM08, ABF^{+}07, AN05, BRC08, Bar05, CWC07, CY09, CL07, CWZL08, CH07a, CBB^{+}02, EAGS01, HHTH00, HASL07, LZ07, MWA^{+}00, NHSC07, PKG06, WZSP04, WZ07, YY01, ZV06, ZX07, ZL07, AAvdG03, BDK^{+}02, CHG00, CC02, CS00, GBHL06, GL02, JLN04, LLS02a, LO04, LS04a, LS04c, LS04b, LS05, LKS05, MO06, PL01, PCL06, PP06, RKM05, SK04]. Dynamically [PvST02, BABD03, CWM01, MM04, QD04].

e-textile [KM06a]. earliest [YJ03].

earliest-first [YJ03]. Early

[EBPG06, Lev07, LP06a, MVG^{+}04].

EaseCAM [RMB05]. Easily

[DM00b, NC03]. Eavesdropping [KM07]. EC/AUED [AB00]. ED [OMM02]. EDAS [Gho01]. EDF [BFG03, ZB09]. Edge [BB03a, yFBC03, HCH01, WTL04]. edges [HC06]. Editor [CDE^{+}00, Gau01, JR02, KP03, LS00, Raj00, Ano01c, Ano02b, Ano07n, Gau02, Har06, Lom07, Lom08a, Pra03a, Pra03b, PL04, Pra04, PL05, Pra05, Pra06a, Pra06b]. Editor-in-Chief [Ano07n, Lom08a, Pra06b]. editorial [LT03, SLG02, UB03]. Editors [AJL02, APV09, BK06, CCdS04, GDM07, HB01b, HC01, mWHP05, KK00b, KMMS09, LML06, PL06, SB05, ST08, SGK08]. effect [HJS04, LKS03]. Effective [KLCV08, MBR^{+}09, SFRV09, BO05, CPN^{+}06, DZZ00, JPJ^{+}04, KPEG04, LJ01, PLK^{+}03, Saa04, XS02]. Effectively [HAD06]. Effectiveness [XCF07]. Effects [BBGM08, JS09, Lev07, TFCW07, BG05]. Efficiency [LF09, TTC09, CCS^{+}04, JWF01, KGMS00, PKG06]. Efficient [ASM06, AC09, Ano05f, BBH06, BCdSFL09, CBT05, CLLL09, Cha09a, CS09b, CH07b, Cil09, CH07c, DZB08, EVDN05, Fio08, GE08, GD03, HvdG02, HWI^{+}02, HX08, JLJ07, KGG08, KS07, LL09, LP07, LLC02b, LCL03, LLWS08, LS08, LS09b, LLLD06, LMF^{+}08, MLB^{+}09, MAA^{+}08, MT06, MSM04, Mud05, MDJM05, NKY08, NZ07, OKLC00, PH08, PCH06, PC07a, PLP08, PPB^{+}07, RM09, RMH03a, RM06, RSQL03, SM03, SC07a, SP07, SHR08, SZZ09, SAYN09, THC^{+}08, Tze04, VT09, WH02, WBW08, WCYR07, XV04, YT07, Yan08, YGB08, YNOJ02, YJ06, YB03, ZMY08, ZW08, dOBNL09, AGG06, BLAA01, Car02, CT05, CC02, Dat05, FDBS05a, FDBS05b, GRV05, HPH^{+}04, Hia00, JS06, KPT04, KLY^{+}05, Kum00, LLC02a, LP06b, LLC03, Low00, MO06, MK06, MKBG00, PKM00, PKEG04b, RMB05, SC05, SLS04, Ste02, SP02, SK01, SSK03, Sun05]. efficient [TA05, YL06a, YD02, YOC^{+}01, Zha03, ZS00]. Efficiently [MKP06]. effort [LR04]. EFSM [DU04]. EFSMs [MNR08]. EKMR [LCL03]. Elastic [BLCA02, CHL09]. election [Sto00]. electronic [SMMMK03]. Elementary [ELMT00, NS09, NKY08, Li04]. Elements [MMPT08, PKL09, BC01a, GMQS02]. Eliminating [KSL05, Lin01]. Elimination [KN08, KMPE02]. Elliptic [AH08, BBGM08, HKM09, LSBV08, LM08, ADMRK02, Mis06, ST03, SSST06]. Elliptic-Curve-Based [LSBV08]. Embedded [AM08, Ano07g, Ano07i, BBGM08, GSS08, KN08, KOA07, LBP08, LV09, PC07b, PPB^{+}07, RK07, SHYV06, SKK^{+}09, ST08, WWC06, XQ08, AAvdG03, BCF^{+}03, BMO04, CH06, CT05, CD04, DK04, FFS02, HWI^{+}02, JBV^{+}05, KGS00, KP03, KPGX05, KGM^{+}05, MT02, MP03, ORM05, PP06, Pie02b, PEP06, SXZ^{+}06, SWP04, TP02a, WB03, XRR^{+}02]. Embedding [NKSG09, HC06, Voy05]. Emergent [Ano06e, Ano06c, Ano06d, GDM07]. Emerging [Ano07d]. Emotional [HWW07]. employing [PP06]. Emulation [BM08b]. En-Route [SX09, TC02]. Enable [XCF07]. enabled [RCA01]. Enabling [OM07, DU04]. Encoded [HY08, YJ00a, KHRR02]. Encoding [CWZL08, CS09b, CHA^{+}09b, SAYN09, Yan08, YB03, ZCW^{+}06]. Encryption [RK06, YW06, BBK^{+}03, LJ05, YGZ05]. End [HASL07, WX06, ASB03, HL01, OM04, RCA01, San06]. end-host [OM04]. end-systems [ASB03]. End-to-End [HASL07, WX06, HL01]. Energy [ACMM07, AZ09, BSPG08, CLLL09, CS08b, GCNS08, ICRSR^{+}09, KYS05, LLLD06, Mud05, Pal05a, PC07a, PPB^{+}07, PKEG04b, RM09, SAYN09, TTC09, WZL08, WYZ08, WRJ07, XQ08, Xie08, ZMY08, ZX07, AGG06, BCF^{+}03, CDV^{+}05, CCS^{+}04, CT05, Dat05, HAK05, HPH^{+}04, JBV^{+}05, KM06a, KLY^{+}05, KGMS00, PKEG04a, PKG06, RMB05, SC05, TA05, VKI^{+}03]. Energy-Aware [Xie08, ZX07, KM06a]. Energy-Delay [XQ08]. Energy-Efficient [PC07a, PPB^{+}07, ZMY08, AGG06, Dat05, HPH^{+}04, KLY^{+}05, SC05, TA05]. engine [Har04]. Engineering [OMG07]. Engines [RG09]. Enhanced [IO03, LSP04, LKTT08, LS04b, Sue09, KMPE02]. Enhancement [Ano04h, KK00a, Uht05]. enhancements [DGZA03, TJB03]. Enhancing [CK06, SM09, ZYZ09, ZC05a]. Enlarging [DVJP07]. enough [YJ00b]. entries [Ata01]. entry [KF04]. Enumeration [PR02b]. Envelope [LKSS06]. Environment [FP09, BJHW00, CN05, LLS02a, LAG^{+}01]. environments [CKBF03, CJDM01, LLSC02, TP04, ZXL02]. EPIC [HJS04]. eQoS [WX06]. Equally [Gol02, LLL01]. equations [BFG01, FDZ03]. equi [RPH01]. equi-continuity [RPH01]. equipped [HSY00]. equivalence [HCC^{+}00]. eRAID [WZL08]. Erasures [CPRS07]. Error [AM07, BCL07, BSH09, BBK^{+}03, BET07, CCL^{+}09, FMRR07, JSW^{+}06, Kim09, KPP09, LMW07, Ngu05, PB07a, SZZ09, YJ06, AFPS01, COP^{+}06, KF04, LVLL06, Lo02, MB03, OMM02, RN04, RB05, Ste02, TTA^{+}02, UF03, UF05, YW06]. Error-Correcting [FMRR07, KPP09, SZZ09]. Error-Detecting [BET07]. Error-detection [Ngu05]. Error-Protecting [Kim09]. Errors [CPRS07, WHZ09, Che05, DH06, HJS04, MBF^{+}04]. establishment [CRJ06, HHTH00, XLN05]. Estimating [PB07a, Ata01]. Estimation [KGG08, MP09, KJD02, MSM04, VKI^{+}03, YWV05]. Ethernet [GCNS08]. Euclidean [TK00]. European [CCSK00b, CCSK00a, CSK^{+}08]. evaluate [MB01]. Evaluating [TSD01, VKI^{+}03, FDZ03]. Evaluation [CCMS02, FFP07, FP09, GSH^{+}08, KS00a, LV07, LV09, WCYR07, ZV06, ASF^{+}01, CHH^{+}03, CM02, CNM^{+}02, EGP03, EKK04, HA05, IB00, KGA01, KF01, KS00b, KL03b, LGML05, PGJ^{+}05, PRO03, RK03, SMSM01, XS02]. even [Wu03]. Event [MP09, CLF05, Gho01, GL03, KI04, KHB02]. Evolutionary [NPB07]. Exact [Ayd07, DZB08, KLS09, LLWS08, SDB03, EKK04, KDB^{+}05]. Examining [MDS02]. examples [LGK01]. exclusion [LWRJ06]. executing [LJ05]. Execution [Ano05f, CBS02, CBT05, ESE05, Gho01, KGA01, MKAP05, PS03, TSC^{+}00]. execution-time [ESE05]. Exercise [MRP^{+}08]. Existence [Imr07]. Existing [HL09]. EXOR [Ano00c, KHP00]. EXOR-Sum-of-Products [KHP00, Ano00c]. Expansion [CKA06, PR02a]. Expected [RK06]. experience [RKFTF03]. Experiences [FN04]. experimental [CWC02, RSQL03]. experiments [Ste02]. explicit [WPP05]. Exploiting [AK00, BSPG08, LX09, SP07, SL04, WYZ08, YSL00, MGT04]. Exploration [GE08, GE09, JS09, SGB08, BFP^{+}06, FFS02, LP06a]. Exploring [PPB^{+}07, PEP06]. explosion [KKKB05]. Exponential [PEB04]. Exponentiation [DKT07, Gra09, BP01, Har04]. Exponentiations [CL06]. Exponents [CL06]. Expression [Lee09]. Expressions [JSM09, SB01]. Extended [FH07b, FH07c, JSM09, WLD06, Mis06, WDL^{+}03, PCC02]. Extending [HL00, MNR08]. extensible [BJHW00]. Extensions [BBGM08, XWC^{+}08, BD05, TJB03]. Extra [Kah04, FB00]. Extra-Precise [Kah04]. Extraction [MBCP07, OS07]. Extreme [Imr07]. extremes [KJD02].

Factor [Sum08, AWY01]. Failure

[DLT07, Red03, TFH07, CAK04, CTA02a, CTA02b, Fet03, HMR02, LFA04]. Failure-detecting [Red03]. Failure-Tolerant [TFH07]. Failures [FS07, HX08, ING09, AN05, FDBS05a, FDBS05b, KJD02]. Fair [HUN07, MS09, NB02, RG05, ST04, YD09, LMS02]. Fairness [YD09]. Fallacies [GMSC09]. Families [SMBS06]. Family [SS00, HMR02]. Fast [Cha09a, EVN04a, FD05, HL02, JLJ07, LR04, LM08, Ngu09, PBB07, PB07b, PF08, RMH03b, YGL07, DJM00, EL02, GVMC^{+}06, Ngu05, TYT01, YJ03, YD02, Zha03]. Fault [Ano04e, Ayd07, BK06, BKM07, CWC07, COP^{+}06, CL09, CSR04, CKC^{+}08, GS00a, Gir06, HY08, HT00, HLTH04, KL08, KLS09, LL08, Lev07, ML08, MP09, PV06, PS03, RMH06, Tsu00, XZP09, YT07, Zha02, ZVT09, ACK^{+}03, BO05, BBR^{+}06, Car02, CP03, CLF05, Con00, Dat05, FN04, GNF^{+}06, Gro04, HSW03, HS04, HCH01, IO03, KKP00, KS00a, KI04, LRJ00, dALB03, LWG01, LC04, LDH06, MME04, MKBG00, NR04, PR02a, PR04c, PGPZ00, RKFTF03, SAKR03, Ste02, TGKL03, Tze02a, Tze02b, Wan03, WDL^{+}03, Wu03, XRR^{+}02, YW04a, YJMS05, YJ00b, YKLM03, YOC^{+}01]. Fault-based [Lev07, YJ00b]. fault-free [HCH01]. Fault-Sensitive [Ayd07].

Fault-Tolerant [CWC07, CL09, HY08, KL08, XZP09, ZVT09, GS00a, HT00, PS03, Tsu00, Zha02, Car02, CLF05, Dat05, FN04, Gro04, HSW03, HS04, IO03, KKP00, KI04, dALB03, LWG01, LDH06, MKBG00, RKFTF03, Tze02a, Tze02b, Wan03, WDL^{+}03, Wu03, XRR^{+}02, YW04a, YJMS05]. Faults [GSS08, HAAvdG06, MRL06, ORM07, TFCW07, UBWF08, AK00, BCGG00, FB00, Fav06, HvdG02, HSW03, HCH01, KHM03, Li05, LMM00, MFR00, MDM04, PR06, TM00, Thi00]. Faulty [PKL09, HC06, HML00]. FCSR [AB05]. FDR [CC03]. Feasibility [Ayd07, HA05]. Feasible [WCLK09, DU04, LS02, ST04]. Feature [ZMM07]. Features [lFB07]. FEC [SSCG06]. feedback [Hey03]. Feistel [JSW^{+}06]. FFT [WKJ07]. fiber [SLZ05a]. Fibonacci [SMSM01]. Field [BSH09, CCdS04, FH07c, FHLM04, Has00, NWA07, NWA08, RMH06, IST06, KB03, RM06, ST03, SSK03, TP04, Wu02a, WHBG02]. Fields [BIN06, FH07b, HCK09, HKM09, JP07, LM08, MMPT08, PDS07, SKG09, Wu08, BS04, GMQS02, LLL01, OKLC00, Sun05, Wu02b, ZV02]. FIFO [PY05a]. FIFO-based [PY05a]. File [BSPG08, LSV00, LBP08, LP07, PCG07, WMZH02, CCK00, LLS02b, LP06b, NC01, TA05, WH03]. Files [EBPG06, TM05]. filter [GCD00, MRT01]. filtered [AB05]. Filtering [KGMS00, ZL07]. Finding [Imr07, LS02, PR00a]. Fine [LVS01, Pal05b]. fine-grain [Pal05b]. Fine-grained [LVS01]. Fingerprint [HBF09]. Finite [BIN06, BSH09, HCK09, Has00, JP07, MMPT08, NWA07, NWA08, WHBG02, Wu08, GMQS02, Hie04, IST06, KB03, KM06b, PR00a, TL02, Wu02a, Wu02b, ZV02]. Finite-Field [BSH09, NWA08]. finite-state [PR00a]. FIR [MRT01]. Firm [LHLL06, DSK00]. first [YJ03]. FITS [CN05]. Five [FH07a, Mon05]. Fixed [BNRB09, DZB08, VT09, Bar04, BB04, DN06, dALB03, OPZ00, XT02, Zac06]. Fixed-Priority [BNRB09, Bar04, dALB03]. fixed-size [OPZ00]. Fixed-Width [VT09]. Flash [HTKL08, KPJ^{+}09, LBP08, LP06b]. Flash-Memory [HTKL08]. flat [ABP^{+}05]. flea [BSN^{+}06]. flea-flicker [BSN^{+}06]. Flexible [CMAB09, LM08, BJHW00, BLCA02, FZM00, HT00, PBWB00]. flicker [BSN^{+}06]. flight [AAS00]. Floating [ACV05, BR09, CHA^{+}09b, EHS09, ES00, Gra09, Kah04, KK09, LB04, TLS09, WSTJ09, XPGP06, BMR04, EP00, LM00, NMLE00, SE04, VLG06, YSL00]. Floating-Point [ACV05, BR09, CHA^{+}09b, EHS09, ES00, Gra09, Kah04, KK09, TLS09, WSTJ09, XPGP06, LB04, BMR04, LM00, NMLE00, SE04, VLG06, YSL00]. flooding [SC05]. floorplanning [KK00a]. Flow [CKA06, GE08, LJS^{+}07, Che05, GG01, WK03, XS02]. fly [Fro00]. FMA [BM08b]. folded [LCD02]. Follow [LLJA07, LCCA02]. follow-me [LCCA02]. Forest [XLSF07]. Form [CH07b, DN06]. Formal [ASMD07, SG07]. Formally [BDL09, Bol09]. Format [CHA^{+}09b, PGK01]. formation [PBC03]. forms [Fal03]. formula [ADMRK02]. Formulae [FH07a, Mon05, WPP05]. Formulas [CO09]. formulations [NASR04]. Forwarding [ANPS07, CS08a, RG09, NMLE00]. foundations [BO05, WT00]. Four [JSW07, MYL^{+}01, JS06]. Four-ary [MYL^{+}01]. Four-Port [JSW07, JS06]. Fourier [ZR07]. FPC [BR09]. FPGA [BM08a, FLW03, HML00, MKS03, NASR04, PDS04, PC07b, RB05, TM04]. FPGA-based [PC07b]. FPGAs [EKK04, RSQL03, SV06, SPS08]. FPU [SST05]. fractional [SSST06]. Fragmented [BP07]. Frame [LS09a, SCK06]. Frame-Based [LS09a, SCK06]. Framework [BHS09, CMAB09, KDM^{+}09, MBG08, PS06b, RG09, SZL08, THC^{+}08, CT05, CSS02, KS00b, MTB^{+}01, PL01, VKI^{+}03]. Free [ABF^{+}07, CRJ07, LMF^{+}08, XZP09, ZW08, HCH01, JJ06, PHA06, Sav05, Sch01, SE05, SW00, Wu03, ZC05a]. Free-Space [ABF^{+}07]. frequencies [CAK04]. frequency [CC03, Uht05]. frequency-directed [CC03]. Frequent [GRV05]. frequently [LCK^{+}01]. Freshness [XHLC08]. friendly [JZ05, WH02]. front [RCA01, San06]. front-end [RCA01, San06]. FSM [PY05b]. full [BI04, LJ01, PR04d, PR06, YW04b]. full-scan [PR04d]. Fully [KLT07, Sue09, Alw06, HV06, PR00b, Zha05]. Function [AK09, HWW07, LV07, LV09, OVB^{+}06, POMB05, SNB07, SLZ05b, YRVS09, DM00a, LGML05, Li04, WRJL05]. Functional [FFP07, WWC06, Car03, PR00a, XP04]. Functions [BWR^{+}07, ELMT00, JPSR07, NS09, NKY08, BCV01, DN06, KSA03, MGZ06, SM03, VD05, WFMSW00]. Fused [BDL09, LB04]. Fusion [LLLD06, PH08]. Future [SBC08]. Fuzzy [ACV05].

G [CMD05]. GAARP [BDBR05]. gain [TP04]. gains [HL03]. Galbraith [HKM09]. Galois [PDS07, SSA09]. GALS [BDBR05, HBH05]. Game [SZL08, HR06, RG05]. game-theoretic [RG05]. Games [GR07]. gate [HR06]. gateways [WTL04]. Gaussian [CCL^{+}09, DHH^{+}06, LLVC04, LVLL06, MBS^{+}08, RM06]. Gbits [HV06]. Gbits/s [HV06]. General [FLW03, Fan02, HL09, LP09, HK00a, JLN04, LJ05, RMH03b, TL02, YSL00, ZP01]. General-Purpose [HL09, LP09]. generalization [AS03c]. Generalized [CHL09, Fal03, GCD00, JSW^{+}06, YJ03, LMS04, Pie02a, Sun04]. Generated [AK09, Voy05]. generating [PR06]. Generation [BBD^{+}08a, BM08a, CWC07, FFJ^{+}06, Gol06, HW07, KB08b, NTA08, Par00, WWC06, XPGP06, Alw06, AGG06, CZM05, CS00, DU04, FFS02, Fuj00b, FS00, NR04, PR00a, PR02a, PR02c, PGPZ00, TSP00]. Generations [SBC08]. Generator [SMS07, AB05, CP03, DCCS01, LLVC04, LVLL06]. Generators [SNB07, GCD00, MMRT06, SGB00]. Generic [FWC02, SWCC00, WD04]. Genetic [HAK05, SC06, SHK06]. Geometric [BF08, CMS08, HSS^{+}08, FCB04]. geometrical [KPDS01]. Geometry [LA05, LLP09, YSL00]. given [PR04c, Sun05]. Global [LR06, YBB00]. Goldschmidt [EIM^{+}00, GS00a]. gossip [GKM03]. gossip-based [GKM03]. GPRS [CD03]. graceful [COP^{+}06]. grain [Pal05b, RTT05]. grained [LVS01]. granularities [KBK03]. Granularity [JS09, Pal05b]. Graph [ACS^{+}09, CBM07, DM00a, JP07, NS09, Fuj00a, SA02, Tsu00]. Graph-Based [ACS^{+}09, JP07, NS09, DM00a]. Graph-Node [Tsu00]. Graphics [LA05, NKY08]. Graphs [CC07, CY09, KL08, SGB08, XWC^{+}08, BCV01, FL05, HCH01, HLTH04, Saa04, WK03]. Gray [ABA07, Jha13]. Grid [CIQC02, DLT07, KLT07, WWH^{+}07, SHK06, DPZ07]. Grid-Oriented [WWH^{+}07]. Grids [KLT07, LZ07, SZL08, ZVT09, CM01]. Group [CWC07, CXP06, KPT04, SH09, LS02, MFC02, WAB^{+}02, Yan02]. Guarantee [MRM07, THC^{+}08, ZX07]. Guaranteed [KB08a, LMV^{+}08, OMG07, ZSS08, LMS02, MKA03]. Guarantees [WX06, FZM00, KBK03, THW03]. guard [KM06b]. guarded [TTA^{+}02]. Guest [JR02, KP03, LT03, LS00, Raj00, SLG02, UB03, AJL02, APV09, BK06, CCdS04, GDM07, Har06, HB01b, HC01, mWHP05, KK00b, KMMS09, LML06, SB05, ST08, SGK08]. Guided [SZL08]. Guidelines [RY05].

H.323 [DLBS03]. halving [FHLM04]. Hamiltonian [BB03a, HC06, HLTH04]. hamiltonicity [HLTH04]. hand [DSK00]. Handheld [NKY08]. Handling [CBS02, GL03, JJ06]. Handshake [CKC^{+}08]. hard [BBL01, BWTE04, CBS02, KL03a, KL04, LMM00, dALB03, PBWB00, WKS^{+}05]. hard-real-time [WKS^{+}05]. Hardening [ORM07]. Hardness [SC07b]. Hardware [ASF^{+}01, Ano07j, Ano07h, CCY00, CMAB09, DKT07, DKV^{+}01, DJJ^{+}08, GPS05, KJM^{+}09, LCLV08, Lee09, Lee12, MKP06, OVB^{+}06, PC07b, SSA09, SGK08, WSTJ09, ZBI^{+}07, ZW08, ZP08a, BBK^{+}03, BP01, BT05, GRV05, JMH02, KT05, KOL02, KP03, KNS01, LwJKW03, LLVC04, LGML05, LVLL06, LRB01, LM00, MAD03, MRS00, OPZ00, OBB^{+}02, PL01, RB05, SXZ^{+}06, SdBF04, SL04, VKI^{+}03, VLG06, YWV05, YKLM03, YW06]. hardware-algorithm [OPZ00]. hardware-based [LLVC04]. Hardware-Efficient [ZW08]. hardware-software [VKI^{+}03]. hardware/software [SXZ^{+}06]. Harley [WPP05]. Harnessing [FFJ^{+}06]. hash [VD05]. hashing [KYS05, JMH02]. Hazard [SM09]. Healing [CD09]. Heterogeneous [QX08, ZMY08, BBP^{+}01, BD05, BFP^{+}06, CS04, HH02, MPP^{+}05, SFJ03]. heterogeneous-connectivity-based [BD05]. Heuristic [CBM07, PS06a]. heuristics [SHK06]. Hidden [FAL06]. Hiding [LP01a, TP02b]. Hierarchical [DPZ07, GMSC09, HJ01, KC01, RKM05, SB07, WYMG09, CT03, FS00, HKA01]. hierarchy [BABD03, DZZ00, FDZ03, LRB01]. High [ANPS07, AH08, ALB00, BP01, BR09, CS08a, CD09, CKC^{+}08, CJDM01, DN05, Hia02, HE05, HDQK09, KNE^{+}00, KG06, KS05, KB08b, LL08, LA05, LG06, Lee09, LLP09, MAA^{+}08, Mel07, MMRT06, PB02, POMB05, PL09b, SSA09, SH09, SPS08, WWH^{+}07, Wil01, WO01, WWSH04, XPGP06, XWC^{+}08, YJ00a, ZMM07, ZP08a, BGL^{+}03, CD04, CYL01, HKL01, Kum00, LY02b, LwJKW03, LG01, Lu05, MM04, ML01, MRS00, ORM05, Par03, Ray06, SKS04, San06, TSP00, TGKL03, WMZH02, XS02, ZK01]. High-Bandwidth [WWH^{+}07, CYL01, LY02b]. High-Fault-Coverage [CKC^{+}08]. High-Level [ZMM07, LG01, MM04]. high-order [Lu05]. High-Performance [AH08, HE05, KB08b, LL08, Lee09, MAA^{+}08, Mel07, PL09b, SSA09, SH09, SPS08, XWC^{+}08, ZP08a, CJDM01, CD04, SKS04, TGKL03, ZK01]. High-Quality [LLP09, TSP00]. High-radix [BP01, Par03]. High-Speed [ANPS07, BR09, CS08a, KNE^{+}00, KG06, KS05, PB02, POMB05, XPGP06, YJ00a, DN05, Hia02, WWSH04, BGL^{+}03, MRS00, San06, XS02]. High-Throughput [LA05, Ray06]. Higher [ZR07, SMM05]. Highly [CWC07, CD09, EP09, MAD03]. hints [ZL04]. Hoc [Ano04h, WCYR07, WDY08, BB03b, CCS^{+}04, DGZA03, LWF03, TNS03, WD04]. hold [Li05]. hold-time [Li05]. holding [TL02]. holistic [KLY^{+}05]. Homing [WYYZ06]. Homogeneous [PL09a]. Horizontal [LF09]. host [OM04, YECV02]. host-based [YECV02]. Hosting [GCN^{+}09]. hot [SAOKM01]. hot-spot [SAOKM01]. Huffman [KKN07]. Human [Ano06e, Ano06c, Ano06d, GDM07, Moo07]. Human-Inspired [Moo07]. Human-Machine [Ano06e, Ano06c, Ano06d, GDM07, Moo07]. HW [CK06]. HW/SW [CK06]. Hybrid [BZ02, LBP08, WKJ07, ZP08b, BBR^{+}06, Che05]. hyperbolic [BBB03]. hypercube [CWC02]. hypercubes [BB03a, KP00, LKF03, LCD02, PKM00, SC04]. hyperelliptic [WPP05]. hypergraphs [Saa04].

I-Cache [CS08b]. I/O [KV02, LSV00, OM04, SD06, TP02a]. IC [BGL^{+}03, SAKR03]. identical [Bar04, KC01, SCZ01]. Identification [YT07]. Identifying [HBF09, Ste02, CHL01]. IEC [YRVS09]. IEEE [Ano04f, Ano04g, Ano04h, Ano06f, Ano06h, Ano06g, Ano07a, Ano07k, Ano07l, Ano07m, Ano09g, Ano09f, CHA^{+}09b, DGZA03, ES00, EP00, LLP09, LLC03, NMLE00, NH06, Pra06b, Raj00, SE04]. II [BC01b, FDBS05b, SK01]. ILP [CN05, LY01, RLJ^{+}09]. ILP-based [CN05]. Image [KOA07, WWH^{+}07, MKS03]. images [TP02b]. immune [YKLM03]. Immunet [PGVB08]. Impact [BGH07, HY08, LG09, KBK03, MS03, MKAP05]. implement [LJ01]. Implementation [ACPP08, CHA^{+}09b, DT05, DJJ^{+}08, Gir06, LCLV08, Lee12, LMV^{+}08, SKJ07, YRVS09, ZSXZ07, AS03a, AS03b, BI04, BBK^{+}03, Che04, CNM^{+}02, EGP03, HW00, LFA04, MKS03, Ngu05, PRO03, SM03, SE04, SLZ05a, YW06, YS01, ZS00]. implementations [BO05, RSQL03, SST05, SK02, VLG06]. implemented [ACK^{+}03, TGKL03]. Implementing [EFX^{+}04, LCCA02, TPB^{+}08, MKBG00]. Implications [LJ05, MDM04, MK07, RVJ^{+}01]. importance [CAK04]. imprecise [AHS06]. Improve [FFJ^{+}06, LX09, LF09, SAT09, XCF07, JZ05, PR02b, WH03, Zha05]. Improved [ACMM07, CO09, HCK09, NBAR08, NKSG09, LMM03]. Improvement [DZ06]. improves [ML00]. Improving [EIM^{+}00, JK09, LJS^{+}07, MR06a, MH01, PG01, SCG08, SXWL04, TTC09, XLN05, YGZ05, YLH05, WTL04]. Impulse [ZFP^{+}01]. In-line [JJ06]. In-Memory [WYZ08]. in-order [BSN^{+}06]. in-transit [FLM^{+}03]. Incompleteness [FFP07]. inconsistent [Gho01]. Incorporates [KC07a]. Incorporating [KK00a, SMN07]. increase [KGMS00, LLVA01]. increasing [PR02b]. Incremental [LOP07, ACCL06, MAMMA03, SSS05]. Independency [BP09]. Independent [PY09, RK07]. independently [Had05]. Index [Ano00b, Ano00d, Ano01b, Ano01d, Ano02a, Ano02c, Ano04b, Ano05a, Ano06b, Ano08a]. indexed [MH01]. indexing [KSL05, MGZ06, PXK^{+}02]. Indirect [KJM^{+}09]. Induced [TFCW07]. indulgent [GR04]. ineffectual [KR04]. Inexpensive [EP09]. InfiniBand [ASMD07]. Information [Ano04f, Ano04g, Ano04i, Ano04j, HWW07, PAW07, Rya04, ET01, Fuj00a, GL05, GR04, Rho03, XP04]. infrastructure [XN06]. inherent [EL02, JPEJ06]. Inherently [ZK01]. Initial [LLWS08, PR00b]. Injection [MRL06, ACK^{+}03, Ste02]. Inner [KK09]. Input [GS09, LS09a, VPG^{+}08, PR02a, PR04d]. Input-Queued [GS09, LS09a]. Inputs [ZDS^{+}07]. Insensitive [Kap09, SMBS06, LPAM04]. inspection [KKKB05]. Inspired [Moo07]. Instability [DTHS09]. Instruction [AF05, BBGM08, CH07a, DVJP07, LF09, TW08, BD05, CT05, CZM05, GYA^{+}03, LCKR03, SS02, WB03]. Instruction-Set [BBGM08]. instructions [KR04, OMM02, YSL00]. Integer [FFLTM09, GSA06, SBAB00, LM00]. Integers [CH07b, MBS^{+}08, RK05, RW08]. Integrated [AD08, JSW07, KOA07, PL09b, SKK^{+}09, ZBS^{+}04, BP06, CN05, SLL^{+}00, VKI^{+}03]. Integration [LX09, PL09a, SXZ^{+}06]. intelligent [KOHC03, LwJKW03, SLT01]. intensive [DK04, SLL^{+}00]. inter [AF05]. inter-PE [AF05]. Interaction [Ano06e, Ano06c, Ano06d, BS08, GDM07, Moo07]. Interactions [OZ06]. InTeRail [KTK06]. interconnect [HR06, PGJ^{+}05]. Interconnection [KL08, KSL08, PGVB08, FB00, LLPC04, MOK04]. Interconnections [Mel07, XWC^{+}08]. Interconnects [BCL07, CKC^{+}08, FML03, KLY^{+}05, YW04b, ZY06]. Interface [OM07, KRP05]. interfaces [BFR01]. interleaved [LC02]. intermittent [BCGG00, KHM03]. Internet [MG02, CM02, CKDS02, EGP03, ET03, LT03, MRY06, MPAS03, NB02, Ros04, RY05, WZX05, WDL^{+}03]. Internet-based [MRY06, Ros04, RY05]. Internode [CXP06]. interplay [MME04]. interpolation [Li04, LGK01, ZV02]. Interpolations [LCLV08]. Interpolator [POMB05]. interrupt [JJ06]. Interval [DLM09, SS00, YRF08, LS04b, LKS05, TN00]. Interval-Based [YRF08]. interval-partitioned [LKS05]. Introducing [Pra06b]. Introduction [APV09, BK06, CCdS04, DLL07, GDM07, HKL01, Har06, mWHP05, KK00b, KMMS09, LML06, Mar08, Mud05, SB05, ST08, SGK08, AJL02, HB01b, JR02, KP03, LS00, Raj00]. intrusion [YECV02]. invalidation [ZXL02]. invariance [CS00]. invariant [MT06]. Inverse [DZ06, ELMT00, NZ07, PB02, SK00, LPS00]. Inverses [AQ08]. Inversion [DH05, FHLM04, GCD00, Sav05, TYT01, WG00, WAB^{+}02, YS03]. Involution [JSW^{+}06]. Involution-Based [JSW^{+}06]. Involutional [JSW^{+}06]. involvement [MYL^{+}01]. IP [CD03, LA03b, LS04b, LKS05, RMB05, SAKR03, SHR08, SZ05]. IP-based [CD03]. IRIS [CJ01]. Irreducible [Cil09, PCH06, HK00a, IST06, RHK03, ZP01]. irredundant [SB01]. Irregular [LHC^{+}08, AN05, CWM01]. iSCSI [WYMG09]. Isolating [PKEG04a]. Isomorph [DBB00].

Isomorph-redundancy [DBB00]. Isomorphic [KYY^{+}03]. Issue [Ano04c, Ano04d, Ano05c, Ano05d, Ano05e, Ano06e, Ano07c, KK00b, SB05, Ano06c, Ano06d, CYL01, GYA^{+}03, LY01, UB03]. Issues [LCL07, CKDS02, MDM04]. iteration [KRCB01]. iterations [SC04]. Iterative [TLS09, WDY08, PGPZ00, Red03, WWSH04].

Java [ISF06, RVJ^{+}01]. jitter [DS00, PP06]. Job [SZL08, SHK06]. Join [Ano06h, Ano06g, Ano07m]. joins [LC02]. Joint [CMD05]. Journal [Lom08b, Lom09]. Journaling [PCG07]. just [Rho03]. just-in-time [Rho03].

Kahan [Bol09]. Karatsuba

[FH07a, Mon05]. Karatsuba-Like

[FH07a, Mon05]. Kernel [BGH07]. Key [DHS08, NHSC07, SH09, KPT04, Lu05, Tze02a, Tze02b]. know [LKF03]. Knowledge [SMN07]. known [Bao04, BMR04]. Koblitz [Lee12, DJJ^{+}08, Has01, VOT08]. KR [Kan05].

L2 [Kim09, SZZ09]. L2/L3 [Kim09]. L3 [Kim09]. Lagrange [BIN06]. Language [FSL07, HW07, SMN07, YF09]. languages [LG01]. LANs [LLC03]. Large [CWC07, CCY00, Cil09, FPFP06, Has00, Kim09, LLWS08, Ray06, ZSXZ07, AS00, Ata01, FS00, HCC^{+}00, LTCH05, Lin01, MH01, SM03, ZZZ04]. Large-capacity [Ray06]. Large-Scale [FPFP06, Lin01]. Last [Bol09]. Latch [ORM07]. Late [MVG^{+}04, QD04]. Latency [AVZ08, CHH^{+}03, CRS09, Kap09, KLCV08, KK09, SMBS06, AF05, HA05, KS00a, LB04, MS02, PG01]. Latency-Aware [KLCV08]. Latency-Insensitive [Kap09]. Lattice [DSV05, SLZ05b, CHL01, LY02a]. Lattice-based [DSV05]. Layered [ZMY08, KHRR02, LLPC04, MG02]. layout [KRCB01, KGM^{+}05]. layout-conscious [KRCB01]. layouts [PKM00]. Leader [Sto00]. Leading [JLZ^{+}09, KM06b]. Leading-One [JLZ^{+}09]. Leakage [CS08b]. Learning [Ano06h, FFJ^{+}06, OMG07, BC01b, OR00]. Least [HLM00, LCK^{+}01, WK03]. Lee [Jha13, ABA07]. Left [HE05, JY00, RK05, KGB05]. Left-to-Right [HE05, JY00, RK05, KGB05]. Length [CCY00, HU09, Ker08, RK06, RK07, AWY01, ABG04, BSM05, CC03, HU02, HU06, PR04c]. Level [BCL07, Cha06, LMW07, XYR^{+}09, ZMM07, ASB03, Ali02, Bar05, HBH05, LG01, MGT04, MM04, RMH05, SD06, WFMSW00, ZC05a]. levels [PEP06, PR02b]. Leveraging [SZZ09]. Lexicographic [SHR08, AVVY05]. LFSR [WBW08]. Library [AK09, DLM09, Ano07a, Ano07k, Ano07l]. Lifetime [LAG^{+}01, BBM^{+}03, PS05, PS06a]. Lifetime-sensitive [LAG^{+}01]. lifetimes [ISF06]. lifting [Che04, RN04]. lifting-based [Che04]. lighting [Har04]. Lightweight [BCL07]. Like [FH07a, Mon05, FDBS05a, FDBS05b]. Limit [CL09, PR04c]. Limitations [KL09, DSK00]. limited [TP02a, YW04b, ZY06]. limits [Pal05a]. Lin [HKM09]. Line [DNVG05, CJ01, JJ06]. Linear [Clu07, PCC02, SPS08, ZP08a, ZP08b, BBP^{+}01, LP01b, RSQL03, Voy05, XTX06]. linear-time [Voy05]. linearly [GD03]. lines [LPAM04, MFR00]. Ling [DN05]. Linguistic [HWW07]. Link [GCNS08, SSCG06, TSS08, VLP^{+}08, AN05, CMD05, LA03b, WDL^{+}03]. link-state [WDL^{+}03]. linking [Tsu00]. Links [VNM07, DLBS03]. Lip [lFB07]. Lip-Motion [lFB07]. List [Ano05g, Ano06a, Ano07b, Ano08e, Ano09a, Ano00a, Ano01a, Ano03, Ano04a]. lived [PKEG04a]. Living [NASK^{+}08]. LNS [CCY00]. Load [CNG^{+}09, JVG07, MS09, ZV06, AS00, BZ02, CP02, HSY00, Kum00, KH00]. load-value [BZ02]. loading [DV04, PR02a]. loads [AE00]. Local [WDY08, TL02]. Locality [CH07a, JDZ07, LX09, LJ04, SZZ09, TSS08, JZ05, LBL01, SXWL04]. Locality-based [LJ04]. Localization [dOBNL09, COP^{+}06]. Localized [Ano04h, ICRSR^{+}09, KB08a, LSS09, PY09, DGZA03]. Locally [CWC02]. Location [GS00b, MP09, SC05, SS09b, CIQC02, Fan02, LCCA02, ZXL02]. Location-aided [SC05]. location-dependent [ZXL02]. lock [JJ06]. lock-up [JJ06]. locking [JLN04, WCKD04]. log [WH03]. log-structured [WH03]. Logarithm [FFLTM09, PEB04]. Logarithmic [CCSK00a, CCSK00b, CSK^{+}08, MR06a, MDJM05, AS03a]. logarithms [AP00]. Logic [AC09, CCdS04, LH08, MRL06, PB07a, SBPV07, CV00, DM00b, HML00, PGPZ00, RTT05, San06, TN00]. Logic-Specified [AC09]. logical [ALMN05]. Long [WHT09, WK06]. Long-Term [WHT09, WK06]. Longest [HCH01]. Look [Has00, MDJM05]. Look-Up [Has00, MDJM05]. Lookahead [CUT00, KHM01]. Lookaside [QD08, JJ06]. Lookup [FFLTM09, KB08b, JPJ^{+}04, RMB05, SZ05, YL06a]. Lookup-Based [KB08b]. Loop [KN08, RLJ^{+}09, TP02a, CHH^{+}00, DNVG05, GRV05, JBV^{+}05, LWG01, PDS04, TSC^{+}00, TL02, WK03]. looping [HB01a]. loops [TSD01]. Loosely [TPB^{+}08]. loss [CLW^{+}03]. Lossless [JHZ01, NYC05]. Low [AVZ08, CHC05, CMCJ04, CH07b, FS07, HE05, KB03, KJM^{+}09, LHJL05, LSLsK07, LMW07, MMPT08, Mel07, NTA08, ORM05, RMH04, RMH05, RHMLL08, SKS04, TTA^{+}02, TLS09, VLP^{+}08, XWC^{+}08, YD09, ZVT09, AS03a, AS03b, CDV^{+}05, CT03, CT05, JBV^{+}05, KL03a, KL04, LW00, RSM^{+}05, Ray06, San06, WWSH04, ZZZ04]. Low-Complexity [RHMLL08, VLP^{+}08, LHJL05, WWSH04]. Low-cost [CMCJ04, TTA^{+}02, Ray06]. low-delay [CT03]. Low-Density [FS07]. Low-Latency [AVZ08]. Low-Power [HE05, TLS09, SKS04, AS03a, AS03b, LW00, RSM^{+}05, San06]. Low-Transition [NTA08]. Low-Weight [CH07b]. Lower [AE00, DV04, PR04d, SW00, ZK01]. lower-power [ZK01]. LPR [RL04]. LRFU [LCK^{+}01]. LRU [JZ05]. LUT [SNB07]. LZ [LW00]. LZW [TM05].

MAC [HSH01, ZW08]. Machine [Ano06e, Ano06c, Ano06d, FFJ^{+}06, GDM07, Moo07, Hie04, LY02a, PR00a, VW05]. Machines [AM07, Fro00, PR00a, PCC02]. Macromodeling [SB07]. Main [TTC09, ASF^{+}01]. Maintaining [CS08b, KLA^{+}03, XHLC08, MO06]. makespan [LZ06]. Making [JZ05]. Malicious [MRL06]. malleable [BKM^{+}06]. Manage [ASMD07]. Management [AM08, ABF^{+}07, AZ09, BVM07, FPFP06, GCN^{+}09, GR07, HUN07, HTKL08, JDZ07, LL08, LG06, LSLsK07, MBG08, MLB^{+}09, MWK^{+}09, NHSC07, NH06, PC07b, SKK^{+}09, SH09, TEG09, ASB03, AHS06, BLCA02, Cao02, CP02, CBB^{+}02, Fan02, Fan03a, GKM03, GBHL06, GL02, HPH^{+}04, LSP04, Lu05, MME04, PCL06, RKM05, SLG02, SMMMK03, ZZ05]. managing [LCC02]. MANETs [WLD06]. manipulation [DM00a, GL03, GD03, iM02]. Manipulations [HL09]. manufacturing [MDM04]. Many [LG09, PKL09, THW03, LCD02]. Many-Core [LG09]. Many-to-Many [PKL09, THW03]. map [LJ01]. mapped [DNVG05]. Mapping [WAU^{+}08, SLT01]. March [BBD^{+}08a, vdGT03]. Market [ET03]. Market-based [ET03]. Markov [Car04, MB01, NYC05]. Markovian [ING09]. Mask [RS08a]. Mask-Based [RS08a]. Masking [PKR04]. Massey [RMH02]. master [LZ06]. master-slave [LZ06]. Mastrovito [HK00a, PDS07, ZP01]. Matching [CCC07, Lee09, CCC06, LTTH04, TM05]. mathematical [BO05]. MATLAB [RB05]. MATLAB-based [RB05]. matrices [BT05]. Matrix [BP07, WS01, LP01b, SP02]. Matrix-Stripe-Cache-Based [BP07]. matroid [RTD00]. matroid-theoretic [RTD00]. maximal [FML03]. maximizing [PR04c]. Maximum [PS05, AWY01, PS06a]. may [TMD05, YJ00b]. MDS [FDBS05a, FDBS05b, FS07]. me [LCCA02]. measure [PR04b]. Measurements [CLE^{+}07]. measures [CAK04, GL02, LTCH05, MR01]. Measuring [JPEJ06, SS02]. Mechanism [DTHS09, KGG08, CP02, CHG00, ICM03, KR04, LLC03]. Mechanisms [EBPG06, HDQK09, BCGG00, Con00, Ste02]. Media [WZ07, WZSP04]. Medium [BIN06, CMD05, ZBC02]. Meetings [PAW07]. Member [Ano09f]. membership [GKM03]. Memoization [ACV05]. Memories [KPP09, LBP08, ZIPL00, vdGT03]. Memory [ACMM07, AGPP09, DZZ00, FPL^{+}08, GE08, Has00, HTKL08, KGG08, KPJ^{+}09, KLCV08, KOA07, MK07, TSS08, TTC09, VOL08, VOT08, WYZ08, ASF^{+}01, AAvdG03, BABD03, BLAA01, BFR01, BW03, CYL01, DSV05, FDZ03, GS00b, GVMC^{+}06, HKL01, JM01, KOHC03, KCHS04, KGMS00, LP01a, LP06b, LRB01, MBF^{+}04, MH01, MS02, MKAP05, NR04, Par04, QD04, SM00, SLT01, TP02a, WB03, Wil01, YGZ05, ZFP^{+}01, ZC05a]. Memory-Based [KPJ^{+}09]. memory-level [ZC05a]. Memory-Link [TSS08]. Mesh [BMlX08, CKC^{+}08, LCL07, LHC^{+}08, TK07, HS04, Jha03, Ros04, Wan03, YD02, Zha03, WS01]. Mesh-Based [LHC^{+}08]. mesh-connected [HS04, YD02, Zha03]. mesh-structured [Ros04]. Meshes [XZP09, MYL^{+}01, Wu03, YW01, Zha02]. Mesochronous [VLP^{+}08]. Message [Ano07n, HC01, LKTT08, HHTH00, ZS00]. Message-Pruning [LKTT08]. messaging [Gro04]. Metastability [PHA06, VPR04]. Method [EP09, HKM09, KT08, NKSG09, PF08, WBW08, AB00, CPN^{+}06, DU04, JSD01, LVLL06, Lo02, MB03, Pie02b, PR02c, Sun04, XTX06, YHIO02, vdGT03]. methodologies [YL06b]. Methodology [LV07, FWCL06, GNF^{+}06, HBH05, YLH05]. Methods [AM07, Gol06, Ker08, WKJ07, dDT05, KGB05, LCL03, SSST06, YW06]. Metric [ABA07, Jha13, MSM02, Pal05b]. Micro [LP09, XW08]. Micro-Scheduler [LP09]. Microarchitectural [YSLH07]. Microarchitecture [MBR^{+}09, VLP^{+}08]. Microarchitectures [ACS^{+}09]. microkernel [AFR02]. microkernel-based [AFR02]. Microphone [PAW07]. micropipelines [CBPC01]. Microprocessor [CCSK00b, GE08, KC07a, LJS^{+}07, TFCW07, BMO04, DH06, CCSK00a, CSK^{+}08]. Microprocessors [KN08, PL09b, ZBS^{+}04]. microsensor [CDV^{+}05]. Minimal [PB04, FB00, KHP00, KGB05, LSV00, PR00a, Wan03]. Minimax [POMB05]. minimization [GD03, NAH02, TP02a, WFMSW00]. Minimize [CS08b, HY08]. Minimizing [KM07, LZ06, HAK05]. Minimum [BRC08, GYA^{+}03, ICRSR^{+}09, KLT07, LKTT08, Ano00c, BDD00, CHH^{+}00, DS00, Par03]. Minimum-Energy [ICRSR^{+}09]. Mirrored [TB06]. MISER [TTC09]. Miss [ZDS^{+}07, FDZ03, JD06, KGM^{+}05]. Misses [MKP06, KSL05, ML00]. mission [MB01]. Mitchell [MR06a]. Mixed [ABA07, BG08, Dad07, HW07, SPS08, LwLH02]. Mixed-Language [HW07]. Mixed-Precision [SPS08]. Mixed-Radix [ABA07, BG08]. MM* [CCC07]. Mobile [CWC07, CRS09, MY07, PLP08, WCYR07, XW08, BB03b, Cao02, EGP03, Fan02, Fan03a, KPDS02, LLS02b, LC02, LLS02a, LLSC02, LA03b, MPAS03, NS03, PMR02, PS03, SLG02, TNS03, Wan04, ZXL02]. mobility [CS04, Fan03a, Lin01]. Möbius [HC06]. Mode [AM07, LS09a, MTHA08, SSA09, Hey03]. Model [ASMD07, BGB08, CC07, CCC07, CT09, HC08a, HC08b, HCL07, HSS^{+}08, ING09, KL09, NH06, SMN07, SS09a, TW08, WAU^{+}08, WRJ07, XZP09, YT07, CLTH04, GS00b, HL01, JLN04, LTTH04, MOK04, MB01, TP04, VC02, Wan03, Wu03]. Model-Driven [WAU^{+}08]. Modeling [Ano07f, CT03, CRS09, DPZ07, yFBC03, GE08, LLW07, MBS^{+}08, SMMMK03, TL02, UBWF08, ZMM07, ZX07, Fan02, GSW02, Gro04, GL03, HJ01, HKA01, KL03b, NYC05, PDS04, PGPZ00, SAOKM01, SAKR03, XV04]. Models [Ano07g, Ano07i, CGS07, PPB^{+}07, TPB^{+}08, AS03c, CU02, Car04, CLW^{+}03, DU04, FLW03]. modern [LRB01]. modes [DKV^{+}01]. modifiable [GL03]. Modified [EVN04b, FS07, ZP01]. modifying [vdGT03]. Modular [BM08a, CBC07, CH07b, DKT07, DZ06, KT08, SBPV07, SK00, SC07a, BP01, Hia00, Hia02, KT05, TK03, VLG06]. Modules [Ano06h]. moduli [CN03]. Modulo [ACS^{+}09, AQ08, EVN03, FFLTM09, KNE^{+}00, PBB07, PB07b, EVN04a, EVN04b, EVDN05, LAG^{+}01, SAJ02, VEN02]. MOLEN [VWG^{+}04]. Monitoring [gLMK07, PF08, VPG^{+}08, WHT09, Che05, GL05, PR04a]. monotone [Wan03]. Monotonic [LLWS08, BG03, BBB03, LMM03]. Montgomery [BP01, DZ06, FH06, HRM09, LHJL05, OS03, PS04, SK00, Sav05, TK03, Wu02b]. MorphoSys [SLL^{+}00]. Motion [lFB07, KGG08, YWV05]. Mounted [CKS^{+}08]. Movement [Fan03a, ZBI^{+}07]. Movement-based [Fan03a]. Moving [LKTT08, PXK^{+}02]. MPLS [OMG07]. MSHR [BW03]. Muller [DN06, Fal03, DM00b, LVLL06]. Multi [PKCD07]. Multi-Rate [PKCD07]. Multiaccess [Par04]. Multibanked [TIVYL09]. Multibit [KS07, LS08, LS09b, SZZ09]. Multibit-Trie [KS07]. Multicast [NHSC07, SC07b, WYYZ06, WLS09, GHP03, KNS01, LS02, Lu05, PY05a, PRO03, THW03]. Multicasting [LLP09, ST04]. multiclass [CSS02, CL01]. Multicomputation [YGL07]. multicomputer [HSY00, HKA01, MJ02]. multicomputer/distributed [HSY00]. multicomputers [LP00, YD02, Zha03]. Multiconstrained [XM07, SS06]. Multicore [SBPV07]. Multidigit [MDJM05]. Multidimensional [Cha09a, HXW09, MDJM05, LLC02b, LCL03]. Multievent [GR07]. multiexponentiations [DJM00]. multigrid [MB03]. Multihop [HDQK09, ICRSR^{+}09, KM07, MS09, ZMY08, CRJ06, PBC03, TNS03]. Multilevel [JDZ07, JVG07, SHYV06, XW08, Che04, Saa04, SA02]. Multimatch [FN09]. Multimedia [ACV05, CCK00, WFP08, HH04, KGM^{+}05, SS02, TJB03, XSMH04]. Multioperand [Dad07, KS05]. Multipartite [dDT05]. multipath [CCK00]. Multiple [HXW09, JPSR07, Jha13, Kag03, NAH02, OMG07, PAW07, SKJ07, TLS09, ZYZ09, ABP^{+}05, AN05, DV04, DM00b, FB00, FDBS05b, HAK05, HMM06, HML00, Li05, LMM00, PEP06, PR02a, Sez04, Thi00, UF03, WK03, YNOJ02, YY01]. Multiple-Distant-Microphone [PAW07].

Multiple-Precision [TLS09].

Multiple-Radix [Jha13]. Multiple-seed [Kag03]. multiple-valued [DM00b]. multiplexer [CLW^{+}03]. Multiplexing [LMV^{+}08]. Multiplication [AH08, BM08c, CO09, CH07b, DKT07, DJJ^{+}08, EHS09, ES00, FH06, FH07c, HKM09, HRM09, Has00, JK09, KT08, KGB07, Lee12, MR06a, NWA07, RMH06, SBAB00, VOT08, BT05, DHH^{+}06, Gol02, KT05, KB03, KGB05, LP01b, Mis06, OS03, RMH03a, RMH03b, RMH04, RM06, TK03, WS01]. multiplication/division [KT05]. Multiplicative [AQ08, DH05, TYT01]. Multiplier [CCL^{+}09, CS09a, Cil09, HE05, KG06, NWA08, PCH06, TLS09, Wu08, YJ00a, CHC05, FD05, HK00a, Hia00, RMH02, SK01, Wu02a, WHBG02, Wu02b]. Multipliers [BM08a, ELMT00, FH07b, KWP06, PDS07, RMPJ08, VT09, EVN04b, EVDN05, EL02, GSA06, IST06, LLL01, LHJL05, OKLC00, RMH05, RHK03, SMM05, Sun04, ZP01]. Multiply [BDL09, DT05, LB04]. Multiply-Accumulate [DT05]. Multiply-Add [BDL09]. multiply-add-fused [LB04]. multipoint [WLD06]. multipolling [LLC03]. Multiprocessor [GE09, KLCV08, PPB^{+}07, BF06, BFP^{+}06, CWM01, IO03, KCHS04, LTCH05, RK03, SW05]. Multiprocessors [AGPP09, SBC08, BG03, BFG03, Bar04, BW03, CHH^{+}03, IB00, MPP^{+}05]. multipurpose [RK03]. Multiresolution [WHT09]. Multiserver [SX09]. multisize [KHM01]. multistage [FB00]. multistate [ZWST03]. Multistep [GR07]. Multithreaded [LG06]. Multithreading [LPS07, RLJ^{+}09, SP07, LVS01, PG01]. Multitier [HASL07]. Multivariate [CL01, YECV02, ZV02]. Multiversion [PC02b]. multizone [MKA03]. mutual [LWRJ06].

NAND [LBP08, LP06b]. Nano [Ano05c, Ano05d, Ano05e, DLL07]. Nanomagnetic [KC07a]. Nanospace [LH08]. Nanowire [CL09, RS08a]. NB [LJ01]. NCQ [MT09]. Near [ACMM07, Li04, CP03]. Near-Memory [ACMM07]. near-perfect [CP03]. Nearest [PC07a]. Need [Tou07]. negotiation [AAS00]. Neighbor [PC07a]. Neighbourhood [Fuj00a]. nests [DNVG05]. NET [FZM00]. nets [GL03, LO04, MB01, PCR01]. Network [APV09, CLE^{+}07, GBD07, KLT07, KRP05, LCL07, LMF^{+}08, NPB07, WHT09, XZP09, ZW08, CU02, FZM00, FCB04, GSW02, GL03, HJ01, HH02, HLLR00, Kan05, LKF03, LTTH04, MJ02, OM04, PGJ^{+}05, Res01, SAKR03, TNS03, XT02, Yan02, YW04a, YCL01, Zac06, ZS00]. network-based [SAKR03]. Network-on-Chip [ZW08, PGJ^{+}05]. Network-Wide [CLE^{+}07]. Networked [BHS09, XQ08]. Networks [AD08, Ano04h, Ano07f, ACPP08, BF08, CLLL09, CWC07, CMS08, CCC07, CXP06, CKC^{+}08, DMS^{+}09, FP09, FPL^{+}08, HC08a, HDQK09, ICRSR^{+}09, JVG07, JSW^{+}06, KM07, KL08, KSL08, LCL07, LMV^{+}08, LSS09, LLP09, LHC^{+}08, LKTT08, LX07, LH09, LLLD06, MY07, Mar08, MAA^{+}08, MBS^{+}08, MR06b, MS09, MP09, NH06, PH08, PGVB08, RM09, SBC08, SX09, VLP^{+}08, WFP08, WLS09, WCYR07, WDY08, WGZ^{+}08, ZV06, ZMY08, ZYZ09, ZSS08, dOBNL09, AE00, BB03b, BP06, CIQC02, CCK00, CLTH04, CCC06, CRJ06, CHH^{+}00, CWC02, CLF05, CCS^{+}04, CSR04, CSS02, CL01, Dat05, DV04, DGZA03, DPIK05, FB00, Fan02, Fan03a, Fan03b, FZM00, FLM^{+}03, GCI06, GNF^{+}06, HHTH00, Kar06, KI04, LWF03, LR06, LLC00, Lin01, LWG01, LLPC04, LDH06, MOK04, MO06, PS05, PS06a, PBC03, PR04a, PV03, RDH^{+}01, SC05, ST04, THWH01]. networks [WD04, XTX06, YJMS05, ZC05b]. Networks- [KSL08]. Networks-on-Chip [ACPP08, CKC^{+}08, FP09, FPL^{+}08, LMV^{+}08, Mar08, SBC08, VLP^{+}08]. Networks-on-Chips [Ano07f]. Neural [NPB07, WLS09, BC01a, BC01b, SAKR03]. Newton [GS00a]. Next [CWC07, AGG06]. Next-Generation [CWC07, AGG06]. no [KJD02]. Node [CT09, HX08, LX07, AN05, Tsu00]. nodes [CDV^{+}05, HC06, ZC05b]. Noise [PA08, HR06, LLVC04, LVLL06]. Noisy [WLS09]. Non [GCI06]. Non-real-time [GCI06]. Nonblocking [Sue09]. Noncongestive [MT09]. Noncooperative [GR07]. nondedicated [GSW02]. Nondeterministic [LPS07, Hie04]. Nondomination [KC07b, Jia04]. Nonideal [MS03]. nonintrusive [GRV05]. Noniterative [CMS08]. nonlinear [BFG01, GCD00]. nonmember [MYL^{+}01]. nonnumeric [JMH02, ML00]. Nonscan [XXF03, PR04d]. nonstandard [BWTE04]. nonstationary [CBB^{+}02]. nonuniform [HK00b, JD06]. NOR [LBP08]. Norm [XM07, TK00]. Normal [CCL^{+}09, FH07c, MMPT08, DHH^{+}06, GPS05, OKLC00, RMH03a, RMH03b, RMH05, RM06, SK01, TYT01]. Normalization [Kor09]. Note [Ano02b, Lom07, Lom08a, Pra04, PL05, Pra05, PL06, Pra06a, Ano01c, CDE^{+}00, Gau01, Gau02, JWF01, Pra03a, Pra03b, PL04, SW00]. Notice [Ano04h, DGZA03]. Novel [GBD07, KB08b, PDS07, Pie02a, RG09, SKJ07, SGB00, ZCWZ08, JZ05, LHLL06, Lu05, WMZH02, WH03, WWSH04]. NP [BWTE04, KLT07, LKTT08]. NP-Complete [KLT07, LKTT08]. NP-hard [BWTE04]. NTRU [OS03]. Null [Yan08]. NUMA [BW03, IB00]. Number [DLM09, FFLTM09, MDJM05, Par00, SMS07, BDD00, BGL^{+}03, KSL05, PR04d, SAJ02, SC04, SW00, SLS04, TL02, WK03]. number-based [KSL05]. Number-Theoretic [Par00]. Numbers [HBF09, SST05, TSP00]. Numerical [KDM^{+}09, SNB07, LLVA01]. numerically [Car02]. NVRAM [KPJ^{+}09].

O [KV02, LSV00, OM04, SD06, TP02a]. OBDD [CAK04, JWF01, WFMSW00]. OBDDs [HLM00, JWF01]. obfuscation [GT06]. object [CKBF03, ISF06]. object-replication [CKBF03]. Objective [WK06]. Objective-optimal [WK06]. Objects [LKTT08, HK00b, KLA^{+}03, PXK^{+}02, RBC^{+}03]. Oblivious [BMlX08, Sch01, Tze04]. occupancy [CHH^{+}03]. Odd [BM08b, KL08, Wu03]. odd-even [Wu03]. OFDM [EFX^{+}04]. Off [SGB08, Fan03a, RB05, ZZZ04]. off-chip [ZZZ04]. offline [HWI^{+}02, KV02]. Offs [KPJ^{+}09, LCLV08, YSLH07, HV06, MEB01, PGJ^{+}05, SD06]. omission [PR02b]. omissive [AK00]. Omura [RMH02]. On-Bound [CX07]. On-Chip [KSL08, LHC^{+}08, MRM07, WHZ09, GRV05, SZ05, TP02a, WO01]. On-Demand [CS08b, LMW07]. On-Disk [WYZ08]. On-line [CJ01]. On-the-fly [Fro00]. onboard [TTA^{+}02]. One [JLZ^{+}09, SLZ05b, CHC05, Che04, LCD02, LLL01, VEN02, YCL01]. one-dimensional [Che04]. one-sided [YCL01]. one-to-many [LCD02]. One-Variable [SLZ05b]. Online [CMAB09, DTHS09, MR06b, RC06, HWI^{+}02, LJ04, PS06a, SWP04]. only [Lo02]. OpenGL [Har04]. Opens [HAAvdG06]. Operand [MR06a, SP07]. operands [PKEG04a]. Operating [LP09, LJS^{+}07, PL09a, SWP04, ASF^{+}01, BCF^{+}03, OM04]. Operation [BKM07, KPDS01, LLS02b]. Operation-Centered [BKM07]. Operation-saving [KPDS01]. Operations [BIN06, BSH09, CLV05, KN08, LA05, WSTJ09, ZP08a, CV00, JMH02, KOL02, LLC02b, LCL03]. Operators [BBD^{+}08b]. opportunities [HL03]. OPT [KV02]. Optical [MR06b, CHG00, HPS02, RDH^{+}01, SMSM01, THWH01, YW04b]. Optimal [ABF^{+}07, AVVY05, ABP^{+}05, AMMMA01, BS04, Bar04, BFG01, BWTE04, BMlX08, CXP06, CV08, CRJ07, FH07c, JY00, KKN07, LCC06, RS08b, ZY06, AE00, BO05, yFBC03, HK00b, KV02, Kan05, dALB03, LML01, LWG01, OPZ00, PC02a, PCL06, PvST02, RMH03a, RK05, SK01, UK01, WK06]. optimality [Li04]. Optimally [Res01]. Optimization [ARSM07, AC09, Ano07f, Fio08, GMSC09, JSM09, KDM^{+}09, LV07, NH06, SZGS09, SC07a, WKJ07, WHZ09, XW08, BBM^{+}03, DLBS03, EAGS01, HR06, ICM03, KM06a, KF01, MTB^{+}01, PL01, WPP05, ZZZ04]. Optimization-Based [NH06]. Optimizations [RCA01, Bar05, BT05, KR01, VKI^{+}03]. Optimized [AK09, DPIK05, JG06, LV09, SE04]. Optimizing [HU06, LGML05]. Optimum [KWP06, YJ06, TMD05]. optoelectronic [WS01]. Orchestrating [LF09]. Orchestration [CLE^{+}07]. Order [MBR^{+}09, PRB09, Yan08, BSN^{+}06, GYA^{+}03, Lu05, MKAP05, NYC05, PR02b, SSS05, YL06a]. Ordered [YGB08, JMH02]. Ordering [SHR08, LLSC02, MWA^{+}00]. orderings [TMD05]. Orders [SS09b]. Organization [TB06, JBV^{+}05, KGM^{+}05]. Organized [Voy08]. organizing [BP06]. Oriented [CLLL09, WWH^{+}07, KLS01, KCKC05, OBB^{+}02, SA02, SV06, vdGT03]. original [ZP01]. orthogonal [BDD00]. OS-Aware [LJS^{+}07]. oscillator [BGL^{+}03]. oscillator-based [BGL^{+}03]. OTIS [WS01]. OTIS-Mesh [WS01]. Out-of-Band [ZSXZ07]. out-of-core [SP02]. Out-of-Order [MBR^{+}09, GYA^{+}03, MKAP05]. Outer [MS09]. Output [JPSR07, BDG03, NR04, PY05a, PKR04, YJ00b]. Outs [MNR08]. Overall [LF09]. overcoming [DH06]. Overflow [SBAB00, GSA06, Lin01, SXZ^{+}06]. Overhead [KN08, Kim09, CK06, ZZZ04]. Overheads [HTKL08]. Overloaded [BVM07, MAMMA03]. overruns [CBS02].

P [Gho01]. P2P [LH09]. PABC [LSLsK07]. PACE [LS04a]. Package [Ano09f]. Packet [ANPS07, CS08a, Cha09a, FN09, LS09a, LS08, LS09b, Mha09, PY09, RG09, YGB08, YD09, ZYZ09, CLW^{+}03, GVMC^{+}06, KKKB05, KL03b, MRS00, NMLE00, PY05a, ZCW^{+}06]. packet-forwarding [NMLE00]. Packet-Mode [LS09a]. Packing [LF09]. Padubidri [Jha03]. page [KHM01, Sez04]. pair [RK05]. Pairing [BBD^{+}08b, PV06, SKG09, GPS05]. Pairing-Based [PV06, GPS05]. pairs [MO06]. pancyclicity [HC06]. Papers [Ano04e, Ano04c, Ano04d, Ano05b, Ano05c, Ano05d, Ano05e, Ano06e, Ano07c, Ano07e, Ano07f, Ano07g, Ano07i, Ano07j, Ano07d, Ano07h, Ano09d, Ano09c, Ano06c, Ano06d]. Paradigm [BCK09, KGA01, NMLE00, Sma03]. Parallel [ANPS07, CPR03, CS08a, CRL00, Cil09, Dad07, FH07b, FPFP06, HRM09, JK09, KNE^{+}00, LL08, LP01b, MTHA08, Mel07, PS04, PCH06, PB07b, RMPJ08, RHK03, RHMLL08, UF05, Wu08, XWC^{+}08, YJ00a, CHC05, CH06, DN05, EVN04a, EL02, FD05, Gho01, HS04, IST06, KV02, LSV00, LLL01, LHJL05, LP00, LPS00, OPZ00, RMH02, RMH04, SLL^{+}00, TB03, Wu02a, ZCW^{+}06]. Parallel-Prefix [KNE^{+}00, PB07b, DN05, EVN04a]. Parallelism [BP09, LF09, MGT04, YSL00, ZC05a]. Parallelization [LOP07, SA02, KHB02]. Parallelizing [MKP06, CWM01]. Parameterized [HSS^{+}08, DNVG05]. parameters [Tze04]. Parametric [KGS00]. Parity [CPRS07, LLW07, VW05, YY01]. Part [FDBS05a, FDBS05b]. Partia [WYYZ06]. Partial [TEG09, XP04, LLSC02, Par03, PY05b]. Partially [Mha09, Bao04, LFA04]. partition [LY02a]. Partitioned [KLS09, MR06b, WH06, BF06, JWF01, LKS05]. partitioned-OBDDs [JWF01]. partitioner [SA02]. Partitioning [CBM07, RSM^{+}05, AS00, Car03, CC03, MGT04, Pie02a, PR02c]. partitionings [BO05]. pass [BSN^{+}06]. Passive [MR06b]. Password [OS07]. Path [BMlX08, CY09, Ker08, KLS09, PKL09, BSM05, HC06, MKAP05, TM00]. Paths [OMG07, AFPS01, HCH01, LCD02, MO06]. Pattern [ABLP07, NTA08, RK07, TM05, CP03, DCCS01, PR02c, PGPZ00, SGB00, XN06]. Patterns [MK07]. PBC [Mha09]. PC [KV02]. PC-OPT [KV02]. PCS [HH02, LLC00]. PDM [RS08b]. PE [AF05]. peak [CP02]. Peer [GKM03, ZSS08, CKBF03, GL05, XLN05]. Peer-to-Peer [ZSS08, GKM03, CKBF03, GL05, XLN05]. Pentanomials [Cil09, PCH06, RHK03]. Perceived [WX06]. Percolation [AD08]. Perfect [Fet03, CP03, WK03]. perfect-rate [WK03]. Performability [Mey01, RK03]. Performance [Ano04h, AH08, BGH07, CM02, CD09, CV08, CD03, CS08b, DLT07, DLBS03, DGZA03, FP09, GSW02, HKA01, HTKL08, HE05, HDQK09, KPJ^{+}09, KBK03, KSL08, KB08b, LL08, LG06, Lee09, LLC00, MAA^{+}08, Mel07, MRM07, PGJ^{+}05, PDS04, PL09b, SSA09, SH09, SPS08, TFH07, WCYR07, XWC^{+}08, ZV06, ZYZ09, ZP08a, ASF^{+}01, BCF^{+}03, CKS^{+}06, CHH^{+}03, CD04, CNM^{+}02, CK06, CJDM01, Fan02, FLM^{+}03, FDZ03, GL02, HKL01, HJ01, HL03, JZ05, KGA01, KS00b, KM03, KL03b, LwJKW03, LMS04, LGK01, LLVA01, MR01, MPP^{+}05, MH01, MOK04, MMRT06, MKAP05, PRO03, SKS04, SS02, TGKL03, Uht05, WMZH02, WH03, Wil01, XS02, XSMH04, YGZ05, YWV05, ZK01, ZBS^{+}04]. Period [AS00, LLWS08]. Period-based [AS00]. Period-Dependent [LLWS08]. Periodic [AZ09, BRC08, KL09, SL06, Tou07, AMMMA01, AMMMA04, Bar04, BDG03, BB04, HSW03, KDB^{+}05]. periods [XR04]. Permutation [JSW^{+}06, MR06b, Dat05, Kan05, MT06, YW04a]. Perspective [TSS08]. perspectives [KS00a]. Petri [GL03, LO04, MB01, PCR01]. Phase [HY08]. phased [KCKC05, MB01, RTT05]. Phonetic [HW07, ZMM07]. Phonetic-Class [ZMM07]. Photonic [GBD07, SBC08]. physical [ACK^{+}03, MVG^{+}04, Rho03]. physically [MH01]. PIM [KGG08]. pinwheel [HL01]. Pipeline [CV08, SC06, TFCW07, WGZ^{+}08, CYL01, KMPE02, NMLE00, PLK^{+}03]. Pipeline-Based [SC06]. Pipelined [AVZ08, BSH09, CCY00, KS07, LS08, Mis06, VT09, WKJ07, YW01, Alw06, HV06, JPJ^{+}04, LP01b, LM00, Ray06, TSD01, ZBS^{+}04]. pipelines [CD04]. Pipelining [MSMS04, PKCD09, Tou07, BSN^{+}06]. Piracy [LH09]. pixel [PLK^{+}03]. PKASSO [PLP08]. PKI [PLP08]. PKI-Based [PLP08]. place [SV06]. Placement [ABF^{+}07, LX07, RG09, Xie08, LML01, RDH^{+}01]. placements [AE00]. Planar [XZP09]. Planning [SKK^{+}09, LP06a]. Platform [ZBI^{+}07]. Platforms [GCN^{+}09, KDM^{+}09, SWP04]. Plugged [LP07]. PMC [AS03c, CC07]. Podcast [Ano08f, Ano08g, Ano09h]. Point [ACV05, BR09, CGS07, CHA^{+}09b, DJJ^{+}08, EHS09, ES00, Gra09, HKM09, Kah04, KGB07, KK09, Lee12, LM08, TLS09, WSTJ09, XPGP06, BMR04, EP00, FHLM04, KGB05, LB04, LM00, NMLE00, SE04, VLG06, YSL00]. Point-to-Point [CGS07]. Points [KLT07]. Polarity [JSM09, DN06]. Policies [FAL06, ZV06, CJ01, LCK^{+}01]. Policy [CX07, PL09a, WCKD04]. Polling [ZMY08]. Pollution [ZL07]. poly [Ali02]. poly-time [Ali02]. polygonal [Wan03, YCL01]. Polymorphic [SS09a, VWG^{+}04]. Polynomial [CO09, CCY00, CH07b, FH06, JSM09, LV07, LCLV08, PCH06, RMH06, Wu08, RMH04, Wu02a]. Polynomial-Based [LV07]. polynomials [CHC05, Gol02, HK00a, LLL01, ZP01]. Port [JSW07, HvdG02, JS06, WT00, YW01]. Portable [GW04, Wan04]. Position [JLZ^{+}09]. Positioning [ZCWZ08]. Positive [Bal04, ABG04]. Possibilities [BBH06]. possible [HSW03]. Postprocessing [Gol06]. Potential [LG09]. pow [LL09]. Power [AMMMA04, CV08, Gir06, GSH^{+}08, Has01, HBF09, HE05, KM07, KSL08, LP06a, LSLsK07, LMW07, MPP^{+}05, MG08, NKY08, PH08, PSZS04, PL09a, PRB09, SL06, TLS09, VT09, VAZSR07, XLSF07, YWV05, ZZ05, AS03a, AS03b, BMO04, BDBR05, Cao02, CT05, CBB^{+}02, DKV^{+}01, GBHL06, JS06, KL03a, KL04, LW00, MME04, MDS02, NAH02, PCL06, RSM^{+}05, RKM05, SKS04, San06, ZK01, ZBS^{+}04]. Power-Aware [LSLsK07, SL06, AMMMA04, LP06a, PSZS04, ZZ05, BDBR05, Cao02]. Power-Efficient [PH08, VT09]. Power-performance [MPP^{+}05]. Power-Scalable [VAZSR07]. Power-Up [HBF09]. Power/Performance [CV08]. Powering [PEB04]. Practical [LZ07, RS08b, XZP09]. practice [BK00]. practices [Rya04]. Pre [HSS^{+}08]. Pre-RTL [HSS^{+}08]. precedence [LP00]. Precise [Kah04]. Precision [BM08c, BR09, LL09, LV09, PB02, SS00, SPS08, TLS09, KM06b, Par03]. Precomputed [VPG^{+}08, PR02c]. Predicate [Ksh07, CK05]. predictability [ML00, ZL04]. Predictable [CKS^{+}06].

Predicting [BSPG08, MP03]. Prediction [BGB08, JLZ^{+}09, KJM^{+}09, LP09, LJS^{+}07, LG09, MKP06, PA08, ZDS^{+}07, GBHL06, GSW02, GG01, ISF06, LY01, LY02b, MGZ06, MGT04, MS02, PSZS04, ZC05a]. Prediction-Based [LP09]. predictor [SSC03]. predictor-directed [SSC03]. predictors [BZ02]. Preemptable [BKM^{+}06]. preemption [Gho01]. preemptive [DRC05]. Prefetch [ZL07, SDT04]. Prefetching [BP09, BGH07, KV02, LwJKW03, LRB01, SSC03, WK06]. Prefix [KNE^{+}00, LKS05, PB07b, SHR08, DN05, EVN03, EVN04a]. Prefixes [CL07, LS04c]. preprocessing [AT05]. prescaling [ML01]. Presence [PKL09, AN05, Fav06, PDS04, SAOKM01, Moo07]. Preserving [YGB08, TMD05]. Pressure [BSPG08]. Prevent [OVB^{+}06]. Preventing [KKKB05, CMCJ04]. Prevention [LH09]. Pricing [LCL07]. primal [XTX06]. Primary [ZVT09]. Primary-Backup [ZVT09]. Prime [BIN06, LM08, KSL05]. Principles [Ano04h, DGZA03]. Priority [BRC08, BNRB09, DZB08, ASL04, Bar04, BB04, DSK00, KDB^{+}05, dALB03, MRS00, WAB^{+}02]. priority-driven [KDB^{+}05]. Proactive [Cao02, HR02, RL04]. Probabilistic [FDZ03, LP00, SB07, TSC^{+}00, Pal05a]. probabilities [CLW^{+}03, SL04]. probability [Had05, LRJ00]. PROBE [CBM07]. PROBE-Based [CBM07]. Problem [AD08, QD08, SC07b, BKM^{+}06, yFBC03, LS02, OR00, RTD00, TPT06, WAB^{+}02]. problems [AKS^{+}03]. Procedures [PR00c, BBK^{+}03]. Process [Kap09, MWK^{+}09, SFJ03, KK00a, LP06a, LVS01]. Process-Variation-Aware [MWK^{+}09]. Processes [SC06]. Processing [JVG07, LC02, PC07a, CRL00, KD02, LLC02a, LLSC02, TJB03, WZX05, YSL00, YY01]. processing/transmission [CRL00]. Processor [BBGM08, KGG08, KOA07, LSBV08, SH09, CRL00, CM01, CZM05, HW00, HMM06, HT00, KYY^{+}03, LP00, RSM^{+}05, ST03, Tsu00, VWG^{+}04, YD02, Zha03]. Processor-In-Memory [KGG08]. Processors [AVZ08, HL09, LV09, NBAR08, SS00, TW08, WKJ07, WWC06, AGG06, CKS^{+}06, CYL01, Gho01, GSG05, HAK05, HV06, HLLR00, JBV^{+}05, KKP00, KR04, KPGX05, KPEG04, LY01, LJ05, MKAP05, PLK^{+}03, QD04, RMH03b, TP02a, WB03, YSL00, YGZ05]. Product [AFM^{+}06, Gra09, HC08a, KK09, MRM07, CLTH04]. production [LAG^{+}01]. Products [KHP00, Ano00c, SB01]. profiling [HJS04, ML00]. Program [GBHL06, KJM^{+}09, ZDS^{+}07, JPEJ06, KF01, SL04, XV04]. Programmable [CCdS04, KKP00]. Programming [Ano07g, Ano07i, FSL07, PPB^{+}07, ST08, XTX06]. Programs [LF09, LLVA01]. Pronunciation [ZMM07]. Proofing [JSW^{+}06]. Propagation [HDQK09, SS09a, HJS04, SCZ01]. Properties [CS00, FFP07, AB05, HLM00, MS03, ZBC02]. Property [WH06, WK03]. Proportional [SCK06, YD09, WZX05]. Proposal [SK02, BBP^{+}01]. Protected [TFCW07]. Protecting [Kim09]. Protection [MG08, WYYZ06, XYR^{+}09, OM04, SXZ^{+}06]. Protocol [Bal04, CWC07, KG07, LMF^{+}08, ZW08, dOBNL09, CMD05, DLBS03, Dat05, DPIK05, GS00b, Gro04, HSH01, Kum00, KPDS02, KM03, LMS02, SC05, SLZ05a, Tze02a, Tze02b, WDL^{+}03, Wu03, ZBC02, PLP08]. Protocols [ABT07, SSCG06, SMBS06, UBWF08, WCYR07, GKM03, HMR02, LJ01, PHA06, RTD00]. Prototype [HXW09]. Provably [DJJ^{+}08, HvAL09, Lee12, SMS07]. Proved [BM08b]. Proven [Bol09]. provides [RBC^{+}03]. Providing [ASMD07, LMS02]. Provisioning [WX06]. proxy [MPAS03, WMZH02]. pruned [JSD01]. Pruning [LKTT08]. Pseudo [ARSM07]. Pseudo-Boolean [ARSM07]. pseudoexhaustive [DCCS01, SGB00]. pseudorandom [AB05]. Publication [DGZA03, Ano04h]. pull [BDK^{+}02]. Purpose [Ano07j, Ano07h, HL09, LP09, SGK08, LJ05, RMH03b, YSL00]. Push [ET01, BDK^{+}02]. Push-based [ET01]. push-pull [BDK^{+}02].

QCA [NPB07, SB07, VOL08]. QoS [ASB03, ASMD07, AHS06, AAS00, HUN07, KBK03, LSP04, LHLL06, MAA^{+}08, SZL08, WX06, XTX06, XM07, ZW08]. QoS-adaptive [ASB03]. QoS-Aware [ZW08]. Quadratic [POMB05]. Quality [BVM07, CKDS02, LLP09, MRM07, SCG08, THC^{+}08, CTA02a, CTA02b, PR04b, SS06, TSP00, THWH01, THW03]. Quality-of-Control [BVM07]. quality-of-service [SS06]. Quantification [JDZ07, LKS03]. quantifying [TP04]. quantization [RB05]. Quantum [CS09a, Imr07, Sue09, VNM07, VOL08, YF09, ZR07]. Quantum-Dot [CS09a, VOL08]. Quasi [SS09a]. Quasi-Species [SS09a]. quasistationarity [Car04]. Queries [HXW09, LC02, PXK^{+}02]. Query [LX09, PC07a, PXK^{+}02, LLC02a]. Queue [ZV06, AWY01, MRS00]. Queue-and-Rate-Adjustment [ZV06]. Queued [GS09, LS09a, PY05a]. Queueing [CMS08]. Queues [CNG^{+}09, TSD01]. Queuing [MT09, NH06, CSS02, CL01, GL02, LMS02]. Queuing-Theoretic [NH06]. quorum [LCC02]. quorum-based [LCC02].

Rabin [FDBS05b]. Rabin-like [FDBS05b]. RACCOOM [GHP03]. Race [OMG07, ZJ08]. Rack [CKS^{+}08]. Rack-Mounted [CKS^{+}08]. Radar [LKSS06]. radars [KCKC05]. Radices [ZR07]. Radio [NH06, SSCG06, CMD05]. Radios [MBG08]. Radix [ABA07, ALB00, BG08, Jha13, LA03a, LN07, PBLM08, BP01, COP^{+}06, ML01, Par03, SMM05]. Radix- [LA03a]. Radix-10 [LN07]. Radix-2 [PBLM08]. RAID [EP09, BP07, FDBS05a, FDBS05b, WZL08, Xie08]. RAID-5 [BP07]. RAID-Structured [Xie08]. Rail [MTHA08, ORM05, Pie02b, SMBY05]. RAM [HAAvdG06]. RAMs [Voy08]. Random [AKS^{+}03, Gol06, HBF09, KGB07, SMS07, ZSS08, BGL^{+}03, Had05, TP04, TSP00, UF03]. randomization [Car04]. Randomized [Ali02, LOP07]. Range [BDK^{+}05, CWZL08, MDJM05, VLG06, ZY06, ZCW^{+}06]. Range-Addressable [MDJM05]. Range-Reduction [BDK^{+}05]. Ranges [Cha06, CL07, LS04c]. Raphson [GS00a]. RaPiD [EFX^{+}04]. rasterization [PLK^{+}03]. Rate [BG03, BBB03, FFJ^{+}06, GCNS08, LLWS08, ML08, PKCD07, PB07a, ZV06, ZDS^{+}07, GHP03, KYL02, KL03b, LMM03, LMS02, WTL04, WZX05, WK03]. rate-based [GHP03, KYL02, LMS02]. Rate-Monotonic [LLWS08, BG03, LMM03]. Rational [LGK01, CL01, CLW^{+}03]. RE [ZCW^{+}06]. Reactivation [KG07]. read [BW03, MS02]. read-after-read [MS02]. Real [AM08, AC09, Ano05f, Ayd07, AZ09, BRC08, BWR^{+}07, BVM07, CHL09, CRJ07, DLM09, DZB08, FSL07, JG06, JS09, KD02, KM00, KCKC05, LKSS06, NASK^{+}08, PC07b, SL06, SKK^{+}09, SC06, THC^{+}08, WCLK09, WZ07, WRJ07, XHLC08, ZB09, ZX07, AS00, AHS06, ACCL06, AAS00, AMMMA01, AMMMA04, BCF^{+}03, BBL01, BDBR05, BJHW00, CBS02, CBT05, CJ01, DSK00, FZM00, GCI06, HHTH00, HSH01, HSW03, HR02, HL02, HL01, JR02, KLA^{+}03, KS00a, KDB^{+}05, KL03a, KL04, KH00, KLS01, KYL02, KKK03, LwLH02, LR04, LWRJ06, LKS03, LMM00, dALB03, LLA04, LHLL06, MAMMA03, MME04, DS00, NS03, Pal05b, PV03, PBWB00, Raj00, RL04, RPH01, SM00, SWP04, WKS^{+}05, WZSP04, WRJL05, XQ06, XR04]. Real-Time [AM08, AC09, Ano05f, Ayd07, AZ09, BRC08, BWR^{+}07, BVM07, CHL09, CRJ07, DZB08, FSL07, JG06, JS09, LKSS06, NASK^{+}08, PC07b, SL06, SKK^{+}09, SC06, THC^{+}08, WCLK09, WZ07, WRJ07, XHLC08, ZB09, ZX07, KD02, KM00, KCKC05, AS00, AHS06, ACCL06, AAS00, AMMMA01, AMMMA04, BCF^{+}03, BBL01, BDBR05, BJHW00, CBS02, CBT05, CJ01, DSK00, FZM00, HHTH00, HSH01, HSW03, HR02, HL02, HL01, JR02, KLA^{+}03, KS00a, KDB^{+}05, KL03a, KL04, KH00, KLS01, KYL02, KKK03, LwLH02, LR04, LWRJ06, LKS03, LMM00, dALB03, LLA04, LHLL06, MAMMA03, MME04, DS00, NS03, Pal05b, PV03, PBWB00, Raj00, RL04, RPH01, SM00, SWP04, WZSP04, WRJL05, XQ06, XR04]. Realistic [WFP08, HvdG02]. realization [CPR03]. realizations [YWV05]. rearrangeable [Kan05, YW04a, YCL01]. receiver [EFX^{+}04]. recently [LCK^{+}01]. Reciprocal [EIM^{+}00, LA03a, PB02]. Reciprocation [ELMT00]. Reclaiming [Ano05f, CBT05]. Recoding [JY00, SSST06]. recodings [SMM05]. Recognition [lFB07, HW07, KOA07]. recommendations [YL06b]. Reconfigurable [ABF^{+}07, KDM^{+}09, SBPV07, SKG09, SdBF04, VT09, ZP08a, ZP08b, BP01, Car03, EFX^{+}04, HPS02, LP01b, LO04, MM04, SKS04, SLL^{+}00, SWP04]. Reconfiguration [JS06, JSW07, LMF^{+}08, TEG09, AN05, Low00, MJ02]. Reconfigured [XLSF07]. Reconstruction [Clu07, XW08]. Recovery [XYR^{+}09, KHM03, KZP05, KKK03, MME04, OBB^{+}02, PMR02, SFJ03, TTA^{+}02, ZC05a]. recovery-free [ZC05a]. recovery-oriented [OBB^{+}02]. rectilinear [Wan03]. rectilinear-monotone [Wan03]. Recurrence [LN07, ALMN05]. Recurring [GCS08]. Recursion [dOBNL09]. Recursive [BCK09, DCCS01]. Red [MRP^{+}08]. reduce [BW03, GYA^{+}03, MGZ06, RSM^{+}05]. Reduced [HU02, HU09, KK09, RMPJ08, ALMN05, Hia02, LB04]. reduced-area [Hia02]. Reducing [BSPG08, GCNS08, Kim09, MS02, PP06, San06, SC04, ZL07, AF05, BMO04, CK06, LA03b]. Reduction [BDL09, BDK^{+}05, KSA03, Ker08, SLZ05b, XLSF07, BO03, BD05, FLW03, ICM03, KGM^{+}05, PR02c, PKEG04a, SDB03, VLG06]. Redundancy [KB08a, WYZ08, DBB00, HSU03, KM06a]. Redundant [HAH08, Kor09, NWA08, AKS^{+}03, CHC05, GS03, KM06b, MSM02, PS04, PGK01, PR06, San03, TM00, WHBG02, EP09]. Reed [DN06, Fal03, DM00b, FDBS05a]. references [KGMS00, MKAP05]. Refinement [VVSA07]. Reflective [RKFTF03, SM00]. regenerative [MB01]. region [CLF05, KI04]. Register [BSPG08, EBPG06, LMW07, Tou07, GYA^{+}03, RSM^{+}05, TA05, TSD01, WK03]. registers [MVG^{+}04]. regrouping [DK04]. Regular [CC07, Lee09, BCV01, CCMS02, MAD03]. Rejuvenation [SAT09]. Rekeying [SH09, VW05]. Related [CLV05, WSTJ09, CV00]. Relation [FMRR07, KR01]. Relations [BCK09]. Relationship [FH06, YOC^{+}01]. Relaxed [KL09, LP01a]. Relay [LX07, SCZ01]. relays [WLD06]. release [MVG^{+}04]. Reliability [AZ09, DPZ07, DLT07, EP09, TB06, WHZ09, CU02, Car02, CAK04, CMD05, SV06, Zha05]. Reliability-Aware [AZ09]. reliability-oriented [SV06]. Reliable [NS03, PV03, PRO03, UB03]. relieving [TNS03]. Remainder [BG08, CO09, YKLM03, Par03]. remapping [RPH01]. Remote [MG02, OS07, BW03, LC02]. Removing [ABLP07]. rename [San06]. rendering [PLK^{+}03]. reorder [KPEG04]. reordering [WH03]. Repair [Ano04h, DGZA03]. repairable [Car02]. Replacement [BGH07, CX07, KS08, BWTE04, CNM^{+}02, HK00b, JD06, JZ05, ZXL02]. Replacing [CNG^{+}09]. rePLay [PL01]. Replica [PBWB00]. replicated [LCC02]. Replication [Zha05, ZVT09, AF05, CKBF03]. Representation [BIN06, CKA06, NWA08, PB07b, CHC05, GMQS02, GS03, KB03, KGB05, LLC02b, PS04, RK05, TPT06, WFMSW00, WHBG02]. Representations [Kor09, NS09, KM06b, PGK01, SSK03, Sun05]. Representing [JP07, JPSR07]. Request [LX09]. requests [CBB^{+}02, KHM01]. required [Par03, SW00]. Requirements [gLMK07]. Rerouting [JSW07]. rescheduling [CS00]. rescheduling-based [CS00]. Reseeding [WBW08]. Reseeding-Based [WBW08]. Reservation [Ano05f, SL06, CBT05, LLA04]. Reservation-Based [Ano05f, CBT05, LLA04]. residual [CU02]. Residue [BG08, VLG06]. Resilience [LZ07]. Resilient [BCdSFL09, GSS08, SHK06]. Resistant [Gir06, MRL06]. resizing [PKG06]. resolution [DSK00, XXF03]. Resource [FPFP06, HUN07, JVG07, LG06, MBG08, MLB^{+}09, NH06, SKJ07, SL06, SKK^{+}09, ASB03, CC03, ET03, GL02, HR02, LWRJ06, MKS03, MG02, OM04, RL04]. Resource-Constrained [MLB^{+}09]. resources [RB05]. Response [Bal04, BNRB09, PKR04]. Response-Time [BNRB09]. responses [BDG03]. Responsive [LP09]. restoration [MKS03]. restricted [CN03]. restructuring [TP02a]. Results [XHLC08, BFG03, CWC02, Ste02]. Retirement [MBR^{+}09]. retrieval [AKS^{+}03, Rho03]. Return [OVB^{+}06]. reuse [HL00, HL03]. Revealed [BBD^{+}08a]. reveals [KJD02]. Reverse [PKCD07]. Reviewers [Ano05g, Ano06a, Ano07b, Ano08e, Ano09a, Ano00a, Ano01a, Ano03, Ano04a]. Revisited [SK00, CMK03, FHLM04]. reward [AMMMA01]. reward-based [AMMMA01]. rewarded [Car04]. Rewriting [VVSA07]. RFID [LSBV08, OS07]. Right [HE05, JY00, KGB05, RK05]. rigor [YLH05]. Rigorous [XRR^{+}02]. Rijndael [SK02]. Ring [GS09, HJ01, KB03, MMRT06]. Rio [NC01]. Risk [KM07, SHK06]. Risk-resilient [SHK06]. RNS [BI04, CN03, Hia00, Hia02, PBB07, VNBE03]. ROBDD [TMD05]. Robin [SCK06, YGB08, YD09]. Robust [WZX05, ZJ08, Ali02]. Robustness [BFG03]. ROC [OBB^{+}02]. ROC-1 [OBB^{+}02]. Root [EIM^{+}00, ELMT00, Kor05, MBCP07, PB02, PBLM08, RHMLL08, LA03a]. Roots [HCK09]. Rotation [ALB00]. Rotation/Vectoring [ALB00]. Round [SCK06, YGB08, YD09]. Round-Robin [SCK06, YGB08, YD09]. Rounded [BM08b, BM08c, BMR04]. Rounding [BM08b, ES00, LL09, Par00, ML01]. Route [Ano04h, SX09, DGZA03, SV06, TC02]. routed [SAOKM01, WT00]. Router [KS07, SHR08, EKK04, LS04c, LS04b, LS05, LKS05, RMB05, SK04]. router-table [LS05, SK04]. Router-Tables [KS07, LS04c, LS04b, LKS05]. routers [NB02, XS02]. Routing [ABF^{+}07, ACPP08, CXP06, LHC^{+}08, LLLD06, MR06b, MS09, Mel07, OMG07, PGVB08, WLS09, XZP09, XWC^{+}08, XM07, BB03b, CCK00, Dat05, EKK04, FLM^{+}03, GNF^{+}06, HS04, KM06a, KNS01, LWG01, LS02, NASR04, PS06a, SDB03, Sch01, SW00, SS06, THW03, Wan03, WDL^{+}03, Wu03, XTX06]. Routing-Conscious [ABF^{+}07]. Row [JSW07, LLW07]. RS [CPRS07]. RSA [BI04, Gir06, YKLM03]. RT [SM00]. RT-CRM [SM00]. RTL [HSS^{+}08, VVSA07]. Run [RK07, CC03]. Run-Length [RK07, CC03]. runahead [MKAP05]. Runtime [TEG09, MTB^{+}01, RVJ^{+}01]. RVR [CU02].

s [HV06]. saboteurs [SAKR03]. Safe [KL09]. Safecharts [HCL07]. Safety [HCL07]. Safety-Critical [HCL07]. Sampling [JLJ07]. SAN [SLZ05a]. SANs [ZSXZ07]. SAT [BBH06, SdBF04]. SAT-Solving [BBH06]. satellite [ET01]. satellite-terrestrial [ET01]. Satisfactions [YRF08]. satisfiability [ASM06, ACCL06]. Saturated [PKCD09]. Saturation [SBAB00]. save [TPT06, UK01]. saving [KPDS01]. Scalable [CH07a, CD09, FPFP06, HXW09, MRS00, RM09, VAZSR07, WYYZ06, XPGP06, ZP08b, BLAA01, GL05, KYS05, MAD03, MKBG00, PXK^{+}02, ST03, TK03]. ScaLAPACK [BBP^{+}01]. Scalar [AH08, VOT08, Mis06]. Scale [CWC07, FPFP06, Sum08, Lin01]. Scaled [CV08]. Scaling [HASL07, LCL07, SW05, ZX07, ZR07, KL03a, KL04, LS04a, PP06]. Scan [HSLN08, XLSF07, XCF07, BO03, BO05, Li05, LC04, NAH02, PR02c, PR04d, PR06, SCZ01, XP04]. Scan-Based [XCF07, BO05, SCZ01]. scan-BIST [LC04]. scatternet [PBC03]. Schduler [YD09]. Schedulability [BB04, DZB08, KM03, LKSS06, LLWS08, ZB09]. Scheduled [KGA01, QD04]. Scheduler [CH07a, GBD07, LP09, SKJ07, YGB08]. Schedules [FSL07]. Scheduling [ACS^{+}09, BWR^{+}07, BNRB09, CLLL09, CHL09, GMSC09, GCS08, GS09, HL01, LKSS06, LZ07, LP09, LCC06, LOP07, LS09a, MEB01, NASK^{+}08, DS00, NBAR08, PKCD07, PY09, QX08, SL06, THC^{+}08, WZ07, WRJ07, XQ06, XHLC08, ZB09, ZVT09, ZX07, ASL04, AMMMA01, AMMMA04, Bar05, BG03, BFG03, Bar04, BF06, BKM^{+}06, BLCA02, CN05, CJ01, DRC05, GCI06, HSW03, HB01a, HH04, ICM03, KMPE02, KHM01, KL03a, KL04, KYL02, KCKC05, LP00, LR04, LWRJ06, LHLL06, LAG^{+}01, MRY06, MAMMA03, NB02, Pal05b, PY05a, PC02a, PBWB00, Ros04, RY05, San03, SHK06, SWP04, TSC^{+}00, WZSP04, WRJL05, ZY06, ZS00]. Scheme [CWZL08, DHS08, HX08, PDS07, RK06, Sum08, Voy08, WFP08, AWY01, BB03b, CNM^{+}02, GVMC^{+}06, LA03b, LLC02b, LCC02, LCL03, MB03, TB03, TA05, WD04, YJMS05, YD02, Zha03]. Schemes [FH07c, PC07a, SZL08, TSS08, BMO04, Fan03b, KL03b, MRT01, Tze04, YB03]. Scientific [CD09]. Scott [HKM09]. Scrambler [Clu07]. screening [MDM04]. SD [ZW08]. SD-MAC [ZW08]. SEA [Xie08]. Search [NKSG09, RM09, WH06, SDB03]. Searching [SLZ05b]. Secion [Ano07i]. SecLib [GSH^{+}08]. Second [PRB09, Yan08, YL06a]. Second-Order [Yan08, YL06a]. Secondary [SMM05]. Section [Ano04e, Ano07e, Ano07f, Ano07g, Ano07d, Ano07h, APV09, BK06, DLL07, GDM07, Har06, KMMS09, LML06, Mar08, Mud05, ST08, SGK08, HKL01, JR02, KP03, LT03, SLG02]. Secure [FPL^{+}08, HvAL09, Lev07, NHSC07, SHYV06, SMS07, LRJ00, Lu05, Tze02a, Tze02b, YGZ05]. Security [Ano08f, Ano08g, Ano09h, DHS08, DMS^{+}09, GSH^{+}08, LSBV08, OVB^{+}06, SXZ^{+}06, MDS02, Rya04, SMBY05, SHK06, XQ06]. security-assured [SHK06]. security-critical [XQ06]. seed [Kag03]. Segment [CL07]. select [EVN03]. select-prefix [EVN03]. selected [Had05]. selecting [PvST02, ZC05b]. Selection [ACPP08, BMlX08, CX07, HWW07, Kor05, LLW07, MK07, yFBC03, HPS02, ML01]. Selective [HSLN08, KKN07]. Self [AGPP09, BP06, CD09, CUT00, DH06, GL03, Kar06, KC07a, MFR00, OZ06, WGZ^{+}08, Ano00c, KHP00, KPGX05, MOK04, Pie02a, Pie02b, SMMMK03, YFKB04]. Self-Adaptive [WGZ^{+}08]. Self-Checking [OZ06, MFR00, Pie02a]. Self-Checkpointing [KC07a]. Self-Healing [CD09]. self-management [SMMMK03]. Self-modifiable [GL03]. Self-organizing [BP06]. self-similar [MOK04]. Self-stabilizing [DH06, Kar06]. self-test [Ano00c, KHP00]. Self-Testing [AGPP09, KPGX05, Pie02b]. Self-Timed [CUT00, YFKB04]. Semantic [LLS02a]. Semantically [PRO03]. semantics [KM00]. Semiconcurrent [AFPS01]. Sensing [CRS09, NASK^{+}08]. Sensitive [Ayd07, LSS09, LAG^{+}01]. Sensitivity [SZGS09]. Sensitivity-Based [SZGS09]. Sensor [AD08, BF08, CLLL09, DMS^{+}09, KLT07, LSS09, LKTT08, LX07, LLLD06, MY07, MP09, PH08, RM09, WHT09, XS07, ZMY08, ZCWZ08, dOBNL09, BP06, CIQC02, CLF05, CSR04, DPIK05, KI04, LWF03, LR06, LDH06, PS06a, PR04a, SC05, ZC05b]. Sensors [KLT07]. Sep [HU09]. separation [RK03]. Sequence [YGB08, Alw06, DU04, PR02a]. Sequences [HU09, HU02, HU06, PR00c, PR02a, PR02b, PR04d, Voy05]. sequencing [GYA^{+}03]. Sequential [BP09, GSS08, PGPZ00, CH06, DBB00, Fro00, Fuj00b, HCC^{+}00, MSMS04, NAH02, PR00c, PR00b, PR02a, PR04c, PR06, RMH05, SC04, TM00, XXF03]. Serial [DH05, HRM09, KWP06, VOL08, LPS00, YY01]. serial-parallel [LPS00]. Series [Ano08f, Ano08g, Ano09h]. Server [WYMG09, ZBI^{+}07, ASB03, KD02, KKK03, MAMMA03]. Servers [CKS^{+}08, HASL07, TTC09, WX06, WZ07, CM02, GL02, HH04, WMZH02, WZX05, WZSP04, YNOJ02]. Service [DPZ07, LSS09, CM02, CTA02a, CTA02b, CBB^{+}02, CKDS02, EGP03, LSV00, LMS02, LMS04, SS06, THWH01, THW03, XL03]. Services [DLT07, MT09, CT03, CD03, CKDS02, HSU03, KL03b, LMS02, MPAS03, XL03]. Session [CP02, KKKB05]. Session-based [CP02]. Set [BBGM08, NKSG09, SHR08, VPG^{+}08, WDY08, Ano00c, BD05, CT05, CN03, KHP00, LCKR03, PB04, PR02c, RZ04, ST04, TMD05]. sets [KS00b, PR04b, PR04d, SS02, WLD06]. setup [Li05]. setup-time [Li05]. SEU [TFCW07]. SEU-Induced [TFCW07]. Seven [FH07a, Mon05]. Seven-Term [FH07a, Mon05]. Several [PAW07]. shaped [FWCL06]. shaping [AVVY05, WTL04]. Shared [AGPP09, BGB08, GS00a]. Shared-Memory [AGPP09]. Sharing [CPRS07, Car03, MG02, RG05]. Sharp [GCS08]. Shift [Kor09, Yan08, TPT06]. shift-unsafe [TPT06]. Shifted [FH06, PCH06]. Shifters [HL09]. shipping [LLS02b]. Short [VOT08, PKEG04a]. short-lived [PKEG04a]. Short-Memory [VOT08]. Shortest [CY09, MO06, XT02, Zac06]. should [LKF03]. side [CMCJ04]. side-channel [CMCJ04]. sided [YCL01]. Sign [PLP08, SLS04]. Sign-On [PLP08]. signaling [LA03b]. Signals [XCF07]. Signature [PB07a, Che05]. Signed [CV00, JY00, COP^{+}06, RK05, SLS04]. signed-binary [RK05]. Signed-Digit [JY00, SLS04]. Silent [LBL01]. silicon [KZP05]. Silver [Ano08f, Ano08g, Ano09h]. SIMD [DT05, Par04, TJB03, TLS09, YSL00]. similar [MOK04]. Similarity [KH00, JPEJ06, KM00, KM03]. Similarity-based [KH00, KM00]. Simple [Gir06, KG06, RK06, YW06, CMCJ04, KR04]. simplex [XTX06]. Simulation [Ano04d, GE08, GE09, Har06, JPSR07, JLJ07, NPB07, SG07, SM09, WWC06, WH06, YL06b, BJHW00, CU02, Gho01, KHB02, MPP^{+}05, YLH05]. Simulation-Based [Ano04d, Har06, WWC06, WH06]. Simulations [YSLH07, LLVC04]. simulators [YL06b]. Simultaneous [HR06, SP07, ZYZ09, PGK01]. Single [PLP08, PB07b, UBWF08, WWH^{+}07, BDG03, MPP^{+}05, PR02a, UF03]. single-chip [MPP^{+}05]. Single-Image [WWH^{+}07]. sinusoidal [Alw06]. sites [CP02]. Six [FH07a, Mon05]. size [BD05, CS00, DNVG05, HLM00, OPZ00, ZZZ04]. sizes [KSA03, Sez04]. sizing [HR06, IC02]. skewed [ABP^{+}05, Sez04]. Slack [AM08]. slave [LZ06]. Slicing [GVB^{+}09, SG07]. slipstream [KR04]. Slowdown [JG06, WZX05]. Small [ELMT00, RK06, GMQS02, JPJ^{+}04, YL06a, Zha02, Zha05, ZV02]. Smart [MG08, BGL^{+}03, MDS02]. smart-card [MDS02]. SmashGuard [OVB^{+}06]. SMT [BG05, CKS^{+}06, CV08]. SMT/CMP [BG05]. SMTs [CKS^{+}06]. SNAP [MP09]. Snapshots [XYR^{+}09]. SoC [BFP^{+}06, LML06]. Society [Ano07a, Ano07k, Ano07l, Ano09g, Ano09f, Ano04f, Ano04g, Ano06f, Ano06h, Ano06g, Ano07m]. SOCs [KTK06, SC07a, XN06, GMSC09, RK07]. Soft [BC01b, WHZ09, DH06, KLA^{+}03, MBF^{+}04]. Software [AGPP09, CS09b, CMAB09, CHA^{+}09b, DHH^{+}06, KPGX05, MBG08, PC07b, SAT09, TP04, Tou07, ACK^{+}03, DKV^{+}01, GPS05, HSW03, HJS04, MS03, MBF^{+}04, PP06, PNRP04, SXZ^{+}06, TTA^{+}02, TGKL03, TSD01, VKI^{+}03, YWV05, RLPV05]. Software-Based [AGPP09, CS09b, CMAB09, KPGX05].

Software-Defined [MBG08]. software-implemented [ACK^{+}03, TGKL03]. Solomon [FDBS05a]. Solomon-like [FDBS05a]. Solution [ARSM07, CS08b, OVB^{+}06, QD08, CSS02, RTD00, TPT06, Voy05]. Solutions [WDY08, CMCJ04, LS02, OR00]. Solve [BCK09]. Solver [SPS08]. solvers [BBP^{+}01, SdBF04]. Solving [BBH06, WAB^{+}02]. Some [BT05, ELMT00, YL06a, Car04, RY05]. Sort [NZ07]. Sorting [RS08b, BB05, OPZ00]. Source [HBF09, RK06, BGL^{+}03, FLM^{+}03]. Sources [PAW07, SMN07]. Space [ABF^{+}07, BCdSFL09, CRJ07, FH07b, GE09, LS08, MS09, WH06, BDG03, CK05, EL02, KRCB01, SDB03, WMZH02]. Space-Efficient [LS08]. Space-Optimal [CRJ07]. spaced [Gol02, LLL01]. Spaces [GE08]. spanning [Res01]. spare [HT00]. spares [YBB00]. Sparse [CL06, LCL03]. Spatial [CH07a, LMV^{+}08, SS09b, WHT09]. Speaker [lFB07, PAW07, ZMM07]. Special [AK09, Ano04e, Ano04c, Ano04d, Ano05c, Ano05d, Ano05e, Ano06e, Ano06c, Ano06d, Ano07c, Ano07e, Ano07f, Ano07g, Ano07i, Ano07j, Ano07d, Ano07h, APV09, BK06, DLL07, GDM07, Har06, KK00b, KMMS09, LML06, Mar08, Mud05, SB05, ST08, SGK08, HKL01, JR02, KP03, LHJL05, LT03, RHK03, SLG02, UB03]. Special-Purpose [Ano07j, Ano07h, SGK08]. Species [SS09a]. Specific [HCK09, BDBR05, CZM05, KKP00]. Specification [AHS06]. Specifications [PF08, BJHW00, PY05b]. Specified [AC09, PR00b]. Specify [MNR08]. Spectral [HWW07, Ker08, Yan08, YL06a]. spectral- [YL06a]. Spectral-Null [Yan08]. spectrum [CH06, LCK^{+}01]. spectrum-based [CH06]. speculation [AGG06, yFBC03, GG01]. Speculative [LOP07, MGT04, TA05]. Speech [Ano06e, Ano06c, Ano06d, lFB07, GDM07, HW07, Moo07, OM07]. Speech-Based [Ano06e, Ano06c, Ano06d, GDM07, Moo07]. Speed [ANPS07, BR09, CS08a, FN09, JK09, KNE^{+}00, KG06, KS05, PB02, POMB05, SC07a, XPGP06, YJ00a, YSLH07, BGL^{+}03, DN05, Hia02, MRS00, ORM05, San06, WWSH04, XS02]. Speeding [KGB07]. speedup [YKLM03]. spill [RSM^{+}05]. spills [GYA^{+}03]. Spoken [SMN07]. sporadic [BF06]. spot [AAvdG03, SAOKM01]. Square [EIM^{+}00, ELMT00, HCK09, Kor05, LA03a, PB02, RHMLL08, CHL01]. Square-root [LA03a]. squarer [Wu02a, Wu02b]. Squaring [HRM09]. SRAM [GVMC^{+}06, HBF09, PL09b, SV06, ZIPL00]. SRAM-based [SV06, ZIPL00]. SRAMs [HvdG02]. SRT [Kor05]. Stabilizing [KG07, DH06, Kar06]. stable [Car02]. stack [KM03]. stage [ET01, YCL01]. stages [FB00]. stalls [BSN^{+}06]. standard [BBK^{+}03, YW06]. Star [MR06b, Fuj00a, HCH01, Res01, THWH01, HX08]. star-coupled [THWH01]. State [HBF09, JS09, Lom08b, Lom09, WAU^{+}08, CS04, Hie04, PR00a, WDL^{+}03, XP04]. states [PR00b]. Static [AAvdG03, PR04d, SGB08, ZL04, KH00, PR00c]. Statistical [GE08, GE09, JLJ07, PRB09, SMN07, ZX07, Hey03, NYC05, YECV02, YLH05]. Statistically [PCL06]. statistics [KJD02]. steering [BBM^{+}03, MSMS04]. Steganography [HvAL09]. Stepwise [VVSA07]. STI [WKS^{+}05]. Stimuli [FFJ^{+}06]. stimulus [MK06]. Stochastic [BC01a, BC01b, KL09, JLN04, KDB^{+}05, MO06, MB01]. Storage [HX08, KPJ^{+}09, WWH^{+}07, WYZ08, WYMG09, XW08, Xie08, AKS^{+}03, CT03, HH04, LP06b, PR02c, RMB05, ZZ05]. storage-based [PR02c]. Store [TIVYL09, LBL01]. stores [LBL01]. STORM [FPFP06]. storms [TNS03]. Strategies [JVG07, LOP07, ZVT09, FN04, LLVA01, MPP^{+}05, PvST02, ZXL02]. Strategy [ACPP08, QX08, XQ08, Xie08, YGL07, ZJ08, HMM06, KYY^{+}03, Red03]. Stream [LX09, SSC03]. streamed [MWA^{+}00]. Streaming [iM02, AVVY05, PCL06]. Streams [DVJP07, JS09, SCG08, WZ07, HHTH00, WZSP04]. strided [SE05]. Stripe [BP07]. Striping [BP09, Xie08]. Striping-Aware [BP09]. Striping-Based [Xie08]. Strongly [HC08a, HC08b, BCV01]. Structural [LO04]. Structure [LKTT08, GR04, Hia00]. Structured [DLT07, Xie08, Ros04, WH03]. Structures [FFLTM09, KN08, CCMS02, Hia02, Kag03, MRT01, SCZ01]. Structuring [BS08]. Student [Ano09f]. Study [BBGM08, KL08, TFCW07, NASR04, Pal05a]. Studying [CKS^{+}08]. style [TJB03]. subarrays [JS06]. subblock [HL03]. subcube [CWC02]. subcube-connected [CWC02]. Subject [Ano00d, Ano01d, Ano02c]. Sublinear [DJJ^{+}08, Lee12]. Sublogarithmic [HPS02]. Subquadratic [FH07b, FH07c, Sun04]. subscribers [TL02]. Substitution [JSW^{+}06]. subsumes [LCK^{+}01]. Subtraction [CCY00]. Success [FFJ^{+}06]. suffix [BB05]. Sum [KHP00, Ano00c, LZ06, SB01]. sum-of-products [SB01]. Sums [BM08b]. superblocks [MEB01]. Supercomputer [MK07, Con00]. supercomputing [GW04]. SuperGrid [Kum00]. Superscalar [NBAR08, TW08, AGG06, GYA^{+}03, KPEG04, PKEG04b, PKG06, ZK01]. Supervisory [PNRP04]. supply [HAK05]. Support [PMR02, PPB^{+}07, TLS09, ZBI^{+}07, ASF^{+}01, BLAA01, DRC05, GT06, HMM06, HL00, JMH02, KCHS04, LCKR03, MPAS03, OM04, OBB^{+}02, Sez04]. Supporting [CLE^{+}07, HXW09, LLP09, WKS^{+}05, AHS06, dALB03, RDH^{+}01, THWH01]. surveillance [BP06, CIQC02]. survey [SdBF04]. survivable [HSU03].

Susceptibility [MBF^{+}04, ORM07]. Sustaining [XL03]. SW [CK06]. Switch [BW03, Mha09, Sue09, FWC02, FLW03, FWCL06, IB00, SMSM01, SWCC00]. Switched [GBD07, MOK04]. Switches [GS09, JSW07, LS09a, PY09, HT00, JS06, MRS00, PY05a]. switching [HSY00, Pal05a, YCL01]. symbol [KF04, MT06]. Symbolic [PCR01, HB01a]. Symmetric [AGPP09, BKM07, KLCV08]. symmetrical [EKK04]. symmetries [Fal03]. symmetry [ASM06, HLM00]. symposium [Raj00]. Synchro [HBH05]. Synchro-tokens [HBH05]. Synchronization [CRJ07, HDQK09, LR06, LLA04, MYL^{+}01]. Synchronized [Ksh07]. Synchronous [HSH01, SGB08, TPB^{+}08, YRVS09, AK00, IM06, LFA04, PR00c, PR00b, PR02a, PR04c, PR06, TM00, XXF03]. Synergy [lFB07, CKS^{+}06]. Synonym [QD08]. Synthesis [NPB07, ZW08, CT05, KGS00, LRJ00, MM04, NC03]. Synthesizable [WAU^{+}08]. System [BCL07, FFLTM09, KL03b, LBP08, LX09, LP07, LF09, MDJM05, NASK^{+}08, SHYV06, WZL08, ZSXZ07, ZCWZ08, ASF^{+}01, Ali02, AS03c, BFP^{+}06, BG05, CKBF03, CC03, CAK04, CT05, HCC^{+}00, IC02, ICM03, KC01, LP06a, LwJKW03, LP01b, LP06b, LHLL06, MFC02, OM04, Par04, PEP06, SAJ02, SXZ^{+}06, SLZ05a, SLL^{+}00, SLS04, WAB^{+}02, Wan04, XRR^{+}02, YFKB04, WMZH02]. System-Level [BCL07, Ali02]. system-on-a-chip [CC03, IC02]. system-on-chip [ICM03, LP06a]. Systematic [BET07, HAH08, ZP01, ABD^{+}04, PEP06, vdGT03]. SystemC [PF08]. Systems [AM08, ARSM07, AC09, Ano04c, Ano05f, Ano05c, Ano05d, Ano05e, Ano06e, Ano06c, Ano06d, Ano07g, Ano07i, BRC08, BGB08, BHS09, BVM07, CGS07, DZB08, DLL07, DRC05, GVB^{+}09, GDM07, HC08b, HCL07, JG06, Kap09, KLCV08, KOA07, LP09, LJS^{+}07, LML06, Mel07, MNR08, NKY08, OM07, PKCD07, PC07b, PL09a, PCG07, PPB^{+}07, QX08, RC06, SMN07, SKK^{+}09, SC06, ST08, VVSA07, WYZ08, WCLK09, XWC^{+}08, XQ08, XW08, Xie08, YT07, ZB09, ZP08b, ASB03, AWY01, ACCL06, ABD^{+}04, AFR02, AAS00, BLAA01, Bar04, BF06, BCF^{+}03, BMO04, BFR01, BBR^{+}06, BBL01, BFG01, BB04, CBS02, CBT05, Cao02, Car02, CT03, CMD05, DSK00, ET01, Fav06, FN04, Fet03, GL02, HKL01, HSY00, HKA01, HR02, HL02, HMM06, HL01, JR02, JLN04, KV02, KHM03, KGS00, KYY^{+}03, KLA^{+}03]. systems [KM06a, KOHC03, KBK03, KCHS04, KDB^{+}05, KP03, KL03a, KL04, KH00, KL03b, LTCH05, LFA04, LSV00, LLS02b, LZ06, LMM00, dALB03, LCC02, LLA04, LO04, MT02, MAMMA03, MME04, MM04, MBF^{+}04, MSM02, MB01, PMR02, PP06, PNRP04, PBWB00, RK03, RVJ^{+}01, RG05, RL04, RKFTF03, RPH01, SdBF04, Sma03, SLG02, SFJ03, SWP04, Sto00, UB03, WH03, WKS^{+}05, Wil01, WRJL05, XLN05, YBB00, ZWST03]. Systems-on-a [Ano04c]. systems-on-a-chip [BBR^{+}06]. Systems-on-Chip [LML06, PPB^{+}07]. systems-on-chips [HMM06]. Systolic [BSH09, CBC07, DH05, Alw06, LLL01, LHJL05, WG00, WWSH04, YS03].

t [CT09]. t-Diagnosability [CT09]. Table [Has00, JMH02, KB08b, WCKD04, dDT05, JPJ^{+}04, KKKB05, LY01, LS05, SK04, XS02]. Table-Based [Has00]. Tables [ASMD07, KS07, MDJM05, LS04c, LS04b, LKS05, MGZ06, YL06a]. Tags [OS07]. TAM [SC07a]. Tamper [JSW^{+}06]. Tang [Jha03]. Target [CLLL09, ZCWZ08, BP06, CIQC02, CSR04]. Target-Oriented [CLLL09]. Task [HLLR00, JG06, KL09, LLA04, QX08, XQ08, Bar05, Bar04, BF06, BKM^{+}06, PC02a]. task-level [Bar05]. Tasks [Ayd07, AZ09, BRC08, CHL09, GCS08, LZ07, MLB^{+}09, PC07b, SL06, ASL04, AMMMA01, AMMMA04, BDBR05, CJ01, HSW03, LP00, LMM00, DS00, SWP04]. Tate [SKG09]. taxonomy [SDT04]. Taylor [CKA06]. TC [Ano04i, Ano04j, Ano09a]. TCAM [CS08a, ANPS07, Cha06, CWZL08, FN09, RMB05, WCKD04, ZCW^{+}06]. TCAM-Based [CS08a, ANPS07, FN09, RMB05, ZCW^{+}06]. TCOT [KPDS02]. TCP [CC02, LP07, WTL04]. Team [MRP^{+}08, MRP^{+}08]. Team/Blue [MRP^{+}08]. Technique [AM08, CMS08, JP07, JPSR07, LLW07, LMW07, MKP06, TEG09, BBR^{+}06, BW03, FLW03, KRCB01, MK06, SLS04, ZC05b]. Techniques [AFM^{+}06, HY08, MDJM05, SAYN09, ACK^{+}03, CU02, DKV^{+}01, GBHL06, KL03a, KL04, LM00, PXK^{+}02, WRJL05]. Technologies [Ano07d, CV08, LCCA02, Raj00]. Technology [OM07]. telephone [Lin01]. telephony [EGP03]. Temperature [CKS^{+}08, HSS^{+}08]. Temperature-Aware [HSS^{+}08]. Temporal [Ksh07, WBW08, WHT09, Car03, KLA^{+}03, THWH01, THW03]. Teraflops [Con00]. Term [FH07a, Mon05, VVSA07, WHT09, WK06]. termination [HHTH00]. terms [BCV01]. Ternary [KPP09]. terrestrial [ET01]. TESH [MJ02]. Test [Ano04c, BBD^{+}08a, CC03, CKC^{+}08, FFS02, GMSC09, IC02, ICM03, KKN07, LL09, LML06, LLWS08, MNR08, NKSG09, NTA08, Par00, RMPJ08, Voy05, VPG^{+}08, WWC06, XPGP06, XLSF07, XCF07, Ano00c, BO03, BDG03, CN05, DCCS01, DU04, Fuj00b, FS00, Had05, HL05, HBH05, KTK06, KHP00, LP06a, MK06, NAH02, PR00a, PR00c, PR02a, PR02b, PR02c, PR04c, PR04b, PR04d, PGPZ00, RZ04, SCZ01, SGB00, XN06, YHIO02]. Test-Data [KKN07]. Testability [XPGP06, FFS02, XXF03]. testable [DM00b, NC03]. tester [ICM03]. Testing [AGPP09, CH06, GMSC09, Hie04, HSLN08, Imr07, LLPC04, MRP^{+}08, MRT01, PY05b, RK07, SC07a, ZIPL00, BJHW00, HWI^{+}02, KJD02, KPGX05, MDM04, Pie02b, PR00b, RTD00, SGB00]. Tests [DZB08, TGKL03, HvdG02, NR04, PR04d, PR06, vdGT03]. text [AT05]. textile [KM06a]. textiles [SMMMK03]. th [MBCP07]. Their [PPB^{+}07, YRF08, Has01, MS03, SSST06, Sma03]. them [MDM04]. theorem [FCB04, BG08, CO09, YKLM03]. Theoretic [NH06, Par00, RG05, RTD00]. theoretical [CWC02]. Theory [PCC02, BK00, FWCL06, GL02, HR06, MRY06]. Thermal [HSS^{+}08]. Thinning [Fan03b, WFP08]. Thread [MGT04]. thread-level [MGT04]. threat [MDS02]. Three [BBD^{+}08b, ES00, FP09, FDBS05a, GPS05, YCL01]. Three-Dimensional [FP09]. three-stage [YCL01]. Threshold [BCGG00, gLMK07, CV00, GL02]. Threshold-based [BCGG00, GL02]. Thresholds [GCS08]. Throttling [LG06]. Throttling-Based [LG06]. Throughput [LA05, LX09, LMV^{+}08, NBAR08, SGB08, TW08, BG05, HV06, MKA03, Ray06]. Throughput-Buffering [SGB08]. Tight [Par03, SW00]. Time [AM08, AC09, Ano05f, Ayd07, AZ09, BRC08, BWR^{+}07, BGB08, BNRB09, BVM07, CHL09, CRJ07, DZB08, FSL07, JG06, JS09, LKSS06, MNR08, NASK^{+}08, PC07b, SL06, SKK^{+}09, SC06, TPB^{+}08, THC^{+}08, WCLK09, WZ07, WRJL05, WRJ07, XHLC08, ZB09, ZX07, AS00, Ali02, AHS06, ACCL06, AAS00, AMMMA01, AMMMA04, BCF^{+}03, BO03, BBL01, BDBR05, BW03, BJHW00, CBS02, CBT05, CK05, CJ01, DSK00, ESE05, FZM00, GS00a, GCI06, HHTH00, HSH01, HSW03, HR02, HL02, HL01, JR02, KHM03, KD02, KLA^{+}03, KS00a, KDB^{+}05, KL03a, KL04, KM00, KH00, KLS01, KYL02, KKK03, KCKC05, LSV00, LwLH02, LR04, Li05, LWRJ06, LKS03, LMM00, dALB03, LLA04, LHLL06, MAMMA03, MME04, DS00, NS03, Pal05b, PGK01, PP06, PV03, PBWB00, Raj00, RL04, Rho03, RPH01, SM00, SLS04, SWP04, TB03]. time [TSC^{+}00, Voy05, WKS^{+}05, WZSP04, XQ06, XR04, ZBC02]. Time-Critical [SC06]. Time-Outs [MNR08]. Time-Shared [BGB08]. time-triggered [KHM03, PP06]. Time/utility [WRJL05]. Timed [CUT00, MNR08, Fet03, HSH01, LMS04, WAB^{+}02, YFKB04, ZBC02, LMS02]. timed-token [HSH01]. timely [VC02]. timeout [KPDS02]. timeout-based [KPDS02]. timer [KBK03]. Times [Ano05f, CBT05, LZ06, PR02a, TL02]. timestamp [LLSC02]. timewheel [MFC02]. Timing [CNG^{+}09, gLMK07, RPH01, UBWF08, YRF08, HL01]. Timing-Centric [CNG^{+}09]. TLB [Sez04]. TLBs [JM01, JJ06]. TLM [PF08]. TMR [GS00a]. Token [GS09, HSH01, LMS04, ZBC02, LMS02]. Token-Ring [GS09]. tokens [HBH05]. Tolerance [Ano04e, BK06, LMM00, SMS07, CSR04, Con00, GNF^{+}06, KS00a, PG01, YOC^{+}01]. tolerances [TGKL03]. Tolerant [CWC07, CL09, HY08, KL08, MP09, TFH07, XZP09, ZVT09, Car02, CLF05, Dat05, FN04, GS00a, Gro04, HSW03, HS04, HT00, IO03, KKP00, KI04, dALB03, LWG01, LDH06, MP03, MKBG00, PS03, RKFTF03, Tsu00, Tze02a, Tze02b, Wan03, WDL^{+}03, Wu03, XRR^{+}02, YW04a, YJMS05, Zha02]. Tolerating [FB00, FS07, QD04, FDBS05a, FDBS05b]. Tool [CKS^{+}08, Fal03]. tools [Rya04]. topologies [AN05]. Topology [CCS^{+}04, PGVB08, XT02, Zac06]. tori [WT00, YW01]. Toroidal [MBS^{+}08]. torus [AE00]. totally [Pie02a]. Toueg [HMR02]. tower [BS04]. TPF [LP07]. TPG [Kag03]. Trace [RLPV05, BGJ^{+}05, JHZ01, LY02b, LJ04]. trace-compression [BGJ^{+}05]. Traces [SG07, YJ06]. tracing [NR04]. Track [BF08, OMG07, HT00]. Tracking [LKTT08, NASK^{+}08, ZCWZ08]. Tractable [PF08]. Trade [KPJ^{+}09, LCLV08, SGB08, YSLH07, Fan03a, HV06, MEB01, PGJ^{+}05, SD06]. Trade-Off [SGB08]. Trade-Offs [KPJ^{+}09, LCLV08, YSLH07, HV06, MEB01, PGJ^{+}05, SD06]. trading [RB05]. Traffic [LHC^{+}08, MS09, OMG07, MOK04, SAOKM01]. Traffic-Balanced [LHC^{+}08]. trails [YECV02]. transaction [KPDS02, KH00, LLSC02, YY01]. Transactions [Pra06b, LwLH02, XR04]. Transfer [LMW07, CS04, Tze04]. Transform [BP07, NZ07, ZR07, Biq05, Che04, JSD01, RZ04]. transformation [BK00, KRCB01, VPR04, YHIO02]. transformations [HPH^{+}04, KPDS01, PDS04]. transformed [GD03]. transforms [LPS00, RN04]. Transient [BSPG08, Car04, ORM07, BCGG00, LMM00, MFR00]. transit [FLM^{+}03]. Transition [NTA08]. translation [EAGS01, JJ06]. Transmission [KM07, KB08a, TB03, ZYZ09, CRL00]. Transparent [KHM03]. Transport [CLV05, GBD07]. transposition [SP02]. Transversal [BF08]. traps [QD04]. Traversal [WH06, PR04a]. Tree [CY09, DLT07, HAH08, HXW09, JVG07, LKTT08, NHSC07, VW05, DV04, Kar06, KNS01, LS05, Lu05, MYL^{+}01, Res01]. tree-based [KNS01, MYL^{+}01]. Tree-Structured [DLT07]. Trees [CL07, LS04b]. Triangular [MY07]. Trie [KS07]. Tries [LS08, LS09b]. Trigger [HSLN08]. Triggered [TPB^{+}08, KHM03, PP06]. Trinomials [RHMLL08, FD05, IST06]. Triple [HX08]. True [HBF09, SMS07]. truly [BGL^{+}03]. Truncated [RC06]. truth [TN00]. Tunable [SAYN09, XQ08, BABD03]. Tuple [BCdSFL09]. turn [Wu03]. Two [AD08, KKK03, MMPT08, MTHA08, TFH07, ZMY08, Ano06f, BSN^{+}06, CHL01, DSK00, ET01, LPS00, NASR04, ORM05, Pie02b, TSP00, WFMSW00, XN06]. Two-Dimensional [AD08, CHL01, LPS00, TSP00]. Two-Disk [TFH07]. Two-Layered [ZMY08]. two-level [WFMSW00]. two-pass [BSN^{+}06]. two-pattern [XN06]. Two-Rail [MTHA08, ORM05, Pie02b]. Two-version [KKK03]. type [SK01].

Ubiquitous [NASK^{+}08]. UCFS [WMZH02]. ultra [CDV^{+}05]. ultra-low [CDV^{+}05]. UML [WAU^{+}08]. Unattended [DMS^{+}09, MY07]. unauthorized [PR04a]. Unaware [JVG07]. uncertain [TSC^{+}00]. Understanding [ML00]. unfolded [WK03]. Unidirectional [ABT07, BET07, DH05, EB09, Tal05]. Unified [ALB00, JP07, NKY08, VKI^{+}03]. uniform [BG03, BFG03, HL02]. unifying [CSS02, LMS02]. Unimodal [WRJ07]. Uniprocessor [JM01, Uht05, KCHS04]. Unit [DT05, KK09, LN07, Har04]. Unithreaded [RLJ^{+}09]. Units [HW07, SBPV07, XPGP06, Car03, EP00, KC01, PKR04]. Universal [AT05, BK00, LPAM04, Sue09, Ano00c, FWC02, KHP00, KYS05, RZ04, SWCC00]. universally [Tze04]. Unix [KBK03]. Unix-based [KBK03]. unknown [PKR04]. Unknowns [WBW08]. unreliable [HMR02, LFA04]. Unroll [KMPE02]. Unroll-based [KMPE02]. unsafe [TPT06]. Unsorted [Imr07]. unstructured [XLN05]. Updatable [SHR08]. Update [WCKD04, LY01, LCCA02, XR04].

Updates [Cha09a, SHR08]. upgrading [TTA^{+}02]. upon [BFG03]. Upper [TK07, HLM00, Par03]. usable [Tze04]. Use [CL06, SZZ09, DK04, PR00b, TSD01]. used [LCK^{+}01]. useful [SM03]. User [ASB03, OM07, GL03, SD06, WH02, WMZH02]. user-friendly [WH02]. User-level [ASB03, SD06]. User-space [WMZH02]. uses [RSQL03]. Using [BIN06, BSH09, BM08b, CO09, CT09, CHA^{+}09b, ELMT00, EBPG06, FH07c, GR07, HR02, HWW07, HCL07, HSS^{+}08, JSM09, JSW^{+}06, KPJ^{+}09, KDM^{+}09, KJM^{+}09, Ksh07, LLWS08, LS08, MGZ06, MR06a, MP09, MRL06, MDJM05, NWA08, PB07a, PAW07, PCH06, POMB05, RMH06, Rho03, SNB07, SG07, SAT09, SHR08, SLZ05b, VVSA07, WLS09, XCF07, Yan08, ZCWZ08, ZVT09, BDG03, BD05, Car04, CC03, CCK00, CAK04, CHC05, CH06, CN03, CLW^{+}03, DHH^{+}06, DLBS03, EVN03, GS00a, GRV05, HR06, HAK05, HLM00, Hie04, HSU03, HL01, JPEJ06, JMH02, KM06a, KB03, KJD02, KSL05, KS00b, LVLL06, LCCA02, OPZ00, PS04, PR02a, PKR04, Ray06, RMH03b, RM06, SAKR03, TYT01, Tsu00, VPR04, VKI^{+}03, Wu02a, WHBG02, YW04b, YOC^{+}01, ZL04]. Utility [BWR^{+}07, CRJ06, WRJ07, LWRJ06, WRJL05]. Utilization [CMK03, WCLK09, ASL04, Bar04, LSP04].

V [FZM00]. V-NET [FZM00]. Validating [SMBS06]. Validation [Ano04d, Har06, SHYV06, Con00, LLSC02]. Value [LG09, MKP06, TSS08, BZ02, yFBC03, GG01, HL00, HL03, LY01, LY02b, LBL01, MGT04, ZL04, ZC05a]. valued [DM00b, SAJ02]. Values [BSPG08, Imr07, LLWS08, PKR04, TN00]. Variable [Ano05f, BWR^{+}07, LCKR03, SS00, SLZ05b, CBT05, NYC05, TMD05]. Variable-Precision [SS00]. variables [RSM^{+}05]. variance [LSV00]. Variation [MWK^{+}09, NBAR08, SC07b].

Variation-Aware [NBAR08]. variational [LML01]. VC [CCK00]. Vector [DT05, KK09, NKY08, VPG^{+}08, PR02b, TK00, Voy05]. Vector/SIMD [DT05]. Vectoring [ALB00]. Vectors [RMPJ08, Had05, KR01, RZ04, SE05]. velocity [PXK^{+}02]. Verifiable [FSL07]. Verification [AC09, CL06, CKA06, FFP07, JPSR07, SG07, VVSA07, ZMM07, NC01, RZ04, WH02]. Verified [BDL09, DLM09]. versatile [FZM00, HW00, HMR02]. version [KKK03, Mis06]. versus [HTKL08, WPP05, YSLH07]. Vertical [LF09]. vertices [CHL01]. Very [ALB00, CCY00, KJM^{+}09, ML01]. Very-High [ALB00, ML01]. VHDL [Gho01, KHB02, SAKR03, WAU^{+}08]. via [CLV05, DN06, MS02, SXZ^{+}06, ZMM07, ZC05a, ZL07]. video [AWY01, CT03, GW04, GL02, KHRR02]. video-on-demand [AWY01, GL02]. View [Kap09]. Violation [Ano04h, DGZA03]. Virtual [KJM^{+}09, QD08, TEG09, JM01, PY05a]. Virtualization [SAT09, ZSXZ07]. Virtualizing [OM04]. visibility [CPN^{+}06]. Vista [OM07]. Visualization [WGZ^{+}08]. Viterbi [Thi00]. VLIW [CS00, GSG05, JBV^{+}05, LLVA01]. VLSI [AS03a, AS03b, Che04, DN05, HAK05, HW00, HL05, HSLN08, JS06, JSW07, KPDS01, Low00, PKM00, SK02, TK00]. VLSI/WSI [Low00]. Voice [HWW07]. VoIP [DLBS03, LCCA02, SCG08]. Voltage [HASL07, KL03a, KL04, VAZSR07, ZX07, Bar05, LS04a, PP06]. Voltage-clock-scaling [KL03a, KL04]. voltages [HAK05]. Volume [XLSF07, BO03, ICM03]. Voting [PH08]. VPC [BGJ^{+}05, KJM^{+}09]. Vulnerability [PR04a].

Wait [CRJ07, PHA06]. Wait-Free [CRJ07, PHA06]. Wake [CH07a]. Wake-Up [CH07a]. Walk [ZSS08]. Walsh [JSD01]. wave [MSMS04, SCZ01]. wavelength [YW04b, ZY06]. wavelet [Che04, RN04]. way [LJ01]. WDDL [GSH^{+}08]. WDM [RDH^{+}01, YW04b, ZY06]. WDMA [THWH01]. WDMA-based [THWH01]. weak [JZ05]. Weakly [BBL01, CKBF03]. Wearable [Sma03, ABD^{+}04, MS03]. Web [BDK^{+}02, CP02, CKDS02, HASL07, PvST02, SX09, TC02, WMZH02, WX06, WK06, XL03]. Weight [CH07b, KGB05, PB04]. Weighted [XCF07, YOC^{+}01]. well [BFG01]. well-conditioned [BFG01]. Wheeler [BK00, YHIO02]. which [MT06]. whole [XV04]. whole-program [XV04]. Wide [CLE^{+}07, WGZ^{+}08, CD04, CYL01, LY01]. Wide-Area [WGZ^{+}08]. wide-issue [LY01]. wideband [KL03b]. Width [LV07, TK07, VT09, Jha03]. WiMax [SCG08]. Window [WZ07, SSST06, WZSP04]. Window-Constrained [WZ07, WZSP04]. windows [RSM^{+}05, OM07]. Wire [FN09]. Wire-Speed [FN09]. Wireless [AD08, CX07, DMS^{+}09, ICRSR^{+}09, KM07, KLT07, LCL07, LSS09, LLP09, LKTT08, LX07, LLLD06, MS09, NH06, PH08, PKCD07, RM09, WFP08, WHT09, WDY08, ZYZ09, ZW08, dOBNL09, CDV^{+}05, CLF05, CCS^{+}04, DLBS03, Dat05, EGP03, ET01, Fan02, Fan03a, Fan03b, GCI06, KI04, LLC02a, LWF03, LT03, LLC03, LDH06, PS05, PS06a, SC05, TL02, TNS03, Wan04, WD04, ZC05b]. wireless-sensor [SC05]. wiring [FML03]. Within [NBAR08]. Within-Die [NBAR08]. Without [Kah04, JM01, MYL^{+}01, WCKD04]. Won’t [MRM07]. Wooley [VT09]. Word [CCY00, Voy08, RMH05, vdGT03]. Word-Length [CCY00]. word-level [RMH05]. Word-Organized [Voy08]. word-oriented [vdGT03]. Wordlength [Fio08, WKJ07]. workload [BLCA02]. Workloads [JLJ07, JZ05]. works [Rya04]. workstation [CJDM01]. wormhole [HS04, MOK04, SAOKM01, Sch01, SW00, WT00]. wormhole-routed [SAOKM01, WT00]. wormhole-switched [MOK04]. Worms [SS09a]. Worst [CC07, SB01, SLZ05b, YD09, ESE05]. Worst-Case [CC07, YD09, ESE05]. worth [DSK00]. Write [KPJ^{+}09, WH03]. Writes [BP07]. wrong [MKAP05]. wrong-path [MKAP05]. WSI [Low00].

x [LL09, WBW08]. X-Block [WBW08]. Xen [GCN^{+}09]. Xen-Based [GCN^{+}09]. XOR [VD05]. XOR-based [VD05]. XPAND [MK06].

year [Ano06f]. yield [KK00a, MJ02, MP03]. yielding [GMQS02].

Z [ABLP07]. Z-Cycles [ABLP07]. Zech [AP00]. Zero [BDG03, PB07b]. Zeroaliasing [BDG03]. ZettaRAM [VAZSR07]. ZRP [BB03b].

References

Atdelzater:2000:QNR

[AAS00] T. F. Atdelzater, E. M. Atkins, and K. G. Shin. QoS negotiation in real-time systems and its application to automated flight control. IEEE Transactions on Computers, 49(11):1170– 1183, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895935.

Al-Ars:2003:SDB

[AAvdG03] Z. Al-Ars and A. J. van de Goor. Static and dynamic behavior of memory cell array spot defects in embedded DRAMs. IEEE Transactions on Computers, 52(3): 293–309, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183945.

Al-Bassam:2000:AMC

[AB00] S. Al-Bassam. Another method for constructing t-EC/AUED codes. IEEE Transactions on Computers, 49(9):964–966, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869327.

Arnault:2005:DPN

[AB05] F. Arnault and T. P. Berger. Design and properties of a new pseudorandom generator based on a filtered FCSR automaton. IEEE Transactions on Computers, 54(11):1374– 1383, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514417.

Anantha:2007:MRG

[ABA07] M. Anantha, B. Bose, and B. F. AlBdaiwi. Mixedradix gray codes in Lee metric. IEEE Transactions on Computers, 56(10):1297–1307, October 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302703. See

comments [Jha13].

Anliker:2004:SAD

[ABD^{+}04] U. Anliker, J. Beutel, M. Dyer, R. Enzler, P. Lukowicz, L. Thiele, and G. Troster. A systematic approach to the design of distributed wearable systems. IEEE Transactions on Computers, 53(8):1017– 1033, August 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1306994.

Ahmadinia:2007:OFS

[ABF^{+}07] A. Ahmadinia, C. Bobda, S. P. Fekete, J. Teich, and J. C. van der Veen. Optimal free-space management and routing-conscious dynamic placement for reconfigurable devices. IEEE Transactions on Computers, 56(5): 673–680, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141239.

Anastasi:2004:CBP

[ABG04] G. Anastasi, A. Bartoli, and G. Giannini. On causal broadcasting with positive acknowledgments and bounded-length counters. IEEE Transactions on Computers, 53(10):1355– 1358, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327584.

Allulli:2007:CRZ

[ABLP07] L. Allulli, R. Baldoni, L. Laura, and S. T. Piergiovanni. On the complexity of removing Zcycles from a checkpoints and communication pattern. IEEE Transactions on Computers, 56 (6):853–858, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4167794.

Ardizzoni:2005:OSD

[ABP^{+}05] E. Ardizzoni, A. A. Bertossi, M. C. Pinotti, S. Ramaprasad, R. Rizzi, and M. V. S. Shashanka. Optimal skewed data allocation on multiple channels with flat broadcast per channel. IEEE Transactions on Computers, 54(5): 558–572, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1407846.

Anantha:2007:APU

[ABT07] M. Anantha, B. Bose, and L. G.

Tallini. ARQ protocols and unidirectional codes. IEEE Transactions on Computers, 56 (4):433–443, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118668.

Andrei:2009:EVO

[AC09] S. Andrei and A. M. K.

Cheng. Efficient verification and optimization of realtime logic-specified systems. IEEE Transactions on Computers, 58(12):1640–1653, December 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5066964.

Andrei:2006:ADR

[ACCL06] S. Andrei, W. N. Chin, A. M. K.

Cheng, and M. Lupu. Automatic debugging of real-time systems based on incremental satisfiability counting. IEEE Transactions on Computers, 55 (7):830–842, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637399.

Arlat:2003:CPS

[ACK^{+}03] J. Arlat, Y. Crouzet, J. Karlsson, P. Folkesson, E. Fuchs, and G. H. Leber. Comparison of physical and softwareimplemented fault injection techniques. IEEE Transactions on Computers, 52(9):1115– 1133, September 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228509.

AbouGhazaleh:2007:NMC

[ACMM07] N. AbouGhazaleh, B. R. Childers, D. Mosse, and R. G. Melhem. Near-memory caching for improved energy consumption. IEEE Transactions on Computers, 56(11):1441–1455, November 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336294.

Ascia:2008:IAN

[ACPP08] Giuseppe Ascia, Vincenzo Catania, Maurizio Palesi, and Davide Patti. Implementation and analysis of a new selection strategy for adaptive routing in networks-on-chip. IEEE Transactions on Computers, 57 (6):809–820, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4459314.

Aleta:2009:AGB

[ACS^{+}09] A. Aleta, J. M. Codina, J. Sanchez, A. Gonzalez, and D. Kaeli. AGAMOS: a graphbased approach to modulo scheduling for clustered microarchitectures. IEEE Transactions on Computers, 58(6): 770–783, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4785457.

Alvarez:2005:FMF

[ACV05] C. Alvarez, J. Corbal, and M. Valero. Fuzzy memoization for floating-point multimedia applications. IEEE Transactions on Computers, 54(7): 922–927, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432675.

Ammari:2008:ICC

[AD08] H. M. Ammari and S. K. Das.

Integrated coverage and connectivity in wireless sensor networks: a two-dimensional percolation problem. IEEE Transactions on Computers, 57(10): 1423–1434, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4492765.

Al-Daoud:2002:NAF

[ADMRK02] E. Al-Daoud, R. Mahmod, M. Rushdan, and A. Kilicman.

A new addition formula for elliptic curves over GF(2^{n}). IEEE Transactions on Computers, 51 (8):972–975, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1024743.

Azizoglu:2000:LBC

[AE00] M. C. Azizoglu and O. Egecioglu. Lower bounds on communication loads and optimal placements in torus networks. IEEE Transactions on Computers, 49(3): 259–266, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=841129.

Aggarwal:2005:IRR

[AF05] A. Aggarwal and M. Franklin. Instruction replication for reducing delays due to inter-PE communication latency. IEEE Transactions on Computers, 54(12):1496–1507, December 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1524932.

Azatchi:2006:AAT

[AFM^{+}06] H. Azatchi, L. Fournier, E. Marcus, S. Ur, A. Ziv, and K. Zohar. Advanced analysis techniques for cross-product coverage. IEEE Transactions on Computers, 55(11):1367– 1379, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705446.

Antola:2001:SED

[AFPS01] A. Antola, F. Ferrandi, V. Piuri, and M. Sami. Semiconcurrent error detection in data paths. IEEE Transactions on Computers, 50(5): 449–465, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926159.

Arlat:2002:DCM

[AFR02] J. Arlat, J.-C. Fabre, and M. Rodriguez. Dependability of COTS microkernelbased systems. IEEE Transactions on Computers, 51(2): 138–163, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980005.

Aragon:2006:CSE

[AGG06] J. L. Aragon, J. Gonzalez, and A. Gonzalez. Control speculation for energyefficient next-generation superscalar processors. IEEE Transactions on Computers, 55(3): 281–291, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1583558.

Apostolakis:2009:SBS

[AGPP09] A. Apostolakis, D. Gizopoulos, M. Psarakis, and A. Paschalis. Software-based self-testing of symmetric shared-memory multiprocessors. IEEE Transactions on Computers, 58(12): 1682–1694, December 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5184820.

Ansari:2008:HPA

[AH08] B. Ansari and M. A. Hasan. High-performance architecture of elliptic curve scalar multiplication. IEEE Transactions on Computers, 57(11):1443– 1453, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599568.

Amirijoo:2006:SMQ

[AHS06] M. Amirijoo, J. Hansson, and S. H. Son. Specification and management of QoS in realtime databases supporting imprecise computations. IEEE Transactions on Computers, 55 (3):304–319, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1583560.

Avresky:2002:GEI

[AJL02] D. Avresky, B. W. Johnson, and F. Lombardi. Guest Editors’ introduction. IEEE Transactions on Computers, 51(2): 97–99, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980048.

Azadmanesh:2000:EOF

[AK00] M. H. Azadmanesh and R. M.

Kieckhafer. Exploiting omissive faults in synchronous approximate agreement. IEEE Transactions on Computers, 49(10): 1031–1042, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888039.

Anand:2009:OCS

[AK09] C. K. Anand and W. Kahl.

An optimized Cell BE special function library generated by Coconut. IEEE Transactions on Computers, 58(8): 1126–1138, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4731241.

Aerts:2003:RRS

[AKS^{+}03] J. Aerts, J. Korst, F. Spieksma, W. Verhaegh, and G. Woeginger. Random redundant storage in disk arrays: complexity of retrieval problems. IEEE Transactions on Computers, 52(9):1210–1214, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228515.

Antelo:2000:VHR

[ALB00] E. Antelo, T. Lang, and J. D. Bruguera. Very-high radix circular CORDIC: Vectoring and unified rotation/ vectoring. IEEE Transactions on Computers, 49(7): 727–739, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863043.

Alippi:2002:RAS

[Ali02] C. Alippi. Randomized algorithms: a system-level, polytime analysis of robust computation. IEEE Transactions on Computers, 51(7): 740–749, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017694.

Antelo:2005:DRD

[ALMN05] E. Antelo, T. Lang, P. Montuschi, and A. Nannarelli. Digit-recurrence dividers with reduced logical depth. IEEE Transactions on Computers, 54 (7):837–851, July 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432667.

Alwan:2006:FPS

[Alw06] N. A. S. Alwan. A fully pipelined systolic array for sinusoidal sequence generation. IEEE Transactions on Computers, 55(5):636–639, May 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613843.

Almukhaizim:2007:CED

[AM07] S. Almukhaizim and Y. Makris.

Concurrent error detection methods for asynchronous burst-mode machines. IEEE Transactions on Computers, 56 (6):785–798, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167789.

Acharya:2008:DSM

[AM08] S. Acharya and R. N. Mahapatra. A dynamic slack management technique for realtime distributed embedded systems. IEEE Transactions on Computers, 57(2):215–230, February 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358241.

Aydin:2001:ORB

[AMMMA01] H. Aydin, R. Melhem, D. Mosse, and P. Mejia-Alvarez. Optimal reward-based scheduling for periodic real-time tasks. IEEE Transactions on Computers, 50(2):111–130, February 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=908988.

Aydin:2004:PAS

[AMMMA04] H. Aydin, R. Melhem, D. Mosse, and P. Mejia-Alvarez. Poweraware scheduling for periodic real-time tasks. IEEE Transactions on Computers, 53(5): 584–600, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275298.

Avresky:2005:DRC

[AN05] D. Avresky and N. Natchev. Dynamic reconfiguration in computer clusters with irregular topologies in the presence of multiple node and link failures. IEEE Transactions on Computers, 54(5): 603–615, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407849.

Anonymous:2000:RL

[Ano00a]  Anonymous.  1999 review-

ers list. IEEE Transactions on Computers, 49(1):95– 96, January 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=822568.

Anonymous:2000:AI

[Ano00b]  Anonymous.  Author in-

dex. IEEE Transactions on Computers, 49(12):1380–1384, December 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895873.

Anonymous:2000:CMU

[Ano00c] Anonymous. Correction to “A minimum universal test set for self-test of EXOR-sum-of-products circuits”. IEEE Transactions on Computers, 49(5):525, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?

                                        tp=&arnumber=859544. See

[KHP00].

Anonymous:2000:SI

[Ano00d]  Anonymous.  Subject in-

dex. IEEE Transactions on Computers, 49(12):1384–1392, December 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895875.

Anonymous:2001:RL

[Ano01a]  Anonymous.  2000 review-

ers list. IEEE Transactions on Computers, 50(1):92– 95, January 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=902756.

Anonymous:2001:AI

[Ano01b]  Anonymous.  Author in-

dex. IEEE Transactions on Computers, 50(12):1377–1380, December 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=970576.

Anonymous:2001:CEN

[Ano01c] Anonymous. Correction to Editor’s note. IEEE Transactions on Computers, 50(7): 765, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=936242.

Anonymous:2001:SI

[Ano01d]  Anonymous.  Subject in-

dex. IEEE Transactions on Computers, 50(12):1380–1388, December 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970577.

Anonymous:2002:AI

[Ano02a]  Anonymous.  Author in-

dex. IEEE Transactions on Computers, 51(12):1463–1466, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1146715.

Anonymous:2002:EN

[Ano02b] Anonymous. Editor’s note.

IEEE Transactions on Computers, 51(1):1–2, January 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=980012.

Anonymous:2002:SI

[Ano02c]  Anonymous.  Subject in-

dex. IEEE Transactions on Computers, 51(12):1466–1472, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1146717.

Anonymous:2003:RL

[Ano03]  Anonymous.  2002 review-

ers list. IEEE Transactions on Computers, 52(1):93– 96, January 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1159756.

Anonymous:2004:RL

[Ano04a]  Anonymous.  2003 review-

ers list. IEEE Transactions on Computers, 53(1):93– 96, January 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1255797.

Anonymous:2004:AI

[Ano04b]  Anonymous.  Annual in-

dex. IEEE Transactions on Computers, 53(12):1636–1648, December 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1347089.

Anonymous:2004:CPSa

[Ano04c] Anonymous. Call for papers for special issue on design and test of systems-on-a chip. IEEE Transactions on Computers, 53 (10):1359, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1327586.

Anonymous:2004:CPSb

[Ano04d] Anonymous. Call for papers for special issue on simulationbased design validation. IEEE Transactions on Computers, 53 (10):1360, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1327587.

Anonymous:2004:CPS

[Ano04e] Anonymous. Call for papers for special section on fault diagnosis and tolerance in cryptography. IEEE Transactions on Computers, 53(11): 1504, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336771.

Anonymous:2004:ICSa

[Ano04f] Anonymous. IEEE Computer Society information. IEEE Transactions on Computers, 53(4):499, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268410.

Anonymous:2004:ICSb

[Ano04g] Anonymous. IEEE Computer Society information. IEEE Transactions on Computers, 53(5):642, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1275303.

Anonymous:2004:NVI

[Ano04h] Anonymous. Notice of violation of IEEE publication principles in “Performance Enhancement of Ad Hoc Networks with Localized Route Repair”. IEEE Transactions on Computers, 53(7):928, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1321053.

Anonymous:2004:TIAa

[Ano04i] Anonymous. TC: Information for authors. IEEE Transactions on Computers, 53(4): 498, April 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268409.

Anonymous:2004:TIAb

[Ano04j] Anonymous. TC: Information for authors. IEEE Transactions on Computers, 53(5): 641, May 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275302.

Anonymous:2005:AI

[Ano05a] Anonymous. 2005 annual index. IEEE Transactions on Computers, 54(12):1631–1648, December 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524943.

Anonymous:2005:CP

[Ano05b]  Anonymous.  Call for pa-

pers. IEEE Transactions on Computers, 54(3):381, March 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1388202.

Anonymous:2005:CPSa

[Ano05c] Anonymous. Call for papers for special issue on nano systems and computing. IEEE Transactions on Computers, 54(6):784, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1424451.

Anonymous:2005:CPSb

[Ano05d] Anonymous. Call for papers for special issue on nano systems and computing. IEEE Transactions on Computers, 54(7):928, July 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432676.

Anonymous:2005:CPSc

[Ano05e] Anonymous. Call for papers for special issue on nano systems and computing. IEEE Transactions on Computers, 54(9): 1184, September 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471679.

Anonymous:2005:CER

[Ano05f] Anonymous. Correction to “Efficient Reclaiming in Reservation-Based Real-Time Systems with Variable Execution Times”. IEEE Transactions on Computers, 54(6):783, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1461364. See

[CBT05].

Anonymous:2005:RL

[Ano05g] Anonymous. Reviewers list. IEEE Transactions on Computers, 54(1):93–96, January 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1362644.

Anonymous:2006:RL

[Ano06a] Anonymous. 2005 reviewers list. IEEE Transactions on Computers, 55(1):93–96, January 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2006:AI

[Ano06b]  Anonymous.  Annual in-

dex. IEEE Transactions on Computers, 55(12):tc06, December 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717396.

Anonymous:2006:CPSb

[Ano06c] Anonymous. Call for papers — special issue on emergent systems, algorithms, and architectures for speech-based humanmachine interaction. IEEE Transactions on Computers, 55 (8):1070, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650205.

Anonymous:2006:CPSc

[Ano06d] Anonymous. Call for papers — special issue on emergent systems, algorithms, and architectures for speech-based humanmachine interaction. IEEE Transactions on Computers, 55 (9):1216, September 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668050.

Anonymous:2006:CPSa

[Ano06e] Anonymous. Call for papers for special issue on emergent systems, algorithms, and architectures for speech-based humanmachine interaction. IEEE Transactions on Computers, 55(7):928, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637409.

Anonymous:2006:ICS

[Ano06f] Anonymous. IEEE Computer Society celebrates two 60-year anniversaries. IEEE Transactions on Computers, 55(8): 1071, August 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650206.

Anonymous:2006:JICb

[Ano06g] Anonymous. Join the IEEE Computer Society! IEEE Transactions on Computers, 55 (10):1312, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683762.

Anonymous:2006:JICa

[Ano06h] Anonymous. Join the IEEE Computer Society — now with 800 course modules for distance learning! IEEE Transactions on Computers, 55(8): 1072, August 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650207.

Anonymous:2007:CAI

[Ano07a]  Anonymous.  180,000 com-

puting articles in the IEEE Computer Society Digital Library. IEEE Transactions on Computers, 56(10):1440, October 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302718.

Anonymous:2007:RL

[Ano07b]  Anonymous.  2006 review-

ers list. IEEE Transactions on Computers, 56(1):139–143, January 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSa

[Ano07c] Anonymous. Call for papers for special issue on computer arithmetic. IEEE Transactions on Computers, 56(1):144, January 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSg

[Ano07d] Anonymous. Call for papers for special section on chips and architectures for emerging technologies and applications. IEEE Transactions on Computers, 56 (10):1438, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSb

[Ano07e] Anonymous. Call for papers for special section on computer arithmetic. IEEE Transactions on Computers, 56(2):287, February 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSc

[Ano07f] Anonymous. Call for papers for special section on networks-on-chips: Modeling, analysis and optimization. IEEE Transactions on Computers, 56(3):432, March 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSd

[Ano07g] Anonymous. Call for papers for special section on programming models and architectures for embedded systems. IEEE Transactions on Computers, 56 (5):720, May 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSh

[Ano07h] Anonymous. Call for papers for special section on specialpurpose hardware for cryptography and cryptanalysis. IEEE Transactions on Computers, 56 (10):1439, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSe

[Ano07i] Anonymous. Call for papers for the special secion on programming models and achitectures for embedded systems. IEEE Transactions on Computers, 56(6):864, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:CPSf

[Ano07j] Anonymous. Call for papers:

Special-purpose hardware for cryptography and cryptanalysis. IEEE Transactions on Computers, 56(7):1008, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2007:ICSa

[Ano07k] Anonymous. IEEE Computer Society Digital Library. IEEE Transactions on Computers, 56 (2):288, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042689.

Anonymous:2007:ICSb

[Ano07l] Anonymous. IEEE Computer Society Digital Library. IEEE Transactions on Computers, 56 (12):1728, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4368207.

Anonymous:2007:JIC

[Ano07m] Anonymous. Join the IEEE Computer Society. IEEE Transactions on Computers, 56 (12):1727, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4368206.

Anonymous:2007:MNE

[Ano07n] Anonymous. Message from the new Editor-in-Chief. IEEE Transactions on Computers, 56(1):1, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016492.

Anonymous:2008:AI

[Ano08a]  Anonymous.  Annual in-

dex. IEEE Transactions on Computers, 57(1):INDEX:1– INDEX:21, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4390000. On-

line only.

Anonymous:2008:BYCb

[Ano08b] Anonymous. Build your career. IEEE Transactions on Computers, 57(10):1439, October 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4607768.

Anonymous:2008:BYCa

[Ano08c] Anonymous. Build your career in computing. IEEE Transactions on Computers, 57(9): 1296, September 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4579705.

Anonymous:2008:BYCc

[Ano08d] Anonymous. Build your career in computing. IEEE Transactions on Computers, 57(11): 1584, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633727.

Anonymous:2008:RL

[Ano08e] Anonymous. Reviewers list.

IEEE Transactions on Computers, 57(1):139–144, January 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2008:SBSa

[Ano08f] Anonymous. Silver Bullet Security Podcast series. IEEE Transactions on Computers, 57 (9):1295, September 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4579704.

Anonymous:2008:SBSb

[Ano08g] Anonymous. Silver Bullet Security Podcast series. IEEE Transactions on Computers, 57 (10):1440, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4607769.

Anonymous:2009:TRL

[Ano09a] Anonymous. 2008 TC reviewers list. IEEE Transactions on Computers, 58(1): 139–144, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4688921.

Anonymous:2009:BYC

[Ano09b] Anonymous. Build your career. IEEE Transactions on Computers, 58(5):720, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4804835.

Anonymous:2009:CPD

[Ano09c] Anonymous. Call for papers:

Dependable computer architecture. IEEE Transactions on Computers, 58(12):1728, December 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2009:CPC

[Ano09d] Anonymous. Call-for-papers on computer arithmetic. IEEE Transactions on Computers, 58 (5):719, May 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2009:CC

[Ano09e] Anonymous. CSDA Certification. IEEE Transactions on Computers, 58(11):1583, November 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5271548.

Anonymous:2009:IIC

[Ano09f] Anonymous. IEEE and IEEE Computer Society 2010 Student Member Package. IEEE Transactions on Computers, 58(11): 1584, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5271549.

Anonymous:2009:ICS

[Ano09g] Anonymous. IEEE Computer Society Computing Now. IEEE Transactions on Computers, 58 (11):1582, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5271547.

Anonymous:2009:SBS

[Ano09h] Anonymous. Silver Bullet Security Podcast series. IEEE Transactions on Computers, 58 (2):288, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4740166.

Akhbarizadeh:2007:TBP

[ANPS07] M. J. Akhbarizadeh, M. Nourani, R. Panigrahy, and S. Sharma. A TCAM-based parallel architecture for high-speed packet forwarding. IEEE Transactions on Computers, 56(1):58– 72, January 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016497. See

comments [CS08a].

Assis:2000:ACZ

[AP00] F. M. Assis and C. E. Pedreira. An architecture for computing Zech’s logarithms in GF(2^{m}). IEEE Transactions on Computers, 49(5): 519–524, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=859543.

Avresky:2009:GEI

[APV09] D. R. Avresky, Harald Prokop, and Dinesh C. Verma. Guest Editors’ introduction: Special section on autonomic network computing. IEEE Transactions on Computers, 58(11):1441– 1443, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5271546.

Arazi:2008:CMI

[AQ08] O. Arazi and Hairong Qi. On calculating multiplicative inverses modulo 2^{m}. IEEE Transactions on Computers, 57(10): 1435–1438, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483503.

Aloul:2007:SOS

[ARSM07] F. A. Aloul, A. Ramani, K. A.

Sakallah, and I. L. Markov. Solution and optimization of systems of pseudo-Boolean constraints. IEEE Transactions on Computers, 56(10):1415– 1424, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302712.

Abdelzaher:2000:PBL

[AS00] T. F. Abdelzaher and K. G. Shin. Period-based load partitioning and assignment for large real-time applications. IEEE Transactions on Computers, 49 (1):81–87, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=822566.

Abed:2003:CVI

[AS03a] K. H. Abed and R. E. Siferd. CMOS VLSI implementation of a low-power logarithmic converter. IEEE Transactions on Computers, 52(11):1421– 1433, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1244940.

Abed:2003:VIL

[AS03b] K. H. Abed and R. E. Siferd. VLSI implementation of a low-power antilogarithmic converter. IEEE Transactions on Computers, 52(9):1221–1228, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1228517.

Araki:2003:DSG

[AS03c] T. Araki and Y. Shibata. (t, k)-diagnosable system: a generalization of the PMC models. IEEE Transactions on Computers, 52(7): 971–975, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214345.

Abdelzaher:2003:ULQ

[ASB03] T. F. Abdelzaher, K. G. Shin, and N. Bhatti. User-level QoSadaptive resource management in server end-systems. IEEE Transactions on Computers, 52 (5):678–685, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197134.

Abali:2001:HCM

[ASF^{+}01] B. Abali, Xiaowei Shen, H. Franke, D. E. Poff, and T. B. Smith. Hardware compressed main memory: operating system support and performance evaluation. IEEE Transactions on Computers, 50(11):1219– 1233, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=966496.

Abdelzaher:2004:UBA

[ASL04] T. F. Abdelzaher, Vivek Sharma, and C. Lu. A utilization bound for aperiodic tasks and priority driven scheduling. IEEE Transactions on Computers, 53(3): 334–350, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261839.

Aloul:2006:ESB

[ASM06] F. A. Aloul, K. A. Sakallah, and I. L. Markov. Efficient symmetry breaking for Boolean satisfiability. IEEE Transactions on Computers, 55(5): 549–558, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613836.

Alfaro:2007:FMM

[ASMD07] F. J. Alfaro, J. L. Sanchez, M. Menduia, and J. Duato. A formal model to manage the InfiniBand arbitration tables providing QoS. IEEE Transactions on Computers, 56(8): 1024–1039, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4264319.

Abel:2005:UTP

[AT05] J. Abel and W. Teahan. Universal text preprocessing for data compression. IEEE Transactions on Computers, 54(5): 497–507, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407841.

Atallah:2001:ELE

[Ata01] M. J. Atallah. On estimating the large entries of a convolution. IEEE Transactions on Computers, 50(3): 193–196, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=910811.

Anastasiadis:2005:OLS

[AVVY05] S. V. Anastasiadis, P. Varman, J. S. Vitter, and Ke Yi. Optimal lexicographic shaping of aggregate streaming data. IEEE Transactions on Computers, 54 (4):398–408, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401859.

Antelo:2008:LLP

[AVZ08] E. Antelo, J. Villalba, and E. L. Zapata. A low-latency pipelined 2D and 3D CORDIC processors. IEEE Transactions on Computers, 57(3): 404–417, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358248.

Aggarwal:2001:MFQ

[AWY01] C. C. Aggarwal, J. L. Wolf, and P. S. Yu. The maximum factor queue length batching scheme for video-on-demand systems. IEEE Transactions on Computers, 50(2): 97–110, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=908987.

Aydin:2007:EFS

[Ayd07] H. Aydin. Exact fault-sensitive feasibility analysis of real-time tasks. IEEE Transactions on Computers, 56(10):1372– 1386, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302709.

Aydin:2009:RAE

[AZ09] H. Aydin and Dakai Zhu. Reliability-aware energy management for periodic real-time tasks. IEEE Transactions on Computers, 58(10):1382– 1397, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4815213.

Balasubramonian:2003:DTM

[BABD03] R. Balasubramonian, D. H. Albonesi, A. Buyuktosunoglu, and S. Dwarkadas. A dynamically tunable memory hierarchy. IEEE Transactions on Computers, 52(10):1243– 1258, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1234523.

Baldoni:2004:RCP

[Bal04] R. Baldoni. Response to comment on “A Positive Acknowledgment Protocol for Causal Broadcasting”. IEEE Transactions on Computers, 53(10): 1358, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1327585.

Bao:2004:CPK

[Bao04]  Feng Bao.  Cryptanalysis

of a partially known cellular automata cryptosystem. IEEE Transactions on Computers, 53(11):1493–1497, November 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336769.

Baruah:2004:OUB

[Bar04] S. K. Baruah. Optimal utilization bounds for the fixedpriority scheduling of periodic task systems on identical multiprocessors. IEEE Transactions on Computers, 53(6): 781–784, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1288553.

Barnett:2005:DTL

[Bar05] J. A. Barnett. Dynamic tasklevel voltage scheduling optimizations. IEEE Transactions on Computers, 54(5): 508–520, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1407842.

Bae:2003:EDH

[BB03a] M. M. Bae and B. Bose. Edge disjoint Hamiltonian cycles in k-ary n-cubes and hypercubes. IEEE Transactions on Computers, 52(10):1271– 1284, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234525.

Beraldi:2003:CSR

[BB03b] R. Beraldi and R. Baldoni. A caching scheme for routing in mobile ad hoc networks and its application to ZRP. IEEE Transactions on Computers, 52(8):1051– 1062, August 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223639.

Bini:2004:SAP

[BB04] E. Bini and G. C. Buttazzo. Schedulability analysis of periodic fixed priority systems. IEEE Transactions on Computers, 53(11):1462– 1473, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336766.

Baron:2005:ASS

[BB05] D. Baron and Y. Bresler.

Antisequential suffix sorting for BWT-based data compression. IEEE Transactions on Computers, 54(4): 385–397, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401858.

Bini:2003:RMA

[BBB03] E. Bini, G. C. Buttazzo, and G. M. Buttazzo. Rate monotonic analysis: the hyperbolic bound. IEEE Transactions on Computers, 52(7): 933–942, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214341.

Benso:2008:MTG

[BBD^{+}08a] A. Benso, A. Bosio, S. Di Carlo, G. Di Natale, and P. Prinetto. March test generation revealed. IEEE Transactions on Computers, 57(12):1704– 1713, December 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4569834.

Beuchat:2008:AAO

[BBD^{+}08b] J.-L. Beuchat, N. Brisebarre, J. Detrey, E. Okamoto, M. Shirase, and T. Takagi. Algorithms and arithmetic operators for computing the ηT pairing in characteristic three. IEEE Transactions on Computers, 57(11):1454– 1468, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4564442.

Bartolini:2008:EIS

[BBGM08] S. Bartolini, I. Branovic, R. Giorgi, and E. Martinelli. Effects of instruction-set extensions on an embedded processor: a case study on elliptic curve cryptography over GF(2^{m}). IEEE Transactions on Computers, 57(5): 672–685, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358294.

Babic:2006:BCN

[BBH06] D. Babic, J. Bingham, and A. J. Hu. B-Cubing: New possibilities for efficient SATsolving. IEEE Transactions on Computers, 55(11):1315– 1324, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705441.

Bertoni:2003:EAD

[BBK^{+}03] G. Bertoni, L. Breveglieri, I. Koren, P. Maistri, and V. Piuri. Error analysis and detection procedures for a hardware implementation of the advanced encryption standard. IEEE Transactions on Computers, 52(4): 492–505, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190590.

Bernat:2001:WHR

[BBL01] G. Bernat, A. Burns, and A. Liamosi. Weakly hard realtime systems. IEEE Transactions on Computers, 50(4): 308–321, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=919277.

Benini:2003:DCS

[BBM^{+}03] L. Benini, D. Bruni, A. Mach, E. Macii, and M. Poncino. Discharge current steering for battery lifetime optimization. IEEE Transactions on Computers, 52(8):985–995, August 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223633.

Beaumont:2001:PHC

[BBP^{+}01] O. Beaumont, V. Boudet, A. Petitet, F. Rastello, and Y. Robert. A proposal for a heterogeneous cluster ScaLA-PACK (dense linear solvers). IEEE Transactions on Computers, 50(10):1052–1070, October 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956091.

Bernardi:2006:NHF

[BBR^{+}06] P. Bernardi, L. M. V. Bolzani, M. Rebaudengo, M. S. Reorda, F. L. Vargas, and M. Violante. A new hybrid fault detection technique for systemson-a-chip. IEEE Transactions on Computers, 55(2): 185–198, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566579.

Brown:2001:SNCa

[BC01a] B. D. Brown and H. C.

Card. Stochastic neural computation. I. computational elements. IEEE Transactions on Computers, 50(9):891–905, September 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954505.

Brown:2001:SNCb

[BC01b] B. D. Brown and H. C.

Card. Stochastic neural computation. II. soft competitive learning. IEEE Transactions on Computers, 50(9):906–920, September 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954506.

Bessani:2009:EBR

[BCdSFL09] A. N. Bessani, M. Correia, J. da Silva Fraga, and Lau Cheuk Lung. An efficient byzantine-resilient tuple space. IEEE Transactions on Computers, 58(8): 1080–1094, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4912196.

Baynes:2003:PEC

[BCF^{+}03] K. Baynes, C. Collins, E. Fiterman, Brinda Ganesh, P. Kohout, C. Smit, T. Zhang, and B. Jacob. The performance and energy consumption of embedded real-time operating systems. IEEE Transactions on Computers, 52(11):1454– 1469, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244943.

Bondavalli:2000:TBM

[BCGG00] A. Bondavalli, S. Chiaradonna, F. Di Giandomenico, and F. Grandoni. Thresholdbased mechanisms to discriminate transient from intermittent faults. IEEE Transactions on Computers, 49(3): 230–245, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841127.

Baneres:2009:RPS

[BCK09] D. Baneres, J. Cortadella, and M. Kishinevsky. A recursive paradigm to solve Boolean relations. IEEE Transactions on Computers, 58(4): 512–527, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624246.

Bakos:2007:LEC

[BCL07] J. D. Bakos, D. M. Chiarulli, and S. P. Levitan. Lightweight error correction coding for system-level interconnects. IEEE Transactions on Computers, 56 (3):289–304, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4079513.

Bernasconi:2001:CBF

[BCV01] A. Bernasconi, B. Codenottl, and J. M. Vanderkam. A characterization of bent functions in terms of strongly regular graphs. IEEE Transactions on Computers, 50(9):984– 985, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954512.

Biswas:2005:CSR

[BD05] P. Biswas and N. D. Dutt. Code size reduction in heterogeneousconnectivity-based DSPs using instruction set extensions. IEEE Transactions on Computers, 54(10):1216– 1226, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501788.

Bhunia:2005:GPA

[BDBR05] Swarup Bhunia, Animesh Datta, Nilanjan Banerjee, and Kaushik Roy. GAARP: a power-aware GALS architecture for real-time algorithmspecific tasks. IEEE Transactions on Computers, 54(6): 752–766, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461362.

Bertolazzi:2000:COD

[BDD00] P. Bertolazzi, G. Di Battista, and W. Didimo. Computing orthogonal drawings with the minimum number of bends. IEEE Transactions on Computers, 49(8): 826–840, August 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=868028.

Bhattacharya:2003:ZAS

[BDG03] B. B. Bhattacharya, A. Dmitriev, and M. Gossel. Zero-aliasing space compaction of test responses using a single periodic output. IEEE Transactions on Computers, 52(12):1646– 1651, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1252860.

Bhide:2002:APP

[BDK^{+}02] M. Bhide, P. Deolasee, A. Katkar, A. Panchbudhe, K. Ramamritham, and P. Shenoy. Adaptive push-pull: disseminating dynamic Web data. IEEE Transactions on Computers, 51 (6):652–668, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009150.

Brisebarre:2005:NRR

[BDK^{+}05] Nicolas Brisebarre, David Defour, Peter Kornerup, Jean-Michel Muller, and Nathalie Revol. A new range-reduction algorithm. IEEE Transactions on Computers, 54(3): 331–339, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0331abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0331.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0331.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388197.
pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388197.
pdf?isnumber=30205&prod=
JNL&arnumber=1388197&arSt=
+331&ared=+339&arAuthor=
Brisebarre%2C+N.%3B+Defour%
2C+D.%3B+Kornerup%2C+P.%
3B+Muller%2C+J.-M.%3B+Revol%
2C+N.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30205;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30205&
arnumber=1388197&count=13&

                                        index=8;   http://ieeexplore.
                                        ieee.org/xpls/references.
                                        jsp?arnumber=1388197.

Boldo:2009:FVA

[BDL09] S. Boldo, M. Daumas, and Ren-Cang Li. Formally verified argument reduction with a fused multiply-add. IEEE Transactions on Computers, 58(8): 1139–1145, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4711042.

Bose:2007:SUE

[BET07] B. Bose, S. Elmougy, and L. G. Tallini. Systematic tunidirectional error-detecting codes over Z_{m}. IEEE Transactions on Computers, 56(7): 876–880, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118666.

Baruah:2006:PMS

[BF06] Sanjoy Baruah and N. Fisher. The partitioned multiprocessor scheduling of deadlineconstrained sporadic task systems. IEEE Transactions on Computers, 55(7):918–923, July 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1637407.

Baumgartner:2008:GTA

[BF08] K. Baumgartner and S. Ferrari. A geometric transversal approach to analyzing track coverage in sensor networks. IEEE Transactions on Computers, 57(8):1113– 1128, August 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483505.

Bianchini:2001:OAW

[BFG01] N. Bianchini, S. Fanelli, and M. Gori. Optimal algorithms for well-conditioned nonlinear systems of equations. IEEE Transactions on Computers, 50 (7):689–698, July 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=936235.

Baruah:2003:RRC

[BFG03] S. Baruah, S. Funk, and J. Goossens. Robustness results concerning EDF scheduling upon uniform multiprocessors. IEEE Transactions on Computers, 52(9):1185–1195, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228513.

Brandolese:2006:ADS

[BFP^{+}06] C. Brandolese, W. Fornaciari, L. Pomante, F. Salice, and D. Sciuto. Affinitydriven system design exploration for heterogeneous multiprocessor SoC. IEEE Transactions on Computers, 55(5): 508–519, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613832.

Benveniste:2001:CMI

[BFR01] C. D. Benveniste, P. A. Franaszek, and J. T. Robinson. Cache-memory interfaces in compressed memory systems. IEEE Transactions on Computers, 50(11):1106– 1116, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966489.

Baruah:2003:RMS

[BG03] S. K. Baruah and J. Goossens. Rate-monotonic scheduling on uniform multiprocessors. IEEE Transactions on Computers, 52 (7):966–970, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1214344.

Burns:2005:ASC

[BG05] J. Burns and J.-L. Gaudiot.

Area and system clock effects on SMT/CMP throughput. IEEE Transactions on Computers, 54(2):141–152, February 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377153.

Bi:2008:MRC

[BG08] Shaoqiang Bi and W. J. Gross.

The mixed-radix Chinese Remainder Theorem and its applications to residue comparison. IEEE Transactions on Computers, 57(12):1624–1632, December 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4585366.

Beltran:2008:NCA

[BGB08] M. Beltran, A. Guzman, and J. L. Bosque. A new CPU availability prediction model for time-shared systems. IEEE Transactions on Computers, 57 (7):865–875, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4445661.

Butt:2007:PIK

[BGH07] A. R. Butt, C. Gniady, and Y. C. Hu. The performance impact of kernel prefetching on buffer cache replacement algorithms. IEEE Transactions on Computers, 56(7): 889–908, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216288.

Burtscher:2005:VTC

[BGJ^{+}05] M. Burtscher, I. Ganusov, S. J. Jackson, J. Ke, P. Ratanaworabhan, and N. B. Sam. The VPC trace-compression algorithms. IEEE Transactions on Computers, 54(11):1329– 1344, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514414.

Bucci:2003:HSO

[BGL^{+}03] M. Bucci, L. Germani, R. Luzzi, A. Trifiletti, and M. Varanonuovo. A high-speed oscillator-based truly random number source for cryptographic applications on a smart card IC. IEEE Transactions on Computers, 52(4): 403–409, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190581.

Bridges:2009:CFC

[BHS09] P. G. Bridges, M. A. Hiltunen, and R. D. Schlichting. Cholla: a framework for composing and coordinating adaptations in networked systems. IEEE Transactions on Computers, 58(11):1456– 1469, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815212.

Bajard:2004:FRI

[BI04] J.-C. Bajard and L. Imbert.

A full RNS implementation of RSA. IEEE Transactions on Computers, 53(6): 769–774, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288551.

Bajard:2006:AOF

[BIN06] J.-C. Bajard, L. Imbert, and C. Negre. Arithmetic operations in finite fields of medium prime characteristic using the Lagrange representation. IEEE Transactions on Computers, 55(9):1167–1177, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668044.

Bique:2005:NCD

[Biq05] S. Bique. New characterizations of 2D discrete cosine transform. IEEE Transactions on Computers, 54(9):1054–1060, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471667.

Brockmeyer:2000:FES

[BJHW00] M. Brockmeyer, F. Jahanlan, C. Heitmeyer, and E. Winner. A flexible, extensible simulation environment for testing real-time specifications. IEEE Transactions on Computers, 49(11):1184– 1201, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895936.

Balkenhol:2000:UDC

[BK00] B. Balkenhol and S. Kurtz. Universal data compression based on the Burrows–Wheeler transformation: theory and practice. IEEE Transactions on Computers, 49(10):1043– 1053, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888040.

Breveglieri:2006:GEI

[BK06] L. Breveglieri and I. Koren. Guest Editors’ introduction: Special section on fault diagnosis and tolerance in cryptography. IEEE Transactions on Computers, 55(9):1073–1074, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668034.

Blazewicz:2006:PMT

[BKM^{+}06] J. Blazewicz, M. Y. Kovalyov, M. Machowiak, D. Trystram, and J. Weglarz. Preemptable malleable task scheduling problem. IEEE Transactions on Computers, 55(4): 486–490, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608009.

Breveglieri:2007:OCA

[BKM07] L. Breveglieri, I. Koren, and P. Maistri. An operationcentered approach to fault detection in symmetric cryptography ciphers. IEEE Transactions on Computers, 56(5): 635–649, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118667.

Barua:2001:CSS

[BLAA01] R. Barua, W. Lee, S. Arnarasinghe, and A. Agarwal. Compiler support for scalable and efficient memory systems. IEEE Transactions on Computers, 50(11):1234– 1247, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966497.

Buttazzo:2002:ESF

[BLCA02] G. C. Buttazzo, G. Lipari, M. Caccamo, and L. Abeni. Elastic scheduling for flexible workload management. IEEE Transactions on Computers, 51 (3):289–302, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=990127.

Beuchat:2008:AGM

[BM08a] J.-L. Beuchat and J.-M. Muller. Automatic generation of modular multipliers for FPGA applications. IEEE Transactions on Computers, 57(12):1600– 1613, December 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4564441.

Boldo:2008:EFC

[BM08b] Sylvie Boldo and Guillaume Melquiond. Emulation of a FMA and correctly rounded sums: Proved algorithms using rounding to odd. IEEE Transactions on Computers, 57 (4):462–471, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358278;   https:
//ens-lyon.hal.science/inria-
00080427/file/odd-rounding.
pdf.

Brisebarre:2008:CRM

[BM08c] N. Brisebarre and J.-M. Muller.

Correctly rounded multiplication by arbitrary precision constants. IEEE Transactions on Computers, 57(2): 165–174, February 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358257.

Busch:2008:OOP

[BMlX08] C. Busch, M. Magdon-lsmail, and Jing Xi. Optimal oblivious path selection on the mesh. IEEE Transactions on Computers, 57(5):660–671, May 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4445660.

Benini:2004:CCC

[BMO04] L. Benini, F. Menichelli, and M. Olivieri. A class of code compression schemes for reducing power consumption in embedded microprocessor systems. IEEE Transactions on Computers, 53(4): 467–482, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268405.

Brisebarre:2004:ACR

[BMR04] N. Brisebarre, J.-M. Muller, and Saurabh Kumar Raina. Accelerating correctly rounded floating-point division when the divisor is known in advance. IEEE Transactions on Computers, 53(8):1069– 1072, August 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306999.

Bini:2009:RTB

[BNRB09] E. Bini, Thi Huyen Chau Nguyen, P. Richard, and S. K. Baruah. A responsetime bound in fixed-priority scheduling with arbitrary deadlines. IEEE Transactions on Computers, 58(2):279–286, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4624248.

Bayraktaroglu:2003:CAC

[BO03] I. Bayraktaroglu and A. Orailoglu. Concurrent application of compaction and compression for test time and data volume reduction in scan designs. IEEE Transactions on Computers, 52(11):1480– 1489, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244945.

Bayraktaroglu:2005:COD

[BO05] I. Bayraktaroglu and A. Orailoglu.

The construction of optimal deterministic partitionings in scan-based BIST fault diagnosis: mathematical foundations and cost-effective implementations. IEEE Transactions on Computers, 54(1):61– 75, January 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362640.

Boldo:2009:KAC

[Bol09] Sylvie Boldo. Kahan’s algorithm for a correct discriminant computation at last formally proven. IEEE Transactions on Computers, 58(2): 220–225, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4663057. See

[Kah04] for the original algorithm.

Blum:2001:HRM

[BP01] T. Blum and C. Paar. Highradix Montgomery modular exponentiation on reconfigurable hardware. IEEE Transactions on Computers, 50(7): 759–764, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936241.

Biswas:2006:SOS

[BP06] P. K. Biswas and S. Phoha. Self-organizing sensor networks for integrated target surveillance. IEEE Transactions on Computers, 55(8):1033– 1047, August 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650200.

Baek:2007:MSC

[BP07] Sung Hoon Baek and Kyu Ho Park. Matrix-stripe-cache-based contiguity transform for fragmented writes in RAID-5. IEEE Transactions on Computers, 56(8):1040–1054, August 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4264320.

Baek:2009:SAS

[BP09] Sung Hoon Baek and Kyu Ho Park. Striping-aware sequential prefetching for independency and parallelism in disk arrays with concurrent accesses. IEEE Transactions on Computers, 58(8):1146– 1152, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4752809.

Burtscher:2009:FHS

[BR09] M. Burtscher and P. Ratanaworabhan. FPC: a high-speed compressor for double-precision floating-point data. IEEE Transactions on Computers, 58 (1):18–31, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4589203.

Balbastre:2008:MDC

[BRC08] P. Balbastre, I. Ripoll, and A. Crespo. Minimum deadline calculation for periodic real-time tasks in dynamic priority systems. IEEE Transactions on Computers, 57(1): 96–109, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358239.

Baktir:2004:OTF

[BS04] S. Baktir and B. Sunar. Optimal tower fields. IEEE Transactions on Computers, 53(10): 1231–1243, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327575.

Bliudze:2008:ACS

[BS08] S. Bliudze and J. Sifakis. The algebra of connectorsstructuring interaction in BIP. IEEE Transactions on Computers, 57(10):1315–1330, October 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4445663.

Bayat-Sarmadi:2009:CED

[BSH09] S. Bayat-Sarmadi and M. A. Hasan. Concurrent error detection in finite-field arithmetic operations using pipelined and systolic architectures. IEEE Transactions on Computers, 58(11):1553–1567, November 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4815219.

Butler:2005:APL

[BSM05] J. T. Butler, T. Sasao, and M. Matsuura. Average path length of binary decision diagrams. IEEE Transactions on Computers, 54(9):1041–1053, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471666.

Barnes:2006:BOS

[BSN^{+}06] R. D. Barnes, J. W. Sias, E. M. Nystrom, S. J. Patel, J. Navarro, and W. W. Hwu. Beating in-order stalls with “flea-flicker” two-pass pipelining. IEEE Transactions on Computers, 55(1):18– 33, January 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545748.

Balkan:2008:PET

[BSPG08] D. Balkan, J. Sharkey, D. Ponomarev, and K. Ghose. Predicting and exploiting transient values for reducing register file pressure and energy consumption. IEEE Transactions on Computers, 57(1): 82–95, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358237.

Boullis:2005:SOH

[BT05] N. Boullis and A. Tisserand.

Some optimizations of hardware multiplication by constant matrices. IEEE Transactions on Computers, 54(10): 1271–1282, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501792.

Buttazzo:2007:QCM

[BVM07] Giorgio Buttazzo, Manel Velasco, and Pau Marti. Qualityof-control management in overloaded real-time systems. IEEE Transactions on Computers, 56(2):253–266, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4042685.

Bhuyan:2003:SMT

[BW03] L. N. Bhuyan and Hujun Wang. Switch MSHR: a technique to reduce remote read memory access time in CC-NUMA multiprocessors. IEEE Transactions on Computers, 52(5): 617–632, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1197128.

Balli:2007:UAR

[BWR^{+}07] U. Balli, H. Wu, B. Ravindran, J. S. Anderson, and E. D. Jensen. Utility accrual realtime scheduling under variable cost functions. IEEE Transactions on Computers, 56(3): 385–401, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079520.

Brehob:2004:ORN

[BWTE04] M. Brehob, S. Wagner, E. Torng, and R. Enbody. Optimal replacement is NP-hard for nonstandard caches. IEEE Transactions on Computers, 53(1): 73–76, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255792.

Burtscher:2002:HLV

[BZ02] M. Burtscher and B. G.

Zorn. Hybrid load-value predictors. IEEE Transactions on Computers, 51(7): 759–774, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017696.

Chang:2004:CSF

[CAK04] Yung-Ruei Chang, S. V. Amari, and Sy-Yen Kuo. Computing system failure frequencies and reliability importance measures using OBDD. IEEE Transactions on Computers, 53(1): 54–68, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255790.

Cao:2002:PPA

[Cao02]  Guohong Cao.  Proactive

power-aware cache management for mobile computing systems. IEEE Transactions on Computers, 51(6): 608–621, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009147.

Carrasco:2002:CEN

[Car02] J. A. Carrasco. Computationally efficient and numerically stable reliability bounds for repairable fault-tolerant systems. IEEE Transactions on Computers, 51(3): 254–268, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=990125.

Cardoso:2003:CTP

[Car03] J. M. P. Cardoso. On combining temporal partitioning and sharing of functional units in compilation for reconfigurable architectures. IEEE Transactions on Computers, 52(10): 1362–1375, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1234532.

Carrasco:2004:TAS

[Car04] J. A. Carrasco. Transient analysis of some rewarded Markov models using randomization with quasistationarity detection. IEEE Transactions on Computers, 53(9):1106–1120, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315605.

Chung:2002:DPM

[CBB^{+}02] Eui-Young Chung, L. Benini, A. Bogliolo, Yung-Hsiang Lu, and G. De Micheli. Dynamic power management for nonstationary service requests. IEEE Transactions on Computers, 51(11):1345– 1361, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047758.

Chen:2007:NSA

[CBC07] Gang Chen, Guoqiang Bai, and Hongyi Chen. A new systolic architecture for modular division. IEEE Transactions on Computers, 56(2): 282–286, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042687.

Chardaire:2007:PBH

[CBM07] P. Chardaire, M. Barake, and G. P. McKeown. A PROBEbased heuristic for graph partitioning. IEEE Transactions on Computers, 56(12):1707– 1720, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358219.

Chatterjee:2000:BAA

[CBP00] M. Chatterjee, S. Banerjee, and D. K. Pradhan. Buffer assignment algorithms on data driven ASICs. IEEE Transactions on Computers, 49(1): 16–32, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=822561.

Choy:2001:NCC

[CBPC01] Chiu-Sing Choy, J. Butas, J. Povazanic, and Cheong-Fat Chan. A new control circuit for asynchronous micropipelines. IEEE Transactions on Computers, 50(9):992– 997, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=954514.

Caccamo:2002:HEO

[CBS02] M. Caccamo, G. Buttazzo, and Lui Sha. Handling execution overruns in hard realtime control systems. IEEE Transactions on Computers, 51 (7):835–849, July 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017703.

Caccamo:2005:ERR

[CBT05] M. Caccamo, G. C. Buttazzo, and D. C. Thomas. Efficient reclaiming in reservation-based real-time systems with variable execution times. IEEE Transactions on Computers, 54(2): 198–213, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377158. See

correction [Ano05f].

Cohen:2002:DAE

[CC02] A. Cohen and R. Cohen. A dynamic approach for efficient TCP buffer allocation. IEEE Transactions on Computers, 51 (3):303–312, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=990128.

Chandra:2003:TDC

[CC03] A. Chandra and K. Chakrabarty.

Test data compression and test resource partitioning for system-on-a-chip using frequencydirected run-length (FDR) codes. IEEE Transactions on Computers, 52(8):1076– 1088, August 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1223641.

Caruso:2007:WCD

[CC07] A. Caruso and S. Chessa. Worst-case diagnosis completeness in regular graphs under the PMC model. IEEE Transactions on Computers, 56(7): 917–924, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216290.

Chang:2006:DMC

[CCC06] Guey-Yun Chang, Gen-Huey Chen, and G. J. Chang. (t; k)- diagnosis for matching composition networks. IEEE Transactions on Computers, 55(1): 88–92, January 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1545754.

Chang:2007:DMC

[CCC07] Guey-Yun Chang, Gen-Huey Chen, and G. J. Chang. (t, k)- diagnosis for matching composition networks under the MM* model. IEEE Transactions on Computers, 56(1): 73–79, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016498.

Cheung:2004:GEI

[CCdS04] P. Y. K. Cheung, G. A.

Constantinides, and J. T. de Sousa. Guest Editors’ introduction: Field programmable logic and applications. IEEE Transactions on Computers, 53(11):1361– 1362, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336758.

Chang:2000:MFA

[CCK00] Pao-Yuan Chang, Deng-Jyi Chen, and K. M. Kavl. Multimedia file allocation on VC networks using multipath routing. IEEE Transactions on Computers, 49(9):971–977, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869329;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=869330.

Chiou:2009:CED

[CCL^{+}09] C. W. Chiou, C.-C. Chang, C.- Y. Lee, T.-W. Hou, and J.-M. Lin. Concurrent error detection and correction in Gaussian normal basis multiplier over GF (2^{m}). IEEE Transactions on Computers, 58(6): 851–857, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4731244.

Caruso:2002:EDA

[CCMS02] A. Caruso, S. Chessa, P. Maestrini, and P. Santi. Evaluation of a diagnosis algorithm for regular structures. IEEE Transactions on Computers, 51(7): 850–865, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1017704.

Cheng:2004:TCA

[CCS^{+}04] M. X. Cheng, M. Cardei, Jinhua Sun, Xiaochun Cheng, Lusheng Wang, Yingfeng Xu, and Ding-Zhu Du. Topology control of ad hoc wireless networks for energy efficiency. IEEE Transactions on Computers, 53(12):1629– 1635, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1347088.

Coleman:2000:AEL

[CCSK00a] J. N. Coleman, E. I. Chester, C. I. Softley, and J. Kadlec. Arithmetic on the European Logarithmic Microprocessor. IEEE Transactions on Computers, 49(7):702–715, July 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863040. See

corrections [CCSK00b].

Coleman:2000:CAE

[CCSK00b] J. N. Coleman, E. I. Chester, C. I. Softley, and J. Kadlec. Corrections to “Arithmetic on the European Logarithmic Microprocessor”. IEEE Transactions on Computers, 49(10): 1152, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888057. See

[CCSK00a].

Chen:2000:PCV

[CCY00] Chichyang Chen, Rui-Lin Chen, and Chih-Huan Yang. Pipelined computation of very large word-length LNS addition/subtraction with polynomial hardware cost. IEEE Transactions on Computers, 49 (7):716–726, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863041.

Chitre:2003:PIB

[CD03] V. A. Chitre and J. N. Daigle.

Performance of IP-based services over GPRS. IEEE Transactions on Computers, 52(6): 727–741, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204829.

Childers:2004:CWC

[CD04] B. R. Childers and J. W. Davidson. Custom wide counterflow pipelines for highperformance embedded applications. IEEE Transactions on Computers, 53(2): 141–158, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261825.

Chen:2009:HSS

[CD09] Zizhong Chen and J. Dongarra. Highly scalable selfhealing algorithms for high performance scientific computing. IEEE Transactions on Computers, 58(11):1512–1524, November 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4799775.

Crovella:2000:EN

[CDE^{+}00] M. Crovella, J. Duato, K. Ebcioglu, L. P. Gewali, P. Montuschi, T. Nanya, K. Schwan, A. Sivasubramaniam, I. G. Tollis, and U. Vishkin. Editor’s note. IEEE Transactions on Computers, 49(10):1009– 1012, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888034.

Cadoli:2006:AC

[CDLS06] M. Cadoli, F. M. Donini, P. Liberatore, and M. Schaerf. kapproximating circuits. IEEE Transactions on Computers, 55 (7):913–917, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637406.

Calhoun:2005:DCU

[CDV^{+}05] B. H. Calhoun, D. C. Daly, Naveen Verma, D. F. Finchelstein, D. D. Wentzloff, A. Wang, Seong-Hwan Cho, and A. P. Chandrakasan. Design considerations for ultra-low energy wireless microsensor nodes. IEEE Transactions on Computers, 54(6):727–740, June 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1424449.

Cameron:2007:AAM

[CGS07] K. W. Cameron, R. Ge, and X.-H. Sun. log_{n} P and log_{3}P: Accurate analytical models of point-to-point communication in distributed systems. IEEE Transactions on Computers, 56 (3):314–327, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079515.

Chen:2006:TES

[CH06] X. Chen and M. S. Hsiao. Testing embedded sequential cores in parallel using spectrumbased BIST. IEEE Transactions on Computers, 55(2): 150–162, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1566576.

Chen:2007:SDI

[CH07a] Chung-Ho Chen and Kuo-Su Hsiao. Scalable dynamic instruction scheduler through wake-up spatial locality. IEEE Transactions on Computers, 56(11):1534– 1548, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4336301.

Chung:2007:LWP

[CH07b] Jaewook Chung and M. Anwar Hasan. Low-weight polynomial form integers for efficient modular multiplication. IEEE Transactions on Computers, 56 (1):44–57, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016496.

Compton:2007:ADA

[CH07c] K. Compton and S. Hauck. Automatic design of area-efficient configurable ASIC cores. IEEE Transactions on Computers, 56 (5):662–672, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141238.

Chang:2006:LTA

[Cha06]  Yeim-Kuan Chang.  A 2-

level TCAM architecture for ranges. IEEE Transactions on Computers, 55(12):1614– 1629, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717392.

Chang:2009:EMP

[Cha09a]  Yeim-Kuan Chang.  Effi-

cient multidimensional packet classification with fast updates. IEEE Transactions on Computers, 58(4):463–479, April 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633349.

Cornea:2009:SII

[CHA^{+}09b] M. Cornea, J. Harrison, C. Anderson, P. Tang, E. Schneider, and E. Gvozdev. A software implementation of the IEEE 754R decimal floating-point arithmetic using the binary encoding format. IEEE Transactions on Computers, 58(2): 148–162, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4674342.

Chang:2005:LCB

[CHC05] Ku-Young Chang, Dowon Hong, and Hyun-Sook Cho. Low complexity bit-parallel multiplier for GF(2^{m}) defined by all-one polynomials using redundant representation. IEEE Transactions on Computers, 54(12):1628– 1630, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524942.

Chen:2004:VIO

[Che04] Pei-Yin Chen. VLSI implementation for one-dimensional multilevel lifting-based wavelet transform. IEEE Transactions on Computers, 53(4): 386–398, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268396.

Chen:2005:CDC

[Che05] Yung-Yuan Chen. Concurrent detection of control flow errors by hybrid signature monitoring. IEEE Transactions on Computers, 54(10):1298– 1313, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501794.

Cohen:2000:OBB

[CHG00] W. E. Cohen, D. W. Hyde, and R. K. Gaede. An optical bus-based distributed dynamic barrier mechanism. IEEE Transactions on Computers, 49(12):1354–1365, December 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895862.

Chen:2000:MDD

[CHH^{+}00] Chiuyuan Chen, F. K. Hwang, Hsien-Sheng Hsiao, Yeh-Hao Chin, and Wei-Pang Yang. The minimum distance diagram of double-loop networks. IEEE Transactions on Computers, 49(9):977–979, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869331.

Chaudhuri:2003:LOB

[CHH^{+}03] M. Chaudhuri, M. Heinrich, C. Holt, J. P. Singh, E. Rothberg, and J. Hennessy. Latency, occupancy, and bandwidth in DSM multiprocessors: a performance evaluation. IEEE Transactions on Computers, 52 (7):862–880, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1214336.

Cohen:2001:CIV

[CHL01] G. D. Cohen, L. Honkala, and A. Lobstein. On codes identifying vertices in the twodimensional square lattice with diagonals. IEEE Transactions on Computers, 50(2): 174–176, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=908992.

Chantem:2009:GES

[CHL09] T. Chantem, Xiaobo Sharon Hu, and M. D. Lemmon. Generalized elastic scheduling for real-time tasks. IEEE Transactions on Computers, 58(4): 480–495, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4626952.

Cilardo:2009:EBP

[Cil09]  A. Cilardo.  Efficient bit-

parallel GF(2^{m}) multiplier for a large class of irreducible pentanomials. IEEE Transactions on Computers, 58(7): 1001–1008, July 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4752811.

Chakrabarty:2002:GCS

[CIQC02] K. Chakrabarty, S. S. Iyengar, Hairong Qi, and Eungchun Cho. Grid coverage for surveillance and target location in distributed sensor networks. IEEE Transactions on Computers, 51(12):1448– 1453, December 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1146711.

Choi:2001:CLS

[CJ01] Kyunghee Choi and Gihyun Jung. Comment on “On-line scheduling policies for a class of IRIS real-time tasks”. IEEE Transactions on Computers, 50 (5):526–528, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926165.

Cuppu:2001:HPD

[CJDM01] V. Cuppu, B. Jacob, B. Davis, and T. Mudge. Highperformance DRAMs in workstation environments. IEEE Transactions on Computers, 50(11):1133–1153, November 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=966491.

Chandra:2005:CBP

[CK05] P. Chandra and A. D. Kshemkalyani. Causality-based predicate detection across space and time. IEEE Transactions on Computers, 54(11):1438–1453, November 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1514422.

Chung:2006:EPH

[CK06] Moo-Kyoung Chung and Chong-Min Kyung. Enhancing performance of HW/SW cosimulation and coemulation by reducing communication overhead. IEEE Transactions on Computers, 55(2):125–136, February 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1566574.

Ciesielski:2006:TED

[CKA06] M. Ciesielski, Priyank Kalla, and S. Askar. Taylor expansion diagrams: a canonical representation for verification of data flow designs. IEEE Transactions on Computers, 55(9):1188–1201, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668046.

Cetintemel:2003:DDP

[CKBF03] U. Cetintemel, P. J. Keleher, B. Bhattacharjee, and M. J. Franklin. Deno: a decentralized, peer-to-peer objectreplication system for weakly connected environments. IEEE Transactions on Computers, 52 (7):943–959, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214342.

Cota:2008:HFC

[CKC^{+}08] E. Cota, F. L. Kastensmidt, M. Cassel, M. Herve, P. Almeida, P. Meirelles, A. Amory, and M. Lubaszewski. A high-fault-coverage approach for the test of data, control and handshake interconnects in mesh networks-on-chip. IEEE Transactions on Computers, 57(9):1202–1215, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487057.

Conti:2002:QSI

[CKDS02] M. Conti, M. Kumar, S. K.

Das, and B. A. Shirazi. Quality of service issues in Internet Web services. IEEE Transactions on Computers, 51(6): 593–594, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1009145.

Cazorla:2006:PPS

[CKS^{+}06] F. J. Cazorla, P. M. W. Knijnenburg, R. Sakellariou, E. Fernandez, A. Ramirez, and M. Valero. Predictable performance in SMT processors: synergy between the OS and SMTs. IEEE Transactions on Computers, 55(7): 785–799, July 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1637396.

Choi:2008:CBT

[CKS^{+}08] Jeonghwan Choi, Youngjae Kim, A. Sivasubramaniam, J. Srebric, Qian Wang, and Joonwon Lee. A CFD-based tool for studying temperature in rack-mounted servers. IEEE Transactions on Computers, 57(8):1129–1142, August 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4479448.

Cuyt:2001:MRA

[CL01] A. Cuyt and R. B. Lenin. Multivariate rational approximants for multiclass closed queuing networks. IEEE Transactions on Computers, 50(11):1279– 1288, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=966500.

Cheon:2006:USA

[CL06] Jung Hee Cheon and Dong Hoon Lee. Use of sparse and/or complex exponents in batch verification of exponentiations. IEEE Transactions on Computers, 55(12):1536–1542, December 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717386.

Chang:2007:DST

[CL07] Yeim-Kuan Chang and Yung-Chieh Lin. Dynamic segment trees for ranges and prefixes. IEEE Transactions on Computers, 56(6): 769–784, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167788.

Chee:2009:LAF

[CL09] Yeow Meng Chee and A. C. H.

Ling. Limit on the addressability of fault-tolerant nanowire decoders. IEEE Transactions on Computers, 58(1):60– 68, January 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4585368.

Calyam:2007:ONW

[CLE^{+}07] P. Calyam, Chang-Gun Lee, E. Ekici, M. Haffner, and N. Howes. Orchestration of network-wide active measurements for supporting distributed computing applications. IEEE Transactions on Computers, 56(12):1629– 1642, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358218.

Chen:2005:CDB

[CLF05] Qingchun Chen, Kam-Yiu Lam, and Pingzhi Fan. Comments on “Distributed Bayesian algorithms for fault-tolerant event region detection in wireless sensor networks”. IEEE Transactions on Computers, 54(9):1182–1183, September 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471678. See

[KI04].

Cai:2009:EET

[CLLL09] Yanli Cai, Wei Lou, Minglu Li, and Xiang-Yang Li. Energy efficient target-oriented scheduling in directional sensor networks. IEEE Transactions on Computers, 58(9):1259–1274, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4798155.

Chang:2004:DCN

[CLTH04] Chien-Ping Chang, Pao-Lien Lai, Jimmy Jiann-Mean Tan, and Lih-Hsing Hsu. Diagnosability of t-connected networks and product networks under the comparison diagnosis model. IEEE Transactions on Computers, 53(12):1582– 1590, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347084.

Cluzeau:2007:RLS

[Clu07]  M. Cluzeau.  Reconstruc-

tion of a linear scrambler. IEEE Transactions on Computers, 56(9):1283–1291, September 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288094.

Cotofana:2005:ARA

[CLV05] Sorin Cotofana, Casper Lageweg, and Stamatis Vassiliadis. Addition related arithmetic operations via controlled transport of charge. IEEE Transactions on Computers, 54(3): 243–256, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0243abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0243.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0243.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388190.
pdf?isnumber=30205&prod=
JNL&arnumber=1388190&arSt=
+243&ared=+256&arAuthor=
Cotofana%2C+S.%3B+Lageweg%
2C+C.%3B+Vassiliadis%2C+
S.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=
1388190&count=13&index=1;
 http://ieeexplore.ieee.
org/xpls/references.jsp?
arnumber=1388190.

Cuyt:2003:CPL

[CLW^{+}03] A. Cuyt, R. B. Lenin, G. Willems, C. Blondia, and P. Rousseeuw. Computing packet loss probabilities in multiplexer models using rational approximation. IEEE Transactions on Computers, 52(5): 633–644, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197129.

Chessa:2001:CAC

[CM01] S. Chessa and P. Maestrini.

Correct and almost complete diagnosis of processor grids. IEEE Transactions on Computers, 50(10):1095–1102, October 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956094.

Chen:2002:PES

[CM02] Xiangping Chen and P. Mohapatra. Performance evaluation of service differentiating Internet servers. IEEE Transactions on Computers, 51(11):1368– 1375, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047760.

Constantinides:2009:FSB

[CMAB09] K. Constantinides, O. Mutlu, T. Austin, and V. Bertacco. A flexible software-based framework for online detection of hardware defects. IEEE Transactions on Computers, 58(8): 1063–1079, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815209.

Chevallier-Mames:2004:LCS

[CMCJ04] B. Chevallier-Mames, M. Ciet, and M. Joye. Low-cost solutions for preventing simple side-channel analysis: sidechannel atomicity. IEEE Transactions on Computers, 53(6): 760–768, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288550.

Chatterjee:2005:JRM

[CMD05] M. Chatterjee, G. D. Mandyam, and S. K. Das. Joint reliability of medium access control and radio link protocol in 3G CDMA systems. IEEE Transactions on Computers, 54(12):1584–1597, December 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524939.

Chen:2003:UBR

[CMK03] Deji Chen, A. K. Mok, and Tei-Wei Kuo. Utilization bound revisited. IEEE Transactions on Computers, 52(3): 351–361, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183949.

Casale:2008:GBN

[CMS08] G. Casale, R. Muntz, and G. Serazzi. Geometric bounds: a noniterative analysis technique for closed queueing networks. IEEE Transactions on Computers, 57(6): 780–794, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4459313.

Conway:2003:NCB

[CN03] R. Conway and J. Nelson. New CRT-based RNS converter using restricted moduli set. IEEE Transactions on Computers, 52 (5):572–578, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197124.

Chin:2005:FII

[CN05] J. Chin and M. Nourani.

FITS: an integrated ILPbased test scheduling environment. IEEE Transactions on Computers, 54(12):1598– 1613, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524940.

Castro:2009:RAL

[CNG^{+}09] F. Castro, R. Noor, A. Garg, D. Chaver, M. C. Huang, L. Piuel, M. Prieto, and F. Tirado. Replacing associative load queues: a timingcentric approach. IEEE Transactions on Computers, 58(4): 496–511, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599576.

Choi:2002:DIP

[CNM^{+}02] Jongmoo Choi, S. H. Noh, Sang Lyul Min, Eun-Yong Ha, and Yookun Cho. Design, implementation, and performance evaluation of a detectionbased adaptive block replacement scheme. IEEE Transactions on Computers, 51(7): 793–800, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1017699.

Cenk:2009:IPM

[CO09] M. Cenk and F. Ozbudak. Improved polynomial multiplication formulas over IF_{2} using Chinese Remainder Theorem. IEEE Transactions on Computers, 58(4):572–576, April 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4674340.

Constantinescu:2000:TSA

[Con00] C. Constantinescu. Teraflops supercomputer: architecture and validation of the fault tolerance mechanisms. IEEE Transactions on Computers, 49(9): 886–894, September 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869320.

Cardarilli:2006:FLE

[COP^{+}06] G. C. Cardarilli, M. Ottavi, S. Pontarelli, M. Re, and A. Salsano. Fault localization, error correction, and graceful degradation in radix 2 signed digitbased adders. IEEE Transactions on Computers, 55(5): 534–540, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613834.

Cherkasova:2002:SBA

[CP02] L. Cherkasova and P. Phaal.

Session-based admission control: a mechanism for peak load management of commercial Web sites. IEEE Transactions on Computers, 51(6): 669–685, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009151.

Chatterjee:2003:PGD

[CP03] Mitrajit Chatterjee and D. K.

Pradhan. A BIST pattern generator design for near-perfect fault coverage. IEEE Transactions on Computers, 52(12):1543–1558, December 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1252851.

Choi:2006:EVC

[CPN^{+}06] Moon-Hee Choi, Woo-Chan Park, F. Neelamkavil, Tack-Don Han, and Shin-Dug Kim. An effective visibility culling method based on cache block. IEEE Transactions on Computers, 55(8):1024–1032, August 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650199.

Campobello:2003:PCR

[CPR03] G. Campobello, G. Patane, and M. Russo. Parallel CRC realization. IEEE Transactions on Computers, 52(10): 1312–1319, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234528.

Cardarilli:2007:AEE

[CPRS07] G. C. Cardarilli, S. Pontarelli, M. Re, and A. Salsano. Analysis of errors and erasures in parity sharing RS codecs. IEEE Transactions on Computers, 56(12):1721–1726, December 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358230.

Channakeshava:2006:UAC

[CRJ06] Karthik Channakeshava, Binoy Ravindran, and E. D. Jensen. Utility accrual channel establishment in multihop networks. IEEE Transactions on Computers, 55(4): 428–442, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608005.

Cho:2007:SOW

[CRJ07] Hyeonjoong Cho, B. Ravindran, and E. D. Jensen. Spaceoptimal, wait-free real-time synchronization. IEEE Transactions on Computers, 56(3): 373–384, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4079519.

Charcranoon:2000:PPC

[CRL00] S. Charcranoon, T. G. Robertazzi, and S. Luryi. Parallel processor configuration design with processing/transmission costs. IEEE Transactions on Computers, 49(9):987–991, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=869332.

Chin:2009:MDL

[CRS09] Tai-Lin Chin, P. Ramanathan, and K. K. Saluja. Modeling detection latency with collaborative mobile sensing architecture. IEEE Transactions on Computers, 58(5): 692–705, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4641910.

Conte:2000:PRS

[CS00] T. M. Conte and S. Sathaye. Properties of rescheduling size invariance for dynamic rescheduling-based VLIW crossgeneration compatibility. IEEE Transactions on Computers, 49 (8):814–825, August 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=868027.

Chanchio:2004:CST

[CS04] K. Chanchio and X.-H. Sun. Communication state transfer for the mobility of concurrent heterogeneous computing. IEEE Transactions on Computers, 53(10):1260– 1273, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1327577.

Chang:2008:CTB

[CS08a] Yeim-Kuan Chang and Cheng-Chien Su. Comments on “A TCAM-Based Parallel Architecture for High-Speed Packet Forwarding”. IEEE Transactions on Computers, 57(4): 574–576, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4433984. See

[ANPS07].

Chung:2008:DSM

[CS08b] Sung Woo Chung and K. Skadron.

On-demand solution to minimize I-cache leakage energy with maintaining performance. IEEE Transactions on Computers, 57(1):7– 24, January 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358227.

Cho:2009:AMD

[CS09a] H. Cho and E. E. Swartzlander. Adder and multiplier design in quantum-dot cellular automata. IEEE Transactions on Computers, 58(6): 721–727, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4760137.

Cho:2009:ESB

[CS09b] Junho Cho and Wonyong Sung. Efficient softwarebased encoding and decoding of BCH codes. IEEE Transactions on Computers, 58(7): 878–889, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4782950.

Coleman:2008:ELM

[CSK^{+}08] J. N. Coleman, C. I. Softley, J. Kadlec, R. Matousek, M. Tichy, Z. Pohl, A. Hermanek, and N. F. Benschop. The European Logarithmic Microprocessor. IEEE Transactions on Computers, 57(4): 532–546, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358243.

Clouqueur:2004:FTC

[CSR04] T. Clouqueur, K. K. Saluja, and P. Ramanathan. Fault tolerance in collaborative sensor networks for target detection. IEEE Transactions on Computers, 53(3):320–333, March 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261838.

Cremonesi:2002:UFA

[CSS02] P. Cremonesi, P. J. Schweitzer, and G. Serazzi. A unifying framework for the approximate solution of closed multiclass queuing networks. IEEE Transactions on Computers, 51(12):1423–1434, December 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1146708.

Chan:2003:MDH

[CT03] S.-H. G. Chan and F. A.

Tobagi. Modeling and dimensioning hierarchical storage systems for low-delay video services. IEEE Transactions on Computers, 52(7): 907–919, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214339.

Cheng:2005:EEI

[CT05] A. C. Cheng and G. S. Tyson.

An energy efficient instruction set synthesis framework for low power embedded system designs. IEEE Transactions on Computers, 54(6): 698–713, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461358.

Chiang:2009:UND

[CT09] Chieh-Feng Chiang and J. J. M. Tan. Using node diagnosability to determine t-diagnosability under the comparison diagnosis model. IEEE Transactions on Computers, 58(2): 251–259, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4609378.

Chen:2002:QSFa

[CTA02a] Wei Chen, S. Toueg, and M. K. Aguilera. On the quality of service of failure detectors. IEEE Transactions on Computers, 51 (1):13–32, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980014.

Chen:2002:QSFb

[CTA02b] Wei Chen, Sam Toueg, and M. K. Aguilera. On the quality of service of failure detectors. IEEE Transactions on Computers, 51(5): 561–580, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1004595.

Cancela:2002:ARS

[CU02] H. Cancela and M. E. Urquhart.

Adapting RVR simulation techniques for residual connectedness network reliability models. IEEE Transactions on Computers, 51(4): 439–443, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=995453.

Cheng:2000:STC

[CUT00] Fu-Chiung Cheng, S. H. Unger, and M. Theobald. Self-timed carry-lookahead adders. IEEE Transactions on Computers, 49 (7):659–672, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863035.

Cotofana:2000:SDA

[CV00] S. Cotofana and S. Vassiliadis. Signed digit addition and related operations with threshold logic. IEEE Transactions on Computers, 49(3): 193–207, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841124.

Chishti:2008:OPP

[CV08] Z. Chishti and T. N. Vijaykumar. Optimal power/ performance pipeline depth for SMT in scaled technologies. IEEE Transactions on Computers, 57(1):69–81, January 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358228.

Chen:2002:LSC

[CWC02] Jianer Chen, Guojun Wang, and Songqiao Chen. Locally subcube-connected hypercube networks: theoretical analysis and experimental results. IEEE Transactions on Computers, 51 (5):530–540, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004592.

Cao:2007:FTG

[CWC07] Jiannong Cao, Guojun Wang, and K. C. C. Chan. A faulttolerant group communication protocol in large scale and highly dynamic mobile nextgeneration networks. IEEE Transactions on Computers, 56 (1):80–94, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016499.

Codrescu:2001:AAC

[CWM01] L. Codrescu, D. S. Wills, and J. Meindl. Architecture of the Atlas chip-multiprocessor: dynamically parallelizing irregular applications. IEEE Transactions on Computers, 50(1): 67–82, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=902753.

Che:2008:DDR

[CWZL08] Hao Che, Zhijun Wang, Kai Zheng, and Bin Liu. DRES: Dynamic range encoding scheme for TCAM coprocessors. IEEE Transactions on Computers, 57 (7):902–915, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378358.

Chen:2007:BSC

[CX07] Hui Chen and Yang Xiao. Onbound selection cache replacement policy for wireless data access. IEEE Transactions on Computers, 56(12):1597– 1611, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358225.

Chen:2006:IDO

[CXP06] Baoxing Chen, Wenjun Xiao, and B. Parhami. Internode distance and optimal routing in a class of alternating group networks. IEEE Transactions on Computers, 55(12):1645– 1648, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1717395.

Chan:2009:SPT

[CY09] E. P. F. Chan and Yaya Yang. Shortest path tree computation in dynamic graphs. IEEE Transactions on Computers, 58 (4):541–557, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4663055.

Cho:2001:HBM

[CYL01] Sangyeun Cho, Pen-Chung Yew, and Gyungho Lee. A highbandwidth memory pipeline for wide issue processors. IEEE Transactions on Computers, 50 (7):709–723, July 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936237.

Clark:2005:ACI

[CZM05] N. T. Clark, H. Zhong, and S. A. Mahlke. Automated custom instruction generation for domain-specific processor acceleration. IEEE Transactions on Computers, 54(10): 1258–1270, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501791.

Dadda:2007:MPD

[Dad07] Luigi Dadda. Multioperand parallel decimal adder: a mixed binary and BCD approach. IEEE Transactions on Computers, 56(10):1320– 1328, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302705.

Lima:2003:OFP

[dALB03] G. de A. Lima and A. Burns.

An optimal fixed-priority assignment algorithm for supporting fault-tolerant hard realtime systems. IEEE Transactions on Computers, 52(10): 1332–1346, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1234530.

Datta:2005:FTP

[Dat05] A. Datta. A fault-tolerant protocol for energy-efficient permutation routing in wireless networks. IEEE Transactions on Computers, 54(11):1409– 1421, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514420.

Das:2000:IRS

[DBB00] D. K. Das, U. K. Bhattacharya, and B. B. Bhattacharya. Isomorph-redundancy in sequential circuits. IEEE Transactions on Computers, 49(9): 992–997, September 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869333.

Dasgupta:2001:CAB

[DCCS01] P. Dasgupta, S. Chattopadhyay, P. P. Chaudhuri, and I. Sengupta. Cellular automatabased recursive pseudoexhaustive test pattern generator. IEEE Transactions on Computers, 50(2):177–185, February 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=908993.

deDinechin:2005:MTM

[dDT05] Florent de Dinechin and Arnaud Tisserand. Multipartite table methods. IEEE Transactions on Computers, 54(3): 319–330, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0319abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0319.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/

t0319.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388196.
pdf?isnumber=30205&prod=
JNL&arnumber=1388196&arSt=
+319&ared=+330&arAuthor=
de+Dinechin%2C+F.%3B+Tisserand%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=
1388196&count=13&index=7;
 http://ieeexplore.ieee.
org/xpls/references.jsp?
arnumber=1388196.

Duggirala:2003:NVI

[DGZA03] R. Duggirala, R. Gupta, Qing-An Zeng, and D. P. Agrawal. Notice of violation of IEEE publication principles “Performance enhancements of ad hoc networks with localized route repair”. IEEE Transactions on Computers, 52(7): 854–861, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214335.

Daneshbeh:2005:CUB

[DH05] Amir K. Daneshbeh and M. Anwar Hasan. A class of unidirectional bit serial systolic architectures for multiplicative inversion and division over GF(2^{m}). IEEE Transactions on Computers, 54(3): 370–380, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0370abs.htm;

                                        http://csdl.computer.org/
                                        dl/trans/tc/2005/03/t0370.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tc/2005/03/
                                        t0370.pdf;   http://ieeexplore.
                                        ieee.org/iel5/12/30205/01388201.
                                        pdf?isnumber=30205&prod=
                                        JNL&arnumber=1388201&arSt=
                                        +370&ared=+380&arAuthor=
                                        Daneshbeh%2C+A.K.%3B+Hasan%
                                        2C+M.A.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=30205&arnumber=
                                        1388201&count=13&index=12;
                                         http://ieeexplore.ieee.
                                        org/xpls/references.jsp?
                                        arnumber=1388201.

Dolev:2006:SSM

[DH06] S. Dolev and Y. A. Haviv. Self-stabilizing microprocessor: analyzing and overcoming soft errors. IEEE Transactions on Computers, 55(4): 385–399, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1608002.

Dahab:2006:SMU

[DHH^{+}06] R. Dahab, D. Hankerson, F. Hu, M. Long, J. Lopez, and A. Menezes. Software multiplication using Gaussian normal bases. IEEE Transactions on Computers, 55(8): 974–984, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650195.

Daza:2008:CSD

[DHS08] V. Daza, J. Herranz, and G. Saez. On the computational security of a distributed key distribution scheme. IEEE Transactions on Computers, 57(8): 1087–1097, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4479446.

Dimitrov:2008:PSP

[DJJ^{+}08] V. S. Dimitrov, K. U. Jarvinen, M. J. Jacobson, W. Chan, and Zhun Huang. Provably sublinear point multiplication on Koblitz curves and its hardware implementation. IEEE Transactions on Computers, 57(11):1469– 1481, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487060. See

comments [Lee12].

Dimitrov:2000:CFA

[DJM00] V. S. Dimitrov, G. A. Jullien, and W. C. Miller. Complexity and fast algorithms for multiexponentiations. IEEE Transactions on Computers, 49(2): 141–147, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833110.

DeLaLuz:2004:ARU

[DK04] V. De La Luz and M. Kandemir. Array regrouping and its use in compiling data-intensive, embedded applications. IEEE Transactions on Computers, 53 (1):1–19, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1255787.

David:2007:HCM

[DKT07] J. P. David, K. Kalach, and N. Tittley. Hardware complexity of modular multiplication and exponentiation. IEEE Transactions on Computers, 56(10):1308–1319, October 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302704.

Delaluz:2001:HST

[DKV^{+}01] V. Delaluz, M. Kandemir, N. Vijaykrishnan, A. Sivasubramaniam, and M. J. Irwin. Hardware and software techniques for controlling DRAM power modes. IEEE Transactions on Computers, 50(11):1154– 1173, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966492.

Das:2003:POV

[DLBS03] S. K. Das, E. Lee, K. Basu, and S. K. Sen. Performance optimization of VoIP calls over wireless links using H.323 protocol. IEEE Transactions on Computers, 52(6): 742–752, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204830.

Dehon:2007:ISS

[DLL07] A. Dehon, C. S. Lent, and F. Lombardi. Introduction to the special section on nano systems and computing. IEEE Transactions on Computers, 56(2):145–146, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042675.

Daumas:2009:VRN

[DLM09] M. Daumas, D. Lester, and C. Muoz. Verified real number calculations: a library for interval arithmetic. IEEE Transactions on Computers, 58(2): 226–237, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4685896.

Dai:2007:PRT

[DLT07] Y.-S. Dai, G. Levitin, and K. S. Trivedi. Performance and reliability of tree-structured grid services considering data dependence and failure correlation. IEEE Transactions on Computers, 56(7): 925–936, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216291.

Dubrova:2000:CGB

[DM00a] E. Dubrova and L. Macchiarulo. A comment on “Graph-based algorithm for Boolean function manipulation”. IEEE Transactions on Computers, 49(11):1290– 1292, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895944.

Dubrova:2000:ETM

[DM00b] E. V. Dubrova and J. C. Muzio. Easily testable multiplevalued logic circuits derived from Reed–Muller circuits. IEEE Transactions on Computers, 49(11):1285–1289, November 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895943.

DiPietro:2009:DSU

[DMS^{+}09] R. Di Pietro, L. V. Mancini, C. Soriente, A. Spognardi, and G. Tsudik. Data security in unattended wireless sensor networks. IEEE Transactions on Computers, 58(11):1500– 1511, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5184811.

Dimitrakopoulos:2005:HSP

[DN05] G. Dimitrakopoulos and D. Nikolos. High-speed parallel-prefix VLSI Ling adders. IEEE Transactions on Computers, 54(2): 225–231, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377160.

Dautovic:2006:CBF

[DN06] S. Dautovic and L. Novak. A comment on “Boolean functions classification via fixed polarity Reed–Muller form”. IEEE Transactions on Computers, 55(8):1067–1069, August 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650204.

DAlberto:2005:LSA

[DNVG05] P. D’Alberto, A. Nicolau, A. Veidenbaum, and Rajesh Gupta. Line size adaptivity analysis of parameterized loop nests for direct mapped data cache. IEEE Transactions on Computers, 54(2): 185–197, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377157.

deOliveira:2009:EDL

[dOBNL09] H. A. B. F. de Oliveira, A. Boukerche, E. F. Nakamura, and A. A. F. Loureiro. An efficient directed localization recursion protocol for wireless sensor networks. IEEE Transactions on Computers, 58 (5):677–691, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4721367.

Durresi:2005:OBP

[DPIK05] A. Durresi, V. K. Paruchuri, S. S. Iyengar, and R. Kannan. Optimized broadcast protocol for sensor networks. IEEE Transactions on Computers, 54(8):1013–1024, August 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453502.

Dai:2007:HMA

[DPZ07] Y.-H. Dai, Y. Pan, and X. Zou. A hierarchical modeling and analysis for Grid service reliability. IEEE Transactions on Computers, 56(5): 681–691, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141240.

Dimitrijevic:2005:SSP

[DRC05] Z. Dimitrijevic, R. Rangaswami, and E. Y. Chang. Systems support for preemptive disk scheduling. IEEE Transactions on Computers, 54(10): 1314–1326, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501795.

Natale:2000:SDR

[DS00] M. Di Natale and J. A.

Stankovic. Scheduling distributed real-time tasks with minimum jitter. IEEE Transactions on Computers, 49(4): 303–316, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844344.

Datta:2000:BHW

[DSK00] A. Datta, S. H. Son, and V. Kumar. Is a bird in the hand worth more than two in the bush? Limitations of priority cognizance in conflict resolution for firm real-time database systems. IEEE Transactions on Computers, 49(5): 482–502, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=859541.

Darte:2005:LBM

[DSV05] A. Darte, R. Schreiber, and G. Villard. Lattice-based memory allocation. IEEE Transactions on Computers, 54(10): 1242–1257, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501790.

Danysh:2005:AIV

[DT05] Albert Danysh and Dimitri Tan. Architecture and implementation of a vector/SIMD multiply-accumulate unit. IEEE Transactions on Computers, 54(3):284– 293, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0284abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0284.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0284.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388193.
pdf?isnumber=30205&prod=
JNL&arnumber=1388193&arSt=

+284&ared=+293&arAuthor=
Danysh%2C+A.%3B+Tan%2C+D.;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30205&arnumber=1388193&count=
13&index=4;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1388193.

Deshpande:2009:OMB

[DTHS09] S. Deshpande, M. Thottan, Tin Kam Ho, and B. Sikdar. An online mechanism for BGP instability detection and analysis. IEEE Transactions on Computers, 58(11):1470– 1484, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5161252.

Duale:2004:MEF

[DU04] A. Y. Duale and M. U. Uyar. A method enabling feasible conformance test sequence generation for EFSM models. IEEE Transactions on Computers, 53 (5):614–627, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275300.

Dharmasena:2004:LBL

[DV04] H. P. Dharmasena and R. Vaidyanathan.

Lower bounds on the loading of multiple bus networks for binary tree algorithms. IEEE Transactions on Computers, 53(12):1535– 1546, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347080.

Desmet:2007:EIS

[DVJP07] L. Desmet, P. Verbaeten, W. Joosen, and F. Piessens. Enlarging instruction streams. IEEE Transactions on Computers, 56(10):1342–1357, October 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302707.

Deng:2006:IMM

[DZ06] Rui Deng and Yujie Zhou. Improvement to Montgomery modular inverse algorithm. IEEE Transactions on Computers, 55(9):1207–1210, September 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668048.

Davis:2008:EES

[DZB08] R. I. Davis, A. Zabos, and A. Burns. Efficient exact schedulability tests for fixed priority real-time systems. IEEE Transactions on Computers, 57(9):1261–1276, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487061.

Du:2000:MHC

[DZZ00] Xing Du, Xiaodong Zhang, and Zhichun Zhu. Memory hierarchy considerations for cost-effective cluster computing. IEEE Transactions on Computers, 49(9):915–933, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869323.

Ebcioglu:2001:DBT

[EAGS01] K. Ebcioglu, E. Altman, M. Gschwind, and S. Sathaye. Dynamic binary translation and optimization. IEEE Transactions on Computers, 50(6): 529–548, June 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=931892.

Elarief:2009:DCA

[EB09] N. Elarief and B. Bose. Diversity combining ARQ over the m(≥ 2)-ary unidirectional channel. IEEE Transactions on Computers, 58(8):1026– 1034, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815216.

Ergin:2006:ERD

[EBPG06] O. Ergin, D. Balkan, D. Ponomarev, and K. Ghose. Early register deallocation mechanisms using checkpointed register files. IEEE Transactions on Computers, 55(9):1153–1166, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668043.

Ebeling:2004:IOR

[EFX^{+}04] C. Ebeling, C. Fisher, Guanbin Xing, Manyuan Shen, and Hui Liu. Implementing an OFDM receiver on the RaPiD reconfigurable architecture. IEEE Transactions on Computers, 53(11):1436– 1448, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336764.

Emako:2003:MAB

[EGP03] B. Emako, R. H. Glitho, and S. Pierre. A mobile agentbased advanced service architecture for wireless Internet telephony: design, implementation, and evaluation. IEEE Transactions on Computers, 52 (6):690–705, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204826.

Erle:2009:DFP

[EHS09] Mark A. Erle, Brian J. Hickmann, and Michael J. Schulte. Decimal floating-point multiplication. IEEE Transactions on Computers, 58(7): 902–916, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4711044.

Ercegovac:2000:IGD

[EIM^{+}00] Miloš D. Ercegovac, Laurent Imbert, David W. Matula, Jean-Michel Muller, and Guoheng Wei. Improving Goldschmidt division, square root, and square root reciprocal. IEEE Transactions on Computers, 49(7): 759–763, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863046.

Eum:2004:CRS

[EKK04] Nak-Woong Eum, Taewhan Kim, and Chong-Min Kyung. CeRA: a router for symmetrical FPGAs based on exact routing density evaluation. IEEE Transactions on Computers, 53 (7):829–842, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321044.

Elia:2002:ISC

[EL02] M. Elia and M. Leone. On the inherent space complexity of fast parallel multipliers for GF(2^{m}). IEEE Transactions on Computers, 51(3): 346–351, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=990131.

Ercegovac:2000:RSR

[ELMT00] Miloš D. Ercegovac, Tomás Lang, Jean-Michel Muller, and Arnaud Tisserand. Reciprocation, square root, inverse square root, and some elementary functions using small multipliers. IEEE Transactions on Computers, 49(7): 628–637, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=863031.

Even:2000:DIC

[EP00] G. Even and W. J. Paul. On the design of IEEE compliant floating point units. IEEE Transactions on Computers, 49 (5):398–413, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=859536.

Elerath:2009:HAM

[EP09] J. G. Elerath and M. Pecht.

A highly accurate method for assessing reliability of Redundant Arrays of Inexpensive Disks (RAID). IEEE Transactions on Computers, 58(3): 289–299, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624244.

Even:2000:CTR

[ES00] G. Even and P.-M. Seidel. A comparison of three rounding algorithms for IEEE floatingpoint multiplication. IEEE Transactions on Computers, 49 (7):638–650, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863033.

Ermedahl:2005:CWC

[ESE05] A. Ermedahl, F. Stappert, and J. Engblom. Clustered worstcase execution-time calculation. IEEE Transactions on Computers, 54(9):1104–1122, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471672.

Ercetin:2001:PBI

[ET01] O. Ercetin and L. Tassiulas. Push-based information delivery in two stage satellite-terrestrial wireless systems. IEEE Transactions on Computers, 50(5):506–518, May 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926163.

Ercetin:2003:MBR

[ET03] O. Ercetin and L. Tassiulas. Market-based resource allocation for content delivery in the Internet. IEEE Transactions on Computers, 52(12):1573– 1585, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252853.

Efstathiou:2005:EDM

[EVDN05] C. Efstathiou, H. T. Vergos, G. Dimitrakopoulos, and D. Nikolos. Efficient diminished-1 modulo 2^{n} + 1 multipliers. IEEE Transactions on Computers, 54(4):491–496, April 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401868.

Efstathiou:2003:MAD

[EVN03] C. Efstathiou, H. T. Vergos, and D. Nikolos. Modulo 2^{n} ± 1 adder design using select-prefix blocks. IEEE Transactions on Computers, 52(11):1399– 1406, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244938.

Efstathiou:2004:FPP

[EVN04a] C. Efstathiou, H. T. Vergos, and D. Nikolos. Fast parallel-prefix modulo 2^{n} + 1 adders. IEEE Transactions on Computers, 53(9):1211–1216, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315614.

Efstathiou:2004:MBM

[EVN04b] C. Efstathiou, H. T. Vergos, and D. Nikolos. Modified Booth modulo 2^{n}−1 multipliers. IEEE Transactions on Computers, 53 (3):370–374, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261842.

Falkowski:2003:CGR

[Fal03] B. J. Falkowski. A comment on “Generalized Reed– Muller forms as a tool to detect symmetries”. IEEE Transactions on Computers, 52(7): 975–976, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1214346.

Frikken:2006:ABA

[FAL06] K. Frikken, M. Atallah, and J. Li. Attribute-based access control with hidden policies and hidden credentials. IEEE Transactions on Computers, 55(10):1259–1270, October 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1683757.

Fang:2002:GMP

[Fan02] Yuguang Fang. General modeling and performance analysis for location management in wireless mobile networks. IEEE Transactions on Computers, 51(10):1169– 1181, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1039843.

Fang:2003:MBM

[Fan03a] Yuguang Fang. Movementbased mobility management and trade off analysis for wireless mobile networks. IEEE Transactions on Computers, 52 (6):791–803, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204834.

Fang:2003:TSC

[Fan03b]  Yuguang Fang.  Thinning

schemes for call admission control in wireless networks. IEEE Transactions on Computers, 52 (5):685–688, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197135;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1197136.

Favalli:2006:DAP

[Fav06] M. Favalli. Diversity analysis in the presence of delay faults affecting duplex systems. IEEE Transactions on Computers, 55 (3):348–352, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1583563.

Fan:2000:TMF

[FB00] C. C. Fan and J. Bruck. Tolerating multiple faults in multistage interconnection networks with minimal extra stages. IEEE Transactions on Computers, 49(9):998–1004, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869334.

Franceschetti:2004:GTN

[FCB04] M. Franceschetti, M. Cook, and J. Bruck. A geometric theorem for network design. IEEE Transactions on Computers, 53 (4):483–489, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268406.

Fan:2005:FBP

[FD05] Haining Fan and Yiqi Dai. Fast bit-parallel GF(2^{n}) multiplier for all trinomials. IEEE Transactions on Computers, 54(4): 485–490, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401867.

Feng:2005:NEMa

[FDBS05a] G.-L. Feng, R. H. Deng, F. Bao, and J.-C. Shen. New efficient MDS array codes for RAID. Part I. Reed–Solomon-like codes for tolerating three disk failures. IEEE Transactions on Computers, 54(9):1071–1080, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471669.

Feng:2005:NEMb

[FDBS05b] G.-L. Feng, R. H. Deng, F. Bao, and J.-C. Shen. New efficient MDS array codes for RAID. Part II. Rabin-like codes for tolerating multiple (>= 4) disk failures. IEEE Transactions on Computers, 54(12):1473– 1483, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524930.

Fraguela:2003:PME

[FDZ03] B. B. Fraguela, R. Doallo, and E. L. Zapata. Probabilistic miss equations: evaluating memory hierarchy performance. IEEE Transactions on Computers, 52 (3):321–336, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183947.

Fetzer:2003:PFD

[Fet03] C. Fetzer. Perfect failure detection in timed asynchronous systems. IEEE Transactions on Computers, 52(2):99–112, February 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176979.

Fine:2006:HML

[FFJ^{+}06] S. Fine, A. Freund, I. Jaeger, Y. Mansour, Y. Naveh, and A. Ziv. Harnessing machine learning to improve the success rate of stimuli generation. IEEE Transactions on Computers, 55(11):1344–1355, November 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705444.

Fit-Florea:2009:DLN

[FFLTM09] A. Fit-Florea, L. Li, M. A. Thornton, and D. W. Matula. A discrete logarithm number system for integer arithmetic modulo 2^{k}: Algorithms and lookup structures. IEEE Transactions on Computers, 58(2): 163–174, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4663061.

Fedeli:2007:PIE

[FFP07] A. Fedeli, F. Fummi, and G. Pravadelli. Properties incompleteness evaluation by functional verification. IEEE Transactions on Computers, 56 (4):528–544, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118675.

Ferrandi:2002:TGT

[FFS02] F. Ferrandi, F. Fummi, and D. Sciuto. Test generation and testability alternatives exploration of critical algorithms for embedded applications. IEEE Transactions on Computers, 51(2): 200–215, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980008.

Fan:2006:RBM

[FH06] H. Fan and M. A. Hasan.

Relationship between GF(2^{m}) Montgomery and shifted polynomial basis multiplication algorithms. IEEE Transactions on Computers, 55(9):1202– 1206, September 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668047.

Fan:2007:CFS

[FH07a] Haining Fan and A. Hasan.

Comments on “Five, Six, and Seven-Term Karatsuba-Like Formulae”. IEEE Transactions on Computers, 56(5): 716–717, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141244. See

[Mon05].

Fan:2007:NAS

[FH07b] Haining Fan and M. Anwar Hasan. A new approach to subquadratic space complexity parallel multipliers for extended binary fields. IEEE Transactions on Computers, 56(2): 224–233, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4042682.

Fan:2007:SCC

[FH07c] Haining Fan and M. Anwar Hasan. Subquadratic computational complexity schemes for extended binary field multiplication using optimal normal bases. IEEE Transactions on Computers, 56(10): 1435–1437, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302715.

Fong:2004:FIP

[FHLM04] K. Fong, D. Hankerson, J. Lopez, and A. Menezes. Field inversion and point halving revisited. IEEE Transactions on Computers, 53(8):1047– 1059, August 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306996.

Fiore:2008:EAW

[Fio08] P. D. Fiore. Efficient approximate wordlength optimization. IEEE Transactions on Computers, 57(11):1561– 1570, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4531730.

Fan:2005:DBG

[FL05] Jianxi Fan and Xiaola Lin.

The t/k-diagnosability of the BC graphs. IEEE Transactions on Computers, 54(2): 176–184, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377156.

Flich:2003:ATB

[FLM^{+}03] J. Flich, P. Lopez, M. P.

Malumbres, J. Duato, and T. Rokicki. Applying in-transit buffers to boost the performance of networks with source routing. IEEE Transactions on Computers, 52(9):1134–1153, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228510.

Fan:2003:GMR

[FLW03] Hongbing Fan, Jiping Liu, and Yu-Liang Wu. General models and a reduction design technique for FPGA switch box designs. IEEE Transactions on Computers, 52(1): 21–30, January 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1159751.

Feng:2003:AAM

[FML03] Wen-Yi Feng, F. J. Meyer, and F. Lombardi. Adaptive algorithms for maximal diagnosis of wiring interconnects. IEEE Transactions on Computers, 52(10):1259– 1270, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234524.

Friedman:2007:AAR

[FMRR07] R. Friedman, A. Mostefaoui, S. Rajsbaum, and M. Raynal. Asynchronous agreement and its relation with errorcorrecting codes. IEEE Transactions on Computers, 56(7): 865–875, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216286.

Felber:2004:ESC

[FN04] P. Felber and P. Narasimhan. Experiences, strategies, and challenges in building faulttolerant CORBA systems. IEEE Transactions on Computers, 53(5):497–511, May 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1275293.

Faezipour:2009:WST

[FN09] M. Faezipour and M. Nourani.

Wire-speed TCAM-based architectures for multimatch packet classification. IEEE Transactions on Computers, 58 (1):5–17, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4609379.

Feero:2009:NCT

[FP09] B. S. Feero and P. P. Pande.

Networks-on-chip in a threedimensional environment: a performance evaluation. IEEE Transactions on Computers, 58 (1):32–45, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4599574.

Frachtenberg:2006:SSR

[FPFP06] E. Frachtenberg, F. Petrini, J. Fernandez, and S. Pakin. STORM: Scalable resource management for large-scale parallel computers. IEEE Transactions on Computers, 55(12): 1572–1587, December 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1717389.

Fiorin:2008:SMA

[FPL^{+}08] L. Fiorin, G. Palermo, S. Lukovic, V. Catalano, and C. Silvano. Secure memory accesses on networks-on-chip. IEEE Transactions on Computers, 57(9):1216–1229, September 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4492766.

Frougny:2000:FAS

[Fro00] C. Frougny. On-the-fly algorithms and sequential machines. IEEE Transactions on Computers, 49(8):859–863, August 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=868030.

Fummi:2000:HTG

[FS00] F. Fummi and D. Sciuto. A hierarchical test generation approach for large controllers. IEEE Transactions on Computers, 49(4): 289–302, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844343.

Fujita:2007:MLD

[FS07] H. Fujita and K. Sakaniwa. Modified low-density MDS array codes for tolerating double disk failures in disk arrays. IEEE Transactions on Computers, 56(4):563–566, April 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118678.

Fischmeister:2007:VLP

[FSL07] S. Fischmeister, O. Sokolsky, and Insup Lee. A verifiable language for programming real-time communication schedules. IEEE Transactions on Computers, 56(11):1505– 1519, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336299.

Fujita:2000:NID

[Fuj00a] S. Fujita. Neighbourhood information dissemination in the star graph. IEEE Transactions on Computers, 49(12):1366– 1370, December 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895865.

Fujiwara:2000:NCS

[Fuj00b] H. Fujiwara. A new class of sequential circuits with combinational test generation complexity. IEEE Transactions on Computers, 49(9):895–905, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=869321.

Fan:2002:CGU

[FWC02] Hongbing Fan, Yu-Liang Wu, and Yao-Wen Chang. Comment on “Generic universal switch blocks”. IEEE Transactions on Computers, 51(1): 93–95, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980020. See

[SWCC00].

Fan:2006:DDT

[FWCL06] H. Fan, Yu-Liang Wu, R. C. C. Cheung, and J. Liu. Decomposition design theory and methodology for arbitraryshaped switch boxes. IEEE Transactions on Computers, 55 (4):373–384, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1608001.

Field:2000:VNV

[FZM00] B. Field, T. F. Znati, and D. Mosse. V-NET: a versatile network architecture for flexible delay guarantees in realtime networks. IEEE Transactions on Computers, 49(8): 841–858, August 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=868029.

Gaudiot:2001:EN

[Gau01] J. Gaudiot. Editor’s note.

IEEE Transactions on Computers, 50(9):863–864, September 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954502.

Gaudiot:2002:EN

[Gau02]  J.-L. Gaudiot.  Editor’s

note. IEEE Transactions on Computers, 51(7):737–739, July 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017693.

Ghosh:2007:NPC

[GBD07] P. Ghosh, K. Basu, and S. K. Das. A novel photonic container switched architecture and scheduler to design the core transport network. IEEE Transactions on Computers, 56(8):1087– 1104, August 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4264323.

Gniady:2006:PCB

[GBHL06] C. Gniady, A. R. Butt, Y. C.

Hu, and Yung-Hsiang Lu. Program counter-based prediction techniques for dynamic power management. IEEE Transactions on Computers, 55(6): 641–658, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1628954.

Golic:2000:GIA

[GCD00] J. D. Golic, A. Clark, and E. Dawson. Generalized inversion attack on nonlinear filter generators. IEEE Transactions on Computers, 49(10): 1100–1109, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888045.

Ganguly:2006:NRT

[GCI06] Samrat Ganguly, Mainak Chatterjee, and R. Izmailov. Nonreal-time content scheduling algorithms for wireless data networks. IEEE Transactions on Computers, 55(7): 893–905, July 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637404.

Govindan:2009:XCC

[GCN^{+}09] S. Govindan, Jeonghwan Choi, A. R. Nath, A. Das, B. Urgaonkar, and A. Sivasubramaniam. Xen and Co.:

Communication-aware CPU management in consolidated Xen-based hosting platforms. IEEE Transactions on Computers, 58(8):1111–1125, August 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815210.

Gunaratne:2008:REC

[GCNS08] C. Gunaratne, K. Christensen, B. Nordman, and S. Suen. Reducing the energy consumption of Ethernet with Adaptive Link Rate (ALR). IEEE Transactions on Computers, 57 (4):448–461, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378356.

Gopalakrishnan:2008:STS

[GCS08] S. Gopalakrishnan, M. Caccamo, and Lui Sha. Sharp thresholds for scheduling recurring tasks with distance constraints. IEEE Transactions on Computers, 57(3): 344–358, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358253.

Gunther:2003:EMM

[GD03] W. Gunther and R. Drechsler. Efficient minimization and manipulation of linearly transformed binary decision diagrams. IEEE Transactions on Computers, 52(9):1196–1209, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228514.

Guido:2007:GEI

[GDM07] Rodrigo Capobianco Guido, Li Deng, and Shoji Makino. Guest Editors’ introduction: Special section on emergent systems, algorithms and architectures for speech-based human-machine interaction. IEEE Transactions on Computers, 56(9):1153–1155, September 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288082.

Genbrugge:2008:MDF

[GE08] D. Genbrugge and L. Eeckhout. Memory data flow modeling in statistical simulation for the efficient exploration of microprocessor design spaces. IEEE Transactions on Computers, 57 (1):41–54, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358235.

Genbrugge:2009:CMD

[GE09] D. Genbrugge and L. Eeckhout. Chip multiprocessor design space exploration through statistical simulation. IEEE Transactions on Computers, 58(12):1668–1681, December 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5010432.

Gonzalez:2001:CFS

[GG01] J. Gonzalez and A. Gonzalez. Control-flow speculation through value prediction. IEEE Transactions on Computers, 50(12):1362–1376, December 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970574.

Ghosh:2001:PAD

[Gho01] S. Ghosh. P^{2}EDAS: asynchronous, distributed event driven simulation algorithm with inconsistent event preemption for accurate execution of VHDL descriptions on parallel processors. IEEE Transactions on Computers, 50(1): 28–50, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=902751.

Gao:2003:RRB

[GHP03] Y. Gao, J. C. Hou, and Sanjoy Paul. RACCOOM: a rate-based congestion control approach for multicast. IEEE Transactions on Computers, 52(12):1521– 1534, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1252849.

Giraud:2006:RIR

[Gir06] C. Giraud. An RSA implementation resistant to fault attacks and to simple power analysis. IEEE Transactions on Computers, 55(9):1116–1120, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1668039.

Ganesh:2003:PPM

[GKM03] A. J. Ganesh, A.-M. Kermarrec, and L. Massoulie. Peerto-peer membership management for gossip-based protocols. IEEE Transactions on Computers, 52(2):139–149, February 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1176982.

Guneysu:2008:CC

[GKN^{+}08] T. Guneysu, T. Kasper, M. Novotny, C. Paar, and A. Rupp. Cryptanalysis with COPACOBANA. IEEE Transactions on Computers, 57(11): 1498–1513, November 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4515858.

Golubchik:2002:BPM

[GL02] L. Golubchik and J. C. S.

Lui. Bounding of performance measures for thresholdbased queuing systems: theory and application to dynamic resource management in video-on-demand servers. IEEE Transactions on Computers, 51 (4):353–372, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=995445.

Guan:2003:SMC

[GL03] Sheng-Uei Guan and Wei Liu. Self-modifiable color Petri nets for modeling user manipulation and network event handling. IEEE Transactions on Computers, 52(7): 920–932, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214340.

Gedik:2005:SPP

[GL05] B. Gedik and L. Liu. A scalable peer-to-peer architecture for distributed information monitoring applications. IEEE Transactions on Computers, 54 (6):767–782, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461363.

Lee:2007:MTC

[gLMK07] Chan gun Lee, A. K. Mok, and P. Konana. Monitoring of timing constraints with confidence threshold requirements. IEEE Transactions on Computers, 56 (7):977–991, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216295.

Geiselmann:2002:NRE

[GMQS02] W. Geiselmann, J. Muller-Quade, and R. Steinwandt. On “A new representation of elements of finite fields GF(2^{m}) yielding small complexity arithmetic circuits”. IEEE Transactions on Computers, 51(12):1460–1461, December 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1146713.

Goel:2009:TSH

[GMSC09] S. Goel, E. J. Marinissen, A. Sehgal, and K. Chakrabarty. Testing of SoCs with hierarchical cores: Common fallacies, test access optimization, and test scheduling. IEEE Transactions on Computers, 58(3): 409–423, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624250.

Gomez:2006:RMA

[GNF^{+}06] M. E. Gomez, N. A. Nordbotten, J. Flich, P. Lopez, A. Robles, J. Duato, T. Skeie, and O. Lysne. A routing methodology for achieving fault tolerance in direct networks. IEEE Transactions on Computers, 55 (4):400–415, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608003.

Gollmann:2002:ESP

[Gol02] D. Gollmann. Equally spaced polynomials, dual bases, and multiplication in F (2^{n}). IEEE Transactions on Computers, 51 (5):588–591, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004597.

Golic:2006:NMD

[Gol06]  J. D. Golic.  New meth-

ods for digital generation and postprocessing of random data. IEEE Transactions on Computers, 55(10): 1217–1229, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1683753.

Granger:2005:HSN

[GPS05] R. Granger, D. Page, and M. Stam. Hardware and software normal basis arithmetic for pairing-based cryptography in characteristic three. IEEE Transactions on Computers, 54 (7):852–860, July 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1432668.

Guerraoui:2004:ISI

[GR04] R. Guerraoui and M. Raynal. The information structure of indulgent consensus. IEEE Transactions on Computers, 53 (4):453–466, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268403.

Gupta:2007:MCM

[GR07] U. Gupta and N. Ranganathan. Multievent crisis management using noncooperative multistep games. IEEE Transactions on Computers, 56(5): 577–589, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141234.

Graillat:2009:AFP

[Gra09] Stef Graillat. Accurate floatingpoint product and exponentiation. IEEE Transactions on Computers, 58(7):994– 1000, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4711041.

Grossman:2004:AMF

[Gro04] J. P. Grossman. Analytically modeling a fault-tolerant messaging protocol. IEEE Transactions on Computers, 53(7): 870–878, July 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321047.

Gordon-Ross:2005:FLD

[GRV05] A. Gordon-Ross and F. Vahid.

Frequent loop detection using efficient nonintrusive onchip hardware. IEEE Transactions on Computers, 54(10): 1203–1215, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501787.

Gallagher:2000:FTN

[GS00a] W. L. Gallagher and E. E. Swartzlander, Jr. Faulttolerant Newton–Raphson and Goldschmidt dividers using time shared TMR. IEEE Transactions on Computers, 49(6): 588–595, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=862218.

Gao:2000:LCN

[GS00b] G. R. Gao and V. Sarkar. Location consistency — a new memory model and cache consistency protocol. IEEE Transactions on Computers, 49(8): 798–813, August 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=868026.

Geiselmann:2003:RRD

[GS03] W. Geiselmann and R. Steinwandt. A redundant representation of GF(q^{n}) for designing arithmetic circuits. IEEE Transactions on Computers, 52 (7):848–853, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214334.

Gourgy:2009:CTR

[GS09] A. Gourgy and T. H. Szymanski. Cooperative token-ring scheduling for input-queued switches. IEEE Transactions on Computers, 58(3): 351–364, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4766374.

Gok:2006:IMO

[GSA06] M. Gok, M. J. Schulte, and M. G. Arnold. Integer multipliers with overflow detection. IEEE Transactions on Computers, 55(8):1062– 1066, August 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650203.

Gibert:2005:DDC

[GSG05] E. Gibert, J. Sanchez, and A. Gonzalez. Distributed data cache designs for clustered VLIW processors. IEEE Transactions on Computers, 54(10): 1227–1241, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501789.

Guilley:2008:SEW

[GSH^{+}08] S. Guilley, L. Sauvage, P. Hoogvorst, R. Pacalet, G. M. Bertoni, and S. Chaudhuri. Security evaluation of WDDL and SecLib countermeasures against power attacks. IEEE Transactions on Computers, 57(11):1482– 1497, November 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4585361.

Gaubatz:2008:SCD

[GSS08] G. Gaubatz, E. Savas, and B. Sunar. Sequential circuit design for embedded cryptographic applications resilient to adversarial faults. IEEE Transactions on Computers, 57(1): 126–138, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358236.

Gong:2002:PMP

[GSW02] Linguo Gong, Xian-He Sun, and E. F. Watson. Performance modeling and prediction of nondedicated network computing. IEEE Transactions on Computers, 51(9):1041–1055, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032624.

Gomathisankaran:2006:ASO

[GT06] Mahadevan Gomathisankaran and A. Tyagi. Architecture support for 3D obfuscation. IEEE Transactions on Computers, 55 (5):497–507, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613831.

Gramoli:2009:SDS

[GVB^{+}09] V. Gramoli, Y. Vigfusson, K. Birman, A.-M. Kermarrec, and R. van Renesse. Slicing distributed systems. IEEE Transactions on Computers, 58(11):1444–1455, November 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5184813.

Garcia-Vidal:2006:DSM

[GVMC^{+}06] J. Garcia-Vidal, M. March, L. Cerda, J. Corbal, and M. Valero. A DRAM/ SRAM memory scheme for fast packet buffers. IEEE Transactions on Computers, 55(5): 588–602, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613839.

Gentile:2004:PVS

[GW04] A. Gentile and D. S. Wills.

Portable video supercomputing. IEEE Transactions on Computers, 53(8):960–973, August 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1306990.

Govindarajan:2003:MRI

[GYA^{+}03] R. Govindarajan, Hongbo Yang, J. N. Amaral, Chihong Zhang, and G. R. Gao. Minimum register instruction sequencing to reduce register spills in out-of-order issue superscalar architectures. IEEE Transactions on Computers, 52 (1):4–20, January 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1159750.

Huang:2005:BLC

[HA05] C. Huang and T. Abdelzaher. Bounded-latency content distribution feasibility and evaluation. IEEE Transactions on Computers, 54(11):1422– 1437, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514421.

Hamdioui:2006:ODF

[HAAvdG06] S. Hamdioui, Z. Al-Ars, and A. J. van de Goor. Opens and delay faults in CMOS RAM address decoders. IEEE Transactions on Computers, 55(12):1630– 1639, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717393.

Hadjicostis:2005:APC

[Had05] C. N. Hadjicostis. Aliasing probability calculations for arbitrary compaction under independently selected random test vectors. IEEE Transactions on Computers, 54(12):1614– 1627, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524941.

Hao:2006:CCB

[HAD06] F. Hao, R. Anderson, and J. Daugman. Combining crypto with biometrics effectively. IEEE Transactions on Computers, 55(9):1081–1088, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668036.

Homma:2008:SAD

[HAH08] N. Homma, T. Aoki, and T. Higuchi. A systematic approach for designing redundant arithmetic adders based on counter tree diagrams. IEEE Transactions on Computers, 57(12):1633– 1646, December 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4569835.

Hariyama:2005:GAM

[HAK05] M. Hariyama, T. Aoyama, and M. Kameyama. Genetic approach to minimizing energy consumption of VLSI processors using multiple supply voltages. IEEE Transactions on Computers, 54(6): 642–650, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461353.

Harris:2004:EUO

[Har04] D. Harris. An exponentiation unit for an OpenGL lighting engine. IEEE Transactions on Computers, 53(3): 251–258, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261833.

Harris:2006:GEI

[Har06] I. G. Harris. Guest Editor’s introduction to the special section on simulation-based design validation. IEEE Transactions on Computers, 55(11):1313– 1314, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705440.

Hasan:2000:LTB

[Has00] M. Anwarul Hasan. Lookup table-based large finite field multiplication in memory constrained cryptosystems. IEEE Transactions on Computers, 49 (7):749–758, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863045.

Hasan:2001:PAA

[Has01] M. A. Hasan. Power analysis attacks and algorithmic approaches to their countermeasures for Koblitz curve cryptosystems. IEEE Transactions on Computers, 50(10): 1071–1083, October 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956092.

Horvath:2007:DVS

[HASL07] T. Horvath, T. Abdelzaher, K. Skadron, and Xue Liu. Dynamic voltage scaling in multitier Web servers with endto-end delay control. IEEE Transactions on Computers, 56 (4):444–458, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118669.

Haynal:2001:ABS

[HB01a] S. Haynal and F. Brewer. Automata-based symbolic scheduling for looping DFGs. IEEE Transactions on Computers, 50 (3):250–267, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=910815.

Hsu:2001:GEI

[HB01b] Wei Hsu and V. Bala. Guest Editors’ introduction. IEEE Transactions on Computers, 50 (6):527–528, June 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=931891.

Holcomb:2009:PSS

[HBF09] D. E. Holcomb, W. P. Burleson, and K. Fu. Power-up SRAM state as an identifying fingerprint and source of true random numbers. IEEE Transactions on Computers, 58(9):1198– 1210, September 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4674345.

Heath:2005:STD

[HBH05] M. W. Heath, W. P. Burleson, and I. G. Harris. Synchrotokens: a deterministic GALS methodology for chip-level debug and test. IEEE Transactions on Computers, 54(12): 1532–1546, December 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1524935.

Hurson:2001:MGE

[HC01] A. R. Hurson and B. Childers.

Message from the Guest Editors. IEEE Transactions on Computers, 50(8):767–768, August 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=946997.

Hsieh:2006:HPE

[HC06] Sun-Yuan Hsieh and Nai-Wen Chang. Hamiltonian path embedding and pancyclicity on the Möbius cube with faulty nodes and faulty edges. IEEE Transactions on Computers, 55 (7):854–863, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637401.

Hsieh:2008:SDP

[HC08a] Sun-Yuan Hsieh and Yu-Shu Chen. Strongly diagnosable product networks under the comparison diagnosis model. IEEE Transactions on Computers, 57(6): 721–732, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4459310.

Hsieh:2008:SDS

[HC08b] Sun-Yuan Hsieh and Yu-Shu Chen. Strongly diagnosable systems under the comparison diagnosis model. IEEE Transactions on Computers, 57(12):1720–1725, December 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4569833.

Huang:2000:AEC

[HCC^{+}00] Shi-Yu Huang, Kwang-Ting Cheng, Kuang-Chien Chen, Chung-Yang Huang, and F. Brewer. AQUILA: an equivalence checking system for large sequential designs. IEEE Transactions on Computers, 49(5): 443–464, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=859539.

Hsieh:2001:LFF

[HCH01] Sun-Yuan Hsieh, Gen-Huey Chen, and Chin-Wen Ho. Longest fault-free paths in star graphs with edge faults. IEEE Transactions on Computers, 50(9):960–971, September 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954510.

Han:2009:ICS

[HCK09] Dong-Guk Han, Dooho Choi, and Howon Kim. Improved computation of square roots in specific finite fields. IEEE Transactions on Computers, 58(2):188–196, February 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4663058.

Hsiung:2007:MCS

[HCL07] Pao-Ann Hsiung, Yean-Ru Chen, and Yen-Hung Lin. Model checking safety-critical systems using Safecharts. IEEE Transactions on Computers, 56 (5):692–705, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141241.

Huang:2009:MPS

[HDQK09] Pai-Han Huang, M. Desai, Xiaofan Qiu, and B. Krishnamachari. On the multihop performance of synchronization mechanisms in high propagation delay networks. IEEE Transactions on Computers, 58 (5):577–590, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4721366.

Huang:2005:HPL

[HE05] Zhijun Huang and Miloš D. Ercegovac. High-performance low-power left-to-right array multiplier design. IEEE Transactions on Computers, 54(3): 272–283, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0272abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0272.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0272.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388192.
pdf?isnumber=30205&prod=
JNL&arnumber=1388192&arSt=
+272&ared=+283&arAuthor=
Zhijun+Huang%3B+Ercegovac%
2C+M.D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=
1388192&count=13&index=3;
 http://ieeexplore.ieee.
org/xpls/references.jsp?
arnumber=1388192.

Heys:2003:ASC

[Hey03] H. M. Heys. Analysis of the statistical cipher feedback mode of block ciphers. IEEE Transactions on Computers, 52(1): 77–92, January 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1159755.

Haung:2002:DCA

[HH02] Yieh-Ran Haung and Jan-Ming Ho. Distributed call admission control for a heterogeneous PCS network. IEEE Transactions on Computers, 51(12):1400–1409, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1146706.

Huang:2004:DSM

[HH04] Yin-Fu Huang and Jiing-Maw Huang. Disk scheduling on multimedia storage servers. IEEE Transactions on Computers, 53 (1):77–82, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255793.

Han:2000:DET

[HHTH00] Ching-Chih Han, Chao-Ju Hou, Kar Sun Tsoi, and S. Ho. Dynamic establishment and termination of real-time message streams in dual-bus networks. IEEE Transactions on Computers, 49(10):1110– 1119, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888046.

Hiasat:2000:NES

[Hia00] A. A. Hiasat. New efficient structure for a modular multiplier for RNS. IEEE Transactions on Computers, 49(2): 170–174, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=833113.

Hiasat:2002:HSR

[Hia02] A. A. Hiasat. High-speed and reduced-area modular adder structures for RNS. IEEE Transactions on Computers, 51 (1):84–89, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980018.

Hierons:2004:TNF

[Hie04] R. M. Hierons. Testing from a nondeterministic finite state machine using adaptive state counting. IEEE Transactions on Computers, 53(10):1330– 1342, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327582.

Hamacher:2001:HRN

[HJ01] V. C. Hamacher and Hong Jiang. Hierarchical ring network configuration and performance modeling. IEEE Transactions on Computers, 50(1):

1–12, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=902749.

Hiller:2004:EPP

[HJS04] M. Hiller, A. Jhumka, and Neeraj Suri. EPIC: profiling the propagation and effect of data errors in software. IEEE Transactions on Computers, 53 (5):512–530, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275294.

Halbutogullari:2000:MMG

[HK00a] A. Halbutogullari and C. K.

Koc. Mastrovito multiplier for general irreducible polynomials. IEEE Transactions on Computers, 49(5): 503–518, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=859542.

Hosseini-Khayat:2000:ORN

[HK00b] S. Hosseini-Khayat. On optimal replacement of nonuniform cache objects. IEEE Transactions on Computers, 49(8): 769–778, August 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=868024.

Han:2001:PMH

[HKA01] Gang Han, R. H. Klenke, and J. H. Aylor. Performance modeling of hierarchical crossbar-based multicomputer systems. IEEE Transactions on Computers, 50(9): 877–890, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954504.

Hadimioglu:2001:ISS

[HKL01] H. Hadimioglu, D. Kaeli, and F. Lombardi. Introduction to the special section on high performance memory systems. IEEE Transactions on Computers, 50(11):1103– 1104, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=966487.

Hankerson:2009:AGL

[HKM09] D. Hankerson, K. Karabina, and A. Menezes. Analyzing the Galbraith–Lin–Scott point multiplication method for elliptic curves over binary fields. IEEE Transactions on Computers, 58(10): 1411–1420, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815218.

Huang:2000:EVR

[HL00] J. Huang and D. J. Lilja.

Extending value reuse to basic blocks with compiler support. IEEE Transactions on Computers, 49(4):331–347, April 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844346.

Hsueh:2001:SRT

[HL01] Chih-Wen Hsueh and Kwei-Jay Lin. Scheduling real-time systems with end-to-end timing constraints using the distributed pinwheel model. IEEE Transactions on Computers, 50 (1):51–66, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=902752.

Hermant:2002:FAU

[HL02] J.-F. Hermant and G. Le Lann. Fast asynchronous uniform consensus in real-time distributed systems. IEEE Transactions on Computers, 51(8): 931–944, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024740.

Huang:2003:BRO

[HL03] Jian Huang and D. J. Lija. Balancing reuse opportunities and performance gains with subblock value reuse. IEEE Transactions on Computers, 52(8): 1032–1050, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1223638.

Hashempour:2005:AAC

[HL05] H. Hashempour and F. Lombardi. Application of arithmetic coding to compression of VLSI test data. IEEE Transactions on Computers, 54(9):1166– 1177, September 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471676.

Hilewitz:2009:NBS

[HL09] Y. Hilewitz and R. B. Lee. A new basis for shifters in general-purpose processors for existing and advanced bit manipulations. IEEE Transactions on Computers, 58(8): 1035–1048, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4721365.

Hsu:2000:TAN

[HLLR00] Tsan-Sheng Hsu, J. C. Lee, D. R. Lopez, and W. A. Royce.

Task allocation on a network of processors. IEEE Transactions on Computers, 49(12):1339– 1353, December 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895858.

Heinrich-Litan:2000:LUB

[HLM00] L. Heinrich-Litan and P. Molitor. Least upper bounds for the size of OBDDs using symmetry properties. IEEE Transactions on Computers, 49(4): 360–368, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844348.

Hsu:2004:FHF

[HLTH04] Hong-Chun Hsu, Tseng-Kuei Li, J. J. M. Tan, and Lih-Hsing Hsu. Fault hamiltonicity and fault Hamiltonian connectivity of the arrangement graphs. IEEE Transactions on Computers, 53(1): 39–53, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255789.

Huang:2000:ADM

[HML00] Wei Kang Huang, F. J. Meyer, and F. Lombardi. An approach for detecting multiple faulty FPGA logic blocks. IEEE Transactions on Computers, 49 (1):48–54, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=822563.

Hopkins:2006:DSS

[HMM06] A. B. T. Hopkins and K. D. McDonald-Maier. Debug support strategy for systems-on-chips with multiple processor cores. IEEE Transactions on Computers, 55(2): 174–184, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1566578.

Hurfin:2002:VFC

[HMR02] M. Hurfin, A. Mostefaoui, and M. Raynal. A versatile family of consensus protocols based on Chandra–Toueg’s unreliable failure detectors. IEEE Transactions on Computers, 51(4): 395–408, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=995450.

Heath:2004:CTE

[HPH^{+}04] T. Heath, E. Pinheiro, J. Hom, U. Kremer, and R. Bianchini. Code transformations for energy-efficient device management. IEEE Transactions on Computers, 53(8):

974–987, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1306991.

Han:2002:SDS

[HPS02] Yijie Han, Yi Pan, and Hong Shen. Sublogarithmic deterministic selection on arrays with a reconfigurable optical bus. IEEE Transactions on Computers, 51(6): 702–707, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009153.

Hegazy:2002:UAB

[HR02] T. Hegazy and B. Ravindran.

Using application benefit for proactive resource allocation in asynchronous real-time distributed systems. IEEE Transactions on Computers, 51(8): 945–962, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024741.

Hanchate:2006:SID

[HR06] N. Hanchate and N. Ranganathan. Simultaneous interconnect delay and crosstalk noise optimization through gate sizing using game theory. IEEE Transactions on Computers, 55(8):1011–1023, August 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650198.

Hariri:2009:BSB

[HRM09] A. Hariri and A. Reyhani-Masoleh. Bit-serial and bitparallel Montgomery multiplication and squaring over GF(2^{m}). IEEE Transactions on Computers, 58(10):1332– 1345, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4912195.

Ho:2004:NAF

[HS04] C.-T. Ho and L. Stockmeyer. A new approach to fault-tolerant wormhole routing for mesh-connected parallel computers. IEEE Transactions on Computers, 53(4): 427–438, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268400.

Han:2001:SBA

[HSH01] Ching-Chih Han, K. G. Shin, and Chao-Ju Hou. Synchronous bandwidth allocation for real-time communications with the timed-token MAC protocol. IEEE Transactions on Computers, 50(5):

414–431, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926157.

Hosseinabady:2008:STS

[HSLN08] Mohammad Hosseinabady, Shervin Sharifi, Fabrizio Lombardi, and Zainalabedin Navabi. A selective trigger scan architecture for VLSI testing. IEEE Transactions on Computers, 57(3): 316–328, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358251.

Huang:2008:APR

[HSS^{+}08] Wei Huang, K. Sankaranarayanan, K. Skadron, R. J. Ribando, and M. R. Stan. Accurate, pre-RTL temperatureaware design using a parameterized, geometric thermal model. IEEE Transactions on Computers, 57(9):1277–1288, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487059.

Hiltunen:2003:BSS

[HSU03] M. A. Hiltunen, R. D. Schlichting, and C. A. Ugarte. Building survivable services using redundancy and adaptation. IEEE Transactions on Computers, 52(2):181–194, February 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1176985.

Han:2003:FTS

[HSW03] Ching-Chih Han, K. G. Shin, and Jian Wu. A fault-tolerant scheduling algorithm for realtime periodic tasks with possible software faults. IEEE Transactions on Computers, 52 (3):362–372, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1183950.

Han:2000:LBM

[HSY00] Ching-Chih Han, K. G. Shin, and Sang Kyun Yun. On load balancing in multicomputer/ distributed systems equipped with circuit or cut-through switching capability. IEEE Transactions on Computers, 49 (9):947–957, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=869325.

Horita:2000:FTP

[HT00] T. Horita and I. Takanami. Fault-tolerant processor arrays based on the 11/2-track switches with flexible spare distributions. IEEE Transactions on Computers, 49(6): 542–552, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=862214.

Hsieh:2008:CFM

[HTKL08] Jen-Wei Hsieh, Yi-Lin Tsai, Tei-Wei Kuo, and Tzao-Lin Lee. Configurable flashmemory management: Performance versus overheads. IEEE Transactions on Computers, 57(11):1571–1583, November 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487056.

Hierons:2002:RLC

[HU02] R. M. Hierons and H. Ural.

Reduced length checking sequences. IEEE Transactions on Computers, 51(9):1111–1117, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1032630.

Hierons:2006:OLC

[HU06] R. M. Hierons and H. Ural. Optimizing the length of checking sequences. IEEE Transactions on Computers, 55(5): 618–629, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613841.

Hierons:2009:CRL

[HU09] R. M. Hierons and H. Ural. Correction to “Reduced Length Checking Sequences” [Sep 02 93–99]. IEEE Transactions on Computers, 58(2): 287, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4626950.

Harada:2007:ARA

[HUN07] F. Harada, T. Ushio, and Y. Nakamoto. Adaptive resource allocation control for fair QoS management. IEEE Transactions on Computers, 56(3): 344–357, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079517.

Hodjat:2006:ATT

[HV06] A. Hodjat and I. Verbauwhede. Area-throughput trade-offs for fully pipelined 30 to 70 gbits/s AES processors. IEEE Transactions on Computers, 55(4): 366–372, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1608000.

Hopper:2009:PSS

[HvAL09] N. Hopper, L. von Ahn, and J. Langford. Provably secure steganography. IEEE Transactions on Computers, 58(5): 662–676, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4663056.

Hamdioui:2002:ETR

[HvdG02] S. Hamdioui and A. J. van de Goor. Efficient tests for realistic faults in dual-port SRAMs. IEEE Transactions on Computers, 51(5): 460–473, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004586.

Hasan:2000:VAA

[HW00] M. A. Hasan and A. G. Wassal. VLSI algorithms, architectures, and implementation of a versatile GF(2^{m}) processor. IEEE Transactions on Computers, 49(10):1064– 1073, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888042.

Huang:2007:GPU

[HW07] Chien-Lin Huang and Chung-Hsien Wu. Generation of phonetic units for mixed-language speech recognition based on acoustic and contextual analysis. IEEE Transactions on Computers, 56(9):1225–1233, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288089.

Hellebrand:2002:EOO

[HWI^{+}02] S. Hellebrand, H.-J. Wunderlich, A. A. Ivaniuk, Y. V. Klimets, and V. N. Yarmolik. Efficient online and offline testing of embedded DRAMs. IEEE Transactions on Computers, 51(7): 801–809, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017700.

Hsia:2007:CFC

[HWW07] Chi-Chun Hsia, Chung-Hsien Wu, and Jian-Qi Wu. Conversion function clustering and selection using linguistic and spectral information for emotional voice conversion. IEEE Transactions on Computers, 56(9):1245–1254, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288091.

Huang:2008:SEC

[HX08] Cheng Huang and Lihao Xu.

STAR : An efficient coding scheme for correcting triple storage node failures. IEEE Transactions on Computers, 57 (7):889–901, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358290.

Hua:2009:BTS

[HXW09] Yu Hua, Bin Xiao, and Jianping Wang. BR-tree: a scalable prototype for supporting multiple queries of multidimensional data. IEEE Transactions on Computers, 58(12):1585– 1598, December 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5184804.

Halak:2008:FTT

[HY08] B. Halak and A. V. Yakovlev.

Fault-tolerant techniques to minimize the impact of crosstalk on phase encoded communication channels. IEEE Transactions on Computers, 57(4): 505–519, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358288.

Iyer:2000:DES

[IB00] R. R. Iyer and L. N. Bhuyan. Design and evaluation of a switch cache architecture for CC-NUMA multiprocessors. IEEE Transactions on Computers, 49(8):779–797, August 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=868025.

Iyengar:2002:TBS

[IC02] V. Iyengar and K. Chakrabarty. Test bus sizing for systemon-a-chip. IEEE Transactions on Computers, 51(5): 449–459, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004585.

Iyengar:2003:TAM

[ICM03] Vikram Iyengar, Krishnendu Chakrabarty, and E. J. Marinissen. Test access mechanism optimization, test scheduling, and tester data volume reduction for systemon-chip. IEEE Transactions on Computers, 52(12):1619– 1632, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252857.

Iguchi-Cartigny:2009:LME

[ICRSR^{+}09] J. Iguchi-Cartigny, P. M. Ruiz, D. Simplot-Ryl, I. Stojmenovic, and C. M. Yago. Localized minimum-energy broadcasting for wireless multihop networks with directional antennas. IEEE Transactions on Computers, 58(1): 120–131, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4585365.

Minato:2002:SBM

[iM02] Shin ichi Minato. Streaming BDD manipulation. IEEE Transactions on Computers, 51 (5):474–485, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004587.

Izumi:2006:CAS

[IM06] T. Izumi and T. Masuzawa.

Condition adaptation in synchronous consensus. IEEE Transactions on Computers, 55 (7):843–853, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637400.

Imre:2007:QET

[Imr07] S. Imre. Quantum existence testing and its application for finding extreme values in unsorted databases. IEEE Transactions on Computers, 56(5): 706–710, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141242.

Iyer:2009:MDM

[ING09] S. M. Iyer, M. K. Nakayama, and A. V. Gerbessiotis. A Markovian dependability model with cascading failures. IEEE Transactions on Computers, 58(9):1238–1249, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4785456.

Izadi:2003:ECA

[IO03] B. A. Izadi and F. Ozguner. Enhanced cluster k-ary ncube, a fault-tolerant multiprocessor. IEEE Transactions on Computers, 52(11):1443– 1453, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244942.

Inoue:2006:PJO

[ISF06] H. Inoue, D. Stefanovic, and S. Forrest. On the prediction of Java object lifetimes. IEEE Transactions on Computers, 55 (7):880–892, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637403.

Imana:2006:BPF

[IST06] J. L. Imana, J. M. Sanchez, and F. Tirado. Bit-parallel finite field multipliers for irreducible trinomials. IEEE Transactions on Computers, 55 (5):520–533, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613833.

Jayapala:2005:CLB

[JBV^{+}05] M. Jayapala, F. Barat, T. Vander Aa, F. Catthoor, H. Corporaal, and G. Deconinck. Clustered loop buffer organization for low energy VLIW embedded processors. IEEE Transactions on Computers, 54(6): 672–683, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1461356.

Jeong:2006:CRA

[JD06] J. Jeong and M. Dubois. Cache replacement algorithms with nonuniform miss costs. IEEE Transactions on Computers, 55 (4):353–365, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1607999.

Jiang:2007:CMB

[JDZ07] Song Jiang, K. Davis, and Xiaodong Zhang. Coordinated multilevel buffer cache management with consistent access locality quantification. IEEE Transactions on Computers, 56 (1):95–108, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016500.

Jejurikar:2006:OSR

[JG06] R. Jejurikar and R. Gupta. Optimized slowdown in realtime task systems. IEEE Transactions on Computers, 55(12):1588–1598, December 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1717390.

Jha:2003:CTP

[Jha03] P. K. Jha. A counterexample to Tang and Padubidri’s claim about the bisection width of a diagonal mesh. IEEE Transactions on Computers, 52 (5):676–677, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1197133.

Jha:2013:CXM

[Jha13] Pranava K. Jha. Comments on “Multiple-Radix Gray Codes in Lee Metric”. IEEE Transactions on Computers, 62(1): 200, January 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [ABA07].

Johnson:2001:LTC

[JHZ01] E. E. Johnson, Jiheng Ha, and M. Baqar Zaidi. Lossless trace compression. IEEE Transactions on Computers, 50(2): 158–173, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=908991.

Jiang:2004:NCC

[Jia04] Jehn-Ruey Jiang. On the nondomination of cohorts coteries. IEEE Transactions on Computers, 53(7):922–923, July 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321051. See

comments [KC07b].

Jaleel:2006:LIH

[JJ06] A. Jaleel and B. Jacob. Inline interrupt handling and lock-up free translation lookaside buffers (TLBs). IEEE Transactions on Computers, 55 (5):559–574, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613837.

Jaberipur:2009:ISP

[JK09] G. Jaberipur and A. Kaivani. Improving the speed of parallel decimal multiplication. IEEE Transactions on Computers, 58(11):1539–1552, November 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5184812.

Joshi:2007:ASS

[JLJ07] A. Joshi, Yue Luo, and L. K. John. Applying statistical sampling for fast and efficient simulation of commercial workloads. IEEE Transactions on Computers, 56(11):1520– 1533, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336300.

Jiang:2004:GSM

[JLN04] Y. Jiang, J. Li, and S. Nishimura. A general stochastic model for dynamic locking in database systems. IEEE Transactions on Computers, 53(3): 308–319, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261837.

Ji:2009:CLO

[JLZ^{+}09] Rong Ji, Zhiqiang Ling, Xianjun Zeng, Bingcai Sui, Liang Chen, Junfeng Zhang, Yingjie Feng, and Gang Luo. Comments on “Leading-One Prediction with Concurrent Position Correction”. IEEE Transactions on Computers, 58(12):1726–1727, December 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4799776.

Jacob:2001:UVM

[JM01] B. Jacob and T. Mudge.

Uniprocessor virtual memory without TLBs. IEEE Transactions on Computers, 50(5): 482–499, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926161.

Jovanov:2002:ANO

[JMH02] E. Jovanov, V. Milutinovic, and A. R. Hurson. Acceleration of nonnumeric operations using hardware support for the Ordered Table Hashing algorithms. IEEE Transactions on Computers, 51(9):1026–1040, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032623.

Jabir:2007:GBU

[JP07] A. M. Jabir and D. K. Pradhan. A graph-based unified technique for computing and representing coefficients over finite fields. IEEE Transactions on Computers, 56(8): 1119–1132, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4264325.

Joshi:2006:MBS

[JPEJ06] Ajay Joshi, Aashish Phansalkar, L. Eeckhout, and L. K. John. Measuring benchmark similarity using inherent program characteristics. IEEE Transactions on Computers, 55(6): 769–782, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628963.

Jeong:2004:CEP

[JPJ^{+}04] Jong-Chul Jeong, Woo-Chan Park, Woong Jeong, Tack-Don Han, and Moon-Key Lee. A cost-effective pipelined divider with a small lookup table. IEEE Transactions on Computers, 53 (4):489–495, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268407.

Jabir:2007:TRM

[JPSR07] A. M. Jabir, D. K. Pradhan, A. K. Singh, and T. L. Rajaprabhu. A technique for representing multiple output binary functions with applications to verification and simulation. IEEE Transactions on Computers, 56(8): 1133–1145, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4264326.

Jensen:2002:GEI

[JR02] E. D. Jensen and B. Ravindran. Guest editors’ introduction to special section on asynchronous real-time distributed systems. IEEE Transactions on Computers, 51(8): 881–882, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024736.

Jigang:2006:RAP

[JS06] W. Jigang and T. Srikanthan.

Reconfiguration algorithms for power efficient VLSI subarrays with four-port switches. IEEE Transactions on Computers, 55 (3):243–253, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1583555.

Jiang:2009:EES

[JS09] Yingxin Jiang and A. Striegel. An exploration of the effects of state granularity through (m, k) real-time streams. IEEE Transactions on Computers, 58 (6):784–798, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4731243.

Jankovic:2001:DDM

[JSD01] D. Jankovic, R. S. Stankovic, and R. Drechsler. Decision diagram method for calculation of pruned Walsh transform. IEEE Transactions on Computers, 50(2):147–157, February 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=908990.

Jankovic:2009:OPE

[JSM09] D. Jankovic, R. S. Stankovic, and C. Moraga. Optimization of polynomial expressions by using the extended dual polarity. IEEE Transactions on Computers, 58(12):1710– 1725, December 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=5184815.

Joshi:2006:TPD

[JSW^{+}06] N. Joshi, J. Sundararajan, K. Wu, Bo Yang, and R. Karri. Tamper proofing by design using generalized involutionbased concurrent error detection for involutional substitution permutation and Feistel networks. IEEE Transactions on Computers, 55(10): 1230–1239, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683754.

Jigang:2007:IRC

[JSW07] Wu Jigang, T. Srikanthan, and Xiaodong Wang. Integrated row and column rerouting for reconfiguration of VLSI arrays with four-port switches. IEEE Transactions on Computers, 56(10):1387– 1400, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302710.

Jingxi:2007:ALD

[JVG07] Jia Jingxi, B. Veeravalli, and D. Ghose. Adaptive load distribution strategies for divisible load processing on resource unaware multilevel tree networks. IEEE Transactions on Computers, 56(7): 999–1005, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216297.

Jain:2001:NCO

[JWF01] J. Jain, I. Wegener, and M. Fujita. A note on complexity of OBDD composition and efficiency of partitioned-OBDDs over OB-DDs. IEEE Transactions on Computers, 50(11):1289–1290, November 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966501.

Joye:2000:OLR

[JY00] M. Joye and Sung-Ming Yen. Optimal left-to-right binary signed-digit recoding. IEEE Transactions on Computers, 49 (7):740–748, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=863044.

Jiang:2005:MLF

[JZ05] S. Jiang and X. Zhang. Making LRU friendly to weak locality workloads: a novel replacement algorithm to improve buffer cache performance. IEEE Transactions on Computers, 54 (8):939–952, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453496.

Kagaris:2003:MST

[Kag03] D. Kagaris. Multiple-seed TPG structures. IEEE Transactions on Computers, 52(12):1633– 1639, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252858.

Kahan:2004:CFP

[Kah04]  W. Kahan.  On the cost

of floating-point computation without extra-precise arithmetic. World-Wide Web document, November 20, 2004. URL

http://www.cs.berkeley.edu/
~wkahan/Qdrtcs.pdf.  See

[Bol09] for a proof of this algorithm for accurate computation of the discriminant needed for the solution of quadratic equations.

Kannan:2005:KBN

[Kan05] Rajgopal Kannan. The KR— Benes network: a controloptimal rearrangeable permutation network. IEEE Transactions on Computers, 54(5): 534–544, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407844.

Kapoor:2009:PAV

[Kap09]  H. K. Kapoor.  A pro-

cess algebraic view of latencyinsensitive systems. IEEE Transactions on Computers, 58 (7):931–944, July 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4711040.

Karaata:2006:SSC

[Kar06] M. H. Karaata. Self-stabilizing clustering of tree networks. IEEE Transactions on Computers, 55(4):416–427, April 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1608004.

Katti:2003:LCM

[KB03] R. Katti and J. Brennan. Low complexity multiplication in a finite field using ring representation. IEEE Transactions on Computers, 52(4): 418–427, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190583.

Khabbazian:2008:LBG

[KB08a] M. Khabbazian and V. K. Bhargava. Localized broadcasting with guaranteed delivery and bounded transmission redundancy. IEEE Transactions on Computers, 57(8): 1072–1086, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4479447.

Kounavis:2008:NTL

[KB08b] M. E. Kounavis and F. L.

Berry. Novel table lookupbased algorithms for highperformance CRC generation. IEEE Transactions on Computers, 57(11):1550–1560, November 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4531728.

Kim:2003:PIC

[KBK03] Jin-Ho Kim, Saewoong Bahk, and Hyogon Kim. Performance impact of coarse timer granularities on QoS guarantees in Unix-based systems. IEEE Transactions on Computers, 52 (1):51–58, January 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1159753.

Koppolu:2001:HDI

[KC01] S. Koppolu and A. Chatterjee. Hierarchical diagnosis of identical units in a system. IEEE Transactions on Computers, 50(2): 186–191, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=908994.

Kothari:2007:ASC

[KC07a] Love Kothari and Nicholas P. Carter. Architecture of a self-checkpointing microprocessor that incorporates nanomagnetic devices. IEEE Transactions on Computers, 56(2): 161–173, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4042677.

Kuo:2007:CNC

[KC07b] Yu-Chen Kuo and Teng-Yi Chiu. Comments on “On the Nondomination of Cohorts Coteries”. IEEE Transactions on Computers, 56(7): 1006–1007, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216298. See

[Jia04].

Kim:2004:ASU

[KCHS04] D. Kim, M. Chaudhuri, M. Heinrich, and E. Speight. Architectural support for uniprocessor and multiprocessor active memory systems. IEEE Transactions on Computers, 53(3): 288–307, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261836.

Kuo:2005:RTD

[KCKC05] Tei-Wei Kuo, Yung-Sheng Chao, Chin-Fu Kuo, and Cheng Chang. Real-time dwell scheduling of componentoriented phased array radars. IEEE Transactions on Computers, 54(1):47–60, January 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1362639.

Kanitkar:2002:RTP

[KD02] V. Kanitkar and A. Delis.

Real-time processing in clientserver databases. IEEE Transactions on Computers, 51(3): 269–288, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=990126.

Kim:2005:ESA

[KDB^{+}05] Kanghee Kim, J. L. Diaz, L. L.

Bello, J. M. Lopez, Chang-Gun Lee, and Sang Lyul Min. An exact stochastic analysis of priority-driven periodic realtime systems and its approximations. IEEE Transactions on Computers, 54(11):1460– 1466, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514424.

Kim:2009:AFA

[KDM^{+}09] Jung Sub Kim, Lanping Deng, P. Mangalagiri, K. Irick, K. Sobti, M. Kandemir, V. Narayanan, C. Chakrabarti, N. Pitsianis, and Xiaobai Sun. An automated framework for accelerating numerical algorithms on reconfigurable platforms using algorithmic/architectural optimization. IEEE Transactions on Computers, 58(12):1654–1667, December 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5010433.

Keren:2008:RAP

[Ker08] O. Keren. Reduction of average path length in binary decision diagrams by spectral methods. IEEE Transactions on Computers, 57(4): 520–531, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358255.

Kistler:2001:CPO

[KF01] T. Kistler and M. Franz. Continuous program optimization: Design and evaluation. IEEE Transactions on Computers, 50 (6):549–566, June 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=931893.

Kaneko:2004:CAA

[KF04] H. Kaneko and E. Fujiwara. A class of M-ary asymmetric symbol error correcting codes for data entry devices. IEEE Transactions on Computers, 53(2): 159–167, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261826.

Kang:2006:SHS

[KG06] J.-Y. Kang and J.-L. Gaudiot. A simple high-speed multiplier design. IEEE Transactions on Computers, 55(10): 1253–1258, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683756.

Karaata:2007:SDR

[KG07] M. H. Karaata and M. G.

Gouda. A stabilizing deactivation/reactivation protocol. IEEE Transactions on Computers, 56(7):881–888, July 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216287.

Kavi:2001:SDE

[KGA01] K. M. Kavi, R. Giorgi, and J. Arul. Scheduled dataflow: execution paradigm, architecture, and performance evaluation. IEEE Transactions on Computers, 50(8):834–846, August 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=947003;   http:
                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=947011.

Khabbazian:2005:NMA

[KGB05] M. Khabbazian, T. A. Gulliver, and V. K. Bhargava. A new minimal average weight representation for left-to-right point multiplication methods. IEEE Transactions on Computers, 54(11):1454–1459, November 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1514423.

Khabbazian:2007:DPC

[KGB07] M. Khabbazian, T. A. Gulliver, and V. K. Bhargava. Double point compression with applications to speeding up random point multiplication. IEEE Transactions on Computers, 56 (3):305–313, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4079514.

Kang:2008:EDD

[KGG08] Jung-Yup Kang, S. Gupta, and J.-L. Gaudiot. An efficient data-distribution mechanism in a processor-in-memory (PIM) architecture applied to motion estimation. IEEE Transactions on Computers, 57(3): 375–388, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358277.

Kulkarni:2005:CCD

[KGM^{+}05] C. Kulkarni, C. Ghez, M. Miranda, F. Catthoor, and H. De Man. Cache conscious data layout organization for conflict miss reduction in embedded multimedia applications. IEEE Transactions on Computers, 54 (1):76–81, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362641.

Kin:2000:FMR

[KGMS00] J. Kin, M. Gupta, and W. H.

Mangione-Smith. Filtering memory references to increase energy efficiency. IEEE Transactions on Computers, 49(1): 1–15, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=822560.

Kang:2000:PDS

[KGS00] Dong-In Kang, R. Gerber, and M. Saksena. Parametric design synthesis of distributed embedded systems. IEEE Transactions on Computers, 49(11):1155–1169, November 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895934.

Kuo:2000:SBL

[KH00] Tei-Wei Kuo and Shao-Juen Ho. Similarity-based load adjustment for static real-time transaction systems. IEEE Transactions on Computers, 49(2):112–126, February 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=833108.

Krishnaswamy:2002:APC

[KHB02] V. Krishnaswamy, G. Hasteer, and P. Banerjee. Automatic parallelization of compiled event driven VHDL simulation. IEEE Transactions on Computers, 51(4): 380–394, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=995448.

Kiniwa:2001:LSR

[KHM01] J. Kiniwa, T. Hamada, and D. Mizoguchi. Lookahead scheduling requests for multisize page caching. IEEE Transactions on Computers, 50(9): 972–983, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954511.

Kandasamy:2003:TRI

[KHM03] N. Kandasamy, J. P. Hayes, and B. T. Murray. Transparent recovery from intermittent faults in time-triggered distributed systems. IEEE Transactions on Computers, 52(2): 113–125, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176980.

Kalay:2000:MUT

[KHP00] U. Kalay, D. V. Hall, and M. A.

Perkowski. A minimal universal test set for self-test of EXORsum-of-products circuits. IEEE Transactions on Computers, 49 (3):267–276, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841130. See

correction [Ano00c].

Kangasharju:2002:DLE

[KHRR02] J. Kangasharju, F. Hartanto, M. Reisslein, and K. W. Ross. Distributing layered encoded video through caches. IEEE Transactions on Computers, 51 (6):622–636, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1009148.

Krishnamachari:2004:DBA

[KI04] B. Krishnamachari and S. Iyengar. Distributed Bayesian algorithms for fault-tolerant event region detection in wireless sensor networks. IEEE Transactions on Computers, 53(3): 241–250, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261832. See

comments [CLF05].

Kim:2009:RAO

[Kim09] Soontae Kim. Reducing area overhead for error-protecting large L2/L3 caches. IEEE Transactions on Computers, 58 (3):300–310, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4626951.

Kaufman:2002:CEU

[KJD02] L. M. Kaufman, B. W. Johnson, and J. B. Dugan. Coverage estimation using statistics of the extremes for when testing reveals no failures. IEEE Transactions on Computers, 51 (1):3–12, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980013.

Kim:2009:VPC

[KJM^{+}09] Hyesoon Kim, J. A. Joao, O. Mutlu, Chang Joo Lee, Y. N. Patt, and R. Cohn. Virtual program counter (VPC) prediction: Very low cost indirect branch prediction using conditional branch prediction hardware. IEEE Transactions on Computers, 58(9):1153–1170, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4731245.

Koren:2000:IYE

[KK00a] I. Koren and Z. Koren. Incorporating yield enhancement into the floorplanning process. IEEE Transactions on Computers, 49(6): 532–541, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=862213.

Koren:2000:GEI

[KK00b] Israel Koren and Peter Kornerup. Guest Editors’ introduction: Special issue on computer arithmetic. IEEE Transactions on Computers, 49 (7):625–627, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=863030.

Kim:2009:FPU

[KK09] Donghyun Kim and Lee-Sup Kim. A floating-point unit for 4D vector inner product with reduced latency. IEEE Transactions on Computers, 58 (7):890–901, July 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4674343.

Kuo:2003:TVB

[KKK03] Tei-Wei Kuo, Yuan-Ting Kao, and Chin-Fu Kuo. Two-version based concurrency control and recovery in real-time client/ server databases. IEEE Transactions on Computers, 52(4): 506–524, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1190591.

Kim:2005:PST

[KKKB05] Hyogon Kim, Jin-Ho Kim, Inhye Kang, and Saewoong Bahk. Preventing session table explosion in packet inspection computers. IEEE Transactions on Computers, 54(2): 238–240, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377163.

Kavousianos:2007:OSH

[KKN07] X. Kavousianos, E. Kalligeros, and D. Nikolos. Optimal selective Huffman coding for testdata compression. IEEE Transactions on Computers, 56(8): 1146–1152, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4264327.

Karri:2000:CAD

[KKP00] R. Karri, K. Kim, and M. Potkonjak. Computer aided design of fault-tolerant application specific programmable processors. IEEE Transactions on Computers, 49(11):1272– 1284, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895942.

Krishna:2003:VCS

[KL03a] C. M. Krishna and Y.-H. Lee.

Voltage-clock-scaling adaptive scheduling techniques for low power in hard real-time systems. IEEE Transactions on Computers, 52(12):1586– 1593, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1252854. See

addendum [KL04].

Kwok:2003:SMP

[KL03b] Yu-Kwong Kwok and V. K. N. Lau. System modeling and performance evaluation of rate allocation schemes for packet data services in wideband CDMA systems. IEEE Transactions on Computers, 52(6): 804–814, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204835.

Krishna:2004:AVC

[KL04] C. M. Krishna and Yann-Hang Lee. Addendum to “Voltageclock-scaling adaptive scheduling techniques for low power in hard real-time systems”. IEEE Transactions on Computers, 53(4):497, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268408. See

[KL03a].

Kim:2008:CSO

[KL08] Jong-Seok Kim and Hyeong-Ok Lee. Comments on “A Study of Odd Graphs as Fault-Tolerant Interconnection Networks”. IEEE Transactions on Computers, 57(6): 864, June 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483502.

Kim:2009:SSA

[KL09] Kanghee Kim and Chang-Gun Lee. A safe stochastic analysis with relaxed limitations on the periodic task model. IEEE Transactions on Computers, 58 (5):634–647, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4674341.

Kao:2003:MTC

[KLA^{+}03] Ben Kao, Kam-Yiu Lam, B. Adelberg, R. Cheng, and T. Lee. Maintaining temporal consistency of discrete objects in soft real-time database systems. IEEE Transactions on Computers, 52(3): 373–389, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183951.

Kim:2008:CEL

[KLCV08] Jongsun Kim, Bo-Cheng Lai, M.-C. F. Chang, and I. Verbauwhede. A cost-effective latency-aware memory bus for symmetric multiprocessor systems. IEEE Transactions on Computers, 57(12):1714– 1719, December 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4531735.

Kuo:2001:AOC

[KLS01] Tei-Wei Kuo, Ming-Chung Liang, and LihChyun Shu. Abort-oriented concurrency control for real-time databases. IEEE Transactions on Computers, 50(7):660–673, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=936233.

Kocan:2009:EPD

[KLS09] F. Kocan, Lun Li, and D. G. Saab. Exact path delay fault coverage calculation of partitioned circuits. IEEE Transactions on Computers, 58(6): 858–864, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4663062.

Ke:2007:CWS

[KLT07] Wei-Chieh Ke, Bing-Hong Liu, and Ming-Jer Tsai. Constructing a wireless sensor network to fully cover critical grids by deploying minimum sensors on grid points is NPcomplete. IEEE Transactions on Computers, 56(5): 710–715, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141243.

Kim:2005:HAD

[KLY^{+}05] E. J. Kim, G. M. Link, K. H. Yum, N. Vijaykrishnan, M. Kandemir, M. J. Irwin, and C. R. Das. A holistic approach to designing energy-efficient cluster interconnects. IEEE Transactions on Computers, 54 (6):660–671, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1461355.

Kuo:2000:RTD

[KM00] Te-Wei Kuo and A. K. Mok.

Real-time data semantics and similarity-based concurrency control. IEEE Transactions on Computers, 49(11):1241– 1254, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895940.

Kuo:2003:SPA

[KM03] Tei-Wei Kuo and A. K. Mok.

Schedulability and performance analysis of the similarity stack protocol. IEEE Transactions on Computers, 52(5): 658–669, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1197131.

Kao:2006:OTS

[KM06a] Jung-Chun Kao and R. Marculescu. On optimization of e-textile systems using redundancy and energy-aware routing. IEEE Transactions on Computers, 55(6): 745–756, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628961.

Kornerup:2006:LGD

[KM06b] P. Kornerup and J.-M. Muller. Leading guard digits in finite precision redundant representations. IEEE Transactions on Computers, 55(5): 541–548, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613835.

Kao:2007:MER

[KM07] Jung-Chun Kao and Radu Marculescu. Minimizing eavesdropping risk by transmission power control in multihop wireless networks. IEEE Transactions on Computers, 56(8): 1009–1023, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4264318.

Kornerup:2009:GEI

[KMMS09] Peter Kornerup, Paolo Montuschi, Jean-Michel Muller, and Eric Schwarz. Guest Editors’ introduction: Special section on computer arithmetic. IEEE Transactions on Computers, 58(2):145–147, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4740165.

Kim:2002:UBC

[KMPE02] Suhyun Kim, Soo-Mook Moon, Jinpyo Park, and K. Ebcioglu. Unroll-based copy elimination for enhanced pipeline scheduling. IEEE Transactions on Computers, 51(9):977– 994, September 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1032620.

Kavvadias:2008:EOO

[KN08] N. Kavvadias and S. Nikolaidis. Elimination of overhead operations in complex loop structures for embedded microprocessors. IEEE Transactions on Computers, 57(2): 200–214, February 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358242.

Kalampoukas:2000:HSP

[KNE^{+}00] L. Kalampoukas, D. Nikolos, C. Efstathiou, H. T. Vergos, and J. Kalamatianos. Highspeed parallel-prefix modulo 2^{n} − 1 adders. IEEE Transactions on Computers, 49(7): 673–680, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=863036.

Kumar:2001:NAH

[KNS01] D. R. Kumar, W. A. Najjar, and P. K. Srimani. A new adaptive hardware treebased multicast routing in kary n-cubes. IEEE Transactions on Computers, 50(7): 647–659, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936232.

Kyo:2007:IMA

[KOA07] S. Kyo, S. Okazaki, and T. Arai. An integrated memory array processor for embedded image recognition systems. IEEE Transactions on Computers, 56 (5):622–634, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141237.

Keen:2003:CCI

[KOHC03] D. Keen, M. Oskin, J. Hensley, and F. T. Chong. Cache coherence in intelligent memory systems. IEEE Transactions on Computers, 52(7): 960–966, July 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214343.

Kim:2002:NHA

[KOL02] Chang Han Kim, Sangho Oh, and Jongin Lim. A new hardware architecture for operations in GF(2^{n}). IEEE Transactions on Computers, 51(1): 90–92, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980019.

Kornerup:2005:DSS

[Kor05] Peter Kornerup. Digit selection for SRT division and square root. IEEE Transactions on Computers, 54(3): 294–303, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0294abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0294.
htm;   http://csdl.computer.

                                        org/dl/trans/tc/2005/03/
                                        t0294.pdf;   http://ieeexplore.
                                        ieee.org/iel5/12/30205/01388194.
                                        pdf?isnumber=30205&prod=
                                        JNL&arnumber=1388194&arSt=
                                        +294&ared=+303&arAuthor=
                                        Kornerup%2C+P.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=30205&
                                        arnumber=1388194&count=13&
                                        index=5;   http://ieeexplore.
                                        ieee.org/xpls/references.
                                        jsp?arnumber=1388194.

Kornerup:2009:CNS

[Kor09]  P. Kornerup.  Correcting

the normalization shift of redundant binary representations. IEEE Transactions on Computers, 58(10):1435– 1439, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4798153.

Kranakis:2000:BAD

[KP00] E. Kranakis and A. Pelc. Better adaptive diagnosis of hypercubes. IEEE Transactions on Computers, 49(10):1013– 1020, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888036.

Koc:2003:GEI

[KP03] C. K. Koc and C. Paar. Guest editors’ introduction to the special section on cryptographic hardware and embedded systems. IEEE Transactions on Computers, 52(4): 401–402, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190580.

Karagianni:2001:OSV

[KPDS01] K. Karagianni, V. Paliouras, G. Diamantakos, and T. Stouraitis. Operation-saving VLSI architectures for 3D geometrical transformations. IEEE Transactions on Computers, 50(6): 609–622, June 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=931896.

Kumar:2002:TTB

[KPDS02] V. Kumar, N. Prabhu, M. H.

Dunham, and A. Y. Seydim. TCOT — a timeout-based mobile transaction commitment protocol. IEEE Transactions on Computers, 51(10):1212– 1218, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039846.

Kucuk:2004:CER

[KPEG04] G. Kucuk, D. V. Ponomarev, O. Ergin, and K. Ghose. Complexity-effective reorder buffer designs for superscalar processors. IEEE Transactions on Computers, 53(6): 653–665, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1288541.

Kranitis:2005:SBS

[KPGX05] N. Kranitis, A. Paschalis, D. Gizopoulos, and G. Xenoulis. Software-based selftesting of embedded processors. IEEE Transactions on Computers, 54(4):461–475, April 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1401865.

Kang:2009:PTO

[KPJ^{+}09] Sooyong Kang, Sungmin Park, Hoyoung Jung, Hyoki Shim, and Jaehyuk Cha. Performance trade-offs in using NVRAM write buffer for flash memorybased storage devices. IEEE Transactions on Computers, 58 (6):744–758, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4731242.

Krishnan:2009:ECC

[KPP09] S. C. Krishnan, R. Panigrahy, and S. Parthasarathy. Errorcorrecting codes for ternary content addressable memories. IEEE Transactions on Computers, 58(2):275–279, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633347.

Kim:2004:GKA

[KPT04] Y. Kim, A. Perrig, and G. Tsudik. Group key agreement efficient in communication. IEEE Transactions on Computers, 53(7): 905–921, July 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321050.

Kandemir:2001:DRV

[KR01] M. Kandemir and J. Ramanujam. Data relation vectors: a new abstraction for data optimizations. IEEE Transactions on Computers, 50(8): 798–810, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=947000;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=947008.

Koppanalil:2004:SMD

[KR04] J. J. Koppanalil and E. Rotenberg. A simple mechanism for detecting ineffectual instructions in slipstream processors. IEEE Transactions on Computers, 53(4): 399–413, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1268397.

Kandemir:2001:LCI

[KRCB01] M. Kandemir, J. Ramanujam, A. Choudhary, and P. Banerjee. A layout-conscious iteration space transformation technique. IEEE Transactions on Computers, 50(12):1321– 1335, December 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=970571.

Kim:2005:NID

[KRP05] H. Kim, S. Rixner, and V. S. Pai. Network interface data caching. IEEE Transactions on Computers, 54(11):1394– 1408, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1514419.

Kim:2000:EFT

[KS00a] Hagbae Kim and K. G. Shin. Evaluation of fault tolerance latency from real-time application’s perspectives. IEEE Transactions on Computers, 49 (1):55–64, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=822564.

Krishnaswamy:2000:FCP

[KS00b] U. Krishnaswamy and I. D.

Scherson. A framework for computer performance evaluation using benchmark sets. IEEE Transactions on Computers, 49(12):1325–1338, December 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895853.

Kenney:2005:HSM

[KS05] R. D. Kenney and M. J.

Schulte. High-speed multioperand decimal adders. IEEE Transactions on Computers, 54 (8):953–963, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453497;   http:
//mesa.ece.wisc.edu/publications/
cp_2005-04.pdf.

Kim:2007:ECP

[KS07] Kun Suk Kim and S. Sahni. Efficient construction of pipelined multibit-trie router-tables. IEEE Transactions on Computers, 56 (1):32–43, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016495.

Kharbutli:2008:CBC

[KS08] M. Kharbutli and Yan Solihin. Counter-based cache replacement and bypassing algorithms. IEEE Transactions on Computers, 57(4): 433–447, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358260.

Karpovsky:2003:RSD

[KSA03] M. G. Karpovsky, R. S. Stankovic, and J. T. Astola. Reduction of sizes of decision diagrams by autocorrelation functions. IEEE Transactions on Computers, 52(5): 592–606, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1197126.

Kshemkalyani:2007:TPD

[Ksh07] A. D. Kshemkalyani. Temporal predicate detection using synchronized clocks. IEEE Transactions on Computers, 56(11):1578–1584, November 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4336305.

Kharbutli:2005:ECM

[KSL05] M. Kharbutli, Y. Solihin, and Jaejin Lee. Eliminating conflict misses using prime numberbased cache indexing. IEEE Transactions on Computers, 54 (5):573–586, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407847.

Kodi:2008:ACB

[KSL08] A. K. Kodi, A. Sarathy, and A. Louri. Adaptive channel buffers in on-chip interconnection networks- A power and performance analysis. IEEE Transactions on Computers, 57(9):1169–1181, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4509423.

Kaihara:2005:HAM

[KT05] M. E. Kaihara and N. Takagi. A hardware algorithm for modular multiplication/ division. IEEE Transactions on Computers, 54(1):12–21 54, January 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362636.

Kaihara:2008:BMM

[KT08] M. E. Kaihara and N. Takagi. Bipartite modular multiplication method. IEEE Transactions on Computers, 57(2):

157–164, February 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358245.

Kagaris:2006:ITA

[KTK06] D. Kagaris, S. Tragoudas, and S. Kuriakose. InTeRail: a test architecture for corebased SOCs. IEEE Transactions on Computers, 55(2): 137–149, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1566575.

Kumar:2000:ESP

[Kum00] A. Kumar. An efficient Super-Grid protocol for high availability and load balancing. IEEE Transactions on Computers, 49(10):1126–1133, October 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888048.

Kallahalla:2002:POO

[KV02] M. Kallahalla and P. J. Varman. PC-OPT: optimal offline prefetching and caching for parallel I/O systems. IEEE Transactions on Computers, 51(11):1333– 1344, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047757.

Kumar:2006:ODS

[KWP06] Sandeep Kumar, T. Wollinger, and C. Paar. Optimum digit serial GF(2^{m}) multipliers for curve-based cryptography. IEEE Transactions on Computers, 55(10):1306– 1311, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683761.

Kuo:2002:CRB

[KYL02] Tei-Wei Kuo, Wang-Ru Yang, and Kwei-Jay Lin. A class of rate-based real-time scheduling algorithms. IEEE Transactions on Computers, 51(6): 708–720, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009154.

Kaps:2005:ESU

[KYS05] J.-P. Kaps, K. Yuksel, and B. Sunar. Energy scalable universal hashing. IEEE Transactions on Computers, 54(12): 1484–1495, December 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1524931.

Kang:2003:ISP

[KYY^{+}03] Moonsoo Kang, Chansu Yu, Hee Yong Youn, Ben Lee, and Myungchul Kim. Isomorphic strategy for processor allocation in k-ary ncube systems. IEEE Transactions on Computers, 52(5): 645–657, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1197130.

Kim:2005:CRC

[KZP05] Suhwan Kim, C. H. Ziesler, and M. C. Papaefthymiou. Charge-recovery computing on silicon. IEEE Transactions on Computers, 54(6):651–659, June 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461354.

Lang:2003:RRS

[LA03a] Tomás Lang and Elisardo Antelo. Radix-4 reciprocal square-root and its combination with division and square root. IEEE Transactions on Computers, 52(9):1100–1114, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1228508.

Lee:2003:NSR

[LA03b] Y. J. Lee and I. F. Akyildiz. A new scheme for reducing link and signaling costs in mobile IP. IEEE Transactions on Computers, 52(6): 706–712, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204827.

Lang:2005:HTC

[LA05] Tomás Lang and Elisardo Antelo. High-throughput CORDIC-based geometry operations for 3D computer graphics. IEEE Transactions on Computers, 54(3): 347–361, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0347abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0347.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0347.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388199.
pdf?isnumber=30205&prod=
JNL&arnumber=1388199&arSt=
+347&ared=+361&arAuthor=
Lang%2C+T.%3B+Antelo%2C+
E.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=
1388199&count=13&index=10;
 http://ieeexplore.ieee.
org/xpls/references.jsp?
arnumber=1388199.

Llosa:2001:LSM

[LAG^{+}01] J. Llosa, E. Ayguade, A. Gonzalez, M. Valero, and J. Eckhardt. Lifetime-sensitive modulo scheduling in a production environment. IEEE Transactions on Computers, 50(3): 234–249, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=910814.

Lang:2004:FPM

[LB04] T. Lang and J. D. Bruguera. Floating-point multiply-add-fused with reduced latency. IEEE Transactions on Computers, 53(8):988–1003, August 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306992;   http:
                                        //www.ac.usc.es/arquivos/
                                        articulos/2004/gac2004-j06.
                                        ps.

Lepak:2001:SSS

[LBL01] K. M. Lepak, G. B. Bell, and M. H. Lipasti. Silent stores and store value locality. IEEE Transactions on Computers, 50(11):1174–1190, November 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966493.

Lee:2008:HFF

[LBP08] Chul Lee, Sung Hoon Baek, and Kyu Ho Park. A hybrid flash file system based on NOR and NAND flash memories for embedded devices. IEEE Transactions on Computers, 57(7): 1002–1008, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4433985.

Lee:2002:PDM

[LC02] Chang-Hung Lee and Ming-Syan Chen. Processing distributed mobile queries with interleaved remote mobile joins. IEEE Transactions on Computers, 51(10):1182–1195, October 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039844.

Liu:2004:CDF

[LC04] C. Liu and Krishnendu Chakrabarty.

Compact dictionaries for fault diagnosis in scan-BIST. IEEE Transactions on Computers, 53 (6):775–780, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288552.

Lin:2002:NQB

[LCC02] Ching-Min Lin, Ge-Ming Chiu, and Cheng-Hong Cho. A new quorum-based scheme for managing replicated data in distributed systems. IEEE Transactions on Computers, 51(12):1442–1447, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1146710.

Ling:2006:ODD

[LCC06] Y. Ling, S. Chen, and C.- Y. J. Chiang. On optimal deadlock detection scheduling. IEEE Transactions on Computers, 55(9):1178–1187, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668045.

Lin:2002:IAL

[LCCA02] Yi-Bing Lin, Hsu-Yung Cheng, Ya-Hsing Cheng, and P. Agrawal. Implementing automatic location update for followme database using VoIP and Bluetooth technologies. IEEE Transactions on Computers, 51(10):1154–1168, October 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039842.

Lai:2002:COM

[LCD02] Cheng-Nan Lai, Gen-Huey Chen, and Dyi-Rong Duh.

Constructing one-to-many disjoint paths in folded hypercubes. IEEE Transactions on Computers, 51(1):33– 45, January 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980015.

Lee:2001:LSP

[LCK^{+}01] Donghee Lee, Jongmoo Choi, Jong-Hun Kim, S. H. Noh, Sang Lyul Min, Yookun Cho, and Chong Sang Kim. LRFU: a spectrum of policies that subsumes the least recently used and least frequently used policies. IEEE Transactions on Computers, 50(12):1352– 1361, December 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970573.

Liu:2003:VIS

[LCKR03] J. Liu, F. Chow, T. Kong, and R. Roy. Variable instruction set architecture and its compiler support. IEEE Transactions on Computers, 52 (7):881–895, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214337.

Lin:2003:EDC

[LCL03] Chun-Yuan Lin, Yeh-Ching Chung, and Jen-Shiuh Liu. Efficient data compression methods for multidimensional sparse array operations based on the EKMR scheme. IEEE Transactions on Computers, 52(12): 1640–1646, December 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1252859.

Lam:2007:APN

[LCL07] R. K. Lam, Dah-Ming Chiu, and J. C. S. Lui. On the access pricing and network scaling issues of wireless mesh networks. IEEE Transactions on Computers, 56(11):1456– 1469, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336295.

Lee:2008:HIT

[LCLV08] Dong-U Lee, R. C. C. Cheung, W. Luk, and J. D. Villasenor. Hardware implementation trade-offs of polynomial approximations and interpolations. IEEE Transactions on Computers, 57(5): 686–701, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4384474.

Luo:2006:DFT

[LDH06] X. Luo, M. Dong, and Y. Huang. On distributed faulttolerant detection in wireless sensor networks. IEEE Transactions on Computers, 55(1): 58–70, January 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545751.

Lee:2009:HAH

[Lee09]  Tsern-Huei Lee.  Hard-

ware architecture for highperformance regular expression matching. IEEE Transactions on Computers, 58(7): 984–993, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4599575.

Lee:2012:CPS

[Lee12] Mun-Kyu Lee. Comments on “Provably Sublinear Point Multiplication on Koblitz Curves and Its Hardware Implementation”. IEEE Transactions on Computers, 61(4):591–592, April 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [DJJ^{+}08].

Leveugle:2007:EAF

[Lev07] R. Leveugle. Early analysis of fault-based attack effects in secure circuits. IEEE Transactions on Computers, 56(10):

1431–1434, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302714.

Lin:2009:OHP

[LF09] Hai Lin and Yunsi Fei. Orchestrating horizontal parallelism and vertical instruction packing of programs to improve system overall efficiency. IEEE Transactions on Computers, 58(9):1211–1220, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4799774.

Larrea:2004:IUF

[LFA04] M. Larrea, A. Fernandez, and S. Arevalo. On the implementation of unreliable failure detectors in partially synchronous systems. IEEE Transactions on Computers, 53(7): 815–828, July 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321043.

Faraj:2007:SLM

[lFB07] M. l. Faraj and J. Bigun. Synergy of lip-motion and acoustic features in biometric speech and speaker recognition. IEEE Transactions on Computers, 56(9):1169–1175, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288084.

Liu:2001:AAC

[LG01] Y. A. Liu and G. Gomez. Automatic accurate cost-bound analysis for high-level languages. IEEE Transactions on Computers, 50(12):1295– 1309, December 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970569.

Lee:2006:TBR

[LG06] S.-W. Lee and J.-L. Gaudiot. Throttling-based resource management in high performance multithreaded architectures. IEEE Transactions on Computers, 55(9):1142–1152, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668042.

Liu:2009:PIV

[LG09] Shaoshan Liu and J.-L. Gaudiot. Potential impact of value prediction on communication in many-core architectures. IEEE Transactions on Computers, 58 (6):759–769, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4782951.

Liu:2001:RIE

[LGK01] C. Liu, W.-B. Gong, and C. M. Krishna. Rational interpolation examples in performance analysis. IEEE Transactions on Computers, 50(9):997– 1003, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=954515.

Lee:2005:OHF

[LGML05] Dong-U. Lee, Altaf Abdul Gaffar, Oskar Mencer, and Wayne Luk. Optimizing hardware function evaluation. IEEE Transactions on Computers, 54(12):1520–1531, December 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1524934.

Lee:2008:LCD

[LH08] S. C. Lee and L. R. Hook. Logic and computer design in nanospace. IEEE Transactions on Computers, 57(7): 965–977, July 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358256.

Lou:2009:CPP

[LH09] Xiaosong Lou and Kai Hwang.

Collusive piracy prevention in P2P content delivery networks. IEEE Transactions on Computers, 58(7):970–983, July 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4775892.

Lin:2008:TBR

[LHC^{+}08] Shu-Yen Lin, Chun-Hsiang Huang, Chih-Hao Chao, Keng-Hsien Huang, and An-Yeu Wu. Traffic-balanced routing algorithm for irregular mesh-based on-chip networks. IEEE Transactions on Computers, 57(9):1156–1168, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487055.

Lee:2005:LCB

[LHJL05] Chiou-Yng Lee, Jenn-Shyong Horng, I-Chang Jou, and Erl-Huei Lu. Low-complexity bitparallel systolic Montgomery multipliers for special classes of GF(2^{m}). IEEE Transactions on Computers, 54(9):1061–1070, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471668.

Liu:2006:FRT

[LHLL06] D. Liu, X. S. Hu, M. D. Lemmon, and Q. Ling. Firm real-time system scheduling based on a novel QoS constraint. IEEE Transactions on Computers, 55(3):320–333, March 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1583561.

Li:2004:NOC

[Li04]  R.-C. Li.  Near optimal-

ity of Chebyshev interpolation for elementary function computations. IEEE Transactions on Computers, 53(6): 678–687, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1288543.

Li:2005:DMH

[Li05] J. C.-M. Li. Diagnosis of multiple hold-time and setuptime faults in scan chains. IEEE Transactions on Computers, 54(11):1467–1472, November 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1514425.

Lin:2001:EOL

[Lin01]  Yi-Bing Lin.  Eliminating

overflow for large-scale mobility databases in cellular telephone networks. IEEE Transactions on Computers, 50(4): 356–370, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=919280.

Li:2001:ACE

[LJ01] Tao Li and L. K. John.

ADir_{p}NB: a cost-effective way to implement full map directorybased cache coherence protocols. IEEE Transactions on Computers, 50(9):921–934, September 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954507.

Luo:2004:LBO

[LJ04] Y. Luo and L. K. John.

Locality-based online trace compression. IEEE Transactions on Computers, 53(6): 723–731, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288547.

Lee:2005:IEC

[LJ05] B. K. Lee and L. K. John.

Implications of executing compression and encryption applications on general purpose processors. IEEE Transactions on Computers, 54(7):

917–922, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1432674.

Li:2007:ABP

[LJS^{+}07] Tao Li, L. K. John, A. Sivasubramaniam, N. Vijaykrishnan, and J. Rubio. OSaware branch prediction: Improving microprocessor control flow prediction for operating systems. IEEE Transactions on Computers, 56(1): 2–17, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016493.

LaForge:2003:WDB

[LKF03] L. E. LaForge, K. F. Korver, and M. Sami Fadali. What designers of bus and network architectures should know about hypercubes. IEEE Transactions on Computers, 52(4): 525–544, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1190592.

Liang:2003:QAE

[LKS03] Ming-Chung Liang, Tei-Wei Kuo, and LihChyun Shu. A quantification of aborting effect for real-time data accesses. IEEE Transactions on Computers, 52(5):670–675, May 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197132.

Lu:2005:PIP

[LKS05] Haibin Lu, K. S. Kim, and S. Sahni. Prefix and intervalpartitioned dynamic IP routertables. IEEE Transactions on Computers, 54(5):545–557, May 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407845.

Lee:2006:SER

[LKSS06] Chang-Gung Lee, Phil-Su Kang, Chi-Sheng Shih, and Lui Sha. Schedulability envelope for real-time radar dwell scheduling. IEEE Transactions on Computers, 55(12):1599– 1613, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717391.

Liu:2008:CMP

[LKTT08] Bing-Hong Liu, Wei-Chieh Ke, Chin-Hsien Tsai, and Ming-Jer Tsai. Constructing a message-pruning tree with minimum cost for tracking moving objects in wireless sensor networks is NP-complete and an enhanced data aggregation structure. IEEE Transactions on Computers, 57(6): 849–863, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4445659.

Lan:2008:AFM

[LL08] Zhiling Lan and Yawei Li. Adaptive fault management of parallel applications for high-performance computing. IEEE Transactions on Computers, 57(12):1647–1660, December 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4531733.

Lauter:2009:ERB

[LL09] C. Q. Lauter and V. Lefevre. An efficient rounding boundary test for pow(x, y) in double precision. IEEE Transactions on Computers, 58(2): 197–207, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4663059.

Lipari:2004:TSR

[LLA04] G. Lipari, G. Lamastra, and L. Abeni. Task synchronization in reservation-based real-time systems. IEEE Transactions on Computers, 53(12):1591– 1601, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347085.

Lin:2000:PAD

[LLC00] Yi-Bing Lin, Wei-Ru Lai, and Rong-Jaye Chen. Performance analysis for dual band PCS networks. IEEE Transactions on Computers, 49(2): 148–159, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833111.

Lee:2002:DAW

[LLC02a] Guanling Lee, Shou-Chih Lo, and A. L. P. Chen. Data allocation on wireless broadcast channels for efficient query processing. IEEE Transactions on Computers, 51(10): 1237–1252, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039850.

Lin:2002:ERS

[LLC02b] Chun-Yuan Lin, Jen-Shiuh Liu, and Yeh-Ching Chung. Efficient representation scheme for multidimensional array operations. IEEE Transactions on Computers, 51(3): 327–345, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=990130.

Lo:2003:EMM

[LLC03] Shou-Chih Lo, Guanling Lee, and Wen-Tsuen Chen. An efficient multipolling mechanism for IEEE 802.11 wireless LANs. IEEE Transactions on Computers, 52(6): 764–778, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204832.

Liu:2007:CFC

[LLJA07] Peng Liu, Jie Li, S. Jajodia, and P. Ammann. Can-follow concurrency control. IEEE Transactions on Computers, 56(10): 1425–1430, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302713.

Lee:2001:BPS

[LLL01] Chiou-Yng Lee, Erl-Huei Lu, and Jau-Yien Lee. Bitparallel systolic multipliers for GF(2^{m}) fields defined by allone and equally spaced polynomials. IEEE Transactions on Computers, 50(5): 385–393, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=926154.

Luo:2006:ADF

[LLLD06] Hong Luo, J. Luo, Y. Liu, and S. K. Das. Adaptive data fusion for energy efficient routing in wireless sensor networks. IEEE Transactions on Computers, 55(10): 1286–1299, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683759.

Li:2009:SHQ

[LLP09] Hui Li, Ming Li, and B. Prabhakaran. On supporting highquality 3D geometry multicasting over IEEE 802.11 wireless networks. IEEE Transactions on Computers, 58(4): 558–571, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4657351.

Liu:2004:TLI

[LLPC04] B. Liu, F. Lombardi, N. Park, and M. Choi. Testing layered interconnection networks. IEEE Transactions on Computers, 53(6):710–722, June 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288546.

Lee:2002:SDB

[LLS02a] K. C. K. Lee, Hong Va Leong, and A. Si. Semantic data broadcast for a mobile environment based on dynamic and adaptive chunking. IEEE Transactions on Computers, 51(10): 1253–1268, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1039851.

Lee:2002:OSM

[LLS02b] Yui-Wah Lee, Kwong-Sak Leung, and M. Satyanarayanan. Operation shipping for mobile file systems. IEEE Transactions on Computers, 51(12):1410– 1422, December 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1146707.

Lee:2002:TPP

[LLSC02] V. C. S. Lee, Kwok-Wa Lam, S. H. Son, and E. Y. M. Chan. On transaction processing with partial validation and timestamp ordering in mobile broadcast environments. IEEE Transactions on Computers, 51(10):1196– 1211, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1039845.

Lopez:2001:CCS

[LLVA01] D. Lopez, J. Llosa, M. Valero, and E. Ayguade. Costconscious strategies to increase performance of numerical programs on aggressive VLIW architectures. IEEE Transactions on Computers, 50(10): 1033–1051, October 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956090.

Lee:2004:GNG

[LLVC04] D.-U. Lee, W. Luk, J. D. Villasenor, and P. Y. K. Cheung. A Gaussian noise generator for hardware-based simulations. IEEE Transactions on Computers, 53(12):1523– 1534, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347079.

Li:2007:CPR

[LLW07] James Chien-Mo Li, Hung-Mao Lin, and Fang-Min Wang. Column parity row selection (CPRS) BIST diagnosis technique: Modeling and analysis. IEEE Transactions on Computers, 56(3): 402–414, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4079521.

Lu:2008:EET

[LLWS08] Wan-Chen Lu, Kwei-Jay Lin, Hsin-Wen Wei, and Wei-Kuan Shih. Efficient exact test for rate-monotonic schedulability using large period-dependent initial values. IEEE Transactions on Computers, 57(5): 648–659, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358279.

Luo:2000:API

[LM00] Z. Luo and M. Martonosi. Accelerating pipelined integer and floating-point accumulations in configurable hardware with delayed addition techniques. IEEE Transactions on Computers, 49(3): 208–218, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=841125.

Longa:2008:FFE

[LM08] Patrick Longa and Ali Miri. Fast and flexible elliptic curve point arithmetic over prime fields. IEEE Transactions on Computers, 57(3):289–302, March 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358259.

Lysne:2008:EDF

[LMF^{+}08] O. Lysne, J. M. Montanana, J. Flich, J. Duato, T. M. Pinkston, and T. Skeie. An efficient and deadlock-free network reconfiguration protocol. IEEE Transactions on Computers, 57 (6):762–779, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4459311.

Ling:2001:VCA

[LML01] Yibei Ling, Jie Mi, and Xiaola Lin. A variational calculus approach to optimal checkpoint placement. IEEE Transactions on Computers, 50(7): 699–708, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=936236.

Lo:2006:GEI

[LML06] Jien-Chung Lo, C. Metra, and F. Lombardi. Guest Editors’ introduction: Special section on design and test of systemson-chip (SoC). IEEE Transactions on Computers, 55(2): 97–98, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566571.

Liberato:2000:TMT

[LMM00] F. Liberato, R. Melhem, and D. Mosse. Tolerance to multiple transient faults for aperiodic tasks in hard real-time systems. IEEE Transactions on Computers, 49(9):906–914, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869322.

Lauzac:2003:IRM

[LMM03] S. Lauzac, R. Melhem, and D. Mosse. An improved ratemonotonic admission control and its applications. IEEE Transactions on Computers, 52 (3):337–350, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183948.

Lenzini:2002:USD

[LMS02] L. Lenzini, E. Mingozzi, and G. Stea. A unifying service discipline for providing rate-based guaranteed and fair queuing services based on the Timed Token protocol. IEEE Transactions on Computers, 51(9):1011–1025, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1032622.

Lenzini:2004:DPA

[LMS04] L. Lenzini, E. Mingozzi, and G. Stea. Design and performance analysis of the generalized timed token service discipline. IEEE Transactions on Computers, 53(7): 879–891, July 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321048.

Leroy:2008:CIS

[LMV^{+}08] A. Leroy, D. Milojevic, D. Verkest, F. Robert, and F. Catthoor. Concepts and implementation of spatial division multiplexing for guaranteed throughput in networks-on-chip. IEEE Transactions on Computers, 57(9):1182–1195, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4527238.

Liang:2007:ECD

[LMW07] Han Liang, Piyush Mishra, and Kaijie Wu. Error correction on-demand: a low power register transfer level concurrent error correction technique. IEEE Transactions on Computers, 56(2): 243–252, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042684.

Lang:2007:RDR

[LN07] Tomas Lang and Alberto Nannarelli. A radix-10 digitrecurrence division unit: Algorithm and architecture. IEEE Transactions on Computers, 56 (6):727–739, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4167785.

Lo:2002:ABO

[Lo02] Jien-Chung Lo. Analysis of a BICS-only concurrent error detection method. IEEE Transactions on Computers, 51(3): 241–253, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=990124.

Llorens:2004:SDC

[LO04] M. Llorens and J. Oliver. Structural and dynamic changes in concurrent systems: reconfigurable Petri nets. IEEE Transactions on Computers, 53(9):1147–1158, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1315608.

Lombardi:2007:EN

[Lom07]  Fabrizio Lombardi.  Edi-

tor’s note. IEEE Transactions on Computers, 56(6):

721–726, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167784.

Lombardi:2008:ECN

[Lom08a] Fabrizio Lombardi. Editorin-Chief’s note. IEEE Transactions on Computers, 57(8): 1009–1011, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4548489.

Lombardi:2008:SJ

[Lom08b] Fabrizio Lombardi. State of the Journal. IEEE Transactions on Computers, 57(1):1– 6, January 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4389998.

Lombardi:2009:SJ

[Lom09] Fabrizio Lombardi. State of the Journal. IEEE Transactions on Computers, 58(1):1– 4, January 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4688920.

Llanos:2007:NSS

[LOP07] D. R. Llanos, D. Orden, and B. Palop. New scheduling strategies for randomized incremental algorithms in the context of speculative parallelization. IEEE Transactions on Computers, 56(6): 839–852, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167793.

Low:2000:ERA

[Low00]  Chor Ping Low.  An ef-

ficient reconfiguration algorithm for degradable VLSI/ WSI arrays. IEEE Transactions on Computers, 49(6): 553–559, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=862215.

Li:2000:PAS

[LP00] K. Li and Y. Pan. Probabilistic analysis of scheduling precedence constrained parallel tasks on multicomputers with contiguous processor allocation. IEEE Transactions on Computers, 49(10):1021– 1030, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888038.

Lee:2001:HRM

[LP01a] Jaejin Lee and D. A. Padua. Hiding relaxed memory consistency with a compiler. IEEE Transactions on Computers, 50 (8):824–833, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=947002;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=947010.

Li:2001:PMM

[LP01b] Keqin Li and V. Y. Pan.

Parallel matrix multiplication on a linear array with a reconfigurable pipelined bus system. IEEE Transactions on Computers, 50(5): 519–525, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926164.

Larsson:2006:PAT

[LP06a] E. Larsson and Z. Peng. Poweraware test planning in the early system-on-chip design exploration process. IEEE Transactions on Computers, 55(2): 227–239, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566582.

Lim:2006:ENF

[LP06b] Seung-Ho Lim and Kyu-Ho Park. An efficient NAND flash file system for flash memory storage. IEEE Transactions on Computers, 55(7): 906–912, July 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1637405.

Lim:2007:TTP

[LP07] Sang Seok Lim and Kyu Ho Park. TPF: TCP plugged file system for efficient data delivery over TCP. IEEE Transactions on Computers, 56(4): 459–473, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118670.

Lee:2009:PBM

[LP09] Jupyung Lee and Kyu Ho Park. Prediction-based microscheduler: Toward responsive scheduling of general-purpose operating systems. IEEE Transactions on Computers, 58 (5):648–661, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4663063.

Lee:2004:UDI

[LPAM04] J. Lee, F. Peper, S. Adachi, and K. Morita. Universal delay-insensitive circuits with bidirectional and buffering lines. IEEE Transactions on Computers, 53(8): 1034–1046, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1306995.

Lim:2000:SPA

[LPS00] Hyesook Lim, V. Piuri, and E. E. Swartzlander, Jr. A serial-parallel architecture for two-dimensional discrete cosine and inverse discrete cosine transforms. IEEE Transactions on Computers, 49(12):1297– 1309, December 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895848.

Leadbitter:2007:NM

[LPS07] P. Leadbitter, D. Page, and N. P. Smart. Nondeterministic multithreading. IEEE Transactions on Computers, 56 (7):992–998, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216296.

Li:2004:FBE

[LR04] Peng Li and B. Ravindran.

Fast, best-effort real-time scheduling algorithms. IEEE Transactions on Computers, 53(9):1159–1175, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1315609.

Li:2006:GCS

[LR06] Qun Li and D. Rus. Global clock synchronization in sensor networks. IEEE Transactions on Computers, 55(2): 214–226, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566581.

Lin:2001:DMM

[LRB01] Wei-Fen Lin, S. K. Reinhardt, and D. Burger. Designing a modern memory hierarchy with hardware prefetching. IEEE Transactions on Computers, 50(11):1202–1218, November 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=966495.

Lakshminarayana:2000:BSF

[LRJ00] G. Lakshminarayana, A. Raghunathan, and N. K. Jha. Behavioral synthesis of fault secure controller/datapaths based on aliasing probability analysis. IEEE Transactions on Computers, 49(9):865–885, September 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869319.

Lombardi:2000:GEI

[LS00] F. Lombardi and M. Sami.

Guest editors’ introduction. IEEE Transactions on Computers, 49(6):529–531, June 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=862212.

Low:2002:FFS

[LS02] Chor Ping Low and Xueyan Song. On finding feasible solutions for the delay constrained group multicast routing problem. IEEE Transactions on Computers, 51(5): 581–588, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004596.

Lorch:2004:PNA

[LS04a] J. R. Lorch and A. J. Smith.

PACE: a new approach to dynamic voltage scaling. IEEE Transactions on Computers, 53 (7):856–869, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321046.

Lu:2004:EIT

[LS04b] Haibin Lu and S. Sahni.

Enhanced interval trees for dynamic IP router-tables. IEEE Transactions on Computers, 53(12):1615–1628, December 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347087.

Lu:2004:DRT

[LS04c] Haibin Lu and Sartaj Sahni. O(log n) dynamic router-tables for prefixes and ranges. IEEE Transactions on Computers, 53 (10):1217–1230, October 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1327574.

Lu:2005:BTD

[LS05] H. Lu and Sartaj Sahni. A B-tree dynamic routertable design. IEEE Transactions on Computers, 54(7): 813–824, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432665.

Lu:2008:PCU

[LS08] Wencheng Lu and S. Sahni. Packet classification using space-efficient pipelined multibit tries. IEEE Transactions on Computers, 57(5): 591–605, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4384473.

Lou:2009:FBP

[LS09a] Jianyu Lou and Xiaojun Shen. Frame-based packetmode scheduling for inputqueued switches. IEEE Transactions on Computers, 58(7): 956–969, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4731240.

Lu:2009:EMT

[LS09b] Wencheng Lu and S. Sahni.

Efficient 2D multibit tries for packet classification. IEEE Transactions on Computers, 58(12):1695–1709, December 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5232806.

Lee:2008:ECB

[LSBV08] Yong Ki Lee, K. Sakiyama, L. Batina, and I. Verbauwhede. Elliptic-curve-based security processor for RFID. IEEE Transactions on Computers, 57(11):1514–1527, November 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4604657.

Lee:2007:PPA

[LSLsK07] M. Lee, Euiseong Seo, Joonwon Lee, and Jin soo Kim. PABC: Power-aware buffer cache management for low power consumption. IEEE Transactions on Computers, 56(4): 488–501, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118672.

Lee:2004:EUB

[LSP04] C.-G. Lee, L. Sha, and Avinash Peddi. Enhanced utilization bounds for QoS management. IEEE Transactions on Computers, 53(2): 187–200, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261828.

Li:2009:LDS

[LSS09] Xu Li, N. Santoro, and I. Stojmenovic. Localized distancesensitive service discovery in wireless sensor and actor networks. IEEE Transactions on Computers, 58(9):1275–1288, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815211.

Lee:2000:FAP

[LSV00] Lin-Wen Lee, P. Scheuermann, and R. Vingralek. File assignment in parallel I/O systems with minimal variance of service time. IEEE Transactions on Computers, 49(2): 127–140, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833109.

Lin:2003:GES

[LT03] Y.-B. Lin and Y.-C. Tseng.

Guest editorial: special section on wireless Internet. IEEE Transactions on Computers, 52(6):689, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204825.

Lai:2005:CDM

[LTCH05] Pao-Lien Lai, J. J. M. Tan, Chien-Ping Chang, and Lih-Hsing Hsu. Conditional diagnosability measures for large multiprocessor systems. IEEE Transactions on Computers, 54(2):165–175, February 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377155.

Lai:2004:DMC

[LTTH04] Pao-Lien Lai, J. J. M. Tan, Chang-Hsiung Tsai, and Lih-Hsing Hsu. The diagnosability of the matching composition network under the comparison diagnosis model. IEEE Transactions on Computers, 53(8): 1064–1069, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306998.

Lu:2005:NHO

[Lu05] H. Lu. A novel high-order tree for secure multicast key management. IEEE Transactions on Computers, 54(2): 214–224, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1377159.

Lee:2007:BWO

[LV07] D.-U. Lee and J. D. Villasenor. A bit-width optimization methodology for polynomial-based function evaluation. IEEE Transactions on Computers, 56(4):567–571, April 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118679.

Lee:2009:OCP

[LV09] Dong-U Lee and J. D. Villasenor. Optimized custom precision function evaluation for embedded processors. IEEE Transactions on Computers, 58 (1):46–59, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4585364.

Lee:2006:HGN

[LVLL06] D.-U. Lee, J. D. Villasenor, W. Luk, and P. H. W. Leong. A hardware Gaussian noise generator using the Box–Muller method and its error analysis. IEEE Transactions on Computers, 55(6): 659–671, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1628955.

Lopes:2001:FGM

[LVS01] L. Lopes, V. T. Vasconcelos, and F. Silva. Finegrained multithreading with process calculi. IEEE Transactions on Computers, 50(8): 852–862, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=947014.

Lin:2000:LPC

[LW00] Kun-Jin Lin and Cheng-Wen Wu. A low-power CAM design for LZ data compression. IEEE Transactions on Computers, 49(10):1139– 1145, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888055.

Li:2003:CWA

[LWF03] Xiang-Yang Li, Peng-Jun Wan, and O. Frieder. Coverage in wireless ad hoc sensor networks. IEEE Transactions on Computers, 52(6): 753–763, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204831.

Liu:2001:OFT

[LWG01] Yu-Liang Liu, Yue-Li Wang, and D. J. Guan. An optimal fault-tolerant routing algorithm for double-loop networks. IEEE Transactions on Computers, 50 (5):500–505, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=926162.

Lee:2003:ICS

[LwJKW03] Jung-Hoon Lee, Seh woong Jeong, Shin-Dug Kim, and C. C. Weems. An intelligent cache system with hardware prefetching for high performance. IEEE Transactions on Computers, 52(5): 607–616, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197127.

Lee:2002:CCM

[LwLH02] V. C. S. Lee, Kwok wa Lam, and Sheung-Lun Hung. Concurrency control for mixed transactions in real-time databases. IEEE Transactions on Computers, 51(7):821–834, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1017702.

Li:2006:UAS

[LWRJ06] P. Li, H. Wu, Binoy Ravindran, and E. D. Jensen. A utility accrual scheduling algorithm for real-time activities with mutual exclusion resource constraints. IEEE Transactions on Computers, 55(4): 454–469, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608007.

Lloyd:2007:RNP

[LX07] E. L. Lloyd and Guoliang Xue.

Relay node placement in wireless sensor networks. IEEE Transactions on Computers, 56 (1):134–138, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016503.

Lee:2009:ESR

[LX09] Rubao Lee and Zhiwei Xu. Exploiting stream request locality to improve query throughput of a data integration system. IEEE Transactions on Computers, 58(10):1356– 1368, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=5089314.

Lee:2001:TBU

[LY01] Sang-Jeong Lee and Pen-Chung Yew. On table bandwidth and its update delay for value prediction on wide-issue ILP processors. IEEE Transactions on Computers, 50(8): 847–852, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=947004;   http:
                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=947012.

Lee:2002:CPL

[LY02a] D. Lee and M. Yannakakis. Closed partition lattice and machine decomposition. IEEE Transactions on Computers, 51(2):216–228, February 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=980009.

Lee:2002:ATC

[LY02b] Sang-Jeong Lee and Pen-Chung Yew. On augmenting trace cache for high-bandwidth value prediction. IEEE Transactions on Computers, 51(9):1074– 1088, September 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1032626.

Leung:2006:MSC

[LZ06] J. Y.-T. Leung and H. Zhao.

Minimizing sum of completion times and makespan in masterslave systems. IEEE Transactions on Computers, 55(8): 985–999, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650196.

Lee:2007:PSB

[LZ07] Y. C. Lee and A. Y. Zomaya.

Practical scheduling of bagof-tasks applications on grids with dynamic resilience. IEEE Transactions on Computers, 56 (6):815–825, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4167791.

Martinez:2008:EDB

[MAA^{+}08] A. Martinez, G. Apostolopoulos, F. J. Alfaro, J. L. Sanchez, and J. Duato. Efficient deadline-based QoS algorithms for high-performance networks. IEEE Transactions on Computers, 57(7): 928–939, July 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4459315.

Mangard:2003:HRS

[MAD03] S. Mangard, M. Aigner, and S. Dominikus. A highly regular and scalable AES hardware architecture. IEEE Transactions on Computers, 52(4): 483–491, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1190589.

Mejia-Alvarez:2003:ISS

[MAMMA03] P. Mejia-Alvarez, R. Melhem, D. Mosse, and H. Aydin. An incremental server for scheduling overloaded real-time systems. IEEE Transactions on Computers, 52(10):1347– 1361, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1234531.

Marculescu:2008:ISS

[Mar08] Radu Marculescu. Introduction to the special section on networks-on-chip. IEEE Transactions on Computers, 57(9): 1153–1155, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4579703.

Mura:2001:MRS

[MB01] I. Mura and A. Bondavalli.

Markov regenerative stochastic Petri nets to model and evaluate phased mission systems dependability. IEEE Transactions on Computers, 50(12):1337– 1351, December 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970572.

Mishra:2003:ABE

[MB03] A. Mishra and P. Banerjee.

An algorithm-based error detection scheme for the multigrid method. IEEE Transactions on Computers, 52(9):1089–1099, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228507.

Montuschi:2007:DDA

[MBCP07] P. Montuschi, J. D. Bruguera, L. Ciminiera, and J.-A. Pieiro. A digit-by-digit algorithm for m th root extraction. IEEE Transactions on Computers, 56(12):1696–1706, December 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358221.

Messer:2004:SCS

[MBF^{+}04] A. Messer, P. Bernadat, G. Fu, D. Chen, Z. Dimitrijevic, D. Lie, D. D. Mannaru, A. Riska, and D. Milojicic. Susceptibility of commodity systems and software to memory soft errors. IEEE Transactions on Computers, 53(12):1557– 1568, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347082.

Marojevic:2008:CRM

[MBG08] V. Marojevic, X. R. Balleste, and A. Gelonch. A computing resource management framework for software-defined radios. IEEE Transactions on Computers, 57(10):1399– 1412, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4528958.

Marti:2009:CEO

[MBR^{+}09] S. P. Marti, J. S. Borras, P. L. Rodriguez, R. U. Tena, and J. D. Marin. A complexity-effective out-of-order retirement microarchitecture. IEEE Transactions on Computers, 58(12):1626– 1639, December 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5161254.

Martinez:2008:MTN

[MBS^{+}08] C. Martinez, R. Beivide, E. Stafford, M. Moreto, and E. M. Gabidulin. Modeling toroidal networks with the Gaussian integers. IEEE Transactions on Computers, 57(8): 1046–1056, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483506.

Muscedere:2005:ETB

[MDJM05] Roberto Muscedere, Vassil Dimitrov, Graham A. Jullien, and William C. Miller. Efficient techniques for binaryto-multidigit multidimensional logarithmic number system conversion using range-addressable look-up tables. IEEE Transactions on Computers, 54(3): 257–271, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.

computer.org/comp/trans/
tc/2005/03/t0257abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0257.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0257.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388191.
pdf?isnumber=30205&prod=
JNL&arnumber=1388191&arSt=
+257&ared=+271&arAuthor=
Muscedere%2C+R.%3B+Dimitrov%
2C+V.%3B+Jullien%2C+G.A.%
3B+Miller%2C+W.C.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30205&
arnumber=1388191&count=13&
index=2;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1388191.

Metra:2004:ICD

[MDM04] C. Metra, S. Di Francescantonio, and T. M. Mak. Implications of clock distribution faults and issues with screening them during manufacturing testing. IEEE Transactions on Computers, 53(5): 531–546, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275295.

Messerges:2002:ESC

[MDS02] T. S. Messerges, E. A. Dabbish, and R. H. Sloan. Examining smart-card security under the threat of power analysis attacks. IEEE Transactions on Computers, 51(5):

541–552, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004593.

Meleis:2001:SSB

[MEB01] W. M. Meleis, A. E. Eichenberger, and I. D. Baev. Scheduling superblocks with boundbased branch trade-offs. IEEE Transactions on Computers, 50 (8):784–797, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=946999;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=947007.

Melhem:2007:LDI

[Mel07]  R. Melhem.  Low diame-

ter interconnections for routing in high-performance parallel systems. IEEE Transactions on Computers, 56(4): 502–510, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118673. See

comments [XWC^{+}08].

Meyer:2001:PAC

[Mey01] J. F. Meyer. Performability of an algorithm for connection admission control. IEEE Transactions on Computers, 50 (7):724–733, July 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936238.

Mishra:2002:TGC

[MFC02] S. Mishra, C. Fetzer, and F. Cristian. The timewheel group communication system. IEEE Transactions on Computers, 51(8):883–899, August 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1024737.

Metra:2000:SCD

[MFR00] C. Metra, M. Favalli, and B. Ricco. Self-checking detection and diagnosis of transient, delay, and crosstalk faults affecting bus lines. IEEE Transactions on Computers, 49(6): 560–574, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=862216.

Monaco:2002:RDC

[MG02] F. J. Monaco and A. Gonzaga. Remote device command and resource sharing over the internet: a new approach based on a distributed layered architecture. IEEE Transactions on Computers, 51 (7):787–792, July 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017698.

Muresan:2008:PCA

[MG08] R. Muresan and S. Gregori. Protection circuit against differential power analysis attacks for smart cards. IEEE Transactions on Computers, 57(11):1540–1549, November 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4585359.

Marcuello:2004:TPV

[MGT04] P. Marcuello, A. Gonzalez, and J. Tubella. Thread partitioning and value prediction for exploiting speculative threadlevel parallelism. IEEE Transactions on Computers, 53(2): 114–125, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261823.

Ma:2006:UIF

[MGZ06] Yi Ma, Hongliang Gao, and Huiyang Zhou. Using indexing functions to reduce conflict aliasing in branch prediction tables. IEEE Transactions on Computers, 55(8): 1057–1061, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650202.

Min:2001:IPL

[MH01] Rui Min and Yiming Hu. Improving performance of large physically indexed caches by decoupling memory addresses from cache addresses. IEEE Transactions on Computers, 50(11):1191–1201, November 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966494.

Mhamdi:2009:PPB

[Mha09] L. Mhamdi. PBC: a partially buffered crossbar packet switch. IEEE Transactions on Computers, 58(11):1568– 1581, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4815222.

Mishra:2006:PCS

[Mis06]  P. M. Mishra.  Pipelined

computation of scalar multiplication in elliptic curve cryptosystems (extended version). IEEE Transactions on Computers, 55(8):1000– 1010, August 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650197.

Maziarz:2002:ARY

[MJ02] B. M. Maziarz and V. K. Jain.

Automatic reconfiguration and yield of the TESH multicomputer network. IEEE Transactions on Computers, 51(8): 963–972, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024742.

Mitra:2006:XET

[MK06] S. Mitra and K. S. Kim.

XPAND: an efficient test stimulus compression technique. IEEE Transactions on Computers, 55(2):163–173, February 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566577.

Murphy:2007:MAP

[MK07] R. C. Murphy and P. M. Kogge.

On the memory access patterns of supercomputer applications: Benchmark selection and its implications. IEEE Transactions on Computers, 56 (7):937–945, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216292.

Michiels:2003:GTM

[MKA03] W. Michiels, J. Korst, and J. Aerts. On the guaranteed throughput of multizone disks. IEEE Transactions on Computers, 52(11):1407– 1420, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244939.

Mutlu:2005:API

[MKAP05] O. Mutlu, H. Kim, D. N. Armstrong, and Y. N. Patt. An analysis of the performance impact of wrong-path memory references on out-of-order and runahead execution processors. IEEE Transactions on Computers, 54(12):1556– 1571, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524937.

Morin:2000:ESA

[MKBG00] C. Morin, A.-M. Kermarrec, M. Banatre, and A. Gefflaut. An efficient and scalable approach for implementing fault-tolerant DSM architectures. IEEE Transactions on Computers, 49(5): 414–430, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=859537.

Mutlu:2006:AVD

[MKP06] O. Mutlu, Hyesoon Kim, and Y. N. Patt. Address-value delta (AVD) prediction: a hardware technique for efficiently parallelizing dependent cache misses. IEEE Transactions on Computers, 55(12):1491– 1508, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717383.

Memik:2003:AFI

[MKS03] S. O. Memik, A. K. Katsaggelos, and M. Sarrafzadeh. Analysis and FPGA implementation of image restoration under resource constraints. IEEE Transactions on Computers, 52 (3):390–399, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183952.

Mowry:2000:UWC

[ML00] T. C. Mowry and C.-K. Luk.

Understanding why correlation profiling improves the predictability of data cache misses in nonnumeric applications. IEEE Transactions on Computers, 49(4): 369–384, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844349.

Montuschi:2001:BVH

[ML01] P. Montuschi and T. Lang. Boosting very-high radix division with prescaling and selection by rounding. IEEE Transactions on Computers, 50(1): 13–27, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=902750.

Maistri:2008:DDR

[ML08] P. Maistri and R. Leveugle. Double-data-rate computation as a countermeasure against fault analysis. IEEE Transactions on Computers, 57(11):1528–1539, November 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4604658.

Marti:2009:DER

[MLB^{+}09] P. Marti, Caixue Lin, S. A. Brandt, M. Velasco, and J. M. Fuertes. Draco: Efficient resource management for resource-constrained control tasks. IEEE Transactions on Computers, 58(1): 90–105, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599571.

Meribout:2004:CAH

[MM04] M. Meribout and M. Motomura. A-combined approach to high-level synthesis for dynamically reconfigurable systems. IEEE Transactions on Computers, 53(12):1508– 1522, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347078.

Melhem:2004:IPM

[MME04] R. Melhem, D. Mosse, and E. Elnozahy. The interplay of power management and fault recovery in realtime systems. IEEE Transactions on Computers, 53(2): 217–231, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261830.

Masuda:2008:LCN

[MMPT08] A. M. Masuda, L. Moura, D. Panario, and D. Thomson. Low complexity normal elements over finite fields of characteristic two. IEEE Transactions on Computers, 57(7): 990–1001, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378359.

Mrugalski:2006:HPD

[MMRT06] G. Mrugalski, M. Mukherjee, J. Rajski, and J. Tyszer. High performance dense ring generators. IEEE Transactions on Computers, 55(1):83– 87, January 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1545753.

Merayo:2008:EES

[MNR08] M. G. Merayo, M. Nunez, and I. Rodriguez. Extending EF-SMs to specify and test timed systems with action durations and time-outs. IEEE Transactions on Computers, 57(6): 835–848, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4433986.

Misra:2006:EDA

[MO06] S. Misra and B. J. Oommen. An efficient dynamic algorithm for maintaining allpairs shortest paths in stochastic networks. IEEE Transactions on Computers, 55(6): 686–702, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628957.

Min:2004:PMW

[MOK04] G. Min and M. Ould-Khaoua.

A performance model for wormhole-switched interconnection networks under selfsimilar traffic. IEEE Transactions on Computers, 53(5): 601–613, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275299.

Montgomery:2005:FSS

[Mon05] Peter L. Montgomery. Five, six, and seven-term Karatsubalike formulae. IEEE Transactions on Computers, 54(3): 362–369, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0362abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0362.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0362.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388200.
pdf?isnumber=30205&prod=
JNL&arnumber=1388200&arSt=
+362&ared=+369&arAuthor=
Montgomery%2C+P.L.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30205&
arnumber=1388200&count=13&
index=11;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1388200. See

comments [FH07a].

Moore:2007:PHI

[Moo07] R. K. Moore. PRESENCE: a human-inspired architecture for speech-based human-machine interaction. IEEE Transactions on Computers, 56(9):1176– 1188, September 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4288085.

Meyer:2003:PDT

[MP03] F. J. Meyer and N. Park. Predicting defect-tolerant yield in the embedded core context. IEEE Transactions on Computers, 52(11):1470–1479, November 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1244944.

Michaelides:2009:SFT

[MP09] M. P. Michaelides and C. G. Panayiotou. SNAP: Fault tolerant event location estimation in sensor networks using binary data. IEEE Transactions on Computers, 58(9):1185–1197, September 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4815217.

McKinley:2003:CPS

[MPAS03] P. K. McKinley, U. I. Padmanabhan, N. Ancha, and S. M.

Sadjadi. Composable proxy services to support collaboration on the mobile Internet. IEEE Transactions on Computers, 52 (6):713–726, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204828.

Meyer:2005:PPS

[MPP^{+}05] B. H. Meyer, J. J. Pieper, J. M. Paul, J. E. Nelson, S. M. Pieper, and A. G. Rowe. Power-performance simulation and design strategies for single-chip heterogeneous multiprocessors. IEEE Transactions on Computers, 54(6): 684–697, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1424448.

Mahevas:2001:BCD

[MR01] S. Mahevas and G. Rubino.

Bound computation of dependability and performance measures. IEEE Transactions on Computers, 50(5): 399–413, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926156.

Mahalingam:2006:IAM

[MR06a] V. Mahalingam and N. Ranganathan. Improving accuracy in Mitchell’s logarithmic multiplication using operand decomposition. IEEE Transactions on Computers, 55(12):1523– 1535, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1717385.

Mei:2006:OPR

[MR06b] Alessandro Mei and R. Rizzi. Online permutation routing in partitioned optical passive star networks. IEEE Transactions on Computers, 55(12):1557– 1571, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1717388.

Monnet:2006:DRC

[MRL06] Y. Monnet, M. Renaudin, and R. Leveugle. Designing resistant circuits against malicious faults injection using asynchronous logic. IEEE Transactions on Computers, 55(9):1104–1115, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1668038.

Metra:2007:WCC

[MRM07] C. Metra, D. Rossi, and T. M. Mak. Won’t on-chip clock calibration guarantee performance boost and product quality? IEEE Transactions on Computers, 56(3): 415–428, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4079522.

Mirkovic:2008:TCD

[MRP^{+}08] J. Mirkovic, P. Reiher, C. Papadopoulos, A. Hussain, M. Shepard, M. Berg, and R. Jung. Testing a collaborative DDoS defense in a Red Team/Blue Team exercise. IEEE Transactions on Computers, 57(8): 1098–1112, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4479443.

Moon:2000:SHP

[MRS00] Sung-Whan Moon, J. Rexford, and K. G. Shin. Scalable hardware priority queue architectures for high-speed packet switches. IEEE Transactions on Computers, 49(11):1215– 1227, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895938.

Mukherjee:2001:TSF

[MRT01] N. Mukherjee, J. Rajski, and J. Tyszer. Testing schemes for FIR filter structures. IEEE Transactions on Computers, 50 (7):674–688, July 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936234.

Malewicz:2006:TTS

[MRY06] G. Malewicz, A. L. Rosenberg, and M. Yurkewych. Toward a theory for scheduling dags in Internet-based computing. IEEE Transactions on Computers, 55(6): 757–768, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628962.

Moshovos:2002:RML

[MS02] A. Moshovos and G. S. Sohi. Reducing memory latency via read-after-read memory dependence prediction. IEEE Transactions on Computers, 51(3): 313–326, March 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=990129.

Martin:2003:NBP

[MS03] T. L. Martin and D. P. Siewiorek. Nonideal battery properties and their impact on software design for wearable computers. IEEE Transactions on Computers, 52(8): 979–984, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223632.

Mei:2009:ROS

[MS09] A. Mei and J. Stefa. Routing in outer space: Fair traffic load in multihop wireless networks. IEEE Transactions on Computers, 58(6): 839–850, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4752812.

Mitra:2002:DDM

[MSM02] S. Mitra, N. R. Saxena, and E. J. McCluskey. A design diversity metric and analysis of redundant systems. IEEE Transactions on Computers, 51 (5):498–510, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004589.

Mitra:2004:EDD

[MSM04] S. Mitra, N. R. Saxena, and E. J. McCluskey. Efficient design diversity estimation for combinational circuits. IEEE Transactions on Computers, 53(11):1483– 1492, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1336768.

Macchiarulo:2004:PSC

[MSMS04] L. Macchiarulo, Shih-Min Shu, and M. Marek-Sadowska. Pipelining sequential circuits with wave steering. IEEE Transactions on Computers, 53(9):1205–1210, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1315613.

Maxion:2002:ADE

[MT02] R. A. Maxion and K. M. C. Tan. Anomaly detection in embedded systems. IEEE Transactions on Computers, 51(2): 108–120, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980003.

Mascella:2006:EAB

[MT06] R. Mascella and L. G. Tallini. Efficient m-ary balanced codes which are invariant under symbol permutation. IEEE Transactions on Computers, 55(8): 929–946, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650192.

Mamatas:2009:DSN

[MT09] L. Mamatas and V. Tsaoussidis. Differentiating services with noncongestive queuing (NCQ). IEEE Transactions on Computers, 58(5): 591–604, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4663054.

Merten:2001:AFR

[MTB^{+}01] M. C. Merten, A. R. Trick, R. D. Barnes, E. M. Nystrom, C. N. George, J. C. Gyllenhaal, and W.-M. W. Hwu. An architectural framework for runtime optimization. IEEE Transactions on Computers, 50 (6):567–589, June 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=931894.

Matakias:2008:CMP

[MTHA08] S. Matakias, Y. Tsiatouhas, T. Haniotakis, and A. Arapoyanni. A current mode, parallel, two-rail code checker. IEEE Transactions on Computers, 57(8):1032–1045, August 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483508.

Mudge:2005:ISS

[Mud05] T. Mudge. Introduction to the special section on energy efficient computing. IEEE Transactions on Computers, 54(6):641, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461352.

Monreal:2004:LAE

[MVG^{+}04] T. Monreal, V. Vinals, J. Gonzalez, A. Gonzalez, and M. Valero. Late allocation and early release of physical registers. IEEE Transactions on Computers, 53(10): 1244–1259, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327576.

McKee:2000:DAO

[MWA^{+}00] S. A. McKee, W. A. Wulf, J. H. Aylor, R. H. Klenke, M. H. Salinas, S. I. Hong, and D. A. B. Weikle. Dynamic access ordering for streamed computations. IEEE Transactions on Computers, 49(11):1255– 1271, November 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895941.

Hwu:2005:GEI

[mWHP05] Wen mei W. Hwu and K. V.

Palem. Guest Editors’ introduction. IEEE Transactions on Computers, 54(10): 1185–1187, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501785.

Mutyam:2009:PVA

[MWK^{+}09] M. Mutyam, Feng Wang, R. Krishnan, V. Narayanan, M. Kandemir, Yuan Xie, and M. J. Irwin. Process-variation-aware adaptive cache architecture and management. IEEE Transactions on Computers, 58(7): 865–877, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4782953.

Ma:2007:ATD

[MY07] Ming Ma and Yuanyuan Yang.

Adaptive triangular deployment algorithm for unattended mobile sensor networks. IEEE Transactions on Computers, 56 (7):946–958, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216293.

Moh:2001:FAT

[MYL^{+}01] Sangman Moh, Chansu Yu, B. Lee, Hee Yong Youn, Dongsoo Han, and Dongman Lee. Four-ary tree-based barrier synchronization for 2D meshes without nonmember involvement. IEEE Transactions on Computers, 50(8): 811–823, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=947001;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=947009.

Nicolici:2002:MSC

[NAH02] N. Nicolici and B. M. Al-Hashimi. Multiple scan chains for power minimization during test application in sequential circuits. IEEE Transactions on Computers, 51(6): 721–734, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009155.

Nam:2008:RTU

[NASK^{+}08] Min-Young Nam, Z. Al-Sabbagh, Jung-Eun Kim, Man-Ki Yoon, Chang-Gun Lee, and Eun Yong Ha. A real-time ubiquitous system for assisted living: Combined scheduling of sensing and communication for real-time tracking. IEEE Transactions on Computers, 57 (6):795–808, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4444352.

Nam:2004:CST

[NASR04] G.-J. Nam, F. Aloul, K. A.

Sakallah, and R. A. Rutenbar. A comparative study of two Boolean formulations of FPGA detailed routing constraints. IEEE Transactions on Computers, 53(6): 688–696, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288544.

Ni:2002:FSI

[NB02] Nan Ni and L. N. Bhuyan.

Fair scheduling in Internet routers. IEEE Transactions on Computers, 51(6): 686–701, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009152.

Ndai:2008:WVA

[NBAR08] P. Ndai, S. Bhunia, A. Agarwal, and K. Roy. Withindie variation-aware scheduling in superscalar processors for improved throughput. IEEE Transactions on Computers, 57 (7):940–951, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4468699.

Ng:2001:DVR

[NC01] W. T. Ng and P. M. Chen. The design and verification of the rio file cache. IEEE Transactions on Computers, 50(4): 322–337, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=919278.

Naidu:2003:SET

[NC03] S. R. Naidu and Vijay Chandru. On synthesis of easily testable (k, K) circuits. IEEE Transactions on Computers, 52(11):1490–1494, November 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1244946.

Nguyen:2005:EDC

[Ngu05] G. D. Nguyen. Error-detection codes: algorithms and fast implementation. IEEE Transactions on Computers, 54(1): 1–11, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1362635.

Nguyen:2009:FC

[Ngu09] G. D. Nguyen. Fast CRCs. IEEE Transactions on Computers, 58(10):1321–1331, October 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5089317.

Niyato:2006:QTO

[NH06] D. Niyato and E. Hossain. A queuing-theoretic and optimizationbased model for radio resource management in IEEE 802.16 broadband wireless networks. IEEE Transactions on Computers, 55(11):1473– 1488, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705455.

Ng:2007:DBK

[NHSC07] W. H. D. Ng, M. Howarth, Z. Sun, and H. Cruickshank. Dynamic balanced key tree management for secure multicast communications. IEEE Transactions on Computers, 56 (5):590–605, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141235.

Nikolos:2009:ISM

[NKSG09] D. Nikolos, D. Kagaris, S. Sudireddy, and S. Gidaros. An improved search method for accumulatorbased test set embedding. IEEE Transactions on Computers, 58 (1):132–138, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4633350.

Nam:2008:PAE

[NKY08] Byeong-Gyu Nam, Hyejung Kim, and Hoi-Jun Yoo. Power and area-efficient unified computation of vector and elementary functions for handheld 3D graphics systems. IEEE Transactions on Computers, 57(4): 490–504, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4432232.

Nielsen:2000:ICF

[NMLE00] Asger Munk Nielsen, David W. Matula, Chung Nan Lyu, and Guy Even. An IEEE compliant floating-point adder that conforms with the pipeline packetforwarding paradigm. IEEE Transactions on Computers, 49 (1):33–47, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=822562.

Neto:2007:NNS

[NPB07] Omar Paranaiba Vilela Neto, Marco Aurelio C. Pacheco, and Carlos R. Hall Barbosa. Neural network simulation and evolutionary synthesis of QCA circuits. IEEE Transactions on Computers, 56(2):191–201, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042679.

Niggemeyer:2004:AGD

[NR04] D. Niggemeyer and E. M. Rudnick. Automatic generation of diagnostic memory tests based on fault decomposition and output tracing. IEEE Transactions on Computers, 53(9):1134– 1146, September 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315607.

Nett:2003:RRT

[NS03] E. Nett and S. Schemmer.

Reliable real-time communication in cooperative mobile applications. IEEE Transactions on Computers, 52(2): 166–180, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176984.

Nagayama:2009:CGB

[NS09] S. Nagayama and T. Sasao.

Complexities of graph-based representations for elementary functions. IEEE Transactions on Computers, 58(1): 106–119, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599569.

Nourani:2008:LTT

[NTA08] M. Nourani, M. Tehranipoor, and N. Ahmed. Low-transition test pattern generation for BIST-based applications. IEEE Transactions on Computers, 57 (3):303–315, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358246.

Namin:2007:CAF

[NWA07] A. H. Namin, Huapeng Wu, and M. Ahmadi. Comb architectures for finite field multiplication in F (2^{m}). IEEE Transactions on Computers, 56 (7):909–916, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4216289.

Namin:2008:NFF

[NWA08] A. H. Namin, Huapeng Wu, and M. Ahmadi. A new finitefield multiplier using redundant representation. IEEE Transactions on Computers, 57 (5):716–720, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4378354.

Nunez-Yanez:2005:CSL

[NYC05] J. L. Nunez-Yanez and V. A.

Chouliaras. A configurable statistical lossless compression core based on variable order Markov modeling and arithmetic coding. IEEE Transactions on Computers, 54(11):1345–1359, November 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514415.

Nong:2007:EAI

[NZ07] Ge Nong and Sen Zhang. Efficient algorithms for the inverse sort transform. IEEE Transactions on Computers, 56(11):1564–1574, November 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4336303.

Oppenheimer:2002:RHS

[OBB^{+}02] D. Oppenheimer, A. Brown, J. Beck, D. Hettena, J. Kuroda, N. Treuhaft, D. A. Patterson, and K. Yelick. ROC-1: hardware support for recovery-oriented computing. IEEE Transactions on Computers, 51(2):100–107, February 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980002.

Oh:2000:ENB

[OKLC00] Sangho Oh, Chang Han Kim, Jongin Lim, and Dong Hyeon Cheon. Efficient normal basis multipliers in composite fields. IEEE Transactions on Computers, 49(10):1133– 1138, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888054.

Okumura:2004:VNE

[OM04] T. Okumura and D. Mosse. Virtualizing network I/O on endhost operating system: operating system support for network control and resource protection. IEEE Transactions on Computers, 53(10):1303– 1316, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327580.

Odell:2007:AUI

[OM07] J. Odell and K. Mukerjee. Architecture, user interface, and enabling technology in Windows Vista’s speech systems. IEEE Transactions on Computers, 56(9):1156–1168, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4288083.

Oommen:2007:RBG

[OMG07] B. J. Oommen, S. Misra, and O.-C. Granmo. Routing bandwidth-guaranteed paths in MPLS traffic engineering: a multiple race track learning approach. IEEE Transactions on Computers, 56(7): 959–976, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4216294.

Oh:2002:EED

[OMM02] N. Oh, S. Mitra, and E. J. Mc-Cluskey. ED^{4}I: error detection by diverse data and duplicated instructions. IEEE Transactions on Computers, 51(2): 180–199, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980007.

Olarlu:2000:OHA

[OPZ00] S. Olarlu, M. C. Pinotti, and Si Qing Zheng. An optimal hardware-algorithm for sorting using a fixed-size parallel sorting device. IEEE Transactions on Computers, 49(12):1310– 1324, December 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895849.

Oommen:2000:CLA

[OR00] B. J. Oommen and T. D. Roberts. Continuous learning automata solutions to the capacity assignment problem. IEEE Transactions on Computers, 49(6):608–620, June 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=862220.

Omana:2005:LCH

[ORM05] M. Omana, D. Rossi, and C. Metra. Low cost and high speed embedded two-rail code checker. IEEE Transactions on Computers, 54(2): 153–164, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1377154.

Omana:2007:LST

[ORM07] M. Omana, D. Rossi, and C. Metra. Latch susceptibility to transient faults and new hardening approach. IEEE Transactions on Computers, 56(9):1255–1268, September 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4288092.

ORourke:2003:ANM

[OS03] C. O’Rourke and B. Sunar.

Achieving NTRU with Montgomery multiplication. IEEE Transactions on Computers, 52 (4):440–448, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190585.

Oren:2007:RPE

[OS07] Y. Oren and A. Shamir. Remote password extraction from RFID tags. IEEE Transactions on Computers, 56(9):1292– 1296, September 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288095.

Ozdoganoglu:2006:SHS

[OVB^{+}06] H. Ozdoganoglu, T. N. Vijaykumar, C. E. Brodley, B. A. Kuperman, and A. Jalote. SmashGuard: a hardware solution to prevent security attacks on the function return address. IEEE Transactions on Computers, 55(10):1271– 1285, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683758.

Oikonomakos:2006:DSC

[OZ06] P. Oikonomakos and M. Zwolinski. On the design of selfchecking controllers with datapath interactions. IEEE Transactions on Computers, 55(11):1423–1434, November 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1705451.

Pujara:2008:CNP

[PA08] P. Pujara and A. Aggarwal. Cache noise prediction. IEEE Transactions on Computers, 57(10):1372– 1386, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4509421.

Palem:2005:EAC

[Pal05a] K. V. Palem. Energy aware computing through probabilistic switching: a study of limits. IEEE Transactions on Computers, 54(9):1123–1137, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471673.

Palis:2005:GMF

[Pal05b] M. A. Palis. The granularity metric for fine-grain real-time scheduling. IEEE Transactions on Computers, 54(12):1572– 1583, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524938.

Parks:2000:NTT

[Par00] M. Parks. Number-theoretic test generation for directed rounding. IEEE Transactions on Computers, 49(7): 651–658, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863034.

Parhami:2003:TUB

[Par03]  B. Parhami.  Tight upper

bounds on the minimum precision required of the divisor and the partial remainder in high-radix division. IEEE Transactions on Computers, 52(11):1509–1514, November 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244949.

Park:2004:MMS

[Par04] Jong Won Park. Multiaccess memory system for attached SIMD computer. IEEE Transactions on Computers, 53(4): 439–452, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268401.

Pardo:2007:SDM

[PAW07] José M. Pardo, Xavier Anguera, and Chuck Wooters. Speaker diarization for multiple-distant-microphone meetings using several sources of information. IEEE Transactions on Computers, 56(9):1212–1224, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288088;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4288090.

Pineiro:2002:HSD

[PB02] J. A. Piñeiro and J. D. Bruguera. High-speed double precision computation of reciprocal, division, square root, and inverse square root. IEEE Transactions on Computers, 51(12):1377–1388, December 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1146704.

Phillips:2004:MWD

[PB04] B. Phillips and N. Burgess. Minimal weight digit set conversions. IEEE Transactions on Computers, 53(6): 666–677, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288542.

Pan:2007:EER

[PB07a] Z. Pan and M. A. Breuer.

Estimating error rate in defective logic using signature analysis. IEEE Transactions on Computers, 56(5): 650–661, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118682.

Patel:2007:FPP

[PB07b] R. A. Patel and S. Boussakta.

Fast parallel-prefix architectures for modulo 2^{n} − 1 addition with a single representation of zero. IEEE Transactions on Computers, 56(11):1484– 1492, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336297.

Patel:2007:FMA

[PBB07] R. A. Patel, M. Benaissa, and S. Boussakta. Fast modulo 2^{n} − (2^{n−2} + 1) addition: a new class of adder for RNS. IEEE Transactions on Computers, 56 (4):572–576, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118680.

Petrioli:2003:CBM

[PBC03] C. Petrioli, S. Basagni, and M. Chlamtac. Configuring BlueStars: multihop scatternet formation for Bluetooth networks. IEEE Transactions on Computers, 52(6): 779–790, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204833.

Pineiro:2008:RDD

[PBLM08] J.-A. Pineiro, J. D. Bruguera, F. Lamberti, and P. Montuschi. A radix-2 digit-by-digit architecture for cube root. IEEE Transactions on Computers, 57 (4):562–566, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4407683.

Poledna:2000:RDF

[PBWB00] S. Poledna, A. Burns, A. Wellings, and P. Barrett. Replica determinism and flexible scheduling in hard real-time dependable systems. IEEE Transactions on Computers, 49(2): 100–111, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833107.

Park:2002:OSA

[PC02a] Chan-Ik Park and Tae-Young Choe. An optimal scheduling algorithm based on task duplication. IEEE Transactions on Computers, 51(4): 444–448, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=995454.

Pitoura:2002:MDB

[PC02b] E. Pitoura and P. K. Chrysanthis. Multiversion data broadcast. IEEE Transactions on Computers, 51(10):1224– 1230, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039848.

Park:2007:EED

[PC07a] Kwangjin Park and Hyunseung Choo. Energy-efficient data dissemination schemes for nearest neighbor query processing. IEEE Transactions on Computers, 56(6): 754–768, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167787.

Pellizzoni:2007:RTM

[PC07b] R. Pellizzoni and M. Caccamo.

Real-time management of hardware and software tasks for FPGA-based embedded systems. IEEE Transactions on Computers, 56(12):1666– 1680, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358220.

Paul:2002:TEL

[PCC02] K. Paul, D. R. Choudhury, and P. P. Chaudhuri. Theory of Extended Linear Machines. IEEE Transactions on Computers, 51(9):1106–1110, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032629.

Piernas:2007:DNJ

[PCG07] Juan Piernas, Toni Cortes, and Jose M. Garcia. The design of new journaling file systems: The DualFS case. IEEE Transactions on Computers, 56(2): 267–281, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4042686.

Park:2006:EBP

[PCH06] Sun-Mi Park, Ku-Young Chang, and Dowon Hong. Efficient bit-parallel multiplier for irreducible pentanomials using a shifted polynomial basis. IEEE Transactions on Computers, 55(9):1211–1215, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1668049.

Pettis:2006:SOD

[PCL06] N. Pettis, L. Cai, and Yung-Hsiang Lu. Statistically optimal dynamic power management for streaming data. IEEE Transactions on Computers, 55 (7):800–814, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637397.

Pastor:2001:SAB

[PCR01] E. Pastor, J. Cortadella, and O. Roig. Symbolic analysis of bounded Petri nets. IEEE Transactions on Computers, 50 (5):432–448, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926158.

Park:2004:PAM

[PDS04] Joonseok Park, P. C. Diniz, and K. R. Shesha Shayee. Performance and area modeling of complete FPGA designs in the presence of loop transformations. IEEE Transactions on Computers, 53(11):1420– 1435, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336763.

Petra:2007:NAG

[PDS07] N. Petra, D. De Caro, and A. G. M. Strollo. A novel architecture for Galois fields GF(2^{m}) multipliers based on Mastrovito scheme. IEEE Transactions on Computers, 56(11):1470– 1483, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336296.

Pineiro:2004:AAL

[PEB04] J. A. Piñeiro, M. D. Ercegovac, and J. D. Bruguera. Algorithm and architecture for logarithm, exponential and powering computation. IEEE Transactions on Computers, 53(9):1085– 1096, September 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315603;   http:
//www.ac.usc.es/arquivos/
articulos/2004/gac2004-j05.
ps.

Pimentel:2006:SAE

[PEP06] A. D. Pimentel, C. Erbas, and S. Polstra. A systematic approach to exploring embedded system architectures at multiple abstraction levels. IEEE Transactions on Computers, 55 (2):99–112, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566572.

Pierre:2008:TFM

[PF08] L. Pierre and L. Ferro. A tractable and fast method for monitoring SystemC TLM specifications. IEEE Transactions on Computers, 57(10): 1346–1356, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4509420.

Parcerisa:2001:ILT

[PG01] J.-M. Parcerisa and A. Gonzalez. Improving latency tolerance of multithreading through decoupling. IEEE Transactions on Computers, 50(10): 1084–1094, October 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956093.

Pande:2005:PED

[PGJ^{+}05] Partha Pratim Pande, C. Grecu, M. Jones, A. Ivanov, and R. Saleh. Performance evaluation and design trade-offs for network-on-chip interconnect architectures. IEEE Transactions on Computers, 54(8): 1025–1040, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453503.

Phatak:2001:CTA

[PGK01] D. S. Phatak, T. Goff, and I. Koren. Constant-time addition and simultaneous format conversion based on redundant binary representations. IEEE Transactions on Computers, 50(11):1267– 1278, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966499.

Psarakis:2000:SFM

[PGPZ00] M. Psarakis, D. Gizopoulos, A. Paschalis, and Y. Zorian. Sequential fault modeling and test pattern generation for CMOS iterative logic arrays. IEEE Transactions on Computers, 49(10): 1083–1099, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888044.

Puente:2008:IDR

[PGVB08] V. Puente, J. A. Gregorio, F. Vallejo, and R. Beivide. Immunet: Dependable routing for interconnection networks with arbitrary topology. IEEE Transactions on Computers, 57(12):1676–1689, December 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4531734.

Pai:2008:PED

[PH08] Hung-Ta Pai and Y. S.

Han. Power-efficient directvoting assurance for data fusion in wireless sensor networks. IEEE Transactions on Computers, 57(2):261–273, February 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358250.

Paynter:2006:MAW

[PHA06] S. E. Paynter, N. Henderson, and J. M. Armstrong. Metastability in asynchronous waitfree protocols. IEEE Transactions on Computers, 55(3): 292–303, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1583559.

Piestrak:2002:CNT

[Pie02a] S. J. Piestrak. Comments on “Novel totally self-checking Berger checker designs based on generalized Berger code partitioning”. IEEE Transactions on Computers, 51(6): 735–736, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009156.

Piestrak:2002:DMC

[Pie02b] S. J. Piestrak. Design method of a class of embedded combinational self-testing checkers for two-rail codes. IEEE Transactions on Computers, 51(2): 229–234, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980010.

Pal:2007:CRA

[PKCD07] S. Pal, S. R. Kundu, M. Chatterjee, and S. K. Das. Combinatorial reverse auction based scheduling in multi-rate wireless systems. IEEE Transactions on Computers, 56(10): 1329–1341, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302706.

Papadantonakis:2009:PSA

[PKCD09] K. Papadantonakis, N. Kapre, S. Chan, and A. DeHon. Pipelining saturated accumulation. IEEE Transactions on Computers, 58(2):208–219, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4585362.

Ponomarev:2004:ISL

[PKEG04a] D. Ponomarev, G. Kucuk, O. Ergin, and K. Ghose. Isolating short-lived operands for energy reduction. IEEE Transactions on Computers, 53(6): 697–709, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288545.

Ponomarev:2004:EEC

[PKEG04b] D. V. Ponomarev, G. Kucuk, O. Ergin, and K. Ghose. Energy efficient comparators for superscalar datapaths. IEEE Transactions on Computers, 53 (7):892–904, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321049.

Ponomarev:2006:DRS

[PKG06] D. Ponomarev, G. Kucuk, and K. Ghose. Dynamic resizing of superscalar datapath components for energy efficiency. IEEE Transactions on Computers, 55(2): 199–213, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566580.

Park:2009:MMD

[PKL09] Jung-Heum Park, Hee-Chul Kim, and Hyeong-Seok Lim. Many-to-many disjoint path covers in the presence of faulty elements. IEEE Transactions on Computers, 58(4): 528–540, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4620105.

Patel:2000:AEV

[PKM00] A. Patel, A. Kusalik, and C. McCrosky. Area-efficient VLSI layouts for binary hypercubes. IEEE Transactions on Computers, 49(2): 160–169, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=833112.

Pomeranz:2004:MUO

[PKR04] I. Pomeranz, Sandip Kundu, and S. M. Reddy. Masking of unknown output values during output response compression by using comparison units. IEEE Transactions on Computers, 53 (1):83–89, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1255794.

Patel:2001:RHF

[PL01] S. J. Patel and S. S. Lumetta. rePLay: a hardware framework for dynamic optimization. IEEE Transactions on Computers, 50(6):590–608, June 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=931895.

Prasanna:2004:ENa

[PL04] V. K. Prasanna and F. Lombardi. Editor’s note. IEEE Transactions on Computers, 53(4):385, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268394.

Prasanna:2005:ENa

[PL05] V. K. Prasanna and F. Lombardi. Editor’s note. IEEE Transactions on Computers, 54 (2):97, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377148.

Prasanna:2006:ENa

[PL06] V. K. Prasanna and F. Lombardi. Editors’ note. IEEE Transactions on Computers, 55(1):1, January 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545746.

Pettis:2009:HAP

[PL09a] N. Pettis and Yung-Hsiang Lu.

A homogeneous architecture for power policy integration in operating systems. IEEE Transactions on Computers, 58 (7):945–955, July 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633348.

Puttaswamy:2009:ISC

[PL09b] K. Puttaswamy and G. H. Loh. 3D-integrated SRAM components for high-performance microprocessors. IEEE Transactions on Computers, 58(10): 1369–1381, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5161253.

Park:2003:EPR

[PLK^{+}03] W.-C. Park, K.-W. Lee, I.- S. Kim, T.-D. Han, and S.- B. Yang. An effective pixel rasterization pipeline architecture for 3D rendering processors. IEEE Transactions on Computers, 52(11):1501– 1508, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244948.

Park:2008:CEP

[PLP08] Ki-Woong Park, Sang Seok Lim, and Kyu Ho Park. Computationally efficient PKI-Based Single Sign-On Protocol, PKASSO for mobile devices. IEEE Transactions on Computers, 57(6): 821–834, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4459312.

Pedregal-Martin:2002:SRM

[PMR02] C. Pedregal-Martin and K. Ramamritham. Support for recovery in mobile systems. IEEE Transactions on Computers, 51(10):1219–1224, October 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039847.

Phoha:2004:SCS

[PNRP04] V. V. Phoha, A. U. Nadgar, A. Ray, and S. Phoha. Supervisory control of software systems. IEEE Transactions on Computers, 53(9):1187–1199, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315611.

Pineiro:2005:HSF

[POMB05] Jose-Alejandro Piñeiro, Stuart F. Oberman, Jean-Michel Muller, and Javier D. Bruguera. High-speed function approximation using a minimax quadratic interpolator. IEEE Transactions on Computers, 54 (3):304–318, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0304abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0304.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0304.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388195.
pdf?isnumber=30205&prod=
JNL&arnumber=1388195&arSt=
+304&ared=+318&arAuthor=
Pineiro%2C+J.-A.%3B+Oberman%
2C+S.F.%3B+Muller%2C+J.-
M.%3B+Bruguera%2C+J.D.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
30205&arnumber=1388195&count=
13&index=6;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1388195.

Phatrapornnant:2006:RJE

[PP06] T. Phatrapornnant and M. J. Pont. Reducing jitter in embedded systems employing a time-triggered software architecture and dynamic voltage scaling. IEEE Transactions on Computers, 55(2): 113–124, February 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1566573.

Poletti:2007:EEM

[PPB^{+}07] F. Poletti, A. Poggiali, D. Bertozzi, L. Benini, P. Marchal, M. Loghi, and M. Poncino. Energyefficient multiprocessor systemson-chip for embedded computing: Exploring programming models and their architectural support. IEEE Transactions on Computers, 56(5): 606–621, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141236.

Pomeranz:2000:FMF

[PR00a] I. Pomeranz and S. M. Reddy.

On finding a minimal functional description of a finite-state machine for test generation for adjacent machines. IEEE Transactions on Computers, 49(1): 88–94, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=822567.

Pomeranz:2000:UFS

[PR00b] I. Pomeranz and S. M. Reddy.

On the use of fully specified initial states for testing of synchronous sequential circuits. IEEE Transactions on Computers, 49(2): 175–182, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833114.

Pomeranz:2000:PSC

[PR00c] I. Pomeranz and S. M. Reddy.

Procedures for static compaction of test sequences for synchronous sequential circuits. IEEE Transactions on Computers, 49(6):596–607, June 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=862219.

Pomeranz:2002:BTS

[PR02a] I. Pomeranz and S. M. Reddy. Built-in test sequence generation for synchronous sequential circuits based on loading and expansion of input sequences using single and multiple fault detection times. IEEE Transactions on Computers, 51(4): 409–419, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=995451.

Pomeranz:2002:ETS

[PR02b] I. Pomeranz and S. M. Reddy. Enumeration of test sequences in increasing chronological order to improve the levels of compaction achieved by vector omission. IEEE Transactions on Computers, 51(7): 866–872, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017705.

Pomeranz:2002:SBB

[PR02c] I. Pomeranz and S. M.

Reddy. A storage-based builtin test pattern generation method for scan circuits based on partitioning and reduction of a precomputed test set. IEEE Transactions on Computers, 51(11):1282–1293, November 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047753.

Phipatanasuphorn:2004:VSN

[PR04a] V. Phipatanasuphorn and Parameswaran Ramanathan. Vulnerability of sensor networks to unauthorized traversal and monitoring. IEEE Transactions on Computers, 53(3): 364–369, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261841.

Pomeranz:2004:MQD

[PR04b] I. Pomeranz and S. M.

Reddy. A measure of quality for n-detection test sets. IEEE Transactions on Computers, 53(11):1497–1503, November 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336770.

Pomeranz:2004:MFC

[PR04c] I. Pomeranz and S. M. Reddy. On maximizing the fault coverage for a given test length limit in a synchronous sequential circuit. IEEE Transactions on Computers, 53(9):1121–1133, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315606.

Pomeranz:2004:STC

[PR04d] I. Pomeranz and S. M. Reddy. Static test compaction for fullscan circuits based on combinational test sets and nonscan input sequences and a lower bound on the number of tests. IEEE Transactions on Computers, 53(12):1569– 1581, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347083.

Pomeranz:2006:GTA

[PR06] I. Pomeranz and S. M. Reddy. On generating tests that avoid the detection of redundant faults in synchronous sequential circuits with full scan. IEEE Transactions on Computers, 55 (4):491–495, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608010.

Prasanna:2003:ENa

[Pra03a]  V. K. Prasanna.  Editor’s

note. IEEE Transactions on Computers, 52(3):257–259, March 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183942.

Prasanna:2003:ENb

[Pra03b]  V. K. Prasanna.  Editor’s

note. IEEE Transactions on Computers, 52(7):833–834, July 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214332.

Prasanna:2004:ENb

[Pra04]  V. K. Prasanna.  Editor’s

note. IEEE Transactions on Computers, 53(12):1505–1507, December 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347077.

Prasanna:2005:ENb

[Pra05]  V. K. Prasanna.  Editor’s

note. IEEE Transactions on Computers, 54(7):785–787, July 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432662.

Prasanna:2006:ENb

[Pra06a]  V. K. Prasanna.  Editor’s

note. IEEE Transactions on Computers, 55(3):241–242, March 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1583554.

Prasanna:2006:INE

[Pra06b] V. K. Prasanna. Introducing the new Editor-in-Chief of the IEEE Transactions on Computers. IEEE Transactions on Computers, 55(12):1489– 1490, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1717382.

Prouff:2009:SAS

[PRB09] E. Prouff, M. Rivain, and R. Bevan. Statistical analysis of second order differential power analysis. IEEE Transactions on Computers, 58(6): 799–811, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4752810.

Pereira:2003:SRM

[PRO03] J. Pereira, L. Rodrigues, and R. Oliveira. Semantically reliable multicast: definition, implementation, and performance evaluation. IEEE Transactions on Computers, 52(2): 150–165, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176983.

Pleisch:2003:FTM

[PS03] S. Pleisch and A. Schiper.

Fault-tolerant mobile agent execution. IEEE Transactions on Computers, 52(2): 209–222, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176987.

Page:2004:PCA

[PS04] D. Page and N. P. Smart.

Parallel cryptographic arithmetic using a redundant Montgomery representation. IEEE Transactions on Computers, 53(11):1474–1482, November 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336767.

Park:2005:MLB

[PS05] J. Park and S. Sahni. Maximum lifetime broadcasting in wireless networks. IEEE Transactions on Computers, 54(9):1081– 1090, September 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471670.

Park:2006:OHM

[PS06a] J. Park and S. Sahni. An online heuristic for maximum lifetime routing in wireless sensor networks. IEEE Transactions on Computers, 55(8): 1048–1056, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1650201.

Phan:2006:FDB

[PS06b] R. C.-W. Phan and M. U. Siddiqi. A framework for describing block cipher cryptanalysis. IEEE Transactions on Computers, 55(11):1402– 1409, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705449.

Parikh:2004:PAB

[PSZS04] D. Parikh, K. Skadron, Y. Zhang, and M. Stan. Power-aware branch prediction: characterization and design. IEEE Transactions on Computers, 53(2): 168–186, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261827.

Pinho:2003:RRT

[PV03] L. M. Pinho and F. Vasques.

Reliable real-time communication in CAN networks. IEEE Transactions on Computers, 52(12):1594–1607, December 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252855.

Page:2006:FAP

[PV06] D. Page and F. Vercauteren. A fault attack on pairing-based cryptography. IEEE Transactions on Computers, 55(9): 1075–1080, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1668035.

Pierre:2002:DSO

[PvST02] G. Pierre, M. van Steen, and A. S. Tanenbaum. Dynamically selecting optimal distribution strategies for Web documents. IEEE Transactions on Computers, 51(6): 637–651, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009149.

Prabhakar:2002:QIV

[PXK^{+}02] S. Prabhakar, Yuni Xia, D. V.

Kalashnikov, W. G. Aref, and S. E. Hambrusch. Query indexing and velocity constrained indexing: scalable techniques for continuous queries on moving objects. IEEE Transactions on Computers, 51(10): 1124–1140, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039840.

Pan:2005:FBM

[PY05a] D. Pan and Y. Yang. FIFObased multicast scheduling algorithm for virtual output queued packet switches. IEEE Transactions on Computers, 54(10):1283–1297, October 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501793.

Petrenko:2005:TPD

[PY05b] A. Petrenko and N. Yevtushenko. Testing from partial deterministic FSM specifications. IEEE Transactions on Computers, 54(9):1154–1165, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471675.

Pan:2009:LIP

[PY09] Deng Pan and Yuanyuan Yang. Localized independent packet scheduling for buffered crossbar switches. IEEE Transactions on Computers, 58(2): 260–274, February 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4599572.

Qiu:2004:TLM

[QD04] X. Qiu and M. Dubois. Tolerating late memory traps in dynamically scheduled processors. IEEE Transactions on Computers, 53(6):732–743, June 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288548.

Qiu:2008:SLB

[QD08] Xiaogang Qiu and M. Dubois.

The synonym lookaside buffer: a solution to the synonym problem in virtual caches. IEEE Transactions on Computers, 57(12):1585–1599, December 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4585360.

Qin:2008:AAT

[QX08] Xiao Qin and Tao Xie. An availability-aware task scheduling strategy for heterogeneous systems. IEEE Transactions on Computers, 57(2):188–199, February 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358217.

Rajkumar:2000:GEI

[Raj00]  R. Rajkumar.  Guest ed-

itor’s introduction: 1997 IEEE real-time technologies and applications symposium. IEEE Transactions on Computers, 49(11):1153–1154, November 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895933.

Ray:2006:LCH

[Ray06] Swapan Kumar Ray. Largecapacity high-throughput lowcost pipelined CAM using pipelined CTAM. IEEE Transactions on Computers, 55(5): 575–587, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613838.

Roy:2005:ATQ

[RB05] Sanghamitra Roy and Prith Banerjee. An algorithm for trading off quantization error with hardware resources for MATLAB-based FPGA design. IEEE Transactions on Computers, 54(7):886–896, July 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432671.

Ren:2003:AAA

[RBC^{+}03] Yansong Ren, D. E. Bakken, T. Courtney, M. Cukier, D. A. Karr, P. Rubel, C. Sabnis, W. H. Sanders, R. E. Schantz, and M. Seri. AQuA: an adaptive architecture that provides dependable distributed objects. IEEE Transactions on Computers, 52(1):31– 50, January 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1159752.

Rajagopal:2006:TOA

[RC06] S. Rajagopal and J. R. Cavallaro. Truncated online arithmetic with applications to communication systems. IEEE Transactions on Computers, 55(10):1240–1252, October 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683755.

Reinman:2001:OED

[RCA01] G. Reinman, B. Calder, and T. Austin. Optimizations enabled by a decoupled frontend architecture. IEEE Transactions on Computers, 50(4):

338–355, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=919279.

Ruan:2001:CPS

[RDH^{+}01] Lu Ruan, Dingzhu Du, Xiaodong Hu, Xiaohua Jia, Deying Li, and Zheng Sun. Converter placement supporting broadcast in WDM optical networks. IEEE Transactions on Computers, 50(7): 750–758, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936240.

Redinbo:2003:FDA

[Red03]  G. R. Redinbo.  Failure-

detecting arithmetic convolutional codes and an iterative correcting strategy. IEEE Transactions on Computers, 52(11):1434–1442, November 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1244941.

Rescigno:2001:OBS

[Res01] A. A. Rescigno. Optimally balanced spanning tree of the star network. IEEE Transactions on Computers, 50(1): 88–91, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=902755.

Rakshit:2005:FBS

[RG05] Sudipta Rakshit and R. K.

Guha. Fair bandwidth sharing in distributed systems: a game-theoretic approach. IEEE Transactions on Computers, 54(11):1384–1393, November 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514418.

Rajan:2009:NCA

[RG09] K. Rajan and R. Govindarajan. A novel cache architecture and placement framework for packet forwarding engines. IEEE Transactions on Computers, 58(8):1009– 1025, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4752813.

Rodriguez-Henriquez:2003:PMB

[RHK03] F. Rodriguez-Henriquez and Ç. K. Koç. Parallel multipliers based on special irreducible pentanomials. IEEE Transactions on Computers, 52(12):1535–1542, December 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1252850.

Rodriguez-Henriquez:2008:LCB

[RHMLL08] F. Rodriguez-Henriquez, G. Morales-Luna, and J. Lopez. Lowcomplexity bit-parallel square root computation over GF(2^{m}) for all trinomials. IEEE Transactions on Computers, 57(4): 472–480, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358282.

Rhodes:2003:UPC

[Rho03] B. Rhodes. Using physical context for just-in-time information retrieval. IEEE Transactions on Computers, 52(8): 1011–1014, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1223636.

Rabah:2003:PEM

[RK03] M. Rabah and K. Kanoun. Performability evaluation of multipurpose multiprocessor systems: the “separation of concerns” approach. IEEE Transactions on Computers, 52(2): 223–236, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1176988.

Ruan:2005:LRO

[RK05] X. Ruan and R. S. Katti.

Left-to-right optimal signedbinary representation of a pair of integers. IEEE Transactions on Computers, 54(2): 124–131, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377151.

Ruan:2006:NSC

[RK06] X. Ruan and R. S. Katti.

A new source coding scheme with small expected length and its application to simple data encryption. IEEE Transactions on Computers, 55(10): 1300–1305, October 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1683760.

Ruan:2007:DIP

[RK07] Xiaoyu Ruan and R. S. Katti.

Data-independent pattern runlength compression for testing embedded cores in SoCs. IEEE Transactions on Computers, 56 (4):545–556, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118676.

Ruiz:2003:RFT

[RKFTF03] J. C. Ruiz, M.-O. Killijian, J.-C. Fabre, and P. Thvenod-Fosse. Reflective fault-tolerant systems: from experience to challenges. IEEE Transactions on Computers, 52(2): 237–254, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176989.

Ren:2005:HAD

[RKM05] Z. Ren, B. H. Krogh, and R. Marculescu. Hierarchical adaptive dynamic power management. IEEE Transactions on Computers, 54(4): 409–420, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401860.

Ravindran:2004:DLP

[RL04] Binoy Ravindran and P. Li. DPR, LPR: proactive resource allocation algorithms for asynchronous real-time distributed systems. IEEE Transactions on Computers, 53(2):201–216, February 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261829.

Raghavan:2009:DLC

[RLJ^{+}09] P. Raghavan, A. Lambrechts, M. Jayapala, F. Catthoor, and D. Verkest. Distributed loop controller for multithreading in unithreaded ILP architectures. IEEE Transactions on Computers, 58(3): 311–321, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624249.

Ramirez:2005:STC

[RLPV05] A. Ramirez, J. L. Larriba-Pey, and M. Valero. Software Trace Cache. IEEE Transactions on Computers, 54(1): 22–35, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362637.

Reyhani-Masoleh:2006:EAA

[RM06] A. Reyhani-Masoleh. Efficient algorithms and architectures for field multiplication using Gaussian normal bases. IEEE Transactions on Computers, 55(1): 34–47, January 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545749.

Rachuri:2009:EES

[RM09] K. K. Rachuri and C. Murthy.

Energy efficient and scalable search in dense wireless sensor networks. IEEE Transactions on Computers, 58(6): 812–826, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4782952.

Ravikumar:2005:EES

[RMB05] V. C. Ravikumar, R. N. Mahapatra, and Laxmi Narayan Bhuyan. EaseCAM: an energy and storage efficient TCAMbased router architecture for IP lookup. IEEE Transactions on Computers, 54(5): 521–533, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1407843.

Reyhani-Masoleh:2002:NCM

[RMH02] A. Reyhani-Masoleh and M. A. Hasan. A new construction of Massey–Omura parallel multiplier over GF(2^{m}). IEEE Transactions on Computers, 51 (5):511–520, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1004590.

Reyhani-Masoleh:2003:EMB

[RMH03a] A. Reyhani-Masoleh and M. A. Hasan. Efficient multiplication beyond optimal normal bases. IEEE Transactions on Computers, 52(4): 428–439, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190584.

Reyhani-Masoleh:2003:FNB

[RMH03b] A. Reyhani-Masoleh and M. A.

Hasan. Fast normal basis multiplication using general purpose processors. IEEE Transactions on Computers, 52(11):1379– 1390, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244936.

Reyhani-Masoleh:2004:LCB

[RMH04] A. Reyhani-Masoleh and M. A.

Hasan. Low complexity bit parallel architectures for polynomial basis multiplication over GF(2^{m}). IEEE Transactions on Computers, 53(8): 945–959, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1306989.

Reyhani-Masoleh:2005:LCW

[RMH05] A. Reyhani-Masoleh and M. A.

Hasan. Low complexity wordlevel sequential normal basis multipliers. IEEE Transactions on Computers, 54(2): 98–110, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377149.

Reyhani-Masoleh:2006:FDA

[RMH06] A. Reyhani-Masoleh and M. A. Hasan. Fault detection architectures for field multiplication using polynomial bases. IEEE Transactions on Computers, 55(9):1089–1103, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1668037.

Rahaman:2008:DRT

[RMPJ08] H. Rahaman, J. Mathew, D. K. Pradhan, and A. M. Jabir. Derivation of reduced test vectors for bit-parallel multipliers over GF(2^{m}). IEEE Transactions on Computers, 57(9): 1289–1294, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4487058.

Redinbo:2004:CED

[RN04] G. R. Redinbo and C. Nguyen. Concurrent error detection in wavelet lifting transforms. IEEE Transactions on Computers, 53(10):1291–1302, October 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327579.

Rosenberg:2004:SMS

[Ros04] A. L. Rosenberg. On scheduling mesh-structured computations for Internet-based computing. IEEE Transactions on Computers, 53(9):1176–1186, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315610.

Ryu:2001:TCR

[RPH01] Minsoo Ryu, Jungkeun Park, and Seongsoo Hong. Timing constraint remapping to achieve time equi-continuity in distributed real-time systems. IEEE Transactions on Computers, 50(12):1310– 1320, December 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=970570.

Rachlin:2008:AMB

[RS08a] E. Rachlin and J. E. Savage. Analysis of mask-based nanowire decoders. IEEE Transactions on Computers, 57(2):175–187, February 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4358247.

Rajasekaran:2008:OPA

[RS08b] S. Rajasekaran and S. Sen. Optimal and practical algorithms for sorting on the PDM. IEEE Transactions on Computers, 57 (4):547–561, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4359929.

Ravindran:2005:PVA

[RSM^{+}05] R. A. Ravindran, R. M. Senger, E. D. Marsman, G. S. Dasika, M. R. Guthaus, S. A. Mahlke, and R. B. Brown. Partitioning variables across register windows to reduce spill code in a low-power processor. IEEE Transactions on Computers, 54(8): 998–1012, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453501.

Rouvroy:2003:EUF

[RSQL03] G. Rouvroy, F.-X. Standaert, J.-J. Quisquater, and J.-D. Legat. Efficient uses of FP-GAs for implementations of DES and its experimental linear cryptanalysis. IEEE Transactions on Computers, 52(4): 473–482, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190588.

Ramalingom:2000:MTS

[RTD00] T. Ramalingom, K. Thulasiraman, and A. Das. A matroidtheoretic solution to an assignment problem in the conformance testing of communication protocols. IEEE Transactions on Computers, 49(4): 317–330, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844345.

Reese:2005:CGP

[RTT05] R. B. Reese, M. A. Thornton, and C. Traver. A coarsegrain phased logic CPU. IEEE Transactions on Computers, 54 (7):788–799, July 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432663.

Radhakrishnan:2001:JRS

[RVJ^{+}01] R. Radhakrishnan, N. Vijaykrishnan, L. K. John, A. Sivasubramaniam, J. Rubio, and J. Sabarinathan. Java runtime systems: characterization and architectural implications. IEEE Transactions on Computers, 50(2): 131–146, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=908989.

Raman:2008:CDI

[RW08] R. Raman and D. S. Wise.

Converting to and from Dilated Integers. IEEE Transactions on Computers, 57(4): 567–573, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358258.

Rosenberg:2005:GSS

[RY05] A. L. Rosenberg and M. Yurkewych. Guidelines for scheduling some common computation-dags for Internet-based computing. IEEE Transactions on Computers, 54 (4):428–438, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1401862.

Ryan:2004:IST

[Rya04] J. J. C. H. Ryan. Information security tools and practices: what works? IEEE Transactions on Computers, 53(8): 1060–1063, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306997.

Radecka:2004:DVT

[RZ04] K. Radecka and Z. Zilic. Design verification by test vectors and arithmetic transform universal test set. IEEE Transactions on Computers, 53(5): 628–640, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275301.

Schweitz:2002:PDO

[SA02] E. A. Schweitz and D. P.

Agrawal. Parallelization domain oriented multilevel graph partitioner. IEEE Transactions on Computers, 51(12):1435– 1441, December 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1146709.

Saab:2004:EMA

[Saa04] Y. G. Saab. An effective multilevel algorithm for bisecting graphs and hypergraphs. IEEE Transactions on Computers, 53 (6):641–652, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1288540.

Saed:2002:NSC

[SAJ02] A. Saed, M. Ahmadi, and G. A. Jullien. A number system with continuous valued digits and modulo arithmetic. IEEE Transactions on Computers, 51(11):1294– 1305, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047754.

Shaw:2003:IBF

[SAKR03] D. B. Shaw, D. Al-Khalili, and C. N. Rozon. IC bridge fault modeling for IP blocks using neural network-based VHDL saboteurs. IEEE Transactions on Computers, 52(10): 1285–1297, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234526.

Sanders:2003:ASR

[San03]  P. Sanders.  Asynchronous

scheduling of redundant disk arrays. IEEE Transactions on Computers, 52(9):1170–1184, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1228512.

Sangireddy:2006:RRL

[San06] Rama Sangireddy. Reducing rename logic complexity for highspeed and low-power front-end architectures. IEEE Transactions on Computers, 55(6): 672–685, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628956.

Sarbazi-Azad:2001:AMW

[SAOKM01] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Analytical modeling of wormhole-routed k-ary ncubes in the presence of hotspot traffic. IEEE Transactions on Computers, 50(7): 623–634, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=936230.

Silva:2009:UVI

[SAT09] L. M. Silva, J. Alonso, and J. Torres. Using virtualization to improve software rejuvenation. IEEE Transactions on Computers, 58(11):1525– 1538, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5184821.

Savas:2005:CFA

[Sav05] E. Savas. A carry-free architecture for Montgomery inversion. IEEE Transactions on Computers, 54(12):1508– 1519, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524933.

Suresh:2009:TEE

[SAYN09] D. C. Suresh, B. Agrawal, Jun Yang, and W. A. Najjar. Tunable and energy efficient bus encoding techniques. IEEE Transactions on Computers, 58(8):1049– 1062, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4798154.

Sasao:2001:WBI

[SB01] T. Sasao and J. T. Butler. Worst and best irredundant sum-of-products expressions. IEEE Transactions on Computers, 50(9):935– 948, September 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=954508.

Schulte:2005:GEI

[SB05] Michael J. Schulte and Jean-Claude Bajard. Guest Editors’ introduction: Special issue on computer arithmetic. IEEE Transactions on Computers, 54 (3):241–242, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/tc/2005/03/
                                        t0241.pdf;   http://csdl.
                                        computer.org/dl/trans/tc/
                                        2005/03/t0241.htm;   http:
                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1388189.

Srivastava:2007:HPM

[SB07] Saket Srivastava and Sanjukta Bhanja. Hierarchical probabilistic macromodeling for QCA circuits. IEEE Transactions on Computers, 56(2):

174–190, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042678.

Schulte:2000:IMO

[SBAB00] M. J. Schulte, P. I. Balzola, A. Akkas, and R. W. Brocato. Integer multiplication with overflow detection or saturation. IEEE Transactions on Computers, 49(7):681–691, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//home.ku.edu.tr/~ahakkas/
publications/overflow.pdf;
http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=863038;   http://
mesa.ece.wisc.edu/publications/
cp_2000-08.pdf.

Shacham:2008:PNC

[SBC08] A. Shacham, K. Bergman, and L. P. Carloni. Photonic networks-on-chip for future generations of chip multiprocessors. IEEE Transactions on Computers, 57(9):1246– 1260, September 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4509424.

Sakiyama:2007:MCB

[SBPV07] K. Sakiyama, L. Batina, B. Preneel, and I. Verbauwhede. Multicore curve-based cryptoprocessor with reconfigurable modular arithmetic logic units over GF(2^{n}). IEEE Transactions on Computers, 56(9):1269–1282, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288093.

Santi:2004:RNS

[SC04] P. Santi and S. Chessa. Reducing the number of sequential diagnosis iterations in hypercubes. IEEE Transactions on Computers, 53(1):89– 92, January 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255796.

Sabbineni:2005:LAF

[SC05] H. Sabbineni and K. Chakrabarty. Location-aided flooding: an energy-efficient data dissemination protocol for wirelesssensor networks. IEEE Transactions on Computers, 54(1): 36–46, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362638.

Sheu:2006:PBG

[SC06] Shiann-Tsong Sheu and Yue-Ru Chuang. A pipelinebased genetic algorithm accelerator for time-critical processes in real-time systems. IEEE Transactions on Computers, 55(11):1435–1448, November 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705452.

Sehgal:2007:ODS

[SC07a] A. Sehgal and K. Chakrabarty.

Optimization of dual-speed TAM architectures for efficient modular testing of SOCs. IEEE Transactions on Computers, 56 (1):120–133, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016502.

Sheu:2007:HAM

[SC07b] Pi-Rong Sheu and Shan-Tai Chen. On the hardness of approximating the multicast delay variation problem. IEEE Transactions on Computers, 56(11):1575–1577, November 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336304.

Sengupta:2008:IQV

[SCG08] S. Sengupta, M. Chatterjee, and S. Ganguly. Improving quality of VoIP streams over WiMax. IEEE Transactions on Computers, 57(2):

145–156, February 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358249.

Schwiebert:2001:DFO

[Sch01]  L. Schwiebert.  Deadlock-

free oblivious wormhole routing with cyclic dependencies. IEEE Transactions on Computers, 50(9):865–876, September 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=954503.

Sarkar:2006:FBP

[SCK06] Arnab Sarkar, P. P. Chakrabarti, and Rajeev Kumar. Framebased proportional roundrobin. IEEE Transactions on Computers, 55(9):1121–1129, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1668040.

Sasidhar:2001:BSB

[SCZ01] K. Sasidhar, A. Chatterjee, and Y. Zorian. Boundary scan-based relay wave propagation test of arrays of identical structures. IEEE Transactions on Computers, 50(10): 1007–1019, October 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956088.

Schaelicke:2006:DTO

[SD06] L. Schaelicke and A. L. Davis.

Design trade-offs for user-level I/O architectures. IEEE Transactions on Computers, 55(8): 962–973, August 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650194.

Schmiedle:2003:ERS

[SDB03] F. Schmiedle, R. Drechsler, and B. Becker. Exact routing with search space reduction. IEEE Transactions on Computers, 52 (6):815–825, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1204836.

Skliarova:2004:RHS

[SdBF04] I. Skliarova and A. de Brito Ferrari. Reconfigurable hardware SAT solvers: a survey of systems. IEEE Transactions on Computers, 53(11):1449– 1461, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336765.

Srinivasan:2004:PT

[SDT04] Viji Srinivasan, E. S. Davidson, and G. S. Tyson. A prefetch taxonomy. IEEE Transactions on Computers, 53(2): 126–140, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261824.

Seidel:2004:DOI

[SE04] P.-M. Seidel and G. Even. Delay-optimized implementation of IEEE floating-point addition. IEEE Transactions on Computers, 53(2): 97–113, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261822.

Seznec:2005:CFA

[SE05] A. Seznec and R. Espasa. Conflict-free accesses to strided vectors on a banked cache. IEEE Transactions on Computers, 54(7):913–916, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1432673.

Seznec:2004:CSM

[Sez04] A. Seznec. Concurrent support of multiple page sizes on a skewed associative TLB. IEEE Transactions on Computers, 53 (7):924–927, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321052.

Ssu:2003:PRH

[SFJ03] Kuo-Feng Ssu, W. K. Fuchs, and H. C. Jiau. Process recovery in heterogeneous systems. IEEE Transactions on Computers, 52(2):126–138, February 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176981.

Santana:2009:DCE

[SFRV09] O. J. Santana, A. Falcon, A. Ramirez, and M. Valero. DIA: a complexity-effective decoding architecture. IEEE Transactions on Computers, 58 (4):448–462, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624251.

Sen:2007:FVS

[SG07] A. Sen and V. K. Garg.

Formal verification of simulation traces using computation slicing. IEEE Transactions on Computers, 56(4): 511–527, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118674.

Srinivasan:2000:NTP

[SGB00] R. Srinivasan, S. K. Gupta, and M. A. Breuer. Novel test pattern generators for pseudoexhaustive testing. IEEE Transactions on Computers, 49(11):1228–1240, November 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=895939.

Stuijk:2008:TBT

[SGB08] S. Stuijk, M. Geilen, and T. Basten. Throughputbuffering trade-off exploration for cyclo-static and synchronous dataflow graphs. IEEE Transactions on Computers, 57(10): 1331–1345, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4483507.

Steinwandt:2008:GEI

[SGK08] Rainer Steinwandt, Willi Geiselmann, and Çetin Kaya Koç. Guest Editors’ introduction to the special section on special-purpose hardware for cryptography and cryptanalysis. IEEE Transactions on Computers, 57(11):1441–1442, November 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633726.

Shoufan:2009:HPR

[SH09] A. Shoufan and S. A. Huss.

High-performance rekeying processor architecture for group key management. IEEE Transactions on Computers, 58(10): 1421–1434, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5156494.

Song:2006:RRH

[SHK06] S. Song, Kai Hwang, and Yu-Kwong Kwok. Risk-resilient heuristics and genetic algorithms for security-assured grid job scheduling. IEEE Transactions on Computers, 55(6): 703–719, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1628958.

Soh:2008:EPU

[SHR08] Sieteng Soh, L. Hiryanto, and S. Rai. Efficient prefix updates for IP router using lexicographic ordering and updatable address set. IEEE Transactions on Computers, 57(1): 110–125, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358233.

Schaumont:2006:MDV

[SHYV06] P. Schaumont, D. Hwang, Shenglin Yang, and I. Verbauwhede. Multilevel design validation in a secure embedded system. IEEE Transactions on Computers, 55(11):1380– 1390, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705447.

Savas:2000:MMI

[SK00] E. Savas and Ç. K. Koç. The Montgomery modular inverse— revisited. IEEE Transactions on Computers, 49(7): 763–766, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=863048.

Sunar:2001:EON

[SK01] B. Sunar and C. K. Koc. An efficient optimal normal basis type II multiplier. IEEE Transactions on Computers, 50(1): 83–87, January 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=902754.

Sklavos:2002:AVI

[SK02] N. Sklavos and O. Koufopavlou. Architectures and VLSI implementations of the AESproposal Rijndael. IEEE Transactions on Computers, 51(12):1454–1459, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1146712.

Sahni:2004:DRT

[SK04] S. Sahni and K. S. Kim.

An O(log n) dynamic routertable design. IEEE Transactions on Computers, 53(3): 351–363, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261840.

Shu:2009:RCA

[SKG09] Chang Shu, Soonhak Kwon, and K. Gaj. Reconfigurable computing approach for Tate pairing cryptosystems over binary fields. IEEE Transactions on Computers, 58(9):1221– 1237, September 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815221.

Sabrina:2007:DAI

[SKJ07] F. Sabrina, S. S. Kanhere, and S. K. Jha. Design, analysis and implementation of a novel multiple resource scheduler. IEEE Transactions on Computers, 56(8): 1071–1086, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4264322.

Shankaran:2009:IPA

[SKK^{+}09] N. Shankaran, J. S. Kinnebrew, X. D. Koutsoukas, Chenyang Lu, D. C. Schmidt, and G. Biswas. An integrated planning and adaptive resource management architecture for distributed real-time embedded systems. IEEE Transactions on Computers, 58(11):1485– 1499, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4799777.

Sangireddy:2004:LPH

[SKS04] Rama Sangireddy, H. Kim, and A. K. Somani. Lowpower high-performance reconfigurable computing cache architectures. IEEE Transactions on Computers, 53(10): 1274–1290, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327578.

Styles:2004:EPB

[SL04] H. Styles and W. Luk. Exploiting program branch probabilities in hardware compilation. IEEE Transactions on Computers, 53(11):1408– 1419, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336762.

Scordino:2006:RRA

[SL06] C. Scordino and G. Lipari. A resource reservation algorithm for power-aware scheduling of periodic and aperiodic realtime tasks. IEEE Transactions on Computers, 55(12):1509– 1522, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1717384.

Srimani:2002:GES

[SLG02] P. K. Srimani, Wang-Chien Lee, and S. K. S. Gupta. Guest editorial: special section on data management systems and mobile computing. IEEE Transactions on Computers, 51(10): 1121–1123, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039839.

Singh:2000:MIR

[SLL^{+}00] H. Singh, Ming-Hau Lee, Guangming Lu, F. J. Kurdahi, N. Bagherzadeh, and E. M. Chaves Filho. MorphoSys: an integrated reconfigurable system for data-parallel and computation-intensive applications. IEEE Transactions on Computers, 49(5): 465–481, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=859540.

Srikanthan:2004:ATE

[SLS04] T. Srikanthan, S. K. Lam, and Mishra Suman. Area-time efficient sign detection technique for binary signed-digit number system. IEEE Transactions on Computers, 53(1): 69–72, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1255791.

Solihin:2001:ACM

[SLT01] Yan Solihin, Jaejin Lee, and J. Torrellas. Automatic code mapping on an intelligent memory architecture. IEEE Transactions on Computers, 50(11):1248–1266, November 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966498.

Shu:2005:DIS

[SLZ05a] Jiwu Shu, Bigang Li, and Weimin Zheng. Design and implementation of an SAN system based on the fiber channel protocol. IEEE Transactions on Computers, 54(4):

439–448, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401863.

Stehle:2005:SWC

[SLZ05b] Damien Stehlé, Vincent Lefèvre, and Paul Zimmermann. Searching worst cases of a one-variable function using lattice reduction. IEEE Transactions on Computers, 54(3):340–346, March 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
tc/2005/03/t0340abs.htm;
http://csdl.computer.org/
dl/trans/tc/2005/03/t0340.
htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/
t0340.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388198.
pdf?isnumber=30205&prod=
JNL&arnumber=1388198&arSt=
+340&ared=+346&arAuthor=
Stehle%2C+D.%3B+Lefevre%
2C+V.%3B+Zimmermann%2C+P.;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30205&arnumber=1388198&count=
13&index=9;   http://ieeexplore.
ieee.org/xpls/references.
jsp?arnumber=1388198.

Shen:2000:RCR

[SM00] Chia Shen and I. Mizumuma. RT-CRM: real-time channel-based reflective memory. IEEE Transactions on Computers, 49(11):1202–1214, November 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=895937.

Sarkar:2003:EIC

[SM03] P. Sarkar and S. Maitra. Efficient implementation of cryptographically useful “large” Boolean functions. IEEE Transactions on Computers, 52 (4):410–417, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1190582.

Shi:2009:ESA

[SM09] Feng Shi and Y. Makris. Enhancing simulation accuracy through advanced hazard detection in asynchronous circuits. IEEE Transactions on Computers, 58(3):394–408, March 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599573.

Smailagic:2003:WCN

[Sma03] A. Smailagic. Wearable computers: a new paradigm in computer systems and their applications. IEEE Transactions on Computers, 52(8): 977–978, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223631.

Suhaib:2006:VFL

[SMBS06] S. Suhaib, D. Mathaikutty, D. Berner, and S. Shukla. Validating families of latency insensitive protocols. IEEE Transactions on Computers, 55(11):1391–1401, November 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705448.

Sokolov:2005:DAD

[SMBY05] D. Sokolov, J. Murphy, A. Bystrov, and A. Yakovlev. Design and analysis of dual-rail circuits for security applications. IEEE Transactions on Computers, 54 (4):449–460, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1401864.

Seidel:2005:SRR

[SMM05] P.-M. Seidel, L. D. McFearin, and D. W. Matula. Secondary radix recodings for higher radix multipliers. IEEE Transactions on Computers, 54(2): 111–123, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377150.

Stanley-Marbell:2003:MAS

[SMMMK03] P. S. Stanley-Marbell, D. Marculescu, R. Marculescu, and P. K. Khosla. Modeling, analysis, and self-management of electronic textiles. IEEE Transactions on Computers, 52(8): 996–1010, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1223635.

Sakti:2007:IKS

[SMN07] S. Sakti, K. Markov, and S. Nakamura. Incorporating knowledge sources into a statistical acoustic model for spoken language communication systems. IEEE Transactions on Computers, 56(9):1199–1211, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4288087.

Sunar:2007:PST

[SMS07] B. Sunar, W. J. Martin, and D. R. Stinson. A provably secure true random number generator with built-in tolerance to active attacks. IEEE Transactions on Computers, 56(1): 109–119, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4016501.

Shum:2001:DEF

[SMSM01] A. Shum, P. M. Melliar-Smith, and L. E. Moser. Design and evaluation of the Fibonacci optical ATM switch. IEEE Transactions on Computers, 50 (5):466–481, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926160.

Sasao:2007:NFG

[SNB07] T. Sasao, S. Nagayama, and J. T. Butler. Numerical function generators using LUT cascades. IEEE Transactions on Computers, 56(6): 826–838, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167792.

Suh:2002:EAC

[SP02] Jinwoo Suh and V. K. Prasanna.

An efficient algorithm for out-of-core matrix transposition. IEEE Transactions on Computers, 51(4):420–438, April 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=995452.

Sharkey:2007:EOA

[SP07] Joseph J. Sharkey and Dmitry V.

Ponomarev. Exploiting operand availability for efficient simultaneous multithreading. IEEE Transactions on Computers, 56(2):208–223, February 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042681.

Sun:2008:HPM

[SPS08] Junqing Sun, G. D. Peterson, and O. O. Storaasli. Highperformance mixed-precision linear solver for FPGAs. IEEE Transactions on Computers, 57(12):1614–1623, December 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4531732.

Schulte:2000:FVP

[SS00] M. J. Schulte and E. E. Swartzlander, Jr. A family of variableprecision, interval arithmetic processors. IEEE Transactions on Computers, 49(5): 387–397, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=859535;   http:
//mesa.ece.wisc.edu/publications/
cp_2000-09.pdf.

Slingerland:2002:MPM

[SS02] N. Slingerland and A. J. Smith. Measuring the performance of multimedia instruction sets. IEEE Transactions on Computers, 51(11):1317– 1332, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047756.

Song:2006:AAM

[SS06] Meongchul Song and S. Sahni.

Approximation algorithms for multiconstrained quality-of-service routing. IEEE Transactions on Computers, 55(5): 603–617, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1613840.

Stephenson:2009:QSM

[SS09a] B. Stephenson and B. Sikdar. A quasi-species model for the propagation and containment of polymorphic worms. IEEE Transactions on Computers, 58(9):1289–1296, September 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815220.

Stocco:2009:SOL

[SS09b] L. J. Stocco and G. Schrack.

On spatial orders and location codes. IEEE Transactions on Computers, 58(3): 424–432, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4626948.

Satoh:2009:HPH

[SSA09] A. Satoh, T. Sugawara, and T. Aoki. High-performance hardware architectures for Galois counter mode. IEEE Transactions on Computers, 58(7): 917–930, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4711043.

Sair:2003:DPD

[SSC03] S. Sair, T. Sherwood, and B. Calder. A decoupled predictor-directed stream prefetching architecture. IEEE Transactions on Computers, 52(3): 260–276, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1183943.

Sarkar:2006:DFA

[SSCG06] J. Sarkar, S. Sengupta, M. Chatterjee, and S. Ganguly. Differential FEC and ARQ for radio link protocols. IEEE Transactions on Computers, 55(11):1458–1472, November 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1705454.

Sunar:2003:CCF

[SSK03] B. Sunar, E. Savas, and C. K.

Koc. Constructing composite field representations for efficient conversion. IEEE Transactions on Computers, 52(11):1391– 1398, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244937.

Satran:2005:OIC

[SSS05] J. Satran, D. Sheinwald, and I. Shimony. Out of order incremental CRC computation. IEEE Transactions on Computers, 54(9):1178–1181, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471677.

Schmidt-Samoa:2006:AFW

[SSST06] K. Schmidt-Samoa, O. Semay, and T. Takagi. Analysis of fractional window recoding methods and their application to elliptic curve cryptosystems. IEEE Transactions on Computers, 55(1):48– 57, January 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545750.

Schwarz:2005:FID

[SST05] E. M. Schwarz, M. Schmookler, and S. D. Trong. FPU implementations with denormalized numbers. IEEE Transactions on Computers, 54(7): 825–836, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1432666.

Satoh:2003:SDF

[ST03] A. Satoh and K. Takano. A scalable dual-field elliptic curve cryptographic processor. IEEE Transactions on Computers, 52 (4):449–460, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1190586.

Sarkar:2004:FBA

[ST04] S. Sarkar and L. Tassiulas. Fair bandwidth allocation for multicasting in networks with discrete feasible set. IEEE Transactions on Computers, 53 (7):785–797, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1321041.

Shukla:2008:GEI

[ST08] Sandeep K. Shukla and Jean-Pierre Talpin. Guest Editors’ introduction to the special section on programming architectures for embedded systems. IEEE Transactions on Computers, 57(10):1297– 1299, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4607767.

Steininger:2002:IEC

[Ste02] A. Steininger. Identifying efficient combinations of error detection mechanisms based on results of fault injection experiments. IEEE Transactions on Computers, 51(2): 235–239, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980011.

Stoller:2000:LEA

[Sto00] S. D. Stoller. Leader election in asynchronous distributed systems. IEEE Transactions on Computers, 49(3): 283–284, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841132.

Sue:2009:EUF

[Sue09] Chuan-Ching Sue. An enhanced universal N × N fully nonblocking quantum switch. IEEE Transactions on Computers, 58(2):238–250, February 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4620106.

Sumanasena:2008:SFC

[Sum08] M. G. B. Sumanasena. A scale factor correction scheme for the CORDIC algorithm. IEEE Transactions on Computers, 57(8):1148–1152, August 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4468700.

Sunar:2004:GMC

[Sun04] B. Sunar. A generalized method for constructing subquadratic complexity GF(2^{k}) multipliers. IEEE Transactions on Computers, 53(9):1097–1105, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1315604.

Sunar:2005:EBC

[Sun05] B. Sunar. An efficient basis conversion algorithm for composite fields with given representations. IEEE Transactions on Computers, 54(8): 992–997, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453500.

Sterpone:2006:NRO

[SV06] L. Sterpone and M. Violante.

A new reliability-oriented place and route algorithm for SRAMbased FPGAs. IEEE Transactions on Computers, 55(6): 732–744, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1628960.

Sheng:2000:NTL

[SW00] Li Sheng and Jie Wu. A note on “A tight lower bound on the number of channels required for deadlock-free wormhole routing”. IEEE Transactions on Computers, 49(9): 1005, September 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869335.

Sassone:2005:SAC

[SW05] P. G. Sassone and D. S. Wills.

Scaling up the Atlas chipmultiprocessor. IEEE Transactions on Computers, 54(1): 82–87, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1362642.

Shyu:2000:GUS

[SWCC00] Michael Shyu, Guang-Ming Wu, Yu-Dong Chang, and Yao-Wen Chang. Generic universal switch blocks. IEEE Transactions on Computers, 49(4): 348–359, April 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=844347. See

comment [FWC02].

Steiger:2004:OSR

[SWP04] C. Steiger, H. Walder, and M. Platzner. Operating systems for reconfigurable embedded platforms: online scheduling of real-time tasks. IEEE Transactions on Computers, 53(11):1393–1407, November 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336761.

Shen:2009:CRW

[SX09] Hong Shen and Shihong Xu. Coordinated en-route Web caching in multiserver networks. IEEE Transactions on Computers, 58(5):605–619, May 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4620107.

Song:2004:IDL

[SXWL04] Yonghong Song, Rong Xu, Cheng Wang, and Zhiyuan Li. Improving data locality by array contraction. IEEE Transactions on Computers, 53(9):1073–1084, September 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1315602.

Shao:2006:SPC

[SXZ^{+}06] Zili Shao, C. Xue, Q. Zhuge, M. Qiu, Bin Xiao, and E. H.- M. Sha. Security protection and checking for embedded system integration against buffer overflow attacks via hardware/software. IEEE Transactions on Computers, 55(4): 443–453, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1608006.

Sun:2005:CIA

[SZ05] X. Sun and Y. Q. Zhao.

An on-chip IP address lookup algorithm. IEEE Transactions on Computers, 54(7): 873–885, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432670.

Sankar:2009:SBO

[SZGS09] S. Sankar, Yan Zhang, S. Gurumurthi, and M. R. Stan. Sensitivity-based optimization of disk architecture. IEEE Transactions on Computers, 58 (1):69–81, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4599570.

Subrata:2008:CGF

[SZL08] R. Subrata, A. Y. Zomaya, and B. Landfeldt. A cooperative game framework for QoS guided job allocation schemes in grids. IEEE Transactions on Computers, 57(10):1413– 1422, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4515857.

Sun:2009:LAL

[SZZ09] Hongbin Sun, Nanning Zheng, and Tong Zhang. Leveraging access locality for the efficient use of multibit error-correcting codes in L2 cache. IEEE Transactions on Computers, 58(10): 1297–1306, October 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4815208.

Tseng:2005:SCS

[TA05] J. H. Tseng and K. Asanovic.

A speculative control scheme for an energy-efficient banked register file. IEEE Transactions on Computers, 54(6): 741–751, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1424450.

Tallini:2005:BCU

[Tal05] L. G. Tallini. Bounds on the capacity of the unidirectional channels. IEEE Transactions on Computers, 54(2): 232–235, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377161.

Tallini:2003:TTA

[TB03] L. G. Tallini and B. Bose.

Transmission time analysis for the parallel asynchronous communication scheme. IEEE Transactions on Computers, 52 (5):558–571, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197123.

Thomasian:2006:MDO

[TB06] A. Thomasian and M. Blaum.

Mirrored disk organization reliability analysis. IEEE Transactions on Computers, 55(12):

1640–1644, December 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1717394.

Tang:2002:CRW

[TC02] Xueyan Tang and S. T. Chanson. Coordinated en-route Web caching. IEEE Transactions on Computers, 51(6): 595–607, June 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1009146.

Taher:2009:VCM

[TEG09] M. Taher and T. El-Ghazawi. Virtual configuration management: a technique for partial runtime reconfiguration. IEEE Transactions on Computers, 58(10):1398–1410, October 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5089315.

Touloupis:2007:SES

[TFCW07] E. Touloupis, J. A. Flint, V. A. Chouliaras, and D. D. Ward. Study of the effects of SEU-induced faults on a pipeline protected microprocessor. IEEE Transactions on Computers, 56(12):1585– 1596, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358223.

Thomasian:2007:PTD

[TFH07] A. Thomasian, G. Fu, and C. Han. Performance of two-disk failure-tolerant disk arrays. IEEE Transactions on Computers, 56(6):799–814, June 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4167790.

Turmon:2003:TTH

[TGKL03] M. Turmon, R. Granat, D. S.

Katz, and J. Z. Lou. Tests and tolerances for high-performance software-implemented fault detection. IEEE Transactions on Computers, 52(5): 579–591, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197125.

Tsai:2008:ERT

[THC^{+}08] Cheng-Han Tsai, Tai-Yi Huang, E. T.-H. Chu, Chun-Hang Wei, and Yu-Che Tsai. An efficient real-time disk-scheduling framework with adaptive quality guarantee. IEEE Transactions on Computers, 57(5): 634–647, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358284.

Thibeault:2000:AVA

[Thi00] C. Thibeault. On the adaptation of Viterbi algorithm for diagnosis of multiple bridging faults. IEEE Transactions on Computers, 49(6): 575–587, June 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=862217.

Tyan:2003:MMM

[THW03] Hung-Ying Tyan, J. C. Hou, and Bin Wang. Many-to-many multicast routing with temporal quality of service guarantees. IEEE Transactions on Computers, 52(6): 826–832, June 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1204837.

Tyan:2001:STQ

[THWH01] Hung-Ying Tyan, J. C. Hou, Bin Wang, and Ching-Chih Han. On supporting temporal quality of service in WDMA-based star-coupled optical networks. IEEE Transactions on Computers, 50(3): 197–214, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=910812.

Torres:2009:SBD

[TIVYL09] E. Torres, P. Ibanez, V. Vinals-Yufera, and J. M. Llaberia. Store buffer design for multibanked data caches. IEEE Transactions on Computers, 58 (10):1307–1320, October 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4815214.

Talla:2003:BMP

[TJB03] D. Talla, L. K. John, and D. Burger. Bottlenecks in multimedia processing with SIMD style extensions and architectural enhancements. IEEE Transactions on Computers, 52(8):1015–1031, August 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223637.

Takagi:2000:VAC

[TK00] N. Takagi and S. Kuwahara.

A VLSI algorithm for computing the Euclidean norm of a 3D vector. IEEE Transactions on Computers, 49(10): 1074–1082, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888043.

Tenca:2003:SAM

[TK03] A. F. Tenca and C. K. Koc. A scalable architecture for modular multiplication based on Montgomery’s algorithm. IEEE Transactions on Computers, 52(9):1215–1221, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1228516.

Tang:2007:UBB

[TK07] K. W. Tang and R. Kamoua. An upper bound for the bisection width of a diagonal mesh. IEEE Transactions on Computers, 56(3): 429–431, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079523.

Tsai:2002:MWL

[TL02] Hsien-Ming Tsai and Yi-Bing Lin. Modeling wireless local loop with general call holding times and finite number of subscribers. IEEE Transactions on Computers, 51(7): 775–786, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1017697.

Tan:2009:LPM

[TLS09] D. Tan, C. E. Lemonds, and Michael J. Schulte. Lowpower multiple-precision iterative floating-point multiplier with SIMD support. IEEE Transactions on Computers, 58(2):175–187, February 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4663060.

Tekumalla:2000:RPD

[TM00] R. C. Tekumalla and P. R.

Menon. On redundant path delay faults in synchronous sequential circuits. IEEE Transactions on Computers, 49(3): 277–282, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841131.

Teifel:2004:ADF

[TM04] J. Teifel and R. Manohar. An asynchronous dataflow FPGA architecture. IEEE Transactions on Computers, 53(11): 1376–1392, November 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1336760.

Tao:2005:PML

[TM05] T. Tao and Amar Mukherjee.

Pattern matching in LZW compressed files. IEEE Transactions on Computers, 54(8): 929–938, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453495.

Teslenko:2005:BSP

[TMD05] M. Teslenko, A. Martinelli, and E. Dubrova. Boundset preserving ROBDD variable orderings may not be optimum. IEEE Transactions on Computers, 54(2): 236–237, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1377162.

Takagi:2000:DIT

[TN00] N. Takagi and K. Nakashima. Discrete interval truth values logic and its application. IEEE Transactions on Computers, 49 (3):219–229, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841126.

Tseng:2003:AAR

[TNS03] Yu-Chee Tseng, Sze-Yao Ni, and En-Yu Shih. Adaptive approaches to relieving broadcast storms in a wireless multihop mobile ad hoc network. IEEE Transactions on Computers, 52 (5):545–557, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197122.

Touati:2007:PRN

[Tou07] S.-A.-A. Touati. On periodic register need in software pipelining. IEEE Transactions on Computers, 56(11):1493– 1504, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336298.

Tembe:2002:LRD

[TP02a] W. Tembe and S. Pande. Loop restructuring for data I/O minimization on limited onchip memory embedded processors. IEEE Transactions on Computers, 51(10):1269– 1280, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039852.

Tseng:2002:DHC

[TP02b] Yu-Chee Tseng and Hsiang-Kuang Pan. Data hiding in 2-color images. IEEE Transactions on Computers, 51(7): 873–878, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017706.

Teng:2004:SCM

[TP04] X. Teng and H. Pham. Software cost model for quantifying the gain with considerations of random field environments. IEEE Transactions on Computers, 53 (3):380–384, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261844.

Tripakis:2008:ISM

[TPB^{+}08] S. Tripakis, C. Pinello, A. Benveniste, A. Sangiovanni-Vincent, P. Caspi, and M. Di Natale. Implementing synchronous models on loosely time triggered architectures. IEEE Transactions on Computers, 57(10): 1300–1314, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4527205.

Tenca:2006:CSR

[TPT06] A. F. Tenca, S. Park, and L. A. Tawalbeh. Carry-save representation is shift-unsafe: the problem and its solution. IEEE Transactions on Computers, 55 (5):630–635, May 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1613842.

Tongsima:2000:PLS

[TSC^{+}00] S. Tongsima, E. H.-M. Sha, C. Chantrapornchai, D. R. Surma, and N. L. Passos. Probabilistic loop scheduling for applications with uncertain execution time. IEEE Transactions on Computers, 49(1): 65–80, January 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=822565.

Tyson:2001:EUR

[TSD01] G. S. Tyson, M. Smelyanskiy, and E. S. Davidson. Evaluating the use of register queues in software pipelined loops. IEEE Transactions on Computers, 50 (8):769–783, August 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=946998;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=947006.

Tomassini:2000:GHQ

[TSP00] M. Tomassini, M. Sipper, and M. Perrenoud. On the generation of high-quality random numbers by two-dimensional cellular automata. IEEE Transactions on Computers, 49(10): 1146–1151, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888056.

Thuresson:2008:MLC

[TSS08] M. Thuresson, L. Spracklen, and P. Stenstrom. Memory-link compression schemes: a value locality perspective. IEEE Transactions on Computers, 57 (7):916–927, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4453820.

Tsuda:2000:FTP

[Tsu00] N. Tsuda. Fault-tolerant processor arrays using additional bypass linking allocated by Graph-Node coloring. IEEE Transactions on Computers, 49 (5):431–442, May 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=859538.

Tai:2002:LCE

[TTA^{+}02] A. T. Tai, K. S. Tso, L. Alkalai, S. N. Chau, and W. H. Sanders. Low-cost error containment and recovery for onboard guarded software upgrading and beyond. IEEE Transactions on Computers, 51(2): 121–137, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980004.

Tolentino:2009:MMI

[TTC09] M. E. Tolentino, J. Turner, and K. W. Cameron. Memory MISER: Improving main memory energy efficiency in servers. IEEE Transactions on Computers, 58(3):336–350, March 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4633346.

Taha:2008:ITM

[TW08] T. M. Taha and D. Scott Wills. An instruction throughput model of superscalar processors. IEEE Transactions on Computers, 57(3):389–403, March 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358262.

Takagi:2001:FAM

[TYT01] N. Takagi, J. Yoshiki, and K. Takagi. A fast algorithm for multiplicative inversion in GF(2^{m}) using normal basis. IEEE Transactions on Computers, 50(5): 394–398, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=926155.

Tzeng:2002:CSF

[Tze02a]  Wen Guey Tzeng.  Cor-

rections to “A secure faulttolerant conference-key agreement protocol”. IEEE Transactions on Computers, 51(12): 1462, December 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1146714. See

[Tze02b].

Tzeng:2002:SFT

[Tze02b]  Wen-Guey Tzeng.  A se-

cure fault-tolerant conferencekey agreement protocol. IEEE Transactions on Computers, 51 (4):373–379, April 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=995447. See

corrections [Tze02a].

Tzeng:2004:EOT

[Tze04] Wen-Guey Tzeng. Efficient 1-out-of-n oblivious transfer schemes with universally usable parameters. IEEE Transactions on Computers, 53(2): 232–240, February 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1261831.

Upadhyaya:2003:GES

[UB03] S. J. Upadhyaya and A. Bondavalli. Guest editorial: special issue on reliable distributed systems. IEEE Transactions on Computers, 52(2): 97–98, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176978.

Uyar:2008:AMC

[UBWF08] M. U. Uyar, S. S. Batth, Yu Wang, and M. A. Fecko. Algorithms for modeling a class of single timing faults in communication protocols. IEEE Transactions on Computers, 57(2):274–288, February 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358229.

Umanesan:2003:CRM

[UF03] G. Umanesan and E. Fujiwara.

A class of random multiple bits in a byte error correcting and single byte error detecting (S_{t}b/EC − S_{b}ED) codes. IEEE Transactions on Computers, 52 (7):835–847, July 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214333.

Umanesan:2005:PDC

[UF05] G. Umanesan and E. Fujiwara.

Parallel decoding cyclic burst error correcting codes. IEEE Transactions on Computers, 54 (1):87–92, January 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1362643.

Uht:2005:UPE

[Uht05]  A. K. Uht.  Uniproces-

sor performance enhancement through adaptive clock frequency control. IEEE Transactions on Computers, 54(2): 132–140, February 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1377152.

Um:2001:OAC

[UK01] Junhyung Um and Taewhan Kim. An optimal allocation of carry-save-adders in arithmetic circuits. IEEE Transactions on Computers, 50(3): 215–233, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=910813.

Venkatesan:2007:ZPS

[VAZSR07] Ravi K. Venkatesan, Ahmed S. Al-Zawawi, Krishnan Sivasubramanian, and Eric Rotenberg. ZettaRAM: a powerscalable DRAM alternative through charge-voltage decoupling. IEEE Transactions on Computers, 56(2):147–160, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042676.

Verissimo:2002:TCB

[VC02] P. Verissimo and A. Casimiro.

The timely computing base model and architecture. IEEE Transactions on Computers, 51 (8):916–930, August 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024739.

Vandierendonck:2005:XBH

[VD05] H. Vandierendonck and K. De Bosschere. XOR-based hash functions. IEEE Transactions on Computers, 54(7): 800–812, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432664.

vandeGoor:2003:SMM

[vdGT03] A. J. van de Goor and I. B. S.

Tlili. A systematic method for modifying march tests for bit-oriented memories into tests for word-oriented memories. IEEE Transactions on Computers, 52(10):1320– 1331, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234529.

Vergos:2002:DOM

[VEN02] H. T. Vergos, C. Efstathiou, and D. Nikolos. Diminishedone modulo 2^{n} + 1 adder design. IEEE Transactions on Computers, 51(12):1389– 1399, December 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1146705.

Vijaykrishnan:2003:EIH

[VKI^{+}03] N. Vijaykrishnan, M. Kandemir, M. J. Irwin, Hyun Suk Kim, Wu Ye, and D. Duarte. Evaluating integrated hardwaresoftware optimizations using a unified energy estimation framework. IEEE Transactions on Computers, 52(1):59– 76, January 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1159754.

Villalba:2006:DRM

[VLG06] J. Villalba, T. Lang, and M. A. Gonzalez. Doubleresidue modular range reduction for floating-point hardware implementations. IEEE Transactions on Computers, 55(3): 254–267, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1583556.

Vitullo:2008:LCL

[VLP^{+}08] F. Vitullo, N. E. L’Insalata, E. Petri, S. Saponara, L. Fanucci, M. Casula, R. Locatelli, and M. Coppola. Lowcomplexity link microarchitecture for mesochronous communication in networks-on-chip. IEEE Transactions on Computers, 57(9):1196–1201, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4527123.

Vergos:2003:DRA

[VNBE03] H. T. Vergos, D. Nikolos, M. Bellos, and C. Efstathiou. Deterministic BIST for RNS adders. IEEE Transactions on Computers, 52(7):896–906, July 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1214338.

VanMeter:2007:CLD

[VNM07] R. Van Meter, K. Nemoto, and W. J. Munro. Communication links for distributed quantum computation. IEEE Transactions on Computers, 56(12):1643–1653, December 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358232.

Vankamamidi:2008:SMQ

[VOL08] V. Vankamamidi, M. Ottavi, and F. Lombardi. A serial memory by quantum-dot cellular automata (QCA). IEEE Transactions on Computers, 57 (5):606–618, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358292.

Vuillaume:2008:SMS

[VOT08] C. Vuillaume, K. Okeya, and T. Takagi. Short-memory scalar multiplication for Koblitz curves. IEEE Transactions on Computers, 57(4):481–489, April 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358286.

Voyiatzis:2005:TVE

[Voy05]  I. Voyiatzis.  Test vector

embedding into accumulatorgenerated sequences: a lineartime solution. IEEE Transactions on Computers, 54(4): 476–484, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1401866.

Voyiatzis:2008:ABS

[Voy08]  I. Voyiatzis.  An ALU-

based BIST scheme for wordorganized RAMs. IEEE Transactions on Computers, 57(5): 577–590, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378355.

Voyiatzis:2008:IVM

[VPG^{+}08] I. Voyiatzis, A. Paschalis, D. Gizopoulos, C. Halatsis, F. S. Makri, and M. Hatzimihail. An input vector monitoring concurrent BIST architecture based on a precomputed test set. IEEE Transactions on Computers, 57(8): 1012–1022, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4479445.

Varma:2004:CMU

[VPR04] P. Varma, B. S. Panwar, and K. N. Ramganesh. Cutting metastability using aperture transformation. IEEE Transactions on Computers, 53(9): 1200–1204, September 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1315612.

Van:2009:PEP

[VT09] Lan-Da Van and Jin-Hao Tu. Power-efficient pipelined reconfigurable fixed-width Baugh– Wooley multipliers. IEEE Transactions on Computers, 58 (10):1346–1355, October 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5156495.

Vasudevan:2007:AVA

[VVSA07] S. Vasudevan, V. Viswanath, R. W. Sumners, and J. A. Abraham. Automatic verification of arithmetic circuits in RTL using stepwise refinement of term rewriting systems. IEEE Transactions on Computers, 56(10): 1401–1414, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4302711.

Volkmer:2005:TPM

[VW05] M. Volkmer and S. Wallner. Tree parity machine rekeying architectures. IEEE Transactions on Computers, 54(4): 421–427, April 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1401861.

Vassiliadis:2004:MPP

[VWG^{+}04] S. Vassiliadis, S. Wong, G. Gaydadjiev, K. Bertels, G. Kuzmanov, and E. M. Panainte. The MOLEN polymorphic processor. IEEE Transactions on Computers, 53(11):1363– 1375, November 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1336759.

Wang:2002:SGP

[WAB^{+}02] Yun Wang, E. Anceaume, F. Brasileiro, F. Greve, and M. Hurfin. Solving the group priority inversion problem in a timed asynchronous system. IEEE Transactions on Computers, 51(8):900–915, August 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1024738.

Wang:2003:RMP

[Wan03] Dajin Wang. A rectilinearmonotone polygonal fault block model for fault-tolerant minimal routing in mesh. IEEE Transactions on Computers, 52 (3):310–320, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183946.

Wang:2004:AWA

[Wan04]  Shiuh-Jeng Wang.  Anony-

mous wireless authentication on a portable cellular mobile system. IEEE Transactions on Computers, 53(10): 1317–1329, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327581.

Wood:2008:MDD

[WAU^{+}08] S. K. Wood, D. H. Akehurst, O. Uzenkov, W. G. J. Howells, and K. D. McDonald-Maier. A model-driven development approach to mapping UML state diagrams to synthesizable VHDL. IEEE Transactions on Computers, 57(10): 1357–1371, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4585363.

Weiss:2003:CBD

[WB03] S. Weiss and S. Beren. Classbased decompressor design for compressed instruction memory in embedded processors. IEEE Transactions on Computers, 52(11):1495– 1500, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1244947.

Wang:2008:XBE

[WBW08] Seongmoon Wang, K. J. Balakrishnan, and Wenlong Wei. X-Block: An efficient LFSR reseeding-based method to block unknowns for temporal compactors. IEEE Transactions on Computers, 57(7): 978–989, July 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378353.

Wang:2004:CCP

[WCKD04] Z. Wang, H. Che, Mohan Kumar, and S. K. Das. CoPTUA: Consistent policy table update algorithm for TCAM without locking. IEEE Transactions on Computers, 53(12):1602– 1614, December 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1347086.

Wang:2009:TCD

[WCLK09] Xiaorui Wang, Yingming Chen, Chenyang Lu, and X. D. Koutsoukos. Towards controllable distributed real-time systems with feasible utilization control. IEEE Transactions on Computers, 58(8): 1095–1110, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4760135.

Wu:2007:DPE

[WCYR07] Weigang Wu, Jiannong Cao, Jin Yang, and M. Raynal.

Design and performance evaluation of efficient consensus protocols for mobile ad hoc networks. IEEE Transactions on Computers, 56(8): 1055–1070, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4264321.

Wu:2004:GDB

[WD04]  J. Wu and F. Dai.  A

generic distributed broadcast scheme in ad hoc wireless networks. IEEE Transactions on Computers, 53(10):1343– 1354, October 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1327583.

Wu:2003:EFT

[WDL^{+}03] Jie Wu, Fei Dai, Xiaola Lin, Jiannong Cao, and Weijia Jia. An extended fault-tolerant linkstate routing protocol in the Internet. IEEE Transactions on Computers, 52(10):1298– 1311, October 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234527.

Wu:2008:ILS

[WDY08] Jie Wu, Fei Dai, and Shuhui Yang. Iterative local solutions for connected dominating set in ad hoc wireless networks. IEEE Transactions on Computers, 57 (5):702–715, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4445662.

Wu:2000:OMB

[WFMSW00] Yu-Liang Wu, Hongbing Fan, M. Marek-Sadowska, and C. K. Wong. OBDD minimization based on two-level representation of Boolean functions. IEEE Transactions on Computers, 49(12):1371– 1379, December 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=895868.

Wang:2008:MRT

[WFP08] Xian Wang, Pingzhi Fan, and Yi Pan. A more realistic thinning scheme for call admission control in multimedia wireless networks. IEEE Transactions on Computers, 57(8): 1143–1147, August 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4459309.

Wang:2000:NSA

[WG00] Chin-Liang Wang and Jyh-Huei Guo. New systolic arrays for C + AB^{2}, inversion, and division in GF(2^{m}). IEEE Transactions on Computers, 49(10): 1120–1125, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=888047.

Wu:2008:SAC

[WGZ^{+}08] Qishi Wu, Jinzhu Gao, Mengxia Zhu, N. S. V. Rao, Jian Huang, and S. S. Iyengar. Selfadaptive configuration of visualization pipeline over widearea networks. IEEE Transactions on Computers, 57(1): 55–68, January 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358234.

Wang:2002:EUF

[WH02] Farn Wang and Pao-Ann Hsiung. Efficient and user-friendly verification. IEEE Transactions on Computers, 51(1):61– 83, January 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=980017.

Wang:2003:NRW

[WH03] Jun Wang and Yiming Hu. A novel reordering write buffer to improve write performance of log-structured file systems. IEEE Transactions on Computers, 52(12):1559–1572, December 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252852.

Wu:2006:NSB

[WH06] Qingwei Wu and M. S.

Hsiao. A new simulationbased property checking algorithm based on partitioned alternative search space traversal. IEEE Transactions on Computers, 55(11):1325–1334, November 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705442.

Wu:2002:FFM

[WHBG02] Huapeng Wu, M. A. Hasan, I. F. Blake, and Shuhong Gao. Finite field multiplier using redundant representation. IEEE Transactions on Computers, 51(11):1306–1316, November 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1047755.

Wang:2009:MST

[WHT09] You-Chiun Wang, Yao-Yu Hsieh, and Yu-Chee Tseng. Multiresolution spatial and temporal coding in a wireless sensor network for long-term monitoring applications. IEEE Transactions on Computers, 58 (6):827–838, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4760136.

Wang:2009:COC

[WHZ09] Shuai Wang, Jie Hu, and S. G. Ziavras. On the characterization and optimization of onchip cache reliability against soft errors. IEEE Transactions on Computers, 58(9):1171– 1184, September 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4785458.

Wilkes:2001:HPM

[Wil01] M. V. Wilkes. High performance memory systems. IEEE Transactions on Computers, 50 (11):1105, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966488.

Wolz:2003:DPR

[WK03] F. Wolz and R. Kolla. Disproving the perfect-rate property of data-flow graphs unfolded by the least common multiple of the number of loop register. IEEE Transactions on Computers, 52(5):688, May 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
tp=&arnumber=1197137.

Wu:2006:OOA

[WK06] B. Wu and A. D. Kshemkalyani.

Objective-optimal algorithms for long-term Web prefetching. IEEE Transactions on Computers, 55(1):2–17, January 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1545747.

Wang:2007:HWO

[WKJ07] Cheng-Yeh Wang, Chin-Bin Kuo, and Jing-Yang Jou. Hybrid wordlength optimization methods of pipelined FFT processors. IEEE Transactions on Computers, 56(8): 1105–1118, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4264324.

Welch:2005:SDH

[WKS^{+}05] B. J. Welch, S. O. Kanaujia, A. Seetharam, D. Thirumalai, and A. G. Dean. Supporting demanding hard-real-time systems with STI. IEEE Transactions on Computers, 54(10): 1188–1202, October 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1501786.

Wu:2006:EMR

[WLD06] J. Wu, Wei Lou, and F. Dai. Extended multipoint relays to determine connected dominating sets in MANETs. IEEE Transactions on Computers, 55 (3):334–347, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1583562.

Wang:2009:DCM

[WLS09] Lipo Wang, Wen Liu, and Haixiang Shi. Delay-constrained multicast routing using the noisy chaotic neural networks. IEEE Transactions on Computers, 58(1):82–89, January 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4585367.

Wang:2002:UNU

[WMZH02] Jun Wang, Rui Min, Yingwu Zhu, and Yiming Hu. UCFS — a novel User-space, high performance, Customized File System for Web proxy servers. IEEE Transactions on Computers, 51(9):1056–1073, September 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032625.

Wilson:2001:HBC

[WO01] K. M. Wilson and K. Olukotun. High bandwidth on-chip cache design. IEEE Transactions on Computers, 50(4): 292–307, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=919276.

Wollinger:2005:CVH

[WPP05] T. Wollinger, J. Pelzl, and C. Paar. Cantor versus Harley: optimization and analysis of explicit formulae for hyperelliptic curve cryptosystems. IEEE Transactions on Computers, 54 (7):861–872, July 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432669.

Wu:2007:UAR

[WRJ07] Haisang Wu, B. Ravindran, and E. D. Jensen. Utility accrual real-time scheduling under the unimodal arbitrary arrival model with energy bounds. IEEE Transactions on Computers, 56(10): 1358–1371, October 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4302708.

Wu:2005:TUF

[WRJL05] Haisang Wu, B. Ravindran, E. D. Jensen, and Peng Li. Time/utility function decomposition techniques for utility accrual scheduling algorithms in real-time distributed systems. IEEE Transactions on Computers, 54(9):1138–1153, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1471674.

Wang:2001:MMO

[WS01] Chih-Fang Wang and S. Sahni. Matrix multiplication on the OTIS-Mesh optoelectronic computer. IEEE Transactions on Computers, 50(7):635–646, July 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=936231.

Wang:2009:HDD

[WSTJ09] Liang-Kai Wang, Michael J. Schulte, J. D. Thompson, and N. Jairam. Hardware designs for decimal floatingpoint addition and related operations. IEEE Transactions on Computers, 58(3): 322–335, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4599577.

Wang:2000:AFB

[WT00] San-Yuan Wang and Yu-Chee Tseng. Algebraic foundations and broadcasting algorithms for wormhole-routed allport tori. IEEE Transactions on Computers, 49(3): 246–258, March 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=841128.

Wei:2004:AIT

[WTL04] Huan-Yun Wei, Shih-Chiang Tsao, and Ying-Dar Lin. Assessing and improving TCP rate shaping over edge gateways. IEEE Transactions on Computers, 53(3):259–275, March 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261834.

Wu:2002:BPF

[Wu02a] Huapeng Wu. Bit-parallel finite field multiplier and squarer using polynomial basis. IEEE Transactions on Computers, 51 (7):750–758, July 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017695.

Wu:2002:MMS

[Wu02b] Huapeng Wu. Montgomery multiplier and squarer for a class of finite fields. IEEE Transactions on Computers, 51 (5):521–529, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004591.

Wu:2003:FTD

[Wu03]  Jie Wu.  A fault-tolerant

and deadlock-free routing protocol in 2D meshes based on odd-even turn model. IEEE Transactions on Computers, 52(9):1154–1169, September 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228511.

Wu:2008:BPP

[Wu08]  Huapeng Wu.  Bit-parallel

polynomial basis multiplier for new classes of finite fields. IEEE Transactions on Computers, 57(8):1023–1031, August 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4487062.

Wen:2006:SBF

[WWC06] C. H.-P. Wen, Li-C Wang, and Kwang-Ting Cheng. Simulationbased functional test generation for embedded processors. IEEE Transactions on Computers, 55(11):1335–1343, November 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705443.

Wang:2007:GOS

[WWH^{+}07] F. Z. Wang, S. Wu, N. Helian, M. A. Parker, Y. Guo, Y. Deng, and V. R. Khare. Grid-oriented storage: a single-image, crossdomain, high-bandwidth architecture. IEEE Transactions on Computers, 56(4): 474–487, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4118671.

Wu:2004:HSL

[WWSH04] Chien-Hsing Wu, Chien-Ming Wu, Ming-Der Shieh, and Yin-Tsung Hwang. High-speed, lowcomplexity systolic designs of novel iterative division algorithms in GF(2^{m}). IEEE Transactions on Computers, 53(3): 375–380, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261843.

Wei:2006:EPC

[WX06] Jianbin Wei and Cheng-Zhong Xu. eQoS: Provisioning of client-perceived end-to-end QoS guarantees in Web servers. IEEE Transactions on Computers, 55(12):1543– 1556, December 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1717387.

Wang:2009:NHD

[WYMG09] Jun Wang, Xiaoyu Yao, C. Mitchell, and Peng Gu. A new hierarchical data cache architecture for iSCSI storage server. IEEE Transactions on Computers, 58(4): 433–447, April 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4624247.

Wang:2006:DHB

[WYYZ06] J. Wang, M. Yang, B. Yang, and S. Q. Zheng. Dual-homing based scalable partia multicast protection. IEEE Transactions on Computers, 55(9):1130– 1141, September 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1668041.

Wang:2008:EMD

[WYZ08] Jun Wang, Xiaoyu Yao, and Huijun Zhu. Exploiting inmemory and on-disk redundancy to conserve energy in storage systems. IEEE Transactions on Computers, 57(6): 733–747, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4479444.

West:2007:CDW

[WZ07] R. West and Y. Zhang. Comments on “Dynamic Window-Constrained Scheduling of Real-Time Streams in Media Servers”. IEEE Transactions on Computers, 56(5): 718–719, May 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4141245. See

[WZSP04].

Wang:2008:ECE

[WZL08] Jun Wang, Huijun Zhu, and Dong Li. eRAID: Conserving energy in conventional diskbased RAID system. IEEE Transactions on Computers, 57 (3):359–374, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358280.

West:2004:DWC

[WZSP04] R. West, Y. Zhang, K. Schwan, and C. Poellabauer. Dynamic window-constrained scheduling of real-time streams in media servers. IEEE Transactions on Computers, 53(6): 744–759, June 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1288549. See

comments [WZ07].

Wei:2005:RPR

[WZX05] J. Wei, X. Zhou, and C.-Z. Xu. Robust processing rate allocation for proportional slowdown differentiation on Internet servers. IEEE Transactions on Computers, 54(8): 964–977, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1453498.

Xiang:2007:UWS

[XCF07] Dong Xiang, Mingjing Chen, and H. Fujiwara. Using weighted scan enable signals to improve test effectiveness of scan-based BIST. IEEE Transactions on Computers, 56(12):1619–1628, December 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358224.

Xiong:2008:DSM

[XHLC08] Ming Xiong, Song Han, Kam-Yiu Lam, and Deji Chen. Deferrable scheduling for maintaining real-time data freshness: Algorithms, analysis, and results. IEEE Transactions on Computers, 57(7):

952–964, July 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4433987.

Xie:2008:SSB

[Xie08] Tao Xie. SEA: a striping-based energy-aware strategy for data placement in RAID-structured storage systems. IEEE Transactions on Computers, 57(6): 748–761, June 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4447660.

Xu:2003:SAW

[XL03] Jun Xu and Wooyong Lee. Sustaining availability of Web services under distributed denial of service attacks. IEEE Transactions on Computers, 52(2): 195–208, February 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176986.

Xiao:2005:IUP

[XLN05] L. Xiao, Yunhao Liu, and L. M. Ni. Improving unstructured peer-to-peer systems by adaptive connection establishment. IEEE Transactions on Computers, 54(9):1091–1103, September 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1471671.

Xiang:2007:RSF

[XLSF07] Dong Xiang, Kaiwei Li, Jiaguang Sun, and H. Fujiwara. Reconfigured scan forest for test application cost, test data volume, and test power reduction. IEEE Transactions on Computers, 56(4): 557–562, April 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118677.

Xue:2007:MQR

[XM07] G. Xue and S. K. Makki. Multiconstrained QoS routing: a norm approach. IEEE Transactions on Computers, 56(6): 859–863, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4118681.

Xu:2006:DIB

[XN06] Qiang Xu and N. Nicolici. DFT infrastructure for broadside two-pattern test of corebased SOCs. IEEE Transactions on Computers, 55(4): 470–485, April 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1608008.

Xiang:2004:PSD

[XP04] Dong Xiang and J. H. Patel.

Partial scan design based on circuit state information and functional analysis. IEEE Transactions on Computers, 53(3): 276–287, March 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1261835.

Xenoulis:2006:TAS

[XPGP06] G. Xenoulis, M. Psarakis, D. Gizopoulos, and A. Paschalis. Testability analysis and scalable test generation for highspeed floating-point units. IEEE Transactions on Computers, 55(11):1449–1457, November 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1705453.

Xie:2006:SSC

[XQ06] T. Xie and X. Qin. Scheduling security-critical real-time applications on clusters. IEEE Transactions on Computers, 55 (7):864–879, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637402.

Xie:2008:EDT

[XQ08] Tao Xie and Xiao Qin. An energy-delay tunable task allocation strategy for collaborative applications in networked embedded systems. IEEE Transactions on Computers, 57(3): 329–343, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358254.

Xiong:2004:DDP

[XR04] M. Xiong and K. Ramamritham. Deriving deadlines and periods for real-time update transactions. IEEE Transactions on Computers, 53(5): 567–583, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1275297.

Xu:2002:RDE

[XRR^{+}02] J. Xu, B. Randell, A. Romanovsky, R. J. Stroud, A. F. Zorzo, E. Canver, and F. von Henke. Rigorous development of an embedded fault-tolerant system based on coordinated atomic actions. IEEE Transactions on Computers, 51(2): 164–179, February 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980006.

Xu:2002:CEF

[XS02] Jun Xu and M. Singhal. Costeffective flow table designs for high-speed routers: architecture and performance evaluation. IEEE Transactions on Computers, 51(9):1089–1099, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1032627.

Xu:2007:AAS

[XS07] Xiaochun Xu and S. Sahni.

Approximation algorithms for sensor deployment. IEEE Transactions on Computers, 56(12):1681–1695, December 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4358215.

Xu:2004:ACP

[XSMH04] Z. Xu, S. Sohoni, R. Min, and Y. Hu. An analysis of cache performance of multimedia applications. IEEE Transactions on Computers, 53(1): 20–38, January 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1255788.

Xue:2002:CSN

[XT02] Guoliang Xue and K. Thulasiraman. Computing the shortest network under a fixed topology. IEEE Transactions on Computers, 51(9):1117–1120, September 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032631;   http:
                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1032632.

See comments [Zac06].

Xiao:2006:QRC

[XTX06] Y. Xiao, Krishnaiyan Thulasiraman, and G. Xue. QoS routing in communication networks: approximation algorithms based on the primal simplex method of linear programming. IEEE Transactions on Computers, 55(7): 815–829, July 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1637398.

Xue:2004:EAA

[XV04] J. Xue and X. Vera. Efficient and accurate analytical modeling of whole-program data cache behavior. IEEE Transactions on Computers, 53 (5):547–566, May 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1275296.

Xie:2008:MMC

[XW08] Tao Xie and Hui Wang. MI-CRO: a multilevel cachingbased reconstruction optimization for mobile storage systems. IEEE Transactions on Computers, 57(10):1386– 1398, October 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4509422.

Xiao:2008:CLD

[XWC^{+}08] Wenjun Xiao, Wenhong Wei, Weidong Chen, Mingxin He, and B. Parhami. Comments on “Low Diameter Interconnections for Routing in High-Performance Parallel Systems,” with connections and extensions to arc coloring of coset graphs. IEEE Transactions on Computers, 57(12):1726– 1728, December 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4624245. See

[Mel07].

Xiang:2003:NDT

[XXF03] Dong Xiang, Yi Xu, and H. Fujiwara. Nonscan design for testability for synchronous sequential circuits based on conflict resolution. IEEE Transactions on Computers, 52(8): 1063–1075, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1223640.

Xiao:2009:DAB

[XYR^{+}09] Weijun Xiao, Qing Yang, Jin Ren, Changsheng Xie, and Huaiyang Li. Design and analysis of block-level snapshots for data protection and recovery. IEEE Transactions on Computers, 58(12):1615– 1625, December 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=5184809.

Xiang:2009:PDF

[XZP09] Dong Xiang, Yueli Zhang, and Yi Pan. Practical deadlockfree fault-tolerant routing in meshes based on the planar network fault model. IEEE Transactions on Computers, 58 (5):620–633, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4674344.

Yang:2002:NCN

[Yan02] Yuanyuan Yang. A new conference network for group communication. IEEE Transactions on Computers, 51(9):995– 1010, September 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032621.

Yang:2008:EEA

[Yan08] Ching-Nung Yang. Efficient encoding algorithm for secondorder spectral-null codes using cyclic bit shift. IEEE Transactions on Computers, 57 (7):876–888, July 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4407684.

Youn:2003:EED

[YB03] Jong-Hoon Youn and B. Bose.

Efficient encoding and decoding schemes for balanced codes. IEEE Transactions on Computers, 52(9):1229–1232, September 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1228518.

Yin:2000:DAS

[YBB00] Meng-Lai Yin, D. M. Blough, and L. Bic. A dependability analysis for systems with global spares. IEEE Transactions on Computers, 49(9): 958–963, September 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869326.

Yen:2001:TSO

[YCL01] Mao-Hsu Yen, Sao-Jie Chen, and S. H. Lan. A threestage one-sided rearrangeable polygonal switching network. IEEE Transactions on Computers, 50(11):1291– 1294, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=966502.

Yoo:2002:FEP

[YD02] B. S. Yoo and C. R. Das. A fast and efficient processor allocation scheme for mesh-connected multicomputers. IEEE Transactions on Computers, 51(1): 46–60, January 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=980016. See

comments [Zha03].

Yuan:2009:FRR

[YD09] Xin Yuan and Zhenhai Duan. Fair round-robin: a low complexity packet schduler with proportional and worst-case fairness. IEEE Transactions on Computers, 58(3): 365–379, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4626953.

Ye:2002:MSA

[YECV02] N. Ye, S. M. Emran, Q. Chen, and S. Vilbert. Multivariate statistical analysis of audit trails for host-based intrusion detection. IEEE Transactions on Computers, 51(7): 810–820, July 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1017701.

Ying:2009:ALD

[YF09] Mingsheng Ying and Yuan Feng. An algebraic language for distributed quantum computing. IEEE Transactions on Computers, 58(6): 728–743, June 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4745631.

Fu:2003:MVS

[yFBC03] Chao ying Fu, J. T. Bodine, and T. M. Conte. Modeling value speculation: an optimal edge selection problem. IEEE Transactions on Computers, 52 (3):277–292, March 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1183944.

Yakovlev:2004:DAS

[YFKB04] A. Yakovlev, S. Furber, R. Krenz, and A. Bystrov. Design and analysis of a self-timed duplex communication system. IEEE Transactions on Computers, 53 (7):798–814, July 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1321042.

Yao:2008:ORR

[YGB08] Jingnan Yao, Jiani Guo, and L. N. Bhuyan. Ordered round-robin: An efficient sequence preserving packet scheduler. IEEE Transactions on Computers, 57(12):1690–1703, December 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4531731.

Yang:2007:FMA

[YGL07] Wu-Chuan Yang, D. J. Guan, and Chi Sung Laih. Fast multicomputation with asynchronous strategy. IEEE Transactions on Computers, 56(2): 234–242, February 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4042683.

Yang:2005:IME

[YGZ05] Jun Yang, Lan Gao, and Youtao Zhang. Improving memory encryption performance in secure processors. IEEE Transactions on Computers, 54(5):630–640, May 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407851.

Yamaguchi:2002:MCT

[YHIO02] T. J. Yamaguchi, Dong Sam Ha, M. Ishida, and T. Ohmi. A method for compressing test data based on Burrows– Wheeler transformation. IEEE Transactions on Computers, 51 (5):486–497, May 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1004588.

Yeh:2000:HSB

[YJ00a] Wen-Chang Yeh and Chein-Wei Jen. High-speed Booth encoded parallel multiplier design. IEEE Transactions on Computers, 49 (7):692–701, July 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=863039.

Yen:2000:CBO

[YJ00b] Sung-Ming Yen and M. Joye.

Checking before output may not be enough against faultbased cryptanalysis. IEEE Transactions on Computers, 49 (9):967–970, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=869328.

Yeh:2003:GEF

[YJ03] Wen-Chang Yeh and Chein-Wei Jen. Generalized earliestfirst fast addition algorithm. IEEE Transactions on Computers, 52(10):1233–1242, October 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1234522.

Yen:2006:OAC

[YJ06] Chia-Chih Yen and Jing-Yang Jou. An optimum algorithm for compacting error traces for efficient design error debugging. IEEE Transactions on Computers, 55(11):1356– 1366, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705445.

Yang:2005:FTD

[YJMS05] J. Yang, Q. Jiang, D. Manivannan, and Mukesh Singhal. A fault-tolerant distributed channel allocation scheme for cellular networks. IEEE Transactions on Computers, 54(5): 616–629, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1407850.

Yen:2003:RSC

[YKLM03] Sung-Ming Yen, Seungjoo Kim, Seongan Lim, and Sang-Jae Moon. RSA speedup with Chinese Remainder Theorem immune against hardware fault cryptanalysis. IEEE Transactions on Computers, 52(4): 461–472, April 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1190587.

Yang:2006:SNE

[YL06a] Ching-Nung Yang and Dong-Jing Lee. Some new efficient second-order spectral- codes with small lookup tables. IEEE Transactions on Computers, 55 (7):924–927, July 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1637408.

Yi:2006:SCA

[YL06b] J. J. Yi and D. J. Lilja.

Simulation of computer architectures: simulators, benchmarks, methodologies, and recommendations. IEEE Transactions on Computers, 55(3): 268–280, March 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1583557.

Yi:2005:ICA

[YLH05] J. J. Yi, D. J. Lilja, and D. M. Hawkins. Improving computer architecture simulation methodology by adding statistical rigor. IEEE Transactions on Computers, 54(11):1360– 1373, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1514416.

Yee:2002:EDA

[YNOJ02] Wai Gen Yee, S. B. Navathe, E. Omiecinski, and C. Jermaine. Efficient data allocation over multiple channels at broadcast servers. IEEE Transactions on Computers, 51(10): 1231–1236, October 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1039849.

Youn:2001:EAB

[YOC^{+}01] Hee Yong Youn, Choong Gun Oh, Hyunseung Choo, Jin-Wook Chung, and Dongman Lee. An efficient algorithmbased fault tolerance design using the weighted data-check relationship. IEEE Transactions on Computers, 50(4): 371–383, April 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=919281.

Yu:2008:IBT

[YRF08] Yue Yu, Shangping Ren, and Ophir Frieder. Interval-based timing constraints their satisfactions and applications. IEEE Transactions on Computers, 57 (3):418–432, March 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358252.

Yoong:2009:SAI

[YRVS09] Li Hsien Yoong, P. S. Roop, V. Vyatkin, and Z. Salcic. A synchronous approach for IEC 61499 function block implementation. IEEE Transactions on Computers, 58(12):1599– 1614, December 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=5235138.

Yu:2001:DID

[YS01] Sungwook Yu and E. E. Swartzlander, Jr. DCT implementation with distributed arithmetic. IEEE Transactions on Computers, 50(9):985–991, September 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954513.

Yan:2003:NSA

[YS03] Z. Yan and D. V. Sarwate. New systolic architectures for inversion and division in GF(2^{m}). IEEE Transactions on Computers, 52(11):1514– 1519, November 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1244950.

Yang:2000:EPG

[YSL00] Chia-Lin Yang, B. Sano, and A. R. Lebeck. Exploiting parallelism in geometry processing with general purpose processors and floating-point SIMD instructions. IEEE Transactions on Computers, 49(9):934– 946, September 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=869324.

Yi:2007:SVA

[YSLH07] J. J. Yi, R. Sendag, D. J. Lilja, and D. M. Hawkins. Speed versus accuracy tradeoffs in microarchitectural simulations. IEEE Transactions on Computers, 56(11):1549– 1563, November 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4336302.

Yang:2007:EFI

[YT07] Xiaofan Yang and Yuan Yan Tang. Efficient fault identification of diagnosable systems under the comparison model. IEEE Transactions on Computers, 56(12):1612– 1618, December 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358226.

Yang:2001:PAA

[YW01] Yuanyuan Yang and Jianchao Wang. Pipelined all-to-all broadcast in all-port meshes and tori. IEEE Transactions on Computers, 50(10):1020– 1032, October 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=956089.

Yang:2004:FTR

[YW04a] Y. Yang and J. Wang. A faulttolerant rearrangeable permutation network. IEEE Transactions on Computers, 53(4): 414–426, April 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1268399.

Yang:2004:DWO

[YW04b] Yuanyuan Yang and Jianchao Wang. Designing WDM optical interconnects with full connectivity by using limited wavelength conversion. IEEE Transactions on Computers, 53(12):1547–1556, December 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1347081.

Yen:2006:SED

[YW06] Chih-Hsu Yen and Bing-Fei Wu. Simple error detection methods for hardware implementation of Advanced Encryption Standard. IEEE Transactions on Computers, 55(6): 720–731, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1628959.

Yang:2005:PPA

[YWV05] S. Yang, W. Wolf, and N. Vijaykrishnan. Power and performance analysis of motion estimation based on hardware and software realizations. IEEE Transactions on Computers, 54 (6):714–726, June 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1461359.

Yeung:2001:DMP

[YY01] K. H. Yeung and T. S. Yum. Dynamic multiple parity (DMP) disk array for serial transaction processing. IEEE Transactions on Computers, 50(9):949–959, September 2001. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=954509.

Zachariasen:2006:CCS

[Zac06] M. Zachariasen. Comment on “Computing the shortest network under a fixed topology”. IEEE Transactions on Computers, 55(6): 783–784, June 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1628964. See

[XT02].

Zhang:2009:SAR

[ZB09] Fengxiang Zhang and A. Burns.

Schedulability analysis for real-time systems with EDF scheduling. IEEE Transactions on Computers, 58(9):1250– 1258, September 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4815215.

Zhang:2002:CTP

[ZBC02] Sijing Zhang, A. Burns, and Tee-Hiang Cheng. Cycle-time properties of the timed token medium access control protocol. IEEE Transactions on Computers, 51(11):1362– 1367, November 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1047759.

Zhao:2007:HSA

[ZBI^{+}07] Li Zhao, L. N. Bhuyan, R. Iyer, S. Makineni, and D. Newell. Hardware support for accelerating data movement in server platform. IEEE Transactions on Computers, 56(6): 740–753, June 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4167786.

Zyuban:2004:IAP

[ZBS^{+}04] V. Zyuban, D. Brooks, Viji Srinivasan, M. Gschwind, Pradip Bose, P. N. Strenski, and P. G. Emma. Integrated analysis of power and performance for pipelined microprocessors. IEEE Transactions on Computers, 53(8): 1004–1016, August 2004. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306993.

Zhou:2005:EML

[ZC05a] H. Zhou and T. M. Conte. Enhancing memory-level parallelism via recovery-free value prediction. IEEE Transactions on Computers, 54(7): 897–912, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1432672.

Zou:2005:DCC

[ZC05b] Yi Zou and K. Chakrabarty.

A distributed coverage- and connectivity-centric technique for selecting active nodes in wireless sensor networks. IEEE Transactions on Computers, 54 (8):978–991, August 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1453499.

Zheng:2006:DRT

[ZCW^{+}06] Kai Zheng, H. Che, Zhijun Wang, Bin Liu, and Xin Zhang. DPPC-RE: TCAM-based distributed parallel packet classification with range encoding. IEEE Transactions on Computers, 55(8):947–961, August 2006. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1650193.

Zhang:2008:NDS

[ZCWZ08] Liqiang Zhang, Qiang Cheng, Yingge Wang, and S. Zeadally. A novel distributed sensor positioning system using the dual of target tracking. IEEE Transactions on Computers, 57(2): 246–260, February 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358244.

Zhong:2007:MRP

[ZDS^{+}07] Y. Zhong, S. G. Dropsho, X. Shen, A. Studer, and C. Ding. Miss rate prediction across program inputs and cache configurations. IEEE Transactions on Computers, 56 (3):328–343, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079516.

Zhang:2001:IMC

[ZFP^{+}01] Lixin Zhang, Zhen Fang, M. Parker, B. K. Mathew, L. Schaelicke, J. B. Carter, W. C. Hsieh, and S. A. Mc-Kee. The impulse memory controller. IEEE Transactions on Computers, 50(11):1117– 1132, November 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=966490.

Zhang:2002:FTM

[Zha02]  Li Zhang.  Fault-tolerant

meshes with small degree. IEEE Transactions on Computers, 51(5):553–560, May 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1004594.

Zhang:2003:CFE

[Zha03]  Lu Zhang.  Comments on

“A fast and efficient processor allocation scheme for mesh-connected multicomputers”. IEEE Transactions on Computers, 52(2):255–256, February 2003. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1176990. See

[YD02].

Zhang:2005:RCS

[Zha05] W. Zhang. Replication cache:

a small fully associative cache to improve data cache reliability. IEEE Transactions on Computers, 54(12):1547– 1555, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1524936.

Zhao:2000:TSB

[ZIPL00] J. Zhao, S. Irrinki, M. Puri, and F. Lombardi. Testing SRAM-based content addressable memories. IEEE Transactions on Computers, 49(10): 1054–1063, October 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=888041.

Zhu:2008:RRA

[ZJ08] Yifeng Zhu and Hong Jiang. RACE: a robust adaptive caching strategy for buffer cache. IEEE Transactions on Computers, 57(1):25–40, January 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358240.

Zyuban:2001:ILP

[ZK01] V. V. Zyuban and P. M. Kogge. Inherently lower-power high-performance superscalar architectures. IEEE Transactions on Computers, 50(3): 268–285, March 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=910816.

Zhao:2004:SCV

[ZL04] Q. Zhao and D. J. Lilja. Static classification of value predictability using compiler hints. IEEE Transactions on Computers, 53(8):929–944, August 2004. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1306988.

Zhuang:2007:RCP

[ZL07] Xiaotong Zhuang and Hsien-Hsin S. Lee. Reducing cache pollution via dynamic data prefetch filtering. IEEE Transactions on Computers, 56(1): 18–31, January 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4016494.

Zhang:2007:SVH

[ZMM07] Shi-Xiong Zhang, Man-Wai Mak, and H. M. Meng. Speaker verification via highlevel feature based phoneticclass pronunciation modeling. IEEE Transactions on Computers, 56(9):1189–1198, September 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4288086.

Zhang:2008:EEM

[ZMY08] Zhenghao Zhang, Ming Ma, and Yuanyuan Yang. Energyefficient multihop polling in clusters of two-layered heterogeneous sensor networks. IEEE Transactions on Computers, 57(2):231–245, February 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4358231.

Zhang:2001:SDO

[ZP01] Tong Zhang and K. K. Parhi.

Systematic design of original and modified Mastrovito multipliers for general irreducible polynomials. IEEE Transactions on Computers, 50(7): 734–749, July 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=936239.

Zhuo:2008:HPD

[ZP08a] Ling Zhuo and V. K. Prasanna. High-performance designs for linear algebra operations on reconfigurable hardware. IEEE Transactions on Computers, 57 (8):1057–1071, August 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4483504.

Zhuo:2008:SHD

[ZP08b] Ling Zhuo and V. K. Prasanna. Scalable hybrid designs for linear algebra on reconfigurable computing systems. IEEE Transactions on Computers, 57(12):1661–1675, December 2008. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4528959.

Zilic:2007:SBA

[ZR07] Zeljko Zilic and Katarzyna Radecka. Scaling and better approximating quantum Fourier transform by higher radices. IEEE Transactions on Computers, 56(2):202–207, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4042680.

Zuberi:2000:DIE

[ZS00] K. M. Zuberi and K. G.

Shin. Design and implementation of efficient message scheduling for controller area network. IEEE Transactions on Computers, 49(2): 182–188, February 2000. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=833115.

Zhong:2008:CGR

[ZSS08] Ming Zhong, Kai Shen, and J. Seiferas. The convergenceguaranteed random walk and its applications in peer-to-peer networks. IEEE Transactions on Computers, 57(5): 619–633, May 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4378357.

Zhang:2007:DIB

[ZSXZ07] Guangyan Zhang, Jiwu Shu, Wei Xue, and Weimin Zheng. Design and implementation of an out-of-band virtualization system for large SANs.

IEEE Transactions on Computers, 56(12):1654–1665, December 2007. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4358222.

Zilic:2002:DMI

[ZV02] Z. Zilic and Z. G. Vranesic. A deterministic multivariate interpolation algorithm for small finite fields. IEEE Transactions on Computers, 51(9):1100– 1105, September 2002. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1032628.

Zeng:2006:DPE

[ZV06] Zeng Zeng and B. Veeravalli. Design and performance evaluation of queue-and-rate-adjustment dynamic load balancing policies for distributed networks. IEEE Transactions on Computers, 55(11):1410– 1422, November 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1705450.

Zheng:2009:DFT

[ZVT09] Qin Zheng, B. Veeravalli, and Chen-Khong Tham. On the design of fault-tolerant scheduling strategies using primarybackup approach for computational grids with low replication costs. IEEE Transactions on Computers, 58(3): 380–393, March 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4626949.

Zhao:2008:SMD

[ZW08] Dan Zhao and Yi Wang. SD-MAC: Design and synthesis of a hardware-efficient collisionfree QoS-aware MAC protocol for wireless network-on-chip. IEEE Transactions on Computers, 57(9):1230–1245, September 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4531729.

Zang:2003:BBA

[ZWST03] X. Zang, D. Wang, H. Sun, and K. S. Trivedi. A BDDbased algorithm for analysis of multistate systems with multistate components. IEEE Transactions on Computers, 52(12):1608–1618, December 2003. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1252856.

Zhong:2007:EAM

[ZX07] X. Zhong and C.-Z. Xu.

Energy-aware modeling and scheduling for dynamic voltage scaling with statistical realtime guarantee. IEEE Transactions on Computers, 56(3): 358–372, March 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=4079518.

Zheng:2002:CIR

[ZXL02] Baihua Zheng, Jianliang Xu, and D. L. Lee. Cache invalidation and replacement strategies for location-dependent data in mobile environments. IEEE Transactions on Computers, 51(10):1141–1153, October 2002. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1039841.

Zhang:2006:OSB

[ZY06] Z. Zhang and Y. Yang. Optimal scheduling in buffered WDM interconnects with limited range wavelength conversion capability. IEEE Transactions on Computers, 55(1): 71–82, January 2006. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1545752.

Zhang:2009:EDP

[ZYZ09] Zhenghao Zhang, Yuanyuan Yang, and Miao Zhao. Enhancing downlink performance in wireless networks by simultaneous multiple packet transmission. IEEE Transactions on Computers, 58(5): 706–718, May 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4657352.

Zhu:2005:PAS

[ZZ05] Q. Zhu and Y. Zhou. Poweraware storage cache management. IEEE Transactions on Computers, 54(5): 587–602, May 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1407848.

Zhang:2004:DOL

[ZZZ04] Zhao Zhang, Zhichun Zhu, and Xiaodong Zhang. Design and optimization of large size and low overhead offchip caches. IEEE Transactions on Computers, 53(7): 843–855, July 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=1321045.